A Complete Bibliography of Publications in Journal of Computational Physics: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

10 February 2022 Version 1.07

Title word cross-reference

0 − Z [PRV89]. 1 [BaB86, CS82, DD87, Eck89, Kli83, Law89, Oli80]. 2 [BHM^{+}83, Eck89, HLAD84, Hu89, Mas87, SW88]. 3 [ABS82, AHH87, BHM^{+}83, BBGN80, CS81a, Eck89, For89, FSD81, GHC^{+}86, GGR89, GRS89, HLAD84, HCH^{+}81, KE87, MD81b, OvOH89, TB89, WMM86, WCR^{+}88, WWM^{+}81]. 600 [For85]. _{4} [Ves82]. Au = λBu [MM89]. Ax = λBx [Kal84]. B [Mon85]. β [TST^{+}86]. d^{2}x/dt^{2} = f(x) [Qin88]. ẍ = f(x) [Gre84a]. E_{ν}(x) [CLM88]. exp(A_{1} + A_{2}X) [Mun81]. F_{λ}(η, x) [Bar82]. G_{λ}(η, x) [Bar82]. H [TST^{+}86]. h^{4} [DH89]. H_{0}^{1} [CBLL86]. I [Hag80]. I_{p} [SGT88]. ∫_{0}^{∞} exp[−s(ψ + y cos ψ − z sin ψ)] dψ [Col81b]. J_{0}(x) [GM88, LM87a]. J_{1}(x) [GM88]. J_{1}(x)Y_{1}(ρx) − J_{1}(ρx)Y_{1}(x) = 0 [Cur82]. J_{ν}(x) [MG89]. L [GS83, Tem80]. λx^{2}/(1 + gx^{2}) [Hau81]. LU [Ker80]. N [Bor86, Can81, LJ81, Tem88b]. ∇ · B [BB80]. ∇ · ∇x = 0 [Car89]. O(v/c) [MK82]. ω_{pe}Δt ≫ 1 [Den81]. P_{n} [Sto84]. ∂^{2} ln |s − t|/∂s∂t [Fre83b]. S [YH88b]. S_{N} [SBBY83]. SU(3) [Hoe84, KKMP86]. SU(n) [GWZ85a]. U(3) [Hoe84]. U(a, x) [SGA81]. u_{t} + uu_{x} = 0 [VSS86]. V [WOZ84]. V (a, x)

[SGA81]. W (a, ±x) [SGA81]. x^{2} + λx^{2}(1 + gx^{2}) [FMC83]. x exp(iπ/4) [Wal84]. y [KL80]. Z [Gla89].

-conjugate [Ker80]. -D [MD81b, BaB86, BBGN80, CS82, HLAD84, KE87, Kli83, Oli80, WMM86]. -dimensional [Eck89, For89, GGR89, Hu89]. -method [Sto84]. -mode-like [TST^{+}86]. -pinch [Gla89]. -states [WOZ84]. -transform [CBLL86]. -wave [LJ81].

1981 [Ano81j].

2 [ZH88]. 205 [KKMP86]. 24 [GKL^{+}87].

Abel [iKI85, Lan82]. Ablowitz [DF88]. above [AEBS86]. Absorbing [KK86, TL81, BBM88, BK88b, IBCS87]. Abstract [Ano89a, Ano89b]. Accelerated [FD82, HH85, SPG89, GRA83, Kre88, Sak81]. Accelerating [Tan86, Col81a, Fer80]. Acceleration [Rig81, Lar88, MB80]. Accessibility [Man86]. accumulated [Chi88]. Accuracy [Bal85, BC87a, BC87b, Rob85, Ske81, BGT85, For86, HRS80, Kan84, KG85, Per85, VR86, ZOWZ88]. Accurate [AB81, FMC83, HM89, RH87, Tan85, BM87a, BM85, HTE81, HEOC87, HC81b, JT86, Kor88, KW87, LBM87, WP83]. ACORN [Wik89]. Acoustic [HL85, BFH86, WMS88]. acoustics [Mos86]. Action [LO86, Hit80]. Action-variable [LO86]. actions [Hoe84]. ad [Ehr81]. adapted [IR87, SN83]. adapting [Hag80]. Adaptive [BO84, BC89, BS82, EW88, HM80b, MD88, Mez87, PVMZ87, SJ88, SOS89, BGMM89, BS83, BB86a, CW89, DO86, DD87, Duk84b, EWR88, FS86, Luc87, Pik87, SSC86, SK86, TF89, VBSS89, vDR89]. adaptive-mesh [Luc87]. adaptively [BR86]. add [Tem88b]. Adding [HN80]. adhesion [Fog84]. ADI [ADG86, CW82, FN80, JT86]. Adiabatic [LM86, Eas87, Gar83]. Adjoint [HM80a, DVW87, McD80, SP87b]. adjusting [HH83]. adsorption [Mer80]. advancing [MSB87]. advantage [Tak81]. advantages [Spe87]. advection [Cho80, Dav80b, RMS85, SB88, Smo84, SC86]. aerofoil [Lyt87]. aerofoil-the [Lyt87]. aerofoils [Boe82]. aerosol [Kai83]. aggregates [Ber83]. aggregation [Fog84]. aided [ER85]. air [RMS85, ZBP84]. airfoil [Kor88]. airfoils [CFG83]. Airy [SC81]. Alfvén [Ota88]. Algebra [EM82, DA86, Var84]. algebraic [MEK86]. algebras [Pap84]. Algorithm [Bor86, And83, ABDV81, AB85, BR80b, Boe82, BLT^{+}81, CBLL86, Cep83, CW89, CW85b, CL81b, CL83, CL84, CL87, CBG86, CP88b, CW81, Die83, EH85, EHD89, Ema81, For89, GR87, HMS89, HCH85, HK84, HR86, Ioa81a, Kal84, Kan81, Keu86, KW85, KW87, LB87, LBO^{+}89, LPS80, Law87, LD89, LBS88, Lon88, Mas87, Med86, MSW87, OW85, Pea83, PW85, PC81b, Ram85b, Ram86, RCK87, SS84a, Sad84, SHB82, Sch80a, SBC84, SGA81, SS89, Smo84, SC86, Son85, SM89, TOO83, Tem80, Tem85, Tem88c, Ter81c, TC87, Tox82, WMM86, Wol85, Yam83, YP88, Zuf89]. Algorithms [CKR87, KH87, OS88, Per81a, SW88, Ter81a, CM82, Cha81, Cho80, CG85, DQ85, GS89, Mak89, MM85a, OF82, Pik87, Pru88, RRS89, SMR89, Tem83b, Tem89, WH87, WOZ84, ZO83]. alloying [SK86]. along [LR81]. Alternative [Adh81a, Sca87]. analog [Mel86]. Analysis [BGR80, CM80, Cul83a, LM87b, MC88a, NN89, SP81, SL85, Abe89, BWY82, BER85b, CB84, CKD87, Chr87, DQS87, Dun81, ER85, For83b, For83a, For84, For86, FB88, Kam81, iKI85, LV89, MM85b, MC88b, ND89, PM80, PC81a, RS85, RV88a, RSR86, Roe86, RMS85, SS88a, Vin89, WCR^{+}88, WD84, Zuf89, LBR81]. analytic [AH88b, BFH86, CD82, CC84, Dav86, Dye84, IA85, Wol85]. Analytical [Che84, CH84, CR84, MHK86, TA84c, TA84a, TA84b, TA88, Afa87, HR86, TA84c]. angle [MLB87]. Angular [Ves82, DA86, Dun85, HG88, WP88]. anharmonic [Boo82a]. animal [FP88a]. anisotropic [GS82, GS83, WP88, WPW89]. annealing [VL84]. Announcement [Ano80a, Ano80b, Ano80c, Ano80d, Ano81a, Ano82a, Ano82b, Ano83a, Ano83b, Ano83c, Ano84a, Ano84b, Ano85a, Ano85b, Ano86a, Ano86b, Ano86c, Ano87a, Ano87b, Ano87c, Ano88a, Ano88b, Ano89c, Ano89d, Ano89e]. Announcements [Ano81b, Ano82c, Ano83d, Ano84c, Ano86d, Ano88c, Ano88d]. annular [Sto84]. anomalous [BK87]. ANTIC [Tam81]. antiquark [ADKM82]. AO [Tak81]. aperture [LM87a]. appear [Ano89a, Ano89b, Ano89t]. appearing [BJ81, Chr88]. appears [BM84]. Application [BaB86, Bon84, CY85, EM82, GP80, Hau82, Her89, KMA89, KM85, KTH88, KG87, Lab87, LBR81, Ohr83, SGT88, Sch82, She86, SC82, Sto84, TL84, WVP82, AHP81, BV82, CW86, CF82, CP88a, FZ88, GM81, GRP82, HRS80, Ioa81b, KK89, Lyn82, Nav87, RSR86, SBBY83, SW81, Tak86, TY87a, CDP80, CA85b, SM84]. Applications [GDM81, MOI81, Dar89, Hen81, HSA89, McN87, MML83, OC83, Paj89, SAR82, SC86, SZH85, Yee87, JWB88, ZBG88]. applied [AH88a, AA89, CSW85, FH81, Gol86b, Jon81, Löt82, Nav83, OG86, Rag82, RS85, Rap80b, RMS85, SH84a, SBT86, SSG81, Zuf89]. approach [BSS82b, CA85a, CA85b, Dav81, FZ87, FS87, GIT87, KR86, KSD81, KTH88, Kli83, Lyn82, MPA88, ML83b, Mru87, Mul89, Rai86b, RS81, SC85, SS82b, Str89, VJ89, VL84]. approaches [Hag81]. approximants [Hau82]. Approximate [Fuc89, Roe81, Bak81, Cat88, DS86a, Gla88a, Gla88b, Mru87, OK86, PC81b, Wil82]. approximate-factorisation [Cat88]. approximate-factorization [PC81b]. approximates [Gre85a]. Approximating [Osg81, Bog88]. Approximation [CDS85, Hal87, IO81, PA86, BK85a, But88, Chr88, DQ83, GP80, HM80a, LM87a, LO87, MJ81, Mor84, SP81, Tak80, TM88, Vat87, Vie83]. approximations [Bal85, Bar80, BWY82, Blo82, BFH86, BPD81, DH89, ER85, GM88, Hof82, KG85, MG89, Pie84, SB88, Shy85, Tre81]. aquatic [FP88a]. Arakawa [ST89a]. arbitrarily [MT86b, TMCS85]. arbitrary [CB89, DH82, DP86, Hoe83, Kam81, KOTK88, Lev88, MNM85, Rob89b, War86, WCR^{+}88]. architecture [HCM89, KS89]. area [HL85]. arguments [TB86, Wal84]. arises [Sid85]. Arising [vdV81, BST85, CG89, HR84b, Hun89, Ixa80, MPPS84, Mos86]. Arnoldi [Nav87]. arrangement [KOTK88]. array [CFH^{+}82]. Articles [Ano82s, Ano84t, Ano85i, Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano80u, Ano80v, Ano80w, Ano80x, Ano80y, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano83s, Ano83t, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano84r, Ano84s, Ano84u, Ano85h, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano85v, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano86u, Ano86v]. articles [Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87u, Ano87v, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano88p, Ano88q, Ano88r, Ano88s, Ano88t, Ano88u, Ano88v, Ano89l, Ano89m, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s]. Artificial [Law88, Noh87, Ros80, Wil80a]. aspect [GH87]. aspects [MM81a, TA84c, TA84a, TA84b, TA88]. assembly [KT86]. Assessment [GG88, SB88]. Associated [OS83b]. asymmetric [KPBRF83]. Asymptotic [Boy84, LMM87, LM89a, Lo85, NR80b, CRS88, Gus80b, Jon81, MM81c, RA81]. atmosphere [Bai80]. atmospheric [Cul89, MGS82]. atom [HC87, Mru87]. atomic [KSD81, MD85, NDAL88, NR84, SC81]. atoms [AEG80, BKH86]. attenuation [Lan82]. Attractions [BH87b]. Au/NaCl [AH88a]. Author [Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano82d, Ano82e, Ano82f, Ano82g, Ano83e, Ano83f, Ano83g, Ano83h, Ano84d, Ano84e, Ano84f, Ano84g, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k]. autoconvolution [DFG81]. automata [CH87]. automatic [Dav83, GP81, HBR82, Moh84, NR80a, Sha84]. automating [TGP80]. automaton [ZH88]. averaged [CPJ89, CBCF80, CFT82, CF82, MS86, VM86, ZC86]. averaging [Kre88]. axial [KP83b, ML85]. axisymmetric [AFLS83, AB81, BS86b, Ber85a, CC87, Gar83, GMT^{+}81, GDM83, Jar81, Mac86, MSK80, Uch83, WG86].

Back [Kre88, Dun85]. Back-averaging [Kre88]. back-projection [Dun85]. background [MSD87]. bag [CL86a]. balance [NMG86]. balanced [NMG86]. ballooning [SB87b]. banded [Tuc89b]. barely [PGBO87]. baroclinic [MRH83]. based [Aug84, Cla87, Flo85, Flo86, IA88, KE87, Kli83, OS88]. basis [Boy87b, DR85, GM82]. bath [Din86]. Baxter [PRV89]. Bayesian [OW85]. be [Mez81]. beam [EM80, GMT^{+}81, HL84, JR82, JHLS89, LLL88, ONM83]. beam-plasma [EM80]. beams [MSD87]. bed [CG89, SS85]. behavior [LR80a]. BEM [KK89]. Bénard [Bue88, CB84, CW84a, CW86, Grö83, WC85b]. benchmark [Ano85w]. Benjamin [CH88]. Bernoulli [Lyt87]. Bessel [Afa89, BPD81, Can81, GM88, MG89, NM80, O’B80, Pie84, PA86, Puo88, RH87, Ter81a, Ter81c, TB86, Wal84]. BETA [BHM^{+}83]. Bethe [ML83a]. better [Osg81, Skö82]. between [CM89, Hal88, Hof82, HW81, Hut84, KGH80, MSK80, YRT80]. bi [MM85b]. bi-conjugate [MM85b]. bias [Mez87]. biased [Dav84, Gol86a, GWZ85b, SG86]. bifurcation [CW86, EG83, KH87]. biharmonic [KF87, LC88, Spa85, Ste84]. biharmonics [BSS82b]. bilinear [GL85]. binary [Ale86, HMS89]. binary-binary [Ale86]. binary-single [Ale86]. binding [Laf89, MJ81]. Birth [ONM83]. Black [J.82, AG80, BST86, Sha81]. black-oil [BST86]. blast [EB81]. Blending [Ste86, Eyr87]. blob [Mei89]. blobs [And86, Ker88]. Block [Van86, BDJS86, BM80b, MSW87]. Block-implicit [Van86]. blocking [HBR82]. blood [Fog84, MP80a, MP89a, PM80, PM89]. blunt [GW80]. bodies [DS86c, GW80, OI87, Ros89]. body [Abe89, AEBS86, CQ89, CP88b, DH82, Eyr84, Eyr87, KTH88, MPPS84]. body-fitted [Abe89]. Boltzmann [Mez87, NDP88, Rei81, Shi81]. bonds [LBO^{+}89]. born [ML83a, ML83b]. boson [LWP89]. bounce [MS86]. bounce-averaged [MS86]. bound [ADKM82, CL80, MES80]. bound-state [CL80]. boundaries [AH85, HN81, IBCS87, KK86, MT86b, MK80, PG88a, Rai86a, Vic86]. Boundary [BS86a, Cou85, Hag87, SSSR83, SA81, TB82, TWM82, VTB89, AM82, Abd87b, And89, Are81, Asc80, AD83, BDR82, BBM88, BT82, BWY82, BK88b, Blo82, BGR80, CO80, CQ87, Coo86, Dav80a, Dav83, DP86, EM81, ET88, Fer88, For86, FRRT87, GDS88, Gar86, Gar78, GP80, Gus80a, Gus82, Hal88, HMI83, HC81a, HO80, HMR82, HGS88, HW81, Hun86, Hun89, IHM81, IO81, Jac89, JY83, KR84, KR86, KG89, Kel84, KRW88, KTH88, Kra80, KS82, LBS88, LC88, Mey81, Mey86, MT84b, MS80, MC88b, Nat89, NMG81, NE82, Nor89, OS83a, Ohr83, OC83, Phi86, PH88, Pom88, QR89, RSR86, RWC89, RM82, RS80, SH89, SS88b, SB84a, Smi81, Smo82, Spa85, Spi88b, Str89, Str80, Sum89, TL81, TMCS85, TL84, TB89, Tho87, Uch83]. boundary [WBL86, Wol81, Zeb84, vDS80]. boundary-conforming [DP86]. Boundary-fitted [TWM82, AD83, HC81a, LBS88, MT84b, MS80, Ohr83, TMCS85, TB89, Uch83]. boundary-layer [TL84]. boundary-orthogonal [Spa85]. boundary-value [Asc80, Dav80a, Dav83, Spi88b]. bounded [KJM80, Law88, Law89]. bounds [Ber83]. box [J.82, Jon81]. branching [EG83, Hal89, LBS88]. break [AEBS86]. break-up [AEBS86]. breaking [Miy86]. breakup [Bro85, KTH88]. Brillouin [Nor80]. Buckley [Pro81a]. Budden [TL81]. bulk [CX86]. Burgers [CSW85, Fle83, Lav88, LJ81, Löt82]. burner [Smo82]. burner-stabilized [Smo82].

C [JDK86]. Calculating [BW81, ADR84, BF86, CKR87, GMT^{+}81, LP83, LWP89, MM85a, MFC^{+}85, MB80, NM80, Pap84, Pru88, Ret82, Rom89, Ros80, Yam83]. Calculation [ADK82, BS81, EG83, GS87b, HK85b, KTT89, MD81a, Bak81, BC87b, BB88, BD82, Bra84, Bra86, Can81, CBLL86, CW82, CL80, Cor82, DW81, FMC83, FP82, Fet84, HCG^{+}85, HGS88, HK81, IMT^{+}89, KD80, Lyn84, Mac89, NOY85, RS81, Rob89b, SHB82, Sch80a, SBBY83, Shu84, Ste81, Sto84, SV89, Tam81]. Calculations [ABS82, DA86, KSD81, LCC^{+}86, ACM88, And83, AB81, AB84, AB85, Bon84, BE85, BR86, Bra88, CQ85, CLJ^{+}89, CHH^{+}86, CW86, Col81a, CDH81, Fer80, GHC^{+}86, GDM83, HCH^{+}81, HCH85, LB87, MW87, Mah82, MJ81, Noh87, PT82, Per81c, PC86, Rai86a, Rai86b, Ram83, RL84, Rei81, RH88b, RS80, SK80, SBM^{+}87, SK83b, SC82, Sho83, Tak81, Tox82, Van84, Vin83, Wil80a, WVP82, YWH85]. Caley [KD80]. Call [Ano87j]. Canonical [Aiz85, IA88]. capabilities [BBGN80]. capillary [LA83]. capturing [SO88, SO89]. Carlo [Abe89, Ada88, BHS87, Bin85, Boo81, Boo82b, BF86, Bro89, CBG86, CDS83, DA86, Dun81, EFCS88, EB80, FC84, Gol86a, GRP82, HM80a, Har81, HP87, HPP^{+}82, HM80b, HSSC83, JR81, KKMP86, KL89, KT86, LM87b, Mez81, PRT83, Pes81, RH88b, RV88b, Sad84, SG86, VL84, Ves82, VR86, Wag87, Wag88, WC85a, WP88, WPW89, Yor86]. carrier [Yor86]. cascade [Ema81]. cascades [Ino83]. case [RL89, ZH88]. cases [WOZ84]. catastrophe [CW84a]. Cauchy [Ioa82, IMT^{+}89, Sid85]. Cauchy-type [Ioa82, IMT^{+}89]. cavities [CW84a, Kum82, LA85, WCR^{+}88]. Cavity [GH87, Gup81, GH86, Phi84a, SKA86, SK83a, SK83b, Str82]. CCl [Ves82]. CDC [KKMP86]. Cebysev [Hos86]. Cell [Wes88, BR86, Bra88, CBL81, CBG86, Hai87, HO80, Hor87b, HSA89, HGM86, LD89, MS83, Riz88, Ste84, Zuf89]. Cell-centered [Wes88]. cells [Duk88]. cellular [CH87, ZH88]. center [GS85b, KO86, SSS^{+}84]. centered [Hof82, Vij86, Wes88]. Certain [vdV81, Mer80, PG88b, SSM83, Spi88a, Spi88b, TA84c, TA84a, TA84b, TA88, TY87b, Wag87, Wag88]. CFL [PC86]. Chain [FFM82, SH84c]. chains [Pes81, SHB82]. change [Lyn82, LS85, PG87]. channel [CA85b, DP86, Flo85, Hor87a, KP89b, LBS88, Pat84]. channels [LC88]. chaos [HS88a]. characteristic [Car80, DO86, MS87b]. characteristics [FS87, Gol86a, Pla81, TS81, VF80]. characterization [Mer80]. characterizing [MD80]. charge [CM82, CDS83, Mor82]. charge-transfer [CDS83]. charged [AD87, Cep83, Mac81, MPA88, MD80]. cheap [Wil81]. Chebychev [McD80]. Chebyshev [Bon84, Boy82, CQ85, DR85, DD89, DQ83, DM85, Fre83a, Fre83b, FT84, HLAD84, KP89b, KPD81, KH84, LA85, Pie84, Tan85, VV82, Zeb84]. Check [FK89]. chemical [Mac89, Tak86]. chemically [CB80, KK81, MGS82]. choice [BK88a, Got88, Gus82, MM81a, MP84, OG86, Tak86]. Choosing [ICP80, Mun81, Pes81]. Christoffel [FZ87, FZ88]. CIP [TNY85, TY87a]. circular [BM87a, For85, IT89, JY83, JY86, LM87a, NB89, ONM83, Pat87]. circulation [BC87b, Kum82, MW84]. class [BB84b, FS89, LR81, Pro82, Pry87]. classical [HC81b, LO86, Rok85]. classification [Bar87]. Clebsch [KTT89, Pap84]. Clenshaw [EHM81]. Closed [Afa87, Afa89, MGM86, Ioa81b, Set84]. closure [VM86]. clotting [Fog84]. cluster [Hai85]. Clusters [BP88, RQ89]. coalescence [Bro85]. coalescence/collision [Bro85]. coalescence/collision-breakup [Bro85]. coarse [HD85]. coarse-mesh [HD85]. coating [SS81]. code [AP88, Asc80, AEG80, BLT^{+}81, CC87, CS81a, CBCF80, CF82, Fre81, FSD81, GRS89, HA89, Hew80, HH86, HSA89, HL84, JL86, JHLS89, KH80, KGH80, MS86, NOY85, ONM83, TLD80, Tam81, Wil81, ZC86, ZGB88, ZBG88, ZBP84, LJ85]. Codes [SHK88, BHM^{+}83, BDR80, CFT82, HMJ86, Hor87b, Ker81, LD89, Mar87, Per83, Pir80, Sha84]. coding [JR81]. coefficient [Bai80, MIK88]. coefficients [Boy84, CDS85, Dil85, EHD89, GT85, KTT89, Lyn84, Pap84, Sca87]. collapse [Man85, Man87]. collapsing [MEK86]. collection [Ano85w, EWR88]. collision [CBG86]. collision-breakup [Bro85]. collisional [Mas83, Per81b]. collisional-radiative [Per81b]. collisionless [Kli87]. Collocation [Mos86, ABW83, Asc80, EWR88, Gau88, KP89b, KP89a, KMA89, MZH85, SP81]. column [Wil80b]. combination [Li89, MK83]. combinations [SC82]. combined [Hei88b]. Combining [BGMT89]. combusting [Kan81]. combustion [BM87b, RD83, Set84]. Coming [BH87b]. Comment [Gus80a]. Comments [EH85]. Compact [DH89, Ros83, AD83, Chr85, GGR89, GP85, Lev82, MD81b, Shu84, Ta’87]. comparative [BK88b, SP87a]. compared [KE87]. Comparison [BHM^{+}83, HMJ86, Hor87a, Mak89, MD85, Sha81, She88, TAJ^{+}81, CL86c, EM80, Fle83, Gup81, MES80, MS87b, NS89, SBT86, TDhLcSb89]. compatibility [HG84]. complementarity [CFH^{+}82]. complete [Wil82]. Complex [Boy85, Dav89, Paj89, Dar89, DA86, Dye84, Ker89, KP86b, Ors80, SH89, SSF83, Tem88c, TB86]. Complex-plane [Dav89]. compliant [Hal88]. composed [SB82]. composite [MM89, ST88]. composition [MM85a]. compound [NR85, YH88a]. compounds [MV83a]. compressibility [HN80, MT81]. compressible [AB85, BT82, BAF84, BaB86, CW82, Dar89, Gau88, GG86, Har80, HK85a, IGW86, KJL86, OK86, PP88a, Pea81, PVMZ87, Pul80, Rei81, SSSR83, Tur87, Van84]. compressible-flow [Van84]. Computation [AH85, Asa87, BK87, BG84, DS86c, Dil85, GK81, GIT87, Har80, HR84b, HT86, JY83, KP85, KP86a, LA85, MMV81, Pri89, BGT85, BR81, Dan80, Duk88, EJL87, Gre85b, Gre88, HO80, HW81, IGW86, Kor88, Mih85, MK86, Nat89, NR80a, PT86, Qua81, QN86, SH84b, Tan89, WMS88, Whe81]. Computational [Baa81, HHS89, KS82, ML83a, ML83b, MM81a, MV83b, Var84, Zab81, AA89, Clo87, ET88, FP88a, FC84, KTH88, Li80, MP89a, NB89, Pap84, PM89, PA89, RE87, SR85, TGP80, YHS86]. Computations [MSK80, DWD83, DH84, For83a, Hal89, HS86, HM89, HM80b, KKMP86, LV89, PVMZ87, Rig81, Vie88]. computed [Riz88]. Computer [BP88, CD80, EM82, ER85, HR86, ST88, AD87, Ano85w, HHN^{+}86, KKMP86, Kra80, KH80, KGH80, McD80, ZMC83]. Computer-aided [ER85]. computers [Ano80z, ERDF83, FWHC86, HA89, HL86, KE87, SMO85]. Computing [Ber83, Cla87, CW81, And86, BC87a, EHD89, GW80, OI87, Puo88, Sak81, SS86, WG86]. concentration [KD88, LW85]. concentrations [SH84b]. concepts [BBG80]. concerning [Vic86]. concurrent [LD89]. condition [Are81, HMI83, HG84, KR86, KRW88, Nat89, NMG81, Ram83, RS84, RS80, TL81, TE86, VTB89]. conditioned [Hag80]. conditioning [Qua81]. Conditions [Cou85, Abd87b, And89, BDR82, BBM88, BT82, BGT85, BWY82, BK88b, Blo82, BGR80, CO80, EM81, Fer88, For86, FRRT87, Gar78, Gus80a, Gus82, Hag87, IO81, JY83, KG89, Kra80, KS82, NR80b, Nor89, OC83, PH88, Pom88, QN86, Rob85, RM82, SS88b, SSSR83, SA81, Tho87]. conduction [JT86, LG85]. conductivity [CL81b, HHN^{+}86]. conductor [ADK82]. conductors [Kam81]. cone [ML85, ML87]. conference [Ano81j, Ano82i]. configuration [MNM85]. configurations [Flo86, HSA89, Ino85]. Confined [MHK86]. confinement [LNS87, NB89]. confluent [Ter81b]. Conformal [Flo85, Flo86, O’B81, Ino83, Ino85, MOI81, MZ80, MZ83]. Conformal-mapping- [Flo86]. Conformal-mapping-based [Flo85]. conformational [SHB82]. conforming [DP86]. congress [Ano87j]. congruential [PP88b]. Conjugate [WVP82, AG80, BDJS86, KE87, Ker80, MM85b, Neu81]. conjunction [Alp81]. CONKUB [Mej86]. connected [BS86a, Vie84, WR86]. connections [Kam81]. consequences [Sch80a]. conservation [BS83, BDS88, BCT89, Car80, CLS89, Dec84, HTE81, Har83a, HH83, Har83b, Hen87, LV89, LS85, MW84, MA84, McD89, Mun88, Nav83, SS82a, Tox83, Vin89, WW89]. Conservative [BKP87, Duk84a, Gre84a, Ram85b, BJ81, Hor87a, LW85, Rai86a]. conserving [CQ89, HPL80, IA88, Qin88, Ste87a]. considerations [AHH87, FS86, SG86]. Consistent [Car89, Col81a, Fer80, HK85b, Jar81, Mez87, Per81a]. constant [BG85a, FR87, GL85]. Constrained [OS83a]. constraint [ADL87, BW81]. constraints [MHM81, Per81c]. Construction [Yee87, AP85, CB89, Hag80, Pet87, TGP80]. contact [PW85, SS80]. containing [HL85]. contaminant [SH84b]. continuation [HK84, Ter81b]. Continued [Bar82, HRT80, Mar81]. Continued-fraction [Bar82]. continuity [Cha81, Eis82b]. Continuous [HW81, Mel86, Mey86, HM80a, HW84, Smi88, VL84]. continuous-slowing-down [HM80a]. Contour [Dri88a, Dri88b, ZOWZ88, ZO83]. contractile [MP89a]. contraction [KP89b, KP89a]. contributions [CM82]. control [BB86a, GE88, JPD86, NR80a, RS81]. control-volume [RS81]. controlled [VV82, ZOWZ88]. controls [Eis82a, Eis82b]. convection [BV82, CW82, CB84, CW84a, CJW85, DO86, GK81, Gau88, Grö83, HKA89, HGM86, JY86, LA85, Mul89, Neu81, Phi84a, Shy85, VF80, Bue88, WC85b]. convection-diffusion [BV82]. convection-dominated [DO86, HKA89, Shy85]. convective [Clo87, GR84, Gla84b, Gup81, MJ87, Zal81]. Convergence [DWD83, GRS89, Hal81, WC85b, ADG86, BR80a, Boy82, Col81a, ER85, Fer80, Gol86a, HH85, LR80b, Nor89, Pan89, Rig81, SB84b, Tan86, Van84, WC85a]. convergent [Ixa80, SP87b]. converging [LR80b, MOFK88]. conversational [Mej86]. convolution [Osg81]. Cooley [Eck89, MES80]. cooling [Law88]. Coordinate [Eis82a, Abe89, ABDV81, ADL87, Boy85, CCCS81, Eis82b, Flo85, Flo86, GDM83, HC81a, KP83b, MFC^{+}85, MT84b, MS80, RM87, Spa85, TMCS85, Tan85, TWM82, Uch83]. coordinates [AD83, BS86b, BS81, CLJ^{+}89, Dav81, DP86, HP87, HCG^{+}85, IM88, KPBRF83, LBS88, Ohr83, Pea81, Rom89, Skö82, WBF86]. core [Gre85a]. corner [Boy86]. corners [Wig88]. corrected [KW87, McD89, MC85, PGBO87, SM89, SMR89]. correction [Kor88, Kor89, MA84, PGBO87, Tal80]. corrections [And86, LWP89]. corrector [TS81]. correlation [Ber88, Hai85]. correlations [PP88b]. Corrigendum [CP87, Tem88a]. Cosserat [BST80]. Cotes [BL88]. Couette [MML83]. Coulomb [Bar82, Mar81, MGM86, MPPS84, TB86]. counterflow [KS87]. counting [HR84b]. couple [JR85]. coupled [Bai80, Ixa80, Moh84, Mru87, Phi84b, Rig87]. Coupling [DGKW84, BG85a, Hoe83, Hoe84, KGH80, Law87, VM86, YRT80]. Cray [GKL^{+}87, ZH88]. Cray-2 [ZH88]. Criteria [Mun81, CR84]. Critical [BG85a, BL85, EG83, PRV89]. cross [AP88, HPL80, ML83a, ML83b, OW85, WL84]. crossflow [DS86c]. crystals [Mac81]. Cubic [TNY85, LJ81, Mac81, Mor84, TY87a, TM88]. cubic-interpolated [TY87a]. cubic-spline [Mor84]. Current [BH87b, Cul83b, BD83, HM82, Kam81, ND89, VE87]. currents [BC87a]. Curtis [EHM81]. Curvature [Cho85, OS88, ZOWZ88]. curvature-dependent [OS88]. curvatures [PA89]. curve [Bal85, BB84b]. curve-fitting [BB84b]. curved [TMCS85, Tha80]. curvilinear [MS80, Spa85, TMCS85, Uch83]. cusp [CW84a]. cusping [Mey81]. cut [OS83b]. CYBER [KKMP86]. cyclic [Ske81]. cycling [ASL82]. cyclotron [Ota88]. cylinder [Ana89, BM87a, For85, IT89, JY83, JY86, MMV81, Pat87, RS84, SGA81]. cylinders [MSK80]. cylindrical [IM88, Pir80, Skö82, SN88, Tam81, Tan85, WBF86]. cylindrically [CB89, KW85, MOFK88].

D [GRS89, MD81b, ABS82, AHH87, BaB86, BHM^{+}83, BBGN80, CS82, CS81a, DD87, FSD81, GHC^{+}86, HLAD84, HCH^{+}81, KE87, Kli83, Law89, Mas87, NB89, Oli80, OvOH89, SW88, TB89, WMM86, WCR^{+}88, WWM^{+}81]. D-shaped [NB89]. damping [AH88b, Cha81, Per87]. DAP [Ada88, PT82]. Darwin [HB87, ZC86]. data [Baa81, BB83, Bow84, CL81a, Chr87, Die83, Har81]. Davidson [HN82, Kos84, Mor89]. decay [BGM87a]. decomposable [Pat86]. decomposition [HG88, MT84b, MT86b]. Decompositions [MvdV81]. Deconvolution [BB83, Die83]. decoupling [Wol85]. decreasing [GM81]. deeply [WPW89]. Defect [Kor88, Kor89]. defects [Mac81]. defined [Hai85]. definite [Smo84, SC86]. deforming [LG80, Lyn82, LS85]. Delaunay [RHOG88]. delay [HS88a]. Delves [IA85]. demand [SBT86]. dense [GKL^{+}87, GS88]. densities [NR84]. density [And85, CX86, Gla89, MJ81, Mor82]. Dependent [MK82, AF86, ADR84, Bai80, BK85b, BF86, CFL86, FW81, Gan85, GS89, HG88, Hu89, KK83, Lar88, LW85, MK80, OA85, OS88, Per81b, PC86, Roe86, SAR82, Tak81, Tho87]. deposition [ONM83, Wie83]. derivations [Gla84a]. derivative [Are81, BV86, DF88, Mru85]. derivatives [AG81, Bar82, HK81, Tan86]. derived [Duk84b, SS84b]. DERPER [HK84]. DERPER-an [HK84]. descent [HH85]. describing [NR84, NB89]. Description [KH80]. designs [Neu83]. Desingularization [Kra86]. determinant [BV86]. determinants [KD80]. Determination [Mer80, BS86b, CX86, CRS88, KH87, KPB87, Mez87, Bar80]. determine [LPS80]. determining [Kai83, PT86, PA89]. Deterministic [NDP88, Her89]. deuterium [NR84]. Development [BS89b, SC86]. devices [HK85b]. DFT [Osg81, Tem88b]. diagonal [PC81b, RCK87]. diagrams [CP88b, KSM82]. diatom [AP88]. diatom-spherical [AP88]. diatomic [Kob85]. dielectric [Rob89b]. Difference [Car89, MSB87, Alv89, AHP81, AG80, BM87a, BCE88, BWY82, BS83, BC87a, BC87b, BKP87, Blo82, BST80, BNS83, BP80, Bue88, CKD87, Chr85, Cul89, Dav84, DFP81, DH89, Dic88, Dyk87, Fle83, Gar86, Gar78, GT85, GP85, GS89, GR82a, GX89, Gus80a, GW80, HTE81, Hei88b, Hof82, HC81b, Hua81, IA88, IM88, KMRM88, KOTK88, LLL88, LG85, Luc87, MC88a, Mac86, MEGH85, MES80, Man85, MS83, MIK88, MNM85, Miy86, MS87b, Ohr81b, Per87, Pic83, Riz88, Roe81, Ros83, RMS85, SS82a, SB88, She86, Shy85, Skö82, SM86, SL85, SW81, SBSH84, Tak80, Tak85, Tha80, Vie83, Vin83, Vin89, WMM86]. difference/Galerkin [Bue88]. differenced [PC82]. differences [AD83, She88]. Differencing [Ker81, BW88b, Dav80b, PC86, Str82, Zal81]. different [Gus80b, Mak89]. Differential [BO84, BM84, MHK86, AF86, BK85a, BSSV88, Boo81, Boo82b, Cat88, CL81a, CL84, CS84, CFL86, DQ83, Dyk87, GDM81, Gre84b, HH88, HS88a, HK84, KPD81, KH87, LW86, LO87, MD86, Moh84, NR85, Oey82, Pom83, Pro82, Pry87, Ram85a, RG88, Rig87, SSM83, SM84, Spi88a, Spi88b, TY87b, TAJ^{+}81, TWM82, Vat87, Vat89, VBSS89, War86, WP83, Wol85, ZBP83]. differential-geometric [War86]. differentiating [Baa81]. differentiation [Pro81b, SH84a]. difficult [Dav80a]. diffraction [MP84, OG86]. diffusion [ADR84, Bai80, Bas84, BGMM89, BB86a, BV82, Fou84, Fre81, GS85a, GR84, Hag87, HC87, Her89, HGM86, Ker81, KS87, Kre88, LW85, MMS86, MO80, MGS82, Mei89, MC85, MC88b, Per81c, Pom88, RSR86, RV88b, RMS85, SS86, SHK88, Smo84, SV89, TS83, VF80, VR86]. diffusion-convection [VF80]. diffusion-type [SS86]. diffusional [LP83]. diffusive [LMM87, LM89a]. digital [ERDF83]. dilatation [MFC^{+}85]. dimension [ADR84, Cho82]. dimensional [AG81, AM82, Ana89, And89, AF86, Aug84, AB84, AB85, Bas83, BK85a, BSSV88, BJ81, Bue88, CA85a, CA85b, CL83, CC84, CPJ89, CY85, CH84, CG89, CLS89, CBL81, Duk88, DH84, DP86, EH85, Eas87, Eck89, Fle83, FP88b, FS87, For89, For84, FB88, FFG87, GGR82, GGR89, Gau88, Gla88b, GG88, Gre85b, Gre88, GX89, Hag81, HHS89, HK85a, HH83, HC87, HK86, HR84a, HMJ86, HH86, HSA89, HGS88, Hu89, Hua81, KD88, KK88, KH80, KGH80, KHT87, KHTR89, LLVL89, LLPS85, LA85, LG81, Lim80, LBS88, MS81, MP84, MEK86, MO80, MP89a, Med86, Mei89, MT84b, MNM85, MTM81, Ohr81b, OG86, PM89, Qua81, Ram85b, Ram86, RM87, RG88, Rei81, Ros89, RM82, SS84a, SK80, SBC84, Sch80b, SM84, SBT86, SSG81]. dimensional [SK86, SS82b, SN88, TS81, TY87a, Tal80, Tan85, Tan89, TOO83, Uch83, VF80, VBSS89, Vin83, WR84, WR86, WC84, ZGB88, ZBG88, ZMC83, vDR89]. dimensions [BDS88, BS82, BF82, BR86, Cat88, HS86, KO86, LM86, Rok89, Sal86, Sco85, SP87b, WOZ84]. diminishing [YWH85]. diode [MS81]. dipole [CM82]. dipoles [NDAL88]. Dirac [dFSS89]. Direct [BS86b, DQ83, HK81, KW85, LCF83, Pul80, RH88b, SB82, Abe89, Bas83, BS89a, CLHP89, Grö83, HL87, Ioa81b, PSV89, Pat86, SS88b, Whi87]. Directional [GE88]. disappearing [BJ81]. discharge [HNK89, SMR89]. discharges [BCE88, TST^{+}86]. discontinuities [GMM80, MC88a, RL89, Sho83]. discontinuous [CLS89, Cul89, Hua81, Law85]. Discrete [CFR84, Cou85, Roe86, Ber85a, BH87a, Dav81, ER85, GM81, IA88, Kan84, MD80, Mor84, SB84a, SP87b, Tem80, WP88]. discretisations [Slo88]. discretised [Iss86]. Discretization [LLPS85, ABDV81, DQS87, KJL86]. discretizations [CH88, For83b, Pat86, Ste84]. discretized [DH82, GRA83, SB82]. disintegration [SE87]. disk [BJ84]. disordered [HHN^{+}86]. dispersion [BST86, For84, Neu81, O’R89, SSS^{+}84]. dispersion-convection [Neu81]. dispersive [BPS85, KRW88, SSC81]. displacement [BD83]. displays [Bow84]. dissipation [Mun88]. dissipative [KP85, KP86a]. distance [MM81c]. distances [CCCS81]. distorting [Gar83]. distortion [ADL87]. distributed [Wik89]. distribution [And86, iKI85, Law85, Mer80]. distributions [CM82, GWZ85b, HS88b, HR84b, Kai83]. disturbance [PG87]. diurnal [Bro80]. Divergence [Tuc89a, Gup81, HW80]. Divergence-free [Tuc89a]. divergent [ML85, ML87]. diverted [HPP^{+}82]. Does [Mil89a]. domain [BP87, Boy82, CH84, LM89b, Mav89, MT84b, MT86b, NMG86, Sch88, SB82, Tan89]. domains [BS86a, CY85, HMR82, KP89a, PSV89, Pat86, RM87, Tha80]. dominant [MSW87]. dominated [DO86, HKA89, Shy85]. donor [Hai87, HGM86]. double [Kra83, Ohr81b]. double-folding [Kra83]. doubling [BM87b, KH87, VJ89]. doubly [Vie84]. down [HM80a]. drift [Mor81, Sco88]. drift-wave [Sco88]. drive [VE87]. Driven [SK83a, CHS88, Gup81, Kum82, SK83b, Str82]. driving [WK89]. drop [SBT86]. drop-on-demand [SBT86]. droplet [O’R89]. dual [CBLL86]. duct [HL85, Mos86, SSG81, WR84]. due [And86, BM80a, CH84, GMT^{+}81, Law88]. Duffing [VV82, VJ89]. Dufort [CR84]. Dukowicz [Kan83]. during [Fog84, SLO87]. DWBA [Dan80]. Dynamic [JPD86, Oli80, JH83, KH80, KGH80, Lub88, Rap80a, Wil80a]. Dynamical [MMR89, Alz84, CW84b, ZO83]. Dynamics [BP88, And83, Ano80z, Ano87j, BAF84, BL82, CGM^{+}86, Clo87, DS86a, Dri88a, Dri88b, ET88, GW85, GH86, GH87, HN81, Hua81, KK83, KH80, KGH80, LBO^{+}89, MHM81, Nav87, O’R84, Pat84, Rap85, Rei81, RE87, Roe86, Sch82, Son85, Spe87, SR85, SMO85, SH84c, Tox82, Tox83, VTB89, Wan83, WH87, WR86, ZOWZ88]. dynamics/surgery [Dri88a]. dynamos [Gla84b].

Earth [Bai80, BG84]. economising [BD83]. eddy [HM82, Hor87a, Kam81, KTH88]. Editor [Ano82h, Edi82]. Effect [JR85, BGR80, BB80, HM89, HO80, TM88]. effective [CP88b, Hoe84, WR84]. effects [CLF84, Mei89, Mor82]. Efficiency [CL87, FC84, HRS80, RD83]. Efficient [CG85, Duk88, GRP82, Lub88, NSJ85, PT86, PG88b, Per81b, SO88, SO89, Bro80, Bro85, GS85b, HKA89, HD85, JR81, JRS86, KKMP86, Kor88, Law87, LA83, Ohr81b, SHB82, SK83a, SH84b, SS89, Wal84, Wig88, Yam83, Zer85]. efficiently [NM80]. eigenfunctions [BW81]. eigenlengths [NR80a]. eigenpairs [GPS88]. eigenproblem [HN88, Nex87]. eigenproblems [Boy85, Pry87, SPG89]. eigensolutions [Kos84, Ret82]. eigenspan [MSW87]. eigensystems [Tan86]. Eigenvalue [Tal80, BM84, CS81b, DR85, DVW87, GP89, GKL^{+}87, GS88, HH88, HN82, Kal84, Ker89, LW86, LO87, Mor89, PG88b, SB84a, YH88a]. eigenvalues [BW81, BD82, Bra84, Bra86, CRS88, CW81, Eck89, EJL87, FMC83, GTD89, GP89, KW86, PT86, Pru88, Sch88]. eigenvectors [CS81b, NV89]. Ekman [RS84]. elastic [NE82, PM89, SKA86, YH88b]. elasticity [Ta’87]. electric [HNK89, HK85b, Mor81]. electrical [HHN^{+}86]. electrode [CL86a]. Electromagnetic [HL87, Wei85, ADK82, BF82, Mar87, MMS86, Mas87, TL81, Tan88, WBF86, ZMC83]. Electron [ASL82, BY85, CBG86, CD80, HM80a, Hew80, HL84, JHLS89, Kli83, Mor81, Wie83, WL84]. electron-field [Hew80]. electronic [Col81a, Fer80, WVP82]. electronics [Phi84b]. electrons [LBR81, SBBY83, SV89, TLD80]. electrostatic [Dec84]. element [ABS82, BW88a, Bas84, BE80, BJ81, CB84, CS82, CL86c, CLS89, CFG83, CM80, Cul82, Cul83a, Cul83c, DD89, DO86, DM85, DQS87, Duk84b, EB83, EW88, Fle83, For83a, Fou84, FW81, FWB^{+}81, GDM81, GHK88, Gla89, Gol86b, GF86, GL85, GIT87, HC87, HGS88, HMW86, JWB88, Kaw87, KJL86, KP86b, KHTR89, LLVL89, LV86, LM89b, LG81, LG80, LPS85, LS85, MES80, Man85, Man87, MK83, Mat86, Mos85, MC88b, Nav83, Pat84, Pat86, Pla81, RV88a, SS81, SPG89, Sch88, Ste87a, SLO87, TL84, WD84, WC85b, ZGB88, ZBG88]. element-finite [Man85]. elementary [MV83a]. Elements [YHS86, DH82, Hut84, KSD81, Lyn82, Mav89, OS83a, Pap84, SMRS85, She88, SHK88, VF80, Vij86]. eliminate [Tak81]. eliminates [GTD89]. Elimination [WD84]. ellipsoids [PW85]. Elliptic [Ten82, BB84a, Boo81, Boo82b, BFT85, BJ84, Cat88, Cha86, DQ85, DM85, Hun86, Kel84, KPD81, Li89, MD88, McD80, PSV89, Phi84b, Phi86, Rig87, Sch80b, SP87b, Wil82, ZWH82, ZWH84]. Elliptic-vortex [Ten82]. elongated [Rag82]. embedding [CJ86]. embeddings [NDAL88]. Emden [FLP80]. empirical [HBR82]. employing [Bai80, Tob86]. enclosed [Phi84a]. enclosure [CHS88]. encountered [Zeb87]. endpoint [Boy86]. Energy [Dec84, Ste87a, Tox83, CQ89, Eck89, HK81, LW89, LNS87, MEGH85, MES80, Mer80, Qin88, Ros80, SHB82, Sch80a, SV89, Wie83]. energy-conserving [CQ89, Qin88]. energy-minimizing [LW89]. energy-momentum [MEGH85]. enforced [Nav83]. enforcing [Ram83]. engine [Ros87]. Enhanced [RWC89]. enhancing [Mah82]. ENO [Har89]. ENSTA [Ano81j]. enters [LO87]. entropy [BKP87, Har83b]. environments [McN87]. Equation [DGKW84, MK82, Abd87b, AI80, Alv89, AOK81, AA89, Are81, BR80a, BGT83, Bel82, BKP87, BK85b, BST85, BE85, BFH86, BS84, Bro85, BGM87b, CSW85, CW89, Cha81, CS84, CRS88, DF88, DH82, DFP81, Eck89, EB80, EW88, EWR88, FFS82, For83a, FR87, FRRT87, FWB^{+}81, GC89, GR84, GM85, GF86, Gre85a, GRA83, GR82a, GR82b, GC88, HLAD84, HM82, HP87, HPBY86, HMM85, HMW85, Her89, HC81b, IBCS87, IHM81, ICP80, Ixa80, IR87, IM88, JT86, iKI85, KF87, Kel83, Kel84, Ker81, KG85, KK88, KK83, Kru82, KW85, LL81, LBR81, Lav88, LW89, Lim80, LG85, LJ81, Löt82, MH89, MS83, MIK88, MS87a, MMS86, MLB87, MGS82, MD80, MD81b, Mer86, MT86b, MJ87, MM81c, Mos86, Neu81, NS89, Ohr83, OS85, OC87, PC82]. equation [Per87, Pic86, PH88, PRV89, PV83, Pro81a, Rai86a, Rap80b, Rei81, RN86, RL89, SS88b, SP81, SH84a, SHK88, Shi81, Ske81, Skö82, Slo88, Spa85, Spi88b, SB87b, SB84b, TA84a, TA84b, TA88, Tal80, Tan85, TY87b, Tem80, TB89, TC87, TM88, Tre81, VF80, Vat89, Vie83, Vie84, Wig88, WK89, vDR89, NDP88]. Equations [BO84, MHK86, MvdV81, AG81, Abd87a, AH88a, Adh81a, Adh81b, AEBS86, ABW83, AHP81, AF86, Aug84, Bai80, BCE88, BDR82, BL85, BWY82, BCG89, BB84a, BB86a, BSSV88, Bog88, BST80, BGR80, Boo81, Boo82b, BB86b, BP87, BV82, BB80, BNS83, BW88b, BGM87b, CDP80, CD82, Car80, Cat88, CCCS81, CA85a, CA85b, CJ86, Cha88, CW85b, CHH^{+}86, CKD87, CL81a, CL83, CC84, CL84, CPJ89, CFL86, Cul83a, Cul83c, Dav84, DR85, DD89, DQ83, DQ85, DM85, Dic83, Dic88, Dru80, DP86, Dyk87, EG86, ER85, EHM81, Eyr87, FN80, Far80, FLP80, Fle83, FP88b, For83b, For84, For86, FB88, Fre83a, Fre83b, FF86, GGR82, GGR89, GDM81, GGS82, GT85, Gil81, GS87a, Gla88a, Gla88b, GS89, Gre84b, Gup81, Hal81, HS88a, HA85, HK86, Hen81, Her89]. equations [Hew80, Hir89, HG84, HK84, HW81, Ioa81a, Ioa81b, Ioa82, Iss86, IA88, Ixa80, Jar81, Kel84, KPD81, KM85, KP86b, Kor89, KG87, KHT87, KH87, Lab87, LM89b, LW86, Li80, LC89, Luc87, LPS85, Lyt87, MEGH85, MZH85, MK83, MM81a, MM81b, McD80, Mej86, MPPS84, Mel88, MZ82, MD86, Moh84, Mor81, MP80b, MP89b, MML83, Mru87, MV85, Mul85, Mül88, Mus81, Nat89, Nav83, NN89, Ng81, Nor89, NMG86, OK86, Oey82, OJJ88, OC83, PSV89, Pat89, PND83, Per81b, Phi84b, Pik87, Pom83, Pro82, Qin88, QQM89, Rag82, Rai86b, Ram85a, RG88, Rig87, Riz88, Rok85, Rok89, Ros83, SS86, SSM83, SP87a, Sch82, Sha81, SM84, SSSR83, Sho86, Soh87, SG88b, Spe87, Spi88a, SW81, SBSH84, Ste87a, Ste87b, SN88, Ta’87, TA84c, TA84a, TA84b, TA88]. equations [Tak80, TNY85, TY87a, TE86, Tan89, TDhLcSb89, TAJ^{+}81, TB82, TWM82, TF89, Tur87, VM86, Van86, Vat87, VBSS89, Wan83, WW89, WMS88, WP83, WCR^{+}88, Wol85, WVP82, WOZ84, YRT80, ZWH82, ZWH84, Zer85, ZBP83, ZOWZ88]. equatorial [GC89]. equidistant [GP85]. equidistribution [CFL86]. equilibria [AB81, ABS82, AHH87, BS86b, FDS81, GIT87, HH85, HHS89, HCG^{+}85, HPL80, MFW88, Shu84]. Equilibrium [LJ85, CS81a, DJT80, GDS88, GRS89, HMJ86, HH86, LCC^{+}86, Mac89, MM85a, Oli80, RC81, RA85, RA87, RQ89, RL84, SBBY83, Ta’87, YNAS86]. equipartition [Bir86]. Equivalence [EM82]. Erratum [Ano80j, Ano81i, Ano84h, Ano84i]. Error [iKI85, Lan82, OS85, Pry87, ACM88, BB86a, CM80, CDS85, KMRM88, NR80a, RV88b, Sch80a]. error-control [NR80a]. Errors [LL81, Noh87, Chi88, Hof82]. essentially [HEOC87, SO88, SO89]. estimates [Cho82, GM82, KMRM88, KS87, Tal80]. Estimating [Vie88]. Estimation [LS87, HM80b, iKI85, Mor84, Pry87, WC85a]. estimator [Dun85]. estimators [Wag88]. etching [VC85]. Euler [CA85a, CA85b, CJ86, Dav84, Dic83, Dic88, ER85, FS89, FB88, Gla88a, Gla88b, LC89, Lyt87, MJ87, MP89b, MV85, Mul85, Mül88, OC83, Pik87, Rai86a, Rai86b, Riz88, Ros83, TB82, WW89, WOZ84, ZOWZ88]. Eulerian [CL86b, HW84, KM80, Neu81, Smi88]. evaluating [BM89, Dan80, Dav89, Lup87, SMRS85, Wil82]. Evaluation [HRT80, Bar82, BV86, CLM88, GM81, GS85b, Hau82, JRS86, Kra83, KPBRF83, LR80b, LR81, MT84a, RH87, SGA81, Sid85, Wag87, Wag88, Wal84]. even [BR80b, GM81]. Event [Bar87, Rap80a]. Evolution [HPL80, HNK89, Kli83, KT86, MZ82, MFW89, NR84, RA81, TA84c, TA84a, TA84b, TA88, Wit84]. evolutional [CL84]. evolutionary [BW88a, BSSV88, CM80, JWB88]. Ewald [AD87]. Exact [Boo81, DS86a, JR82, KG89, SS82a, WH87]. example [Lab87, NMG86]. exchange [OC87, Sad84]. excitation [AA89, SA81]. excited [NDAL88]. Existence [Vat89]. expansion [AI80, DG88, Dil85, Eyr84, Fre83a, Fre83b, Hoe83, Hoe84, KPD81, KH84, LCC^{+}86, MMV81, Mat86, McL80, Pat84, Sca87, Tan85]. expansions [Boy86, CFR84, Gus80b, HRT80]. expectation [WC85a]. Expeditious [Whe81]. experience [Baa81, HMM85]. experiment [BK87]. experimental [BB83, Die83]. Experiments [Cul83c, MV85, ABW83, Ano80z, Kra80, OW85]. Explicit [CQ89, AM82, Alp81, Alv89, BM85, FWCW86, GHC^{+}86, HC81b, MJ87, PC82, Qin88, SS82a, SMR89, Yee87]. explosions [FH81, Gar83]. explosives [She86]. Exponential [Mun81, Vie83, Bal85, CLM88, Chr88, GS85b, Lev82, Ter81b]. exponential-type [GS85b]. Exponentially [SS84b]. Expressions [LR80a, Afa87, Afa89, MGM86]. extended [Dri88b, KH84]. Extension [DF88, LW86, FK89, LC88]. exterior [Mer86]. extinction [GS87b]. extracted [WMM86]. extraction [Whe81, WWM^{+}81]. Extrapolated [HGS88]. Extrapolating [Mon85]. extrapolation [Eck89]. extreme [Ret82].

FACR [Tem80]. factor [Cat88, Cha86, Tem83b, Tem85, Tem88c, Tem89]. factored [SSSR83]. factorisation [Bak81, Cat88]. factorizable [Laf89]. Factorization [GWZ85a, AOK81, OK86, PC81b]. factors [LR80b]. Falkner [SH84a, Sum89]. false [Str82]. falsi [FR87]. family [Pom83]. Far [BT82, Lar88]. Fast [Pat86, SGT88, Sch80b, SS88b, Tem83a, WL84, vDR89, BB84a, BK85b, CBLL86, Cha86, FP88b, GR87, KS81, KW85, MT81, PSV89, PRT83, Pic83, SS86, CFR84, Tem83c]. feature [BER85b]. featuring [Cat88]. Feautrier [Mih85, MK86]. Feir [CH88]. Fermi [CS84, FLP80, FRRT87, Kru82]. fermion [LWP89]. fermions [BBG^{+}87]. ferromagnetic [MMS86]. Feynman [ML85, ML87]. FFT [Dil85, EHD89, Osg81, Pic86, PH88, Tem83b, Tem85, Tem88c, Tem89]. fibers [MP89a, PM89]. fibre [Bel82]. fibre-reinforced [Bel82]. FICE [HW84]. field [AB81, And86, BT82, DG88, FSD81, Hai85, HK85b, Hew80, HSA89, KPB87, MMS86, Mas87, NM80, NTK80, Ram83, Sco88, SBAM82]. field-reversed [AB81, FSD81, SBAM82]. fields [HNK89, Hai85, MFW89, Mor81, OI87, Per81a, SMR89, Tuc89a, ZMC83]. filamentation [Kli87]. filled [RS84]. film [AH88a, Hun89]. filter [FT84]. filtering [Alp81, Den81]. Fin [Kui88]. Finding [MFW88, WBL86, Dye84, QQM89]. finer [BLT^{+}81]. Finite [Bas84, BK85a, BST80, Car89, CB84, CC84, DFP81, EB83, FW81, Gol86b, GW80, Kaw87, KJL86, LG81, LG80, LPS85, MEGH85, MNM85, Miy86, MC88b, NY82, RV88a, Sch88, SLO87, Tak85, VF80, Alv89, ABS82, BW88a, BCE88, BS83, BC87a, BC87b, BE80, BKP87, Blo82, BJ81, BNS83, Bue88, CKD87, CS82, Chr85, CL86c, CLS89, CFG83, CM80, Cul82, Cul83a, Cul83c, Cul89, DH82, DD89, DO86, DH89, DM85, DQS87, Duk84b, Dyk87, EW88, Fle83, For83a, Fou84, FWB^{+}81, Gar86, GDM81, GT85, Gla89, GP85, GS89, GF86, GIT87, GR82a, HTE81, HC87, Hei88b, Hof82, HMW86, JWB88, KMRM88, KOTK88, LLVL89, LLL88, LV86, LM89b, Lev88, LG85, Luc87, MC88a, Mac86, MES80, Man85, Man87, MK83, Mat86, MP89b, Mos85, Nav83]. finite [OS83a, Ohr81b, PC82, Pat86, Per87, Pic83, Pla81, Ros83, RMS85, SS81, SS82a, SPG89, SB88, She86, She88, SHK88, Shy85, SM86, SL85, SW81, SBSH84, Ste87a, Tak80, TL84, Tha80, Vij86, Vin83, Vin89, WW89, WMM86, WD84, WC85b, ZGB88, ZBG88]. Finite-difference [DFP81, GW80, MEGH85, Miy86, Alv89, BCE88, BC87a, BC87b, BKP87, Blo82, DH89, Gar86, GR82a, Hof82, Luc87, RMS85, SS82a, SW81, Tak80, Tha80, Vin83, Vin89]. Finite-dimensional [BK85a]. Finite-element [EB83, FW81, KJL86, BE80, CS82, CL86c, Mat86, Nav83, Pla81, TL84, WC85b]. Finite-sized [NY82]. finite-volume [Vin89]. first [ABW83, Ano87j, BR80a, ML83a, ML83b, MP80b, NR84, Wal84, Wil82, YRT80, ZBP83]. first-born [ML83a, ML83b]. first-order [MP80b, ZBP83]. fitted [Abe89, AD83, DH82, HC81a, LBS88, MT84b, MS80, Ohr83, TMCS85, TB89, TWM82, Uch83]. fitting [BB84b, MES80, MM88, MP89b]. five [AG80, HTE81]. five-point [AG80, HTE81]. fixed [Ana89, CH84]. Flame [Cho80, KS87]. flames [GS87b, OA85, Smo82]. flanged [CDP80]. flight [Yor86]. FLIP [BR86]. flooding [CS82]. flow [Abe89, Ana89, Ano80z, Ano82i, BM87a, Bak81, Ber85a, Ber83, BE85, Bra88, BD82, Bra84, Bra86, BR81, CH84, CHS88, CL86c, CDH81, CFG83, Dic83, DP86, EB81, EM81, Fer88, FS87, For85, FFG87, FW81, FF86, GGS82, GC87, GS87b, GMM81, GG86, GX89, GH87, GW80, Hal88, HD85, Hor87a, IT89, Iss86, JY83, JR85, Kan81, KP89b, KP89a, KK81, Kel83, KTH88, KH80, KGH80, Leo80, LG81, LV89, LBS88, Löt82, LC88, LG80, Lyt87, Mac86, Mah82, MP80a, MP89a, MZ83, MSK80, Mey81, MSB87, MML83, MTM81, O’B81, OI87, PG88a, Pat87, Pat84, Pea81, PVMZ87, PM80, PM89, Pul80, RC81, RA85, ROS85, RA87, RM87, RH84, RH88a, Rig81, Rob89a, RWC89, RM82, SS81, SB88, SC82, SSG81, Shy85, Soh87, SS84b]. flow [SW84, SZH85, Sum89, Tak85, TL84, Tan89, Ten82, Tut86, Uch83, Van84, Vij86, Whi87, WR84, WC84, YRT80]. flow- [FS87]. flows [AH85, And85, And89, Asa87, AD83, BDR82, BT82, BaB86, Ba89, BE80, Boe82, BR86, BS89b, CA85b, CPJ89, CB80, CBL81, Cul89, DS86b, Dar89, DS86c, Fer87, Flo85, Flo86, GGR89, GG88, GP81, Gre85b, Gre88, GH86, HMI83, Har80, HL85, HR84a, HP87, HW84, Hu89, Hun89, IGW86, Ker88, Keu86, KMA89, Kor88, KHTR89, LL81, Li80, Mac89, MGS82, MT81, MK80, Pan89, PP88a, PL86, PJ83, Qua81, QN86, RS81, RH88b, Ros80, Ros89, Sak81, SK83a, SH84b, SS85, SS82b, TR86, Vie88]. fluctuation [KD88]. fluid [Ada88, Ano80z, Ano81j, Ano87j, BM89, BAF84, BST80, BR86, Clo87, Duk80, FF86, GMM80, HN81, Hua81, Iss86, JH83, KH80, KGH80, Lam81, Mah82, MZ83, MK80, MTM81, Nav87, NY82, NY83, O’R84, Pat84, Per83, PM89, ROS85, RV88a, RE87, Sak81, SS88a, Sch82, Sch80b, SBT86, SH84b, Spe87, SR85, TLD80, Tak85, TS81, VTB89, Wan83, Wil80a, WR86, WC84, YP88, Zer85]. fluid-structure [Sch80b, TS81]. fluidized [CG89]. fluids [CO80, SH84c]. Flux [McD89, MC85, Mül88, SW81, ADR84, CLJ^{+}89, CM89, Clo87, Dic88, Dun85, Gan85, Gar83, HCG^{+}85, HPL80, KW87, KP83b, Mac89, MA84, MMR89, MS87b, Noh87, PG87, PGBO87, Pom88, Rom89, SM89, SMR89, Sto84, YRT80]. flux-conserving [HPL80]. Flux-corrected [McD89, MC85, PGBO87, SM89, SMR89]. flux-difference [Dic88]. flux-limited [Pom88]. Flux-vector [Mül88]. fluxes [LvLS89, MC88a]. Fock [QQM89]. Focusing [BNS83, CH88, OG86]. Fokker [AH88a, AOK81, BKP87, FWB^{+}81, Her89, LLPS85, MS86, MD80, Wie83]. folding [Kra83]. follower [Mej86]. force [Gol86a, SG86, TLD80]. force-biased [Gol86a, SG86]. forced [For86, JY86, KP85, KP86a]. Forces [BP88, vDR89]. Ford [FK89]. forecast [TGP80]. foreword [Ano82h, Edi82]. form [Har83b, MGM86, PC81b, Qin88, Sca87, SBBY83, Tuc89b, Wal84, WW89, WG86]. formation [MK86, Smi81, YH88b]. formed [FDS81]. forms [Gup81, Hor87a, MM81b]. formula [Aiz85, Eck89, Ioa82, ZBP84]. formulas [HC81b]. Formulation [Rob89a, Bas84, FF86, GS89, HK86, Laf89, LM89b, Lyt87, Spe87, SBSH84, WR84]. formulations [LV89, OS88, Vin89]. Forthcoming [Ano82s, Ano84t, Ano85i, Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano80u, Ano80v, Ano80w, Ano80x, Ano80y, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano83s, Ano83t, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano84r, Ano84s, Ano84u, Ano85h, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano85v, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano86u, Ano86v]. forthcoming [Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87u, Ano87v, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano88p, Ano88q, Ano88r, Ano88s, Ano88t, Ano88u, Ano88v, Ano89l, Ano89m, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s]. found [TDhLcSb89]. Four [Wil81, Ste84]. Fourier [CFR84, Tem83c, AI80, BK88a, Can81, GS85b, GC88, Hal81, HPBY86, Hei88b, Kli83, KK83, Kra83, LR80a, Lyn84, NM80, NS89, Pic83, SS88b, SM84, Tem83a]. Fourth [KJM80, Gup84, HTE81, Ano81j]. Fourth-order [KJM80, Gup84]. Fowler [FLP80]. fraction [Bar82, HRT80]. Fractional [Ste81, BP87, GM88, KM85, LM87a, MG89]. fractional-like [LM87a]. fractional-step [KM85]. fractions [Mar81]. FRAM [Cha81]. FRAM-Nonlinear [Cha81]. Frame [MK82]. France [Ano81j]. Frankel [CR84]. Fredholm [BR80a, Hag80]. Free [Yor86, AH85, Asa87, BY85, BE85, BG84, CL86b, ET88, For89, FW81, GDS88, GS85a, HN81, HK81, Hun86, Hun89, Keu86, LC88, Mey81, MSB87, PJ83, SS80, TLD80, Tuc89a]. free-boundary [ET88, GDS88]. Free-flight [Yor86]. free-surface [Asa87, BE85, For89, MSB87]. frequencies [LO86, WMS88]. frequency [ADR84, BFH86, SS88a]. frequency-dependent [ADR84]. frog [AA89, CR84, SM86]. Front [CGM^{+}86, DW81, Gre84b, Löt82]. Fronts [BM87b, OS88, She87]. fuel [BM87b]. Full [CLJ^{+}89, CHH^{+}86, DH82, HW84]. full-implicit-continuous-Eulerian [HW84]. Full-wave [CLJ^{+}89]. Fully [BSS82a, GP81, Smo84, SK86]. function [Boy84, CFR84, Dav86, DQ83, Dye84, EJL87, GM81, Hau82, JRS86, LM87a, MM88, MG89, O’R84, PT86, Paj89, PW85, Puo88, SSS^{+}84, ST88, SC81, Ter81c]. functional [Hoe83, MB80, Wag87, Wag88]. functions [Afa87, Afa89, ADL87, Bar82, Ber88, Boy86, Boy87a, Boy87b, BPD81, CD82, CRS88, Die83, Eyr87, Gre84b, GS85b, GM88, IA85, JR82, Law85, MMV81, Mar81, O’B80, OS83b, PC81a, Pie84, PA86, Pri89, RH87, SN83, SGA81, Ter81a, Ter81b, TB86, Vin83, Wal84, WK89]. fundamental [KF87]. Further [SC86, TY87b]. fuselage [SC82]. fusion [Wil80b]. future [Ano89a, Ano89b].

galaxy [Sel83]. Galerkin [AEBS86, ABW83, Bue88, CLS89, DO86, GS89, GF86, MD86, MP80b, MS87b, Nav83, RMS85, SSC81, SBSH84, Ste87a, Tre81, VJ89]. GAMM [Ano81j]. gamma [Sto84]. gap [DQ85]. gas [Abe89, CGM^{+}86, CW84b, DS86a, GW85, HR84a, HMS89, KK81, KS89, McL80, O’R89, Pul80, Rei81, RH88b, Roe86, SMR89, Tak86]. gas-dynamical [CW84b]. gasdynamic [SW81]. gases [CG85, Gla88a, Gla88b, GG88, HA85, LvLS89, MM85a, Mül88]. gauge [Hoe83, KKMP86, ML87]. gauges [ML85]. Gauss [Mar87]. Gaussian [LR80b, Pri89, RH87, Shi81, Ter81a]. General [FZ87, OF82, BCT89, BS81, DVW87, Duk84a, Duk85, DCA89, GDM81, GM82, HP87, LD89, Lup87, MEK86, Nor80, OC83, Pir80, Smi88, SB84b, CA85a]. Generalised [MP80b]. generalization [HN82]. Generalizations [ST89a]. Generalized [BaB86, Ter81b, Ba89, CLM88, CS81b, FWHC86, GP89, GKL^{+}87, GS88, Hau82, Ioa82, Kal84, Mor89, Pri89, Ram85b, Ram86]. generate [Hos86]. generated [MNM85]. generating [ADL87, CP88b, Hoe83, Sal86, Spa85, Wik89]. Generation [Hai85, Uch83, And89, BM87a, BSS82b, CW85a, Eis82a, Eis82b, Flo85, Flo86, FK89, FWHC86, GE88, HC81a, Ino83, Ino85, KL89, MEK86, MD88, MT84b, MS80, Ohr81a, Ohr81b, Pea83, Per81a, She86, Sid85, Ste86, TMCS85, TB89, War86, Yor86, vDS80]. generator [ABDV81, KS81, PHD89]. generators [CM86, Hal89, PP88b]. geodesic [Sha81, Tha80]. geomagnetism [CP88a]. Geometric [LB87, War86]. geometries [KP89a, Ors80, OC83, Tuc89a]. geometry [CLJ^{+}89, CHH^{+}86, KP86b, KW85, SM84, SN88, WG86]. geophysical [Nav87]. given [Hai85]. glaciers [YHS86]. Glimm [FH81]. Global [PC81a, ST89b, Hew80, ZMC83]. glow [Bal85, BCE88]. goal [Ng81]. Godunov [BDS88, BCT89, BAF84, Duk85, DCA89, GW85, Vij86]. Godunov-type [BAF84]. Goldstone [CP88b]. Gordan [KTT89, Pap84]. Gottlieb [FT84]. governed [Li80]. GPST [CL87]. gradient [AG80, BDJS86, Ker80, MM85b, ROS85, SHB82, WVP82]. gradient-type [BDJS86]. gradients [GDM81, KE87]. grading [EW88, EWR88]. Graetz [KH84]. grain [Bas84]. Gram [Tob86]. Graphic [Bow84]. graphs [FK89]. gravitational [Bow84]. gravity [BP80, CS82, GC89, Lim80, Pro81a]. gravity-wave [BP80]. Green [JRS86]. grey [Lar88]. Grid [Bor86, CH88, GS85a, Ino83, Ino85, Pik87, RA81, Abd87a, Abd87b, BSS82b, BS83, BK88a, BSSV88, BLT^{+}81, CLF84, Far80, GE88, HH83, LB87, MEK86, MD88, MT84a, MT86a, NY82, Pet87, Pic86, Pro81b, Rai86b, SJ88, Sco85, SOS89, Ste86, VBSS89, War86]. Grid-free [GS85a]. grids [ADL87, CFR84, Cul82, DD87, Gar83, Got88, Lev88, MM89, Neu81, PH88, Pik87, TGP80, TB89]. Grimm [JDK86]. Group [Tre83, BGMT89, KSD81]. groups [Bar80, GWZ85a]. Growth [RQ89]. Guidelines [MvdV81, Pes81]. Guiding [KO86, SSS^{+}84]. Guiding-center [SSS^{+}84]. Gustafsson [Tre83]. Gyrokinetic [Lee87].

half [SKA86]. half-complex [Tem88c]. half-space [SKA86]. Hall [HM89]. Hamada [BS86b]. Hamilton [Hit80, OS88]. Hamiltonian [Aiz85, HR86, IA88, KPB87]. Hamiltonian-conserving [IA88]. Hamiltonians [HBR82]. Hankel [BK85b, GM81]. hard [PW85]. harmonic [CB89, DG88, Dil85, EHD89, SN83, SB87a]. harmonically [CFG83]. harmonics [BS84]. Hartree [QQM89]. Hausdorff [Cho82]. having [DS86c]. heart [MP80a, MP89a, PM80, PM89]. Heat [LS85, Din86, FF86, GMT^{+}81, HK83, IM88, JT86, LG85, Noh87, PG87, She86, SS89]. heatbath [KKMP86]. heated [Bon84, Mac86]. heating [BM80a, VE87]. heavy [Kra83]. heavy-ion [Kra83]. Hele [AH85, Mey81]. HELIAC [SM84]. helical [DG88, HHS89]. helically [GDS88]. Helmholtz [BGT83, HLAD84, JRS86, MS83]. HELP [Sch80a]. HERA [BHM^{+}83]. Hermite [Boy84, GP80, Pri89]. Hermitian [HN88, Nex87, SMRS85]. heterogeneous [SV89]. hexagonal [Pic86, PH88]. hexahedral [Duk88]. High [ASIO86, BL88, BM85, BFH86, Eis82b, Gar86, GGS82, Har83a, HC81b, IM88, Lev88, MIK88, MA84, SS88a, Zal81, CJW85, Gla89, GW85, HEOC87, Kob85, MS83, Mor82, Mun88, Pat86, SV89, Ten82, Tur80, WW89, Yam83]. high-density [Mor82]. High-frequency [BFH86, SS88a]. High-order [ASIO86, BL88, MA84, GW85, Pat86, Tur80]. High-Re [GGS82]. high-resolution [WW89]. Higher [BM87a, BCT89, GM88, BDS88, GH87, Kui88, Lup87, MD81b, OF82, Rob89a]. Higher-order [BM87a, Lup87, MD81b, OF82]. highly [Dav89, Ixa80, JT86]. hoc [Ehr81]. Hodgkin [CCCS81]. hole [Sha81]. homogeneous [Adh81b, Bas83, KD88, Moh84, YP88]. hopscotch [BR80b]. horseshoe [Tan89]. hot [SB87b]. Hugenholtz [CP88b]. Huxley [CCCS81]. Hybrid [BHS87, MSD87, Mru87, FSD81, FWCW86, GIT87, Har82, Hew80, HSA89, KTH88, Mas83, MM85b, VE87, ZC86]. hydrocodes [HN80]. hydrodynamic [Boy85, Ker81, SHK88, TR86]. hydrodynamically [Vie88]. Hydrodynamics [BC89, BDR80, For89, FWCW86, FOF88, GM82, LM86, Smi88]. Hyperbolic [BO84, Cou85, RH84, RH88a, AM82, BL85, BWY82, BCT89, CQ87, Cul82, DQS87, Gar78, Gus80a, Gus80b, Gus82, HTE81, Har83a, HH83, Hen87, JT86, MA84, McD89, Mor81, MP80b, Mun88, TNY85, TY87a, TE86, Tho87, Tur80, Vat87]. hyperbolic-type [TNY85]. hypergeometric [Ter81b]. Hypersonic [RH88b]. hypersphere [NSJ85]. hyperthermia [LPS85]. Hyperviscosity [PP88a]. hysteresis [HSV88].

ICE [MT81, RD83, Smi88]. ICE-like [RD83]. Ideal [GDM83, HG86, BW88b, MZ83, Pul80, WG86]. ideal-gas [Pul80]. Identification [HSV88]. II [Abd87a, ACM88, Bar80, CA85b, CS81b, GMM80, HTE81, JWB88, KP86a, KGH80, LM89a, ML83b, MP89a, MK86, ML87, SO89, TA84a, VSS86, ZBG88, ZWH84]. III [CLS89, HEOC87, TA84b]. imbedding [KR84]. Immersed [PM89]. Implementation [FLP80, HMS89, PHD89, Tem85, ZBP84, Ada88, CW81, HR86, HMW86, KHTR89, O’R89, PT82, SO88, SO89]. Implementing [Hit80]. Implicit [Alp81, BKLT83, BDR82, CLF82, Cul89, Mas81, Mas83, Oey82, YWH85, AM82, AHP81, ADR84, AB85, BWY82, BSS82a, BF82, BDR80, BM80b, BF86, Bro89, Bro85, CFT82, CLF84, CLHP89, CBG86, FS89, FC84, FS86, FWCW86, GHC^{+}86, HK85a, HS86, HM89, HTE81, HL87, HW84, Hu89, HL84, IGW86, KR89, LCF83, Lev82, Mas87, MZ82, MV85, PGBO87, PC86, PC81b, SBM^{+}87, Sco88, She86, SSSR83, Smi88, Smo84, SL85, SM89, SMR89, Tan88, TB82, TR86, Van86, WBF86, Yee87]. implicit-explicit [FWCW86]. implicitly [Iss86]. Improved [BFT85, FR87, RA85, Din86, GPS88, IHM81, LC89, OS85, Rom89, Son85, TE86]. Improvement [EFCS88, CL87, Mat86]. improvements [Bas83, Wil81]. improves [WC85a]. Improving [ADG86, BR80a, FC84]. impulse [CCCS81]. in-place [Tem85, Tem88c]. inadequate [CL81a, Nol85]. Including [DGKW84]. Inclusion [RA87]. Incomplete [MvdV81, Ker80, Ter81c]. Incompressible [PG88a, Abd87a, And89, AB84, BCG89, BE80, BP87, BS89b, CPJ89, CBL81, Fer87, FB88, FW81, GGR89, GGS82, GS89, Gre85b, Gre88, GG86, HMI83, HN80, IGW86, KM85, KHT87, KHTR89, LG81, MK80, MTM81, Nat89, PM89, PL86, QN86, Riz88, Ros89, RWC89, Soh87, SG88b, SBSH84, SN88, Tak85, Tan89, Ten82, TF89, Tur87, Tut86]. incorporating [Mar87, VF80]. Incorporation [Mei89]. increasing [BKP87, KT86]. independent [Hal89, HG88]. index [Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano82d, Ano82e, Ano82f, Ano82g, Ano83e, Ano83f, Ano83g, Ano83h, Ano84d, Ano84e, Ano84f, Ano84g, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k]. induced [CM82, SS85]. inductance [HGS88]. induction [But88]. inductive [CF82]. inelastic [AP88, MT86a]. inertia [Hew80]. Inertial [MK82]. Inertial-Frame [MK82]. infinite [ANAN84, Boy87a, Boy87b, GP89, HRT80, KL80, Mav89, WC85b]. infinite-domain [Mav89]. infinity [Kru82]. Influence [Blo82, Nor89, PC86]. inherent [AA89, Gol86a]. inhomogeneous [CL86a, SW85, Wol81]. Initial [Cou85, NTK80, Pom88, Str80, AM82, Bow84, Dav81, DD87, NR80b, Rob85, ZOWZ88]. Initial-Boundary-Value [Cou85]. initio [MW87, MD81a, Tak81]. injection [FDS81, GMT^{+}81, Law88]. inlet [Ino85]. inner [KR86]. innovation [Zab81]. instabilities [BM80a, Boy85, CH88, HT86, SKS82, SM86, Slo88, Smi81, Sto86, SLO87, SA81]. instability [Bra88, BNS83, EM80, HSA89, MZ83, MJ87, Ota88, Sid85, Try88, VSS86, Zuf89]. integrable [LM87a]. Integral [BGM87b, HM82, QN86, Adh81a, Adh81b, AEBS86, ABW83, BS86a, Bal85, BR80a, Bel82, BST85, CDP80, Col81b, CRS88, EHM81, Eyr87, EW88, Fet84, Fre83a, HA85, IHM81, Ioa81a, Ioa81b, Ioa82, iKI85, Kel83, Kel84, Lim80, LC88, MGM86, Mer86, Mos86, Nav83, RH87, Rok85, Rok89, SP87a, Sch82, SH89, Sid85, SB87b, Str89, TDhLcSb89, Zer85]. integral-equation [Bel82]. integrals [Afa87, Afa89, CLM88, Cla87, Dan80, Dav89, DFG81, FP82, FZ88, Hal87, JR82, KK89, LR80b, LR81, Lup87, MD81a, ML85, ML87, Pri89, Puo88, Tak81, Ter81a, Wag87, Wag88, Wil82]. integrands [Dav89]. Integrated [Chr87]. integrating [Alz84, Ber88, LR81, TC87, Wol85]. integration [Aiz85, ACM88, BL88, CLF82, EB80, GHC^{+}86, GR82a, GR82b, KK89, KK88, KW86, Kru82, LL81, Neu83, NSJ85, Nor80, OS85, Per81b, SSM83, Sha81, Ves82, Wol81, ZBP83, ZBP84]. integrations [Dri88b, Pea83]. integro [Pro82, Ram85a, TY87b, Vat89]. integro-differential [Pro82, Ram85a, TY87b, Vat89]. integrodifferential [Fre83b, MPPS84]. Interacting [BP88]. Interaction [Hau81, CW85a, Hal88, Kra83, LLL88, MV83a, MPPS84, SS84a]. interactions [CDS85, Mac81, MMR89, Sch80b, TS81]. Interactive [BR81]. interesting [BB86b]. interface [Fou84, GMM81, MC88a, Wes88]. interfaces [BaB86, Li89, Smi88]. interior [BW81, Boy85, Mer86]. intermediate [IT89]. internal [GP81, O’B81, PL86, Soh87]. International [Ano82i]. interpolated [TNY85, TY87a]. Interpolating [Ber88, KOTK88]. interpolation [Hor87b, Ioa82, KW86, Mon85, SW88, OvOH89]. interpolations [ASIO86, BLT^{+}81]. interpretation [Tre83]. interval [Boy87a, Boy87b]. intrinsic [FDS81]. introduction [HL86]. Invariance [Vic86, Aiz85, Alz84]. invariant [GWZ85a, GWZ85b, KR84, Lab87, MT81]. invariants [Nav83]. Inverse [ADKM82, Dun81, CL83, CL84, DS86b, DJT80, DH84, GIT87, Hag81, HH86, Tob86]. Inversion [DGKW84, CL87, DFG81, DS89, Kai83, Lan82, NMG81, SP87b]. investigation [IT89, SE87, TM88]. inviscid [Fer88, FS87, GG88, Lav88, LvLS89, Pul80, SW81, SS84b]. involving [Afa87, Afa89, ABW83, AHH87, CD82, Mac81, MGM86, MK80, RH87, Sid85, SMR89]. Ion [WWM^{+}81, CDS85, FSD81, FDS81, Kra83, MSD87, Ota88, Whe81]. ion-cyclotron [Ota88]. ionic [Mac81, MPA88]. ionised [SBBY83]. ionosphere [YRT80]. ions [WMM86]. irregular [Cul82, PSV89, Pik87, SW88, TGP80]. Ising [Har81, HSSC83, HK81, Lub88, Sad84]. islands [MFW88]. Isolas [KSM82]. isoparametric [KP86b]. isothermal [JH83]. isotherms [Mer80]. isotope [Ema81]. isotropic [KD88]. issues [Ano89a, Ano89b]. iteration [Boe82, RA85]. iterations [SPG89]. Iterative [Adh81b, EHM81, Hir89, RRS89, Tob86, vdV81, Adh81a, BGT83, BDJS86, CSW85, CL84, CX86, DJT80, Duk85, FLP80, GPS88, Ioa81a, IGW86, KE87, Kre88, MSW87, NV89, Ret82, RL89, Tan86]. IV [TA88].

Jacobi [OS88]. Jacobian [ST89a]. jet [BST80, Lim80, SBT86, Chr87]. jets [SE87]. joint [HP87]. Jones [RQ89]. Josephson [MB80]. junctions [MB80].

Kadanoff [HK81]. Kansa [RD83]. Kantorovich [Rag82]. keeping [Osg81]. Keith [Kil87]. Keller [Jon81]. Kelvin [BL82, Hen81, MW84]. Kernel [GM82, Fre83a, Fre83b, HS88b, OC87]. kernels [BM85, Ioa82]. kind [ABW83, BR80a, Wil82]. Kinetic [Ber85a, Gla84a, Mus81, Sch80a]. kinetics [Bro80, KK81, Sad84]. KKR [Che84]. KKR-method [Che84]. known [OA85, WH87]. Korteweg [AI80, AA89, HPBY86, MS87a, NS89, Slo88, TA84b, TA88]. Kreiss [Tre83]. Kutta [CLS89, Wil80c].

L [Sch88]. L-shaped [Sch88]. Ladik [DF88]. Lagrangian [Aug84, BP87, Clo87, FWCW86, FOF88, Ker81, LB87, Neu81, Oli80, Per83, SHK88, SS85]. Laguerre [GM81, Mav89, Ram85a]. lake [BC87a]. lakes [BS81, Rag82]. Laminar [Pat84, Ana89, Ano82i, GC87, GS87b, HK83, HD85, JY83, JY86, KP89b, OA85, Smo82, Tan89, vDS80]. Lanczos [BBG^{+}87, CW81, Kal84, MSW87]. Lanczos-type [Kal84]. Langevin [EB80]. Laplace [Are81, BS84, RN86, Skö82, Wig88]. Laplacian [Sch88, TB89]. Large [Ker89, BJ84, CKR87, CS81b, CW81, Fuc89, GPS88, GDM81, GKL^{+}87, GS88, HN82, Hir89, HN88, Hor87a, KTH88, Kos84, LBO^{+}89, LCF83, MM81c, NV89, Nex87, O’B80, OJJ88, Pea83, Ret82, Riz88, SPG89, SGT88]. large-eddy [KTH88]. Large-scale [Ker89, Riz88]. largely [MT81]. Larson [Mon81]. laser [Bon84, CW85a, CW80, DKD85, DKWD89, SS84a]. laser-molecule [CW85a]. laser-plasma [SS84a]. lasers [BY85]. Lattice [Mac81, Ber88, BGM87a, HA85, HMS89, Hoe83, KKMP86, KS89, PT82, ZMC83, dFLR85]. lattice-gas [KS89]. lattices [CKR87]. law [BKP87, Hit80, Mar87, MW84]. laws [BS83, BDS88, BCT89, CLS89, HTE81, Har83a, HH83, Har83b, Hen87, LV89, MA84, McD89, Mun88, Vin89]. Lax [BLT^{+}81]. layer [GHK88, Hal88, HO80, HW81, KTH88, MM81c, NE82, Sum89, TL84, WC85b, vDS80]. layered [SS88a]. LBI [KG87]. LCAO [Tak81]. lead [SA81]. leap [AA89, CR84, SM86]. leap-frog [AA89, CR84, SM86]. leapfrog [CH88]. Least [Mun81, BB84b, CFG83, MM88]. Least-Squares [Mun81, BB84b]. leftmost [GPS88]. Legendre [Afa87, Afa89, Hit80, OS83b, Sca87, TE86]. Leibbrandt [ML87]. length [CH87, Van84]. Lennard [RQ89]. Lennard-Jones [RQ89]. letter [Kan83]. level [Eis82b, SC85]. levels [Kob85]. Leverett [Pro81a]. lie [Mil89a, Pap84]. lifting [Lyt87]. light [ML85, ML87]. light-cone [ML85, ML87]. like [KM80, LM87a, RD83, TST^{+}86]. limitations [BS84]. limited [ADR84, HN80, MS87b, Pom88]. limiter [Clo87, YNAS86]. limiting [WOZ84]. Limits [TST^{+}86, GS87b, Vie88]. Line [Hei88a, BF86, KPB87, MK86, RL84, RMS85, SS80]. Linear [MvdV81, Zuf89, vdV81, Aiz85, ABW83, Alz84, BDJS86, CFH^{+}82, CR84, DFP81, DO86, EG86, FN80, FK89, Hir89, MIK88, Mat86, Per87, PP88b, PC86, Rap80b, SGT88, SB84a, SB84b, TL84, Var84, WVP82]. linearised [Gla88a, Gla88b]. linearization [Ohr81b]. Linearized [Mun81, BM80b, For83b, For84, FSD81, JL86, KJL86, LL81]. linearly [Tak81]. lines [BB86a, Car80, DKD85, DKWD89, Ohr81a, Str80]. Liouville [Boy85, EJL87, LBM87, MV83b, PT86, Pru88]. liquid [Duk80, SE87, TDhLcSb89]. liquids [Gol86a, SG86]. List [Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano80u, Ano80v, Ano80w, Ano80x, Ano80y, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano82s, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano83o, Ano83p, Ano83q, Ano83r, Ano83s, Ano83t, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano84r, Ano84s, Ano84t, Ano84u, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano85v, Ano86k, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano86u, Ano86v]. List [Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87u, Ano87v, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano88p, Ano88q, Ano88r, Ano88s, Ano88t, Ano88u, Ano88v, Ano89l, Ano89m, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s, FK89]. Liu [Kos84]. loadings [NTK80]. Local [BC89, And86, BV82, BV86, Cha86, CLS89, Got88, OC87, SJ88]. localization [SW88]. Locating [Dav86, IA85]. locomotion [FP88a]. Log [Mru85, TL84]. Log-derivative [Mru85]. log-linear [TL84]. logarithmic [Fre83a, MT84a, MT86a]. Logical [Bor86]. Long [PP88b, CCCS81, HS86, Mac81, ZBP84]. long-range [Mac81, ZBP84]. look [VSS86]. lookback [ZMC83]. loop [LWP89]. loops [HSV88]. Low [Wil80c, BCE88, Bra88, Cul83a, Cul83c, IT89, PV83, QQM89, Tur87]. low-lying [QQM89]. low-order [Cul83a, Cul83c]. low-speed [Bra88]. Low-storage [Wil80c]. low-temperature [PV83]. lower [VE87]. LTE [SC85]. LU [OK86]. lumping [Nav83]. lying [QQM89]. Lyness [IA85]. Lytton [Lyt87].

M [BBGN80]. Macroscale [Tan88]. Macroscopic [EM80, Har82]. MACSYMA [Lo85]. magnetic [DG88, KPBRF83, KPB87, MFW88, MFW89, NM80, NTK80, Per81a, Ram83, RG88, Sco88]. magnetized [BKLT83, Tan88, Wei85]. Magneto [MHK86, CF82]. magneto-inductive [CF82]. Magneto-Plasma [MHK86]. Magnetohydrodynamic [BLT^{+}81, AB85, BS86b, BB80, HH85, HHS89, HK85a, HS86, HM89, HW84, Hu89, SK80, SBC84, SBM^{+}87, Sto86, TLD80]. magnetohydrodynamics [BW88b, KR89, VTB89]. magnetostatic [KO86]. maintaining [LBO^{+}89, Per83]. Mandelstam [ML87]. manipulation [Gla84a, KTT89, SR85]. manipulator [Lo85]. many [CP88b, Tak86]. many-body [CP88b]. Mapping [SB87a, CY85, Flo85, Flo86, Ino83, Ino85, MOI81, MZ80, RL83]. mappings [FZ87, FZ88, O’B81]. maps [MZ83]. marching [PL86, Soh87]. Markov [Pes81]. mask [BS89b]. Mass [Fou84]. massive [MH89]. Massively [BP88]. master [PV83, AH88a]. matching [AH88b, Lo85, RS84]. material [Bel82, BaB86, Hal88, MC88a, WHW87]. materials [MMS86]. mathematical [Fog84, Lav88, Zab81]. matrices [Bog88, CW81, GPS88, GWZ85a, GWZ85b, NV89, Ret82, Tuc89b]. Matrix [vdV81, AOK81, CS81b, DVW87, DS89, Fuc89, GP89, Hut84, IMT^{+}89, KSD81, KOTK88, Kos84, KHTR89, Lab87, NR85, OJJ88, Pap84, PG88b, Pes81, SMRS85, SHB82, YH88a, YH88b]. maximally [IR87]. Maximum [CH87]. Maximum-length [CH87]. Maxwell [CB89, CW85b, LM89b, LPS85, WCR^{+}88]. Maxwellian [WL84]. MCSCF [BKH86]. mean [BB88]. means [GRP82, MMV81, Mil89b]. measure [GWZ85a]. mechanical [CKD87, MPPS84, SAR82, WC85a]. mechanics [Ano81j, BK85a, Kan84, PW85]. mechanism [BNS83]. Media [MK82, CB80, GK81, HHN^{+}86, Mih85, MK86, PG88a, Rob89a, SW85, ST88]. medicine [DKWD89]. medium [EB81, PG87, SBBY83, She87, SV89]. melt [SLO87]. membranes [MPA88]. memoriam [Kil87, JDK86]. Menger [KD80]. Mesh [BO84, BC89, AF86, Aug84, AP85, BCE88, CW82, CFL86, DH82, Duk84b, Eis82a, EW88, EWR88, Fou84, Fre81, FOF88, GP85, Hag80, HD85, KM80, KOTK88, LW89, LJ81, Luc87, MC85, Per83, Wil81]. mesh-one [Aug84]. meshes [Duk84a, FS89, Hof82, Ram85b, Ram86, Sal86, SW88]. metal [Bas84]. metallic [LBR81]. metals [MV83a]. Method [CW84b, MHK86, Abe89, And85, And86, ABDV81, ADL87, Aug84, Baa81, Bai80, Bak81, BBG^{+}87, BCE88, BGT83, BGMM89, BS83, BST86, BDS88, BCG89, Bel82, BaB86, BC87a, BE80, Ber85a, BB86a, Bin85, Bog88, Bon84, BJ81, BP87, BE85, BF82, BR86, BR81, BDR80, BS89b, BP80, Bue88, BG85b, Car80, CW85a, CM86, CW82, CJ86, CS82, CC84, CX86, CL86b, Che84, CL80, CLS89, Col81a, CBL81, Cur82, CR84, DS86b, Dar89, Dav83, DKD85, DKWD89, DF88, DJT80, DH82, DO86, DW81, Dic83, Dic88, Din86, Duk84b, Duk85, DCA89, Eas87, Eck89, Ehr81, EFCS88, ET88, FN80, FFS82, Fer80, FC84, FH81, Flo85, Flo86, Fog84, FP88b, FW81, FR87, FWHC86, FWCW86, GTD89, Gau88, GDM81, GGS82, GHK88, GS87a, Gla84b, GMM81, Got88]. method [GF86, GP81, Gre85a, GRA83, GG86, GC88, GX89, Hai87, HS88b, HH85, HK85a, HS86, HK86, HC81a, HPBY86, Her89, Hew80, HN82, HN81, HD85, HGS88, HMW86, HK81, HGM86, IBCS87, IA85, Ixa80, IR87, Ixa87, Jac89, JT86, Jon81, KR86, Kai83, KP83a, KF87, KK89, Kaw87, Kel83, Kel84, KKMP86, Ker80, KM85, Kli83, Kli87, KK88, KP86b, KOTK88, KK83, Kos84, Kre88, KH80, KHT87, KHTR89, LLVL89, LBR81, LP83, LM87b, Lar88, Law85, LW85, LW86, Lev82, Lim80, LO87, Löt82, LC88, Luc87, Mac86, Mac89, Mah82, MZH85, MM89, Man87, Mar87, MM81a, MP84, MP80a, McD80, McD89, MP89a, Mel86, Mez81, MT84b, MM85b, MD86, Moh84, Mon81, MG83, MML83, Mos85, Mru85, NM80, Nav87, NR85, ND89, O’R84, OK86, OI87]. method [Ohr81a, Ohr83, OG86, OS85, PSV89, Pat84, PM89, Pol87, PA89, PL86, Pro82, PJ83, Puo88, Qin88, RS85, Ram85a, RC81, Ram83, ROS85, Ret82, Rob85, RCK87, Rom89, Ros80, Ros89, RL89, SS81, SMRS85, SSM83, SH89, SBBY83, SH84a, SM84, SSSR83, Shi86, SB84a, SC82, SSG81, SP87b, Smi81, Spa85, Ste86, Ste87b, Sto84, SV89, Str80, SN88, SH84c, Ta’87, Tak81, Tak86, Tak85, TS81, TNY85, TY87a, TE86, Tan86, Ten82, Tha80, TGP80, TB89, Tob86, VBSS89, WP83, Wie83, Wig88, Wik89, Wil82, WC85a, WC85b, WVP82, WP88, WG86, YH88a, YNAS86, YRT80, Zeb84, Zer85, ZBP84, Lyt87]. Methods [BM89, MZ80, vdV81, ADG86, AH88a, ABW83, Alz84, Ano81j, Ano82i, AG80, BW88a, Bar87, BL88, BL85, BM85, BCT89, BK88a, BSSV88, BV82, Boy82, Boy85, Boy87b, BFT85, BGMT89, CQ85, CQ87, Cha88, CL81a, CHT84, Cla87, CL86c, CM80, Cul82, Cul89, Dav89, EB83, For84, FS86, FD82, FT84, Gar86, GS85a, GM82, Gol86b, Gre84a, GW80, Hal81, HH83, Hei88a, Hei88b, Hir89, HMR82, IHM81, ICP80, IM88, JWB88, Jon85, KE87, KP89b, KP89a, KPD81, Kra83, KH84, Kui88, LLPS85, LLL88, Leo80, Lev88, LWP89, LCC^{+}86, MEGH85, MES80, Man85, MS83, MIK88, MK83, MD88, MO80, MA84, Mon89, Mor89, MP80b, MS87b, MV85, NR80b, NS89, Ors80, OF82, PP88a, Per85, Pes81, Pic86, PH88, Pul80, Rig87, Rob89b, Sal86, SSC81]. methods [Sch82, Sco85, SG88a, Sha81, SAR82, Smo82, ST89b, SW81, Ste81, SW84, SZH85, SSF83, TDhLcSb89, TAJ^{+}81, Try89, Tur80, Tur87, Tut86, Var84, Wan83, ZWH82, ZWH84, Zeb87, ACM88]. metric [DJT80]. Metropolis [Gol86a, SG86]. MHD [AB81, AB84, BHM^{+}83, CHH^{+}86, CC87, CS81a, EH85, Eas87, GHC^{+}86, GRS89, GDM83, GIT87, HG86, HMJ86, HH86, HT86, HW84, JR85, KJL86, LCC^{+}86, Oli80, SKS82, WG86]. microscopic [EM80]. Miller [Ter81c]. minimal [CQ85]. minimizing [LW89]. minimum [Tem88b]. minimum-add [Tem88b]. mirror [ABS82, MS86, SBAM82]. mixed [AG81, Bue88, LM89b, Mac86, Man85, OA85, Tem83a, Tem83c]. mixed-radix [Tem83a, Tem83c]. mixing [Hos86]. mixture [MM85a]. MO [Tak81]. mode [HSA89, SB87b, TST^{+}86]. Model [Mun81, BC87b, Ber85a, BPS85, CW80, CG89, Duk80, FP88a, Fog84, For83a, Gar83, GC89, Gla89, Gla84b, HPP^{+}82, KD88, KO86, Kli87, LV86, Lee87, MH89, MS81, MRH83, Nav83, NB89, NMG86, O’R89, Ohr81b, SKS82, TL84, TS83, War86, Wit84]. Modeling [PM80, JPD86, Kli83, Pet87, VE87, ZMC83]. modelling [Cul89, Eas87, MS87b, RRS89]. Models [SW84, BH87a, CBG86, Cul83b, EM80, HRS80, HK81, KG85, Pla81, RH84, RH88a, Roe86, SBT86, SC81, TGP80, VM86]. modes [AA89, BS81, BK87, CD80, FFM82, HCH^{+}81, WD84, WG86, Zeb87]. Modification [Kos84, IA85]. modified [BE85, BS89a, Dye84, EHD89, GTD89, GM85, HK81, LBO^{+}89, MS87a, MLB87, SBBY83, TA88, Wal84]. modulational [Slo88]. modules [Tem88b]. moduli [CX86]. Molecular [BP88, MW87, MHM81, Rap85, SMO85, And83, KD80, KK83, LBO^{+}89, Pri89, Rap80a, Son85, SH84c, TAJ^{+}81, Tox82, Tox83, WH87, WVP82]. molecule [CW85a, Mru87]. molecules [LBO^{+}89, RQ89, Son85, SH84c]. moment [Mas81, Mas83, WBF86, ML83b]. Moment-theory [ML83b]. moments [Gan85]. momentum [Dec84, DA86, HG88, MEGH85]. monoenergetic [Gan85]. Monotonic [Bor86, LB87]. Monte [Gol86a, Abe89, Ada88, BHS87, Bin85, Boo81, Boo82b, BF86, Bro89, CBG86, CDS83, DA86, Dun81, EFCS88, EB80, FC84, GRP82, HM80a, Har81, HP87, HPP^{+}82, HM80b, HSSC83, JR81, KKMP86, KL89, KT86, LM87b, Mez81, PRT83, Pes81, RH88b, RV88b, Sad84, SG86, VL84, Ves82, VR86, Wag87, Wag88, WC85a, WP88, WPW89, Yor86]. motion [CL86a, CKD87, Tha80]. motions [Lam81, RV88a]. Moving [BW88a, JWB88, MK82, AF86, BSSV88, CCCS81, Duk84b, Fou84, GDM81, Gla89, HMW86, MK86, MC88b, NY82, OI87, PG88a, Smi81, Smi88, VBSS89, WHW87]. MP/24 [GKL^{+}87]. Muller [Dye84]. Multi [JR81, Riz88, Sco85, BS86a, CL84, HM80b, MO80, SC85, SJ88, TY87a, Tan89, Wag88]. Multi-cell [Riz88]. multi-connected [BS86a]. multi-dimensional [MO80, TY87a]. multi-domain [Tan89]. Multi-grid [Sco85, SJ88]. multi-level [SC85]. multi-parameter [CL84, HM80b]. Multi-spin [JR81]. multi-step [Wag88]. multichannel [Mel86]. multicode [NR84]. multicomponent [LW85, Rob89a]. multidimensional [BK88a, HW84, Kan81, KG87, MD81b, Roe86, Smo84, SC86, Wil80a]. Multigrid [CW89, Hei88b, Kor89, Mul85, Ta’87, Boe82, BDR80, BFT85, BGMT89, DH82, GGS82, GRA83, Hei88a, Hun86, J.82, Kor88, LBS88, Lon88, Mul89, Ohr83, OF82, Sch82, SC82, SZH85, TF89, Van86, Wes88, ZWH82, ZWH84]. Multigroup [GS82, WHW87, Wil80b, Cla87, GS83, HC87]. multilevel [MM89]. multiminima [Eck89]. multiphase [CDH81]. Multiple [CL86a, Ema81, Hal89, SW85, BDS88, GDM81, HL86, KTH88, MM88, NE82, SH84c]. multiple-body [KTH88]. multiple-processor [HL86]. Multiple-water-bag [CL86a]. multiplicative [HR84b]. multiplicity [Yam83]. multiply [WR86]. multipole [CM82, RN86]. multiregion [MS86]. Multislab [GS83]. multistep [Aiz85, Alz84, Hu89]. Multitasking [ZH88]. multivariate [Ter81a].

nacelle [Ros87]. NaCl [AH88a]. Nanbu [BS89a]. national [Ano87j]. Natural [CB84, CJW85, Phi84a, Ioa82, LA85]. Navier [AG81, Abd87a, AHP81, BDR82, BCG89, BGR80, BP87, Cha88, CC84, CPJ89, Cul83a, Cul83c, DH89, FF86, GGR82, GGR89, GGS82, GS89, Gup81, Hal81, HK86, HG84, KM85, KP86b, Kor89, KHT87, Li80, Lon88, Luc87, MZH85, MML83, Nat89, Ng81, Nor89, OK86, Rap80b, Ros83, RS80, SSSR83, Soh87, SG88b, SBSH84, SN88, Tak80, Tan89, TB82, TF89, Tre81, VM86, Van86]. NBS [Ano85w]. Near [Bor86, CL86a, HL85, LB87, LR80a, McL80, PRV89, SS80]. nearly [Gol86b, HC81a, ROS85, TR86]. nearly-implicit [TR86]. neglect [BD83]. Neighbor [Bor86, LB87]. Neoclassical [LNS87]. nerve [CCCS81]. Nesting [Tem89]. network [AHP81]. Neumann [Abd87b, KP83a, Mer86, SB82]. neutral [AEG80, CDS85, GMT^{+}81, HC87, HPP^{+}82, HSA89, ONM83, Tam81, WPW89, ZC86]. neutral-ion [CDS85]. neutral-particle [HPP^{+}82]. neutron [CRS88, Gan85, GF86]. Newton [BL88, Boe82, Mel86, SSG81, Zer85]. Newtonian [BE80]. no [CW81]. nodal [HD85, Kaw87]. node [Got88, Mos85, ZOWZ88]. noise [CW85b]. noisy [Baa81, BB83, Nol85]. Non [vdV81, Abd87a, Abd87b, BE80, CFR84, Clo87, DFP81, DVW87, Duk85, Dye84, EG86, FRRT87, GP85, HEOC87, IGW86, KG89, Mac81, Mac89, Mez87, Mor81, MC85, ONM83, OC87, Pan89, PSV89, Per87, PC86, Ret82, Rob85, SC85, SO88, SO89, SMR89, SB84b]. non-analytic [Dye84]. non-Boltzmann [Mez87]. non-circular [ONM83]. non-cubic [Mac81]. non-equidistant [GP85]. non-equilibrium [Mac89]. non-iterative [Duk85, IGW86]. non-Lagrangian [Clo87]. non-linear [DFP81, EG86, Per87, PC86, SB84b]. non-local [OC87]. non-LTE [SC85]. non-Newtonian [BE80]. non-oscillatory [HEOC87, SO88, SO89]. non-reflecting [KG89]. non-self-adjoint [DVW87]. non-separable [PSV89]. non-smooth [Rob85]. non-spherical [FRRT87]. non-staggered [Abd87a, Abd87b]. Non-symmetric [vdV81, Ret82]. non-uniform [CFR84, Mor81, MC85, Pan89, SMR89]. nonaxisymmetric [HK85b, SKS82]. Noncharacteristic [Cou85]. noncircular [HG86]. nonconforming [Li89]. nonconservation [MM81b]. Nonequilibrium [LV89]. Nonlinear [Cha81, FFM82, LLL88, MM88, MHK86, Ohr81a, SK80, AH88b, AB84, BL85, BST85, BPS85, BB86b, BNS83, CW89, CL86b, CG89, DF88, DO86, FC84, GHC^{+}86, GM85, HMM85, HMW85, HCH^{+}81, HL84, IBCS87, KP83a, Kel84, LM87b, LA83, LBS88, Lon88, MMS86, Mej86, MZ82, MJ87, MNM85, OW85, Pat89, PC81a, RSR86, RL89, SSC81, SSC86, SM86, Ste87b, TA84c, TA84a, TA84b, TA88, TY87a, TY87b, TDhLcSb89, TM88, VSS86, Vat89, WP83, Wit84, Wol85, dFSS89]. Nonlinear-multiple-function [MM88]. nonlinearly [BM84, LO87]. nonneutral [HB87]. Nonoscillatory [Lav88]. nonperiodic [Tuc89a]. Nonperturbative [Pol87]. nonphysical [Str82]. Nonpolynomial [Pic83]. nonrational [Gre84b]. nonrectangular [HMR82]. nonreflecting [RS80]. Nonrelativistic [BKH86]. nonself [McD80]. nonself-adjoint [McD80]. nonstaggered [Got88]. nonstationary [BY85]. nonsteady [EB81]. nonsymmetric [HN82, RRS89]. nonuniform [Far80, Hof82, NTK80, NY82, Pro81b]. nonvariational [CC87, GRS89]. nonzero [BB80]. normal [Are81, BS81, FFM82, WG86]. normalization [Che84]. Note [Bra84, BD83, Tem83b, BB88, CSW85, Mon81, Pro81a]. Notes [Ano89t]. NOVA [CC87]. nuclear [QQM89]. nucleation [AH88a]. Number [Ten82, CM86, For85, GH87, HO80, KS81, KL89, KT86, PHD89, Pea83, PC86]. Numbers [Wik89, CJW85, CH87, Hos86, IT89]. Numerical [Abd87a, Abd87b, AFLS83, ABW83, Ana89, Bai80, BBG80, Bel82, BY85, BPS85, CDP80, CCCS81, CHH^{+}86, CL81a, CB80, CHT84, Cho82, Dru80, DH84, DP86, EB81, EM81, EB80, GHK88, Gla84b, Gus80b, HMM85, HW80, IT89, JH83, Jon85, KR84, Kru82, KPBRF83, KP83b, KPB87, LWP89, LR81, MH89, MS81, MM81b, MOFK88, MGS82, MPPS84, MT86b, Mor81, NE82, OC87, PND83, Phi84b, Pir80, QR89, RC81, RM87, RG88, Shi86, SE87, Sho83, Sho86, SBAM82, Shu84, Sid85, SAR82, Spi88a, Spi88b, Sto86, TMCS85, Try88, VE87, VC85, War86, WMS88, Wol81, WK89, YH88a, YNAS86, ACM88, AH88b, Ano81j, Ano82i, Are81, BCE88, Bas83, BL85, BGT85, BWY82, BST86, BKH86, BM80a, BGR80, BB80, BFH86, Bro80, But88, CW85b, CL81b, CL83]. numerical [CC84, CL84, CW80, CG89, CW84a, Cor82, Cul83b, Dav80a, DW81, Duk80, Fog84, FP88b, FB88, FFG87, GM81, GGR82, GGR89, GR84, GMM81, Gre84a, Grö83, GR82a, GR82b, GG86, Gup81, Gus82, Hag81, HRS80, HRT80, Hen87, HD85, HGM86, Hun86, Hun89, Hut84, IHM81, Ioa81a, Ioa81b, JRS86, KP83a, KF87, Kli83, KK88, Kum82, LL81, LBR81, LMM87, LM89a, Lim80, LA83, LV89, LR80b, MPA88, MEGH85, MS87a, MMS86, McL80, MOI81, MZ80, MS80, MJ81, MK80, MML83, Mun88, Mus81, Neu81, Neu83, NR80b, NMG86, O’R84, O’R89, OS85, Pan89, PSV89, Pea81, Per81c, PM80, Phi86, Pro81b, PJ83, Qin88, Ram83, RA87, Rob89a, Roe86, SMRS85, SJ88, SK83a, SGA81, Sha81, SKS82, SBT86, SOS89, Smi81, Smi88, Ste87a, SS82b, SV89]. numerical [SN88, TA84c, TA84a, TA84b, TA88, TDhLcSb89, TB89, TAJ^{+}81, TWM82, TM88, TR86, Tre81, VSS86, Vic86, Wan83, WP83, WR84, WR86, WC84, YP88, Zer85, ZBP83, TA84a, TA84b, TA88, WOZ84]. numerical-analytic [BFH86]. numerically [ADL87]. Numerov [BG85b, Eck89, IR87, Ixa87, LW86, MES80, Nav83].

obstacle [HR84a, RH88b]. obtained [BHM^{+}83]. obtaining [Kos84, NV89]. OCCAM [PHD89]. Occur [MvdV81, OJJ88]. ocean [BH87a, GC89, HRS80, LV86, MRH83]. oceans [BS81]. October [Ano81j]. odd [BR80b]. odd-even [BR80b]. ODE [BH87b, KK81, Sha84]. off [Wig88]. oil [BST86, Sco85]. One [CLS89, Rei81, Aug84, ADR84, BSSV88, BJ81, CG89, DO86, DH84, Fle83, Gol86b, GG88, Hag81, HH83, Hua81, KK88, Kos84, LLVL89, LLPS85, MH89, MS81, Pom83, SS84a, SBT86, SSG81, Tal80, VF80, VBSS89, Vin83, ZGB88, ZBG88]. one- [Fle83]. One-dimensional [CLS89, Rei81, BSSV88, BJ81, CG89, DH84, GG88, Hag81, HH83, Hua81, KK88, LLVL89, LLPS85, MS81, SS84a, SBT86, SSG81, Tal80, VF80, VBSS89, Vin83, ZGB88, ZBG88]. one-pair [MH89]. one-parameter [Pom83]. one-way [Gol86b]. Onsager [Vie83, Vie84]. Onset [HS88a]. Open [CO80, Fer88, HRS80, HMI83, KRW88, MRH83, Nat89, Nor89, Set84]. operator [BK85a, Coo86, Gla88b, GC88, HPBY86, Iss86, JRS86, Lev82, Vie83]. operator-splitting [Iss86]. Operators [Car89, DH89, LLPS85, Pol87, Pom83, Rap80b, SMRS85, SP87b, Ter81b]. optically [LMM87, LM89a]. optics [WWM^{+}81]. Optimal [AG81, Bar80, BK88a, Neu83, Kaw87]. optimisation [Cat88]. optimization [AFLS83, Boy82, CLHP89, GPS88, HA89, ND89, VL84]. optimized [Lup87]. Optimizing [dFLR85]. optimum [Van84]. Orbit [CFT82, DGKW84, CBCF80, CF82, DWD83, KSD81, MV83a, ZC86]. Orbit-averaged [CFT82, CF82, ZC86]. orbitals [ACM88]. orbits [FDS81, Wil80b]. Order [Bor86, ASIO86, BM87a, BL88, BM85, BDS88, BCT89, BCG89, BAF84, CB89, Can81, Cul83a, Cul83c, DM85, ERDF83, GS83, Gar86, GW85, GM88, Gup84, HH88, HTE81, HEOC87, HC81b, IM88, JY83, KJM80, Kui88, LW86, Lup87, MS83, MIK88, MG89, MA84, MD81b, Moh84, MP80b, NY83, O’B80, OF82, Pat86, Ram86, Rob89a, Ste84, TAJ^{+}81, TB86, Tur80, Vie83, War86, YRT80, Zal81, ZBP83]. ordered [AG80]. orders [Wal84]. ordinary [DQ83, HK84, KH87, NR85, Spi88a]. ordinate [SB84a]. ordinates [Ber85a, MD80]. origin [MFC^{+}85]. ORMEC [HH86]. Ornstein [HA85]. Orr [Dru80, MM81c, NR80b]. Orthogonal [Boy87a, Pet87, RL83, ADL87, BS81, CFR84, CY85, Dav81, HC81a, HP87, MS80, Per83, RM87, Spa85, TB89]. orthogonalization [Tob86]. orthonormalization [Dav83, Mey86]. oscillating [Ana89, CFG83]. oscillations [BM87b, BG84]. oscillator [Boo82a, SN83, VV82, VJ89]. oscillatory [BM87a, Dav89, FP82, GK81, HEOC87, MT84a, SO88, SO89]. other [CM80, GWZ85a]. outer [Jon81]. outflow [Hag87, RS80]. overcoming [Kli87]. overdetermined [SH89]. Overlap [PW85]. overlapping [MT84b, MT86b]. overrelaxation [Dic83].

Padé [Adh81a, Hau82]. pair [Lam81, MH89, Phi84b]. pairing [WC85a]. Pairwise [BP88]. papers [Ano87j, Ano89a, Ano89b]. Parabolic [CW84b, ADG86, BB86a, DO86, Hun89, Li80, MMV81, MIK88, MK83, MLB87, SGA81, She88, Spi88b, Wan83]. parabolized [BDR82, CM89]. Parallel [BP88, KL89, Ano85w, Hal89, HMS89, KHTR89, Lub88]. parameter [BM84, CL84, EG83, HH88, HM80b, LO87, Pom83, Roe81]. parameters [Bar80, LPS80, SBBY83, Ves82]. Parametric [AA89, SH84a]. Parasitic [MS87a]. Paris [Ano81j]. Parker [Mel88]. Parker-type [Mel88]. part [ZBP84, KH80, KGH80]. Partial [BO84, GL85, AF86, BSSV88, Boo81, Boo82b, Cat88, CL84, CFL86, Dyk87, GDM81, Gre84b, Hai87, IMT^{+}89, KPD81, RC81, RA85, RA87, Rig87, SSM83, SM84, Spi88b, TWM82, VBSS89, WP83, Wol85, ZBP83]. partially [Pri89, RS84]. Particle [HA89, Law89, ASIO86, ASL82, AP88, BKLT83, BST85, BR86, Bra88, BLT^{+}81, BBGN80, CBCF80, CFT82, CF82, CLF84, CLHP89, CBG86, Den81, Duk80, GM82, GMT^{+}81, GR87, HPP^{+}82, Her89, Hor87b, HSA89, JL86, KO86, KPBRF83, KP83b, LL81, Lam81, LCF83, Law88, Lee87, LD89, LM86, Mas81, MD80, Mez81, Mon81, MG83, Mon89, NTK80, NDP88, NY82, NY83, NOY85, SS89, SB87b, TLD80, TNY85, TY87a, Tan88, WHW87, Wil81]. particle-fluid [Duk80]. particle-in-cell [BR86, Bra88, CBG86, Hor87b, HSA89, LD89]. particle-mesh [Wil81]. Particles [BP88, DP86, FP88b, KT86, WPW89, YP88]. past [CL86c, For85, HR84a, IT89, RH88b, Whi87]. patched [Rai86b]. patched-grid [Rai86b]. path [Mej86]. path-follower [Mej86]. pattern [Smi81]. PDE [vdV81]. PDE-Problems [vdV81]. PDEs [Cha86]. PDF [HP87]. penetrating [WPW89]. penetration [Mon89]. percolation [HCM89]. perfect [ADK82, GG88]. Performance [CLHP89, MM85a, HMS89, KE87, NB89]. Period [VJ89, Aki84, BM87b, KH87]. period-doubling [KH87]. Periodic [PH88, Pom83, BP87, BFT85, Flo86, HK84, JRS86, KP85, KP86a, Kra86, NMG86, RS85, SM84]. permanent [CM82]. perturbation [BD82, Bra84, Bra86, KR86]. perturbational [SS88a]. perturbative [ICP80, Ixa80]. perturbed [Mez81]. petroleum [LS87]. Petrov [DO86, GF86, SSC81]. Phase [PG87, Ada88, BST86, Bin85, CB84, CL80, EB81, GMM81, LS87, Löt82, Lup87, Lyn82, LS85, PT86, Paj89, RH84, RH88a, Rob89a, SW84, TR86, YH88b]. Phase-change [PG87]. phases [BJ81]. phenomena [Har82]. photon [WL84]. photon-Maxwellian [WL84]. Physical [Per81c]. PIC [CDH81, Mar87, Mas83]. PIC-hybrid [Mas83]. piecewise [FR87, GL85, CW84b]. pinch [Gla89]. pipe [Bra86]. pipelines [HK83]. pivots [Ker80]. place [Tem85, Tem88c]. placement [ZOWZ88]. planar [SLO87]. Planck [AH88a, AOK81, BKP87, FWB^{+}81, Her89, LLPS85, MS86, MD80, Wie83]. plane [CDP80, Dav89, IHM81, KL80]. planning [LPS85]. Plasma [BM80a, MHK86, NMG81, RL84, ASL82, AHH87, BF82, BD83, CL86a, CLF82, CDS83, EH85, EM80, Gla84a, Har82, Hew80, HL87, Jar81, Kli83, Kli87, Kre88, LLL88, Law88, Law89, LBM87, LD89, MSD87, Mas87, MM85b, MFW88, Mus81, NTK80, RRS89, RL89, SS84a, SBAM82, SA81, TL81, TS83, WBF86, Whe81, WMM86, Wil80b, YRT80]. plasmas [AEG80, BKLT83, Bir86, CC87, GDS88, HC87, HPP^{+}82, HB87, KJM80, KPBRF83, LCF83, Mas81, Mas83, MS86, MFW89, Mor82, NDAL88, NB89, ONM83, Per81a, Rob89b, Tam81, Tan88, Wei85, Wie83, ZGB88, ZBG88]. platelet [Fog84]. point [AG80, Ber88, Gar86, GS87b, GM88, HTE81, Jac89, MG89, Mil89a, PRV89, QR89, SS88a, Ske81, ST88, TL81, WBL86, vDR89]. point-source [SS88a]. pointed [DS86c]. points [BK88a, EG83, Hai85, KOTK88, KH87, LR81, MD86, PA86, WBL86]. Pointwise [BER85b]. Poiseuille [BD82, Bra84, Bra86, JR85]. Poisson [Abd87b, BB84a, BE85, Cha86, Far80, Gup84, KE87, KJM80, KW85, MD81b, MT86b, Ohr83, Pat86, Pic86, PH88, RL89, SB82, SS88b, Ske81, Tan85, Tem80, WMM86, WK89, vDR89]. polar [IM88]. pollutants [RMS85, ZBP84]. Polyatomic [ACM88]. polygon [Mil89a]. polyhedra [Med86]. polylogarithm [Cla87]. polylogarithm-based [Cla87]. polymer [SHB82]. Polynomial [Boy86, Boy82, DR85, PG88b, RN86, Sca87]. polynomials [GM81, Hit80, LA85, Mav89, Tan85]. porous [CB80, GK81, JH83, Löt82, PG88a, Rob89a, She87]. positional [JPD86]. positive [SP87b, Smo84, SC86]. posteriori [KMRM88]. Potential [Bak81, Whi87, ABS82, Boe82, CH84, DH82, FMC83, FR87, Gre85b, Gre88, HG88, IHM81, KH80, KGH80, MES80, MGM86, Mru85, Mru87, PJ83, Rok85, SH89, SC82, SZH85, Tut86, Ves82, WR84, WR86]. potential-vorticity [Tut86]. potentials [BG85a, BG85b, Eck89, Ixa87, Kra83]. PPM [CW84b]. Practical [FS86, MvdV81, AHH87, SB88, SG86, Tur80]. practice [RE87]. pre [OA85]. pre-mixed [OA85]. precise [Eis82a, Eis82b]. precision [LV86]. Preconditioned [CQ85, Tur87, AG80, BDJS86, MM85b]. preconditioning [DD89, DM85, Mor89]. predictability [Vie88]. predicting [GP81]. prediction [CH84, Cul83b, SOS89, Ste87a]. Predictions [FF86, Ano80z]. predictor [TS81]. predictor-corrector [TS81]. Preface [CGH84]. premixed [GS87b, Smo82]. preprogrammed [HPL80]. prescription [ML87]. presence [FP88b, Pro81a, Ros87]. preservation [Kan84]. preserving [Sho83, YRT80]. Pressure [ROS85, Abd87b, BCE88, GL85, QN86, RS81, RH84, RH88a, RWC89]. pressurized [Sch80b]. prime [Tem83b, Tem85, Tem88c, Tem89]. primitive [Abd87a, GC89, HK86, Ste87a, Van86]. Princeton [LJ85]. principal [Sid85]. Principles [BBGN80]. printing [SBT86]. probabilities [LP83]. probability [ST88]. Problem [BaB86, ADKM82, BR80b, Bel82, Ba89, BP80, CB84, CS82, CQ89, CW86, CG85, CS81b, DKD85, DKWD89, DJT80, DVW87, DW81, GDS88, GP89, GP80, Gup81, Hag87, Hau82, Kal84, KP83a, Ker80, Kli87, KH84, LLVL89, Law87, LA83, Lon88, MPPS84, Mel86, Mer86, Mon89, Nav87, NR80b, NE82, Rap80a, Rob85, SB82, She88, Sid85, Smi81, Spi88b, SL85, Str82, Uch83, VC85, Wol81]. Problems [Cou85, MvdV81, vdV81, AM82, Asc80, BW88a, BGMM89, BK85a, BJ81, BB84b, BS82, BFT85, BM84, Bro80, BJ84, CL83, CL84, CL86b, Coo86, CFH^{+}82, Cul82, Dav80a, Dav83, DO86, DQS87, DD87, DH84, EG83, EJL87, EB83, ET88, Eyr84, Fer88, For89, Fuc89, FF86, Gar86, GKL^{+}87, GS88, GG86, Hag80, Hag81, HRS80, HH88, HM82, HK83, Hei88b, HN82, HMR82, HD85, Hua81, Hun86, Hun89, IHM81, Jac89, JWB88, KR84, KR86, Kan81, Kel83, Ker89, KS87, KTH88, KK86, KH80, KGH80, LMM87, Lar88, LM89a, Li89, LO87, Lyn82, MP84, MO80, Mey86, MV83b, Mor89, Mru85, Mul89, MD85, O’B81, OS83a, Ors80, PG88b, Pea81, Phi86, QR89, RA81, RD83, RSR86, SSC86, SC85, Sch82, SH89, SB88, Shi81, SB84a, Shy85]. problems [SAR82, SMR89, Ste84, Str80, Tut86, VTB89, Wes88, WBL86, WR84, WR86, YH88a, Zeb84, Zeb87]. procedure [Che84, FC84, HBR82, Li80, MM88, MT81, MV83b, MB80, Pap84, Shi81, Smi88]. procedures [HM82, HM80b]. process [Dyk87, HR84b]. processes [BY85, BGM87a, Chr88, ERDF83]. processing [Din86, KL89]. processor [CFH^{+}82, HMS89, HL86, HSSC83, PRT83]. product [GS85b, RH87, TM88, Wil80b]. production [Dav81]. products [Osg81]. profile [OA85]. Program [KKMP86, dFLR85]. programming [HL86, Lav88, Ng81]. programs [KTT89]. progress [Bin85, Cul83b]. progressive [BB88]. projection [BCG89, CLS89, Dun85, FD82]. projector [EFCS88]. propagating [OS88]. Propagation [BH87a, GT85, BBM88, Chi88, Cho80, EB81, Gol86b, IMT^{+}89, KK86, Lev88, Lon80, SL85, Wei85]. propagative [PG87]. propagator [AH88a]. propagators [CKR87]. properties [AB81, Eis82a, Gol86a, GRS89, HG86, LB87, LS87, ML83a, O’R89, Rob89b, SS82a, ST89b, WC85b]. property [Kan84]. proposed [HG86]. prospects [Cul83b]. prosthetic [PM80]. protium [NR84]. protonosphere [YRT80]. Prüfer [PT86, Paj89]. PSEC [LJ85]. Pseudo [Car80, Hal89, Hua81, MTM81, Pat89, BGMM89, Dav80b, Hos86, ND89, Pea83, SP87a, SH84a, TNY85, TY87a, Wik89]. pseudo-current [ND89]. pseudo-particle [TNY85]. Pseudo-random [Hal89, Hos86, Wik89]. Pseudo-spectral [MTM81, Pat89, BGMM89, SH84a]. Pseudo-unsteady [Hua81]. pseudo-upstream [Dav80b]. pseudocompressibility [RCK87]. pseudodifferential [Bog88]. pseudoperiodic [RS85]. pseudopotentials [MD81a]. Pseudospectral [HK83, BK88a, Cor82, DD89, DM85, GC88, HLAD84, HK86, HPBY86, Jac89, KHT87, KHTR89, Mac86, McD89, NS89, SBC84, ST89b, TE86]. pseudospectral/finite [Mac86]. pulse [SKA86]. purpose [HHN^{+}86, HSSC83].

Q2R [ZH88]. QCD [BGM87a, dFLR85]. QN3D [HSA89]. Quadratic [VR86, Cul82, HR86]. Quadrature [FZ88, CS84, EHM81, Hut84, Ioa81b, Lup87, Shi81]. quadrilateral [Duk84a]. quantal [MES80]. quantum [BK85a, Cep83, CKD87, EW88, GRP82, MPPS84, Pol87, SAR82, WC85a]. quantum-mechanical [SAR82]. quark [ADKM82, CKR87]. quark-antiquark [ADKM82]. Quasi [KKMP86, MHK86, Per83, GIT87, GM88, HSA89, KP85, KP86a, MG89, Oli80, SP87a, ZC86]. Quasi-Analytical [MHK86]. quasi-equilibrium [Oli80]. quasi-fractional [GM88, MG89]. Quasi-heatbath [KKMP86]. quasi-inverse [GIT87]. Quasi-Lagrangian [Per83]. quasi-neutral [HSA89, ZC86]. quasi-periodic [KP85, KP86a]. quasigeostrophic [MRH83]. Quasineutral [Har82]. quaternion [Ves82]. quaternions [Rap85]. quench [DW81]. quenched [KKMP86]. Quiet [Sel83, Law88]. quotients [IA88].

Racah [KTT89]. radiation [AH88b, ADR84, Cla87, EM81, FC84, IO81, KL80, Mih85, MK86, VTB89]. radiational [KRW88]. radiative [Bon84, GS87a, LM87b, Lar88, MK82, Per81b, SC85]. radioisotope [Bas84]. radix [Tem83a, Tem83c]. Random [Cha88, CHS88, GWZ85b, Ber83, CM86, Cep83, CH87, ERDF83, FC84, GS85a, GS87a, Got88, GWZ85a, Hai85, Hal89, Hos86, KS81, KL89, MM81a, MP84, OG86, PHD89, Pea83, Rob85, SP87a, SS89, Sum89, Tak86, Wik89]. random-particle [SS89]. Random-vortex [CHS88]. range [BG85a, Mac81, PP88b, ZBP84]. Rapid [Rok85, Rok89, Bir86]. rapidly [RS84, SP87b, WK89]. Rarefied [Abe89]. Rarnshaw [Kan83]. rate [ADG86, CDS85]. rates [Bir86, GMT^{+}81, PV83]. ratio [GRP82, GH87, Wil82]. Rational [BPD81, Boy87a, Boy87b, Chr88, YH88b]. Rattle [And83]. Rayleigh [Bue88, CB84, CJW85, Grö83, Ker88, MZ83, Try88, Zuf89]. Raymond [JDK86]. rays [Sto84]. Re [GGS82]. reacting [CB80, GP81, KK81, MGS82]. reaction [BGMM89, BB86a, CDS83, MC88b, PV83, RSR86]. reaction-diffusion [BB86a, RSR86]. reactions [JH83, Mac89, SK86, Yam83]. reactive [Ba89, EB81, Mru87, Tak86]. reactor [NR84, Sch80b]. real [CG85, Gla88a, Gla88b, Kos84, LvLS89, Mül88, NV89, Ret82, Tem83a, Tem88c]. real-symmetric [Kos84, NV89]. real/half [Tem88c]. real/half-complex [Tem88c]. recirculating [GC87, IGW86]. recirculation [BR81]. recombination [LP83]. reconnection [Dri88b]. Reconstruction [Gan85]. rectangles [SB82]. rectangular [CH84, CHS88, Kum82, SB87a]. rectangularly [Pat86]. Recurrence [Mar81]. Recursive [CW85a, CP88b, FWHC86]. red [AG80]. red-black [AG80]. Redesign [Ros87]. reduce [ACM88]. reduced [FWB^{+}81, GC89, KR89]. reducing [BST86, HH88]. reduction [GRP82, HGM86]. reductions [Ske81]. reference [Ves82]. Refinement [BO84, BC89, SJ88, SOS89]. reflecting [KG89]. reflection [Vic86]. regimes [HW81, LMM87, LM89a]. region [CL80]. Regional [Boo82b]. regions [LG80, Man86, Mer86, Vie84, WR86]. register [KS81]. regula [FR87]. regular [MD86, Pic86]. Regularization [ZO83, LS87, Mer80]. reinforced [Bel82]. related [MW84, Shi81]. relation [CM89, MB80]. relations [Mar81, WPW89]. Relationship [Hof82]. Relativistic [MT86a, ADKM82, EM80, HL84, JHLS89, Man85, Man87, MS86, MT84a, Pir80, Pol87, Wei85]. relaxation [BM87b, BV82, Cha86, Eas87, Hei88a, Mul85, OF82, Rai86b, RL84]. reliability [HCH85, Tre81]. remapping [Duk84a]. remark [Ioa81b]. remarks [RD83]. remeshing [PVMZ87]. remote [CL81b, OW85]. Removal [GP89, Zeb87]. renormalization [Bar80, BGMT89]. renormalized [LW86]. reorthogonalization [CW81]. Reply [Kan83, RD83]. represent [Gre84b]. representation [ABS82, CB89, KM80, SB87a, Vat87, Wil80b]. representations [AHH87, Col81b, Dyk87]. requirements [Grö83]. reservoir [BST86, BDJS86, FS86, LS87, SJ88, Sco85]. residual [CQ85]. residue [CW85a, FWHC86]. Resistive [MFW89, HCH^{+}81, HT86, KJL86, Sto86, TS83]. Resolution [LLVL89, GMM80, Grö83, Har83a, Har89, Lev88, Mun88, Vie88, WW89]. resolving [Nol85]. resonance [BL85]. resonances [CH88, MFC^{+}85]. response [KH80, KGH80, MB80, OJJ88, Pla81, SS88a]. restrain [GC88, HPBY86]. restricted [QQM89]. restrictions [ZBP83]. resultant [KW86]. results [Aug84, BHM^{+}83, ST88, WOZ84]. retrieval [OW85]. Reversal [DGKW84, Alz84]. reversed [AB81, FSD81, HSA89, SBAM82]. Review [BH87b, Bin85, TWM82]. revisited [Hau81, RH88a]. Reynolds [CPJ89, For85, GH87, HO80, IT89, Ten82]. Reynolds-averaged [CPJ89]. rezoning [Duk84a, Per83, Ram85b, Ram86]. rf [BCE88]. Richardson [Hen81]. Riemann [BaB86, Ba89, CG85, Duk85, Gla88a, Gla88b, GG88, Roe81]. rigid [LBO^{+}89, Son85]. ring [FDS81]. ringing [Bra88]. rings [FSD81]. ripple [BHS87]. RLW [GC88]. Roberts [Kil87]. Roe [Lyt87, SB84b]. Roe/Lytton [Lyt87]. roll [Kra86, Try89]. roll-up [Kra86, Try89]. root [Pol87]. rotated [Tem88b]. Rotating [WOZ84, CO80, Har80, IT89, Lon88, Mac86, MSK80, RS84]. rotation [Kob85]. rotational [BS81, Hor87a]. rotationally [Dav84]. routine [Dye84]. routines [Hor87b]. RSF [EH85, HCH85]. rules [Afa89, FZ88, Ioa81b]. Runge [CLS89, Wil80c].

S [KGH80]. saddle [LR81]. sampling [Mez87, Nor80, WP88]. SARLAC [JHLS89]. Satisfy [Car89]. saturated [GK81]. saturation [BM80a]. scalable [KS89]. scalar [BDS88, GR84, McD89, SB84b, WG86]. scale [HS86, Ker89, Pea83, Riz88]. scales [Gus80b]. Scaling [KK89, ROS85, Str82]. Scattering [SKA86, AEBS86, Ale86, AP88, ADK82, BL88, CL80, DH84, Eyr84, EW88, GS82, GS83, HG88, LLVL89, ML83a, ML83b, MT86a, Mel86, Mru85, Mru87, Paj89, Rok89, SW85, TAJ^{+}81, Tob86, WMS88, WPW89, YH88b]. SCF [ACM88]. scheduling [Rap80a]. scheme [AA89, ADR84, BAF84, BS89a, BW88b, BF86, Bro85, Cat88, Cor82, CFG83, Dav80b, Dav84, Dri88b, GW85, GP85, HTE81, Hen87, HKA89, HW84, Hu89, Hut84, IGW86, KW86, Lam81, LC89, LG85, Lup87, MZ82, MD81b, MJ87, MP89b, NV89, NN89, Neu81, NY83, Ohr81b, Per87, RA85, RA87, SS82a, She86, SH84b, Skö82, SS85, SB84b, Ta’87, TR86, Vij86, WW89, WL84, ZMC83]. schemes [AM82, AG80, BKP87, BPS85, BM80b, CSW85, Cat88, CM89, CKD87, CQ89, Chr85, Cul83a, Cul83c, FS89, GHC^{+}86, Gar78, GGR89, GR82a, Gus80a, HTE81, Har83a, HEOC87, Har89, Hen81, Hua81, MD85, Mun88, Oey82, OC83, OF82, Pic83, RD83, Roe81, Ros83, SO88, SO89, SM86, SS84b, Ste87a, Wil80c, YWH85, Yee87, dFSS89]. Schmidt [Tob86]. Schrödinger [BK85b, BGM87b, CW89, DF88, DFP81, Eck89, FFS82, FR87, GRA83, GR82a, GR82b, HMM85, HMW85, IBCS87, ICP80, Ixa80, IR87, KK88, KK83, LW89, OS85, OC87, Pat89, Per87, TA84a, Tal80, TC87, TM88]. Schwarz [FZ87, FZ88]. Schwinger [MH89]. science [McN87]. Second [NY83, BCG89, BAF84, DM85, LW86, Man86, Moh84, Ram86, TAJ^{+}81, War86, YRT80]. Second-order [NY83, BCG89, BAF84, DM85, LW86, Moh84, Ram86, TAJ^{+}81, War86]. secondary [CA85b, GS88]. section [AP88]. sections [HPL80, ML83a, ML83b, WL84]. Secular [DGKW84]. segmentation [BSS82b]. seismic [SGT88]. seismology [Aki84, CHT84]. Selected [RE87, For84]. selection [GWZ85a, GWZ85b, Mez81, Sha84]. selective [Nav83]. Self [HH83, Jar81, MHK86, Mez87, Tem83c, Bai80, BM80a, But88, Col81a, DVW87, Fer80, HK85b, LBR81, Per81a, SP87b, Tem85, Tem88c]. self-adjoint [SP87b]. Self-Confined [MHK86]. Self-consistent [Jar81, Mez87, Col81a, Fer80, HK85b, Per81a]. self-diffusion [Bai80]. self-heating [BM80a]. self-induction [But88]. Self-sorting [Tem83c, Tem85, Tem88c]. self-supporting [LBR81]. Semi [HK85a, HS86, KR89, SBM^{+}87, Boy87a, BDR80, HM89, HBR82, Ioa81b, KJL86, KL80, PSV89, She86]. semi-closed [Ioa81b]. semi-direct [PSV89]. semi-discretization [KJL86]. semi-empirical [HBR82]. Semi-implicit [HK85a, HS86, KR89, SBM^{+}87, BDR80, HM89, She86]. semi-infinite [Boy87a, KL80]. semiconductor [KG87, MV83a]. semiconductors [MEGH85, MV83a, NDP88, Yor86]. sensing [CL81b, OW85]. Sensitivity [RSR86, BER85b, PC81a]. separable [BFH86, Eyr84, PSV89, SP87b]. separated [EB81]. Separating [SS80, vDS80]. separation [DS86c, Ing84]. sequence [Hal89, KS81]. sequences [CH87, SP87a, Wik89]. series [ANAN84, Boy84, Boy86, Col81b, HRT80, Mat86, Mor84, Pie84, VV82]. series-expansion [Mat86]. set [BB86b, CM82, Tem88b]. Sets [MvdV81]. Several [CS81b, Kos84, Ret82]. shake [And83, LBO^{+}89]. shallow [Aug84, FN80, For83b, For84, For86, Gil81, Hen81, MM81a, MM81b, Nav83, NN89, Sho86]. shallow-water [FN80, Hen81, Nav83, NN89]. Shape [Smi81, DP86, Sho83, SS80, WCR^{+}88]. shaped [MT86b, NB89, Sch88]. shapes [Kam81]. Shaw [AH85, Mey81]. shear [CX86, GHK88, MM81c]. sheared [Sco88]. sheet [KS87, Kra86, Try89]. shell [Mac86, Tak81]. Shift [MFC^{+}85, CL80, KS81]. shift-register [KS81]. shifted [Hit80]. shifts [YH88b]. ship [MSB87, Ohr81b]. ships [MNM85]. Shock [BC89, MG83, BR80b, BSS82a, GM85, KM80, MP84, MOFK88, MS87b, MP89b, OG86, SO88, SO89, SSG81]. shock-capturing [SO88, SO89]. shock-like [KM80]. shocks [Cor82, HL85, Noh87, Ros80, WC84]. shooting [NR80b, NE82, Pru88]. Short [Aki84, BG85a]. Sielecki [Hen81]. signal [LPS80]. signals [RS85]. Similarity [WPW89, HW81]. SIMM1 [JL86]. Simple [DD87, WH87, BSSV88, Kai83, LP83, Pap84, QQM89, SSC86, Tak81]. Simplified [Ram86, Duk84b]. simulate [YRT80]. simulated [VL84]. simulating [Kre88, KT86, Lam81]. Simulation [Ale86, CBL81, Fer87, Ker88, Try89, ASIO86, Abe89, ASL82, AD87, BKLT83, Bas83, BHS87, BST86, BDJS86, BY85, BS89a, BF82, BD83, CL86a, Cep83, CHS88, CLF82, CLF84, CLHP89, CD80, CDS83, EH85, ERDF83, EB80, FS86, FFG87, GS85a, GC87, GHK88, GRP82, Grö83, Har81, Har82, HK85a, Hew80, HL87, HB87, HSSC83, Hor87a, HCM89, JH83, Keu86, KO86, KJM80, KL89, KP83b, LCF83, Law85, Law89, Lee87, Leo80, LBM87, LD89, LG80, Lyn82, LS85, MSD87, Mas81, Mas83, Mas87, MOFK88, MNM85, Miy86, MG83, MD85, MTM81, NTK80, PG87, PRT83, Pul80, Rap80a, Rap85, RH88b, RWC89, RMS85, Sad84, SJ88, SBC84, Sco85, Sco88, SB88, SBAM82, ST88, SS89, Son85, Spi88a, SLO87, Sum89, Tan88, WBF86, WC84, Yor86, ZGB88, ZBG88]. simulation [dFLR85, CCCS81, Ota88]. Simulations [Tha80, BM89, BBGN80, CW84b, Dec84, Den81, FS89, Gla84b, Gol86a, GRP82, GR87, JR81, KS89, LBO^{+}89, Law88, LG81, Lub88, NMG81, NDP88, Sel83, SG86, SA81, SH84c, TLD80, TL81, Try88, Vij86, WP88, WPW89, YP88]. Simultaneous [Bir86, Can81, KGH80, BB86b, CX86, EG86, MM88, SPG89]. simultaneously [Kos84]. Sinc [EJL87, Boy86]. SING [KH80, KGH80]. Single [MS83, Ste84, Ale86, HR84b]. single-stage [HR84b]. singular [BS82, BG85b, CDP80, FZ88, Fre83a, Fre83b, Gar86, Ioa81a, Ioa81b, Ioa82, Ixa87, KR84, KR86, KK89, Kel83, MD86, OS83a, PG88b, Pro82, Spi88b, Wan83, WBL86]. Singularities [EM82, Boy86, Kel84, LR80a, Phi86, SSF83, Wig88]. singularity [Boy85, KH80, KP83b, OF82, vDS80]. sinusoidal [LPS80, Whi87]. site [Hai85]. site-to-site [Hai85]. sixth [Vie83]. size [HS88b, Kai83, iKI85]. sized [NY82]. sizes [ICP80]. Skan [SH84a, Sum89]. slender [Ros89]. slight [And85]. Slip [SS80]. slow [Kel83]. slowing [HM80a]. slowly [LR80b]. small [GR84, Smo84, Tem88b]. small- [Tem88b]. smooth [GG86, Rob85]. Smoothing [CLF84, CP88a, Baa81]. software [GS88]. solenoid [DG88]. solenoidal [Ram83]. solenoids [NM80]. solid [BM87b, BJ84, JH83, Lon80, MK80, Phi84b, SK86]. solid-fluid [JH83]. solid-solid [SK86]. solidification [CB84, Cho85, EB83, Smi81, Str89, SLO87]. solitons [MS87a]. Solution [AH88a, AOK81, BB84a, BST85, CD82, Cha86, DKD85, DKWD89, Far80, FFS82, HA85, Iss86, Kel84, Lon88, Lyt87, MK82, Ng81, OJJ88, SH89, SHK88, Smo82, SB87b, Vie84, vdV81, AI80, Adh81a, Adh81b, AEBS86, Alv89, AHP81, Ana89, Are81, BDR82, BL85, Ber85a, Bir86, BK85b, BST80, Boo81, Boo82b, BB80, BS84, Bro80, Bro85, CDP80, Car80, CCCS81, CW85b, CS84, CG85, CFH^{+}82, Cul82, Dav80a, DD89, DQ83, DM85, DQS87, Dru80, DP86, EB81, EHM81, For83b, FWB^{+}81, GGR89, GKL^{+}87, GRA83, Gus80b, HK86, Hen87, Hir89, HG84, HD85, HR86, Hun86, Hun89, IHM81, Ioa81a, Ioa81b, Ixa80, Jon81, KR84, KF87, Kel83, KPD81, KP86b, KK83, KG87, KH84, KHT87, KHTR89, KSM82, KW85, Lab87, Lav88, Lim80, LA83, LJ81, LBS88]. solution [Luc87, LPS85, MH89, MS81, MK83, MS87a, MMS86, McL80, MGS82, MD80, MPPS84, MT81, MT86b, MSB87, Moh84, MK80, MM81c, Mor81, Mru87, MD85, Mus81, NR80b, NE82, NS89, OA85, OC87, PSV89, Pat89, PND83, Pea81, Phi84b, Pic86, Pik87, PRV89, Pro82, QR89, RRS89, RM87, RG88, Riz88, Rob89a, Rok85, Rok89, SP87a, SGT88, SB82, SS88b, SSSR83, Shi81, SB84a, Sho86, Smi81, Soh87, SG88b, SL85, Tan85, TWM82, TM88, VV82, Van84, Van86, Vat89, VC85, WP83, WR84, WR86, WVP82, WK89, YH88a, Zeb84, Zer85]. Solutions [CPJ89, Gil81, Abd87a, Abd87b, AH88b, BM87a, BR80a, Bel82, Blo82, BGR80, CB80, DFP81, DW81, EW88, FLP80, Fle83, Fuc89, GGS82, GT85, Gre84b, GG86, Gup81, HK83, HP87, HK84, Hua81, Jar81, KP85, KP86a, KF87, Kli87, KM80, LMM87, LM89a, MES80, Pan89, Pom83, QQM89, RN86, SK83b, SK86, Str82, TB82, Tre81, VJ89, Vat87, WOZ84]. solve [BGM87b, HN88, Nex87, RL89, Tak85]. solver [BDJS86, Duk85, Gla88a, Gla88b, Gup84, KE87, KJM80, Whe81]. Solvers [BH87b, BB84a, Cha86, GG88, HLAD84, Pat86, Roe81, Sch80b]. Solving [Asc80, BJ84, DGKW84, Eck89, Eyr87, MHK86, MvdV81, Nol85, BCE88, Cat88, CA85a, CA85b, CW89, CL81a, CL83, CL84, CC87, CL86c, Cur82, Dav83, DJT80, ET88, EWR88, FP88b, For84, FR87, GS87a, GF86, Hag81, Hew80, ICP80, KR86, KK81, Law87, Li89, MEGH85, McD80, Mel88, Per87, Pro81a, Ram85a, Ros89, SAR82, TNY85, TDhLcSb89, TAJ^{+}81, Tur87, Zeb87]. Some [AHH87, Hag81, LV86, Man85, Pla81, SG86, Afa87, Afa89, Bin85, Bog88, CY85, Cul83a, Cul83c, Ter81a]. Sommerfeld [Dru80, MM81c, NR80b]. Sommerfeld-type [Dru80]. sonic [HL85]. SOR [BE85, Ehr81, Rig87]. sorting [Tem83c, Tem85, Tem88c]. source [GMT^{+}81, RMS85, SS88a]. sources [Gar86, Hoe83, OS85]. Space [Mor82, CRS88, DO86, Kli87, Neu81, PL86, SKA86]. Space-charge [Mor82]. space-marching [PL86]. space-time [Neu81]. spacing [Got88]. Sparse [vdV81, SGT88]. sparsity [CW85a]. Spatial [Grö83, CLF84, KG85]. spatially [Pan89]. special [Hau82, HHN^{+}86, HSSC83]. special-purpose [HSSC83]. species [Tak86]. Spectral [Boy87b, DQ85, KP89a, LJ85, MO80, Mor84, Ors80, SZH85, ZWH82, ZWH84, AHH87, Baa81, BM89, BGMM89, BH87a, BFT85, CQ85, CQ87, FFS82, FT84, GTD89, Gau88, GX89, HLAD84, Hei88a, HH86, IBCS87, KP89b, KMA89, KP86b, LO87, MZH85, Mav89, MML83, MTM81, PP88a, Pat84, Pat89, RS85, SH84a, Ste87b, SSF83, Tan89, Zeb87, dFSS89, ML83a]. spectral-difference [GX89]. Spectrum [MK86]. speed [Bra88, OS88, Tur87]. SPH [MG83]. sphere [AP85, CL86c, Dri88a, Gil81, HS88b, iKI85, Neu83, Tha80]. spheres [Ber83, JH83]. Spherical [Mar81, O’B80, AP88, BS84, CB89, DQ85, Dil85, EHD89, FRRT87, KS82, Mac86, Neu83, Ves82]. spherically [Man85, Man87]. Spin [DGKW84, MV83a, EFCS88, HSSC83, JR81, KSD81, Lub88, Sad84]. spin-exchange [Sad84]. Spin-orbit [MV83a, KSD81]. Spline [AEBS86, EWR88, ASIO86, Asc80, Die83, LJ81, Mor84]. Spline-collocation [EWR88, Asc80]. Spline-Galerkin [AEBS86]. Splines [Eyr84, CP88a, Mon85, Sho83]. Split [IBCS87, dFSS89]. Split-step [IBCS87, dFSS89]. Splitting [LvLS89, Rap80b, AG81, CM89, Coo86, Dic88, Gar83, Gla88b, HN88, Iss86, LBR81, MK83, Mül88, Nex87, SW81]. splitting-up [LBR81, MK83]. spontaneous [vDS80]. sprays [AFLS83, Duk80]. spreading [Gre85a]. SPUDNUT [AEG80]. Spurious [SK83b, GTD89, WD84, Zeb87]. square [Pol87]. Squares [Mun81, BB84b, CFG83, MM88]. SSD [GKL^{+}87]. Stability [AM82, BWY82, CKD87, CF82, Dyk87, KG85, PC82, ZBP83, AB81, BHM^{+}83, BGR80, BL82, CC87, CS81a, CFL86, CR84, EH85, FB88, FSD81, GT85, GDM83, HRS80, HMW85, HMJ86, KMA89, LCC^{+}86, Mah82, Man86, MM85b, RV88a, TE86, Tre83, Zeb87]. stability-enhancing [Mah82]. stabilized [Smo82]. stable [JT86]. stage [HR84b]. Staggered [Got88, Abd87a, Abd87b, BCE88, CW82, SS88b]. staggered-mesh [BCE88]. stagnation [GS87b]. standard [KE87, OS85]. star [Ale86]. starting [KS87]. starts [Sel83]. state [ADG86, BJ84, CL80, ER85, Hua81, Lav88, MES80, Nor89, Phi84b, Shy85, Tak80, Van84, WOZ84, YWH85]. states [ADKM82, MJ81, WOZ84]. static [SS80, Mih85]. stationary [BD82, Bra84, Bra86, Fer88, Rap80b]. Statistical [O’R89, PW85, RV88b]. statistics [KD88]. Steady [AD83, For85, JY86, SSG81, WOZ84, ADG86, Boe82, CB84, CA85a, CA85b, CJ86, Dic83, Dic88, ER85, EB83, GK81, GR84, GW85, GW80, Hua81, IT89, JY83, Kor89, Lav88, LG81, MP84, MP89b, Nat89, Nor89, Sak81, Shy85, SS82b, Tak80, Van84, YWH85]. Steady-state [WOZ84, Hua81, Lav88, Shy85, Tak80, Van84, YWH85]. steep [Gre84b]. steepest [HH85]. Stefan [BJ81, DKD85]. stellar [Gla84b]. Stellarator [LCC^{+}86, BBG80, HCG^{+}85]. stellarators [BHS87, LNS87]. step [BP87, Dyk87, For83b, IBCS87, ICP80, KM85, KK88, LCF83, LG81, Mah82, MSW87, Ste81, SH84c, Van84, Wag88, dFSS89]. stepping [Got88]. stepsize [ZBP83, ZBP84]. stereological [iKI85]. Stiff [BH87b, Dav83, KK81]. Stiffness [PV83, Sha84, Bog88]. stimulated [Chr88]. stochastic [DWD83, DS89, EWR88, Gol86a, KD88, Spi88a]. stochasticity [FDS81]. Stokes [AG81, Abd87a, AHP81, BDR82, BCG89, BGR80, BP87, Cha88, CC84, CPJ89, Cul83a, Cul83c, DD89, DH89, DP86, FP88b, FF86, GGR82, GGR89, GGS82, GS89, Gup81, Hal81, HK86, HG84, KP89a, KM85, KP86b, Kor89, KHT87, Li80, Lon88, Luc87, MZH85, MML83, Nat89, Ng81, Nor89, OK86, Rap80b, Ros83, RS80, SSSR83, Soh87, SG88b, SBSH84, SN88, Tak80, Tan89, TB82, TF89, Tre81, VM86, Van86]. storage [GS88, Wil80c]. storm [TGP80]. strain [Mei89]. strategies [CFL86, Dan80, Tem89]. strategy [Nor80, QQM89]. stratified [GHK88, HMI83, HR84a, MTM81, RH88b, SW85, SS82b]. stream [Ana89, DQ83]. streamfunction [GS89]. streamline [Pea81]. Streamlined [HB87]. Streamwise [Gre85b, Gre88]. stress [SC81]. stresses [JR85]. stretching [Vin83]. string [LR81]. Strong [FDS81, Hoe83, Hoe84, ADL87, Noh87, WC84]. strongly [GM81, Har80, Mor81]. structural [KGH80]. structure [BM80b, Col81a, Fer80, Har81, Sch80b, Tak81, TS81, WVP82]. structured [CW85a]. structures [JRS86, KD80, KH80, KGH80]. Studies [VSS86, AFLS83, Bai80, BBG80, FSD81, HL84, KPBRF83, ML83a, ML83b, Sto86]. Study [SS81, AH88b, BL85, Bin85, BK88b, But88, CG89, CW84a, GGR82, GR84, HRT80, HHS89, Kum82, SP87a, SG88a, Shy85, TY87b, ZH88]. studying [Fog84]. Sturm [Boy85, EJL87, MV83b, PT86, Pru88]. Sturmian [DR85]. sub [ASL82]. sub-cycling [ASL82]. subcell [Har89]. subcritical [DS86b]. Subgrid [GMM80]. subject [PG87]. subsonic [RS80]. subtracting [Wig88]. Successive [NDAL88, RL84]. suitable [Pea83]. sum [AD87, Afa89]. summation [vDR89]. sums [Afa87, Mac81]. Sundström [Tre83]. SUNION [SS84a]. SUNION-an [SS84a]. Supercomputer [McN87, BJ84]. superconducting [MMR89]. superconvergent [MC88a]. supercritical [Ros87]. superfluid [But88]. Supersonic [FS87, BDR82, DS86c, GW85, Li80, MP84, RM82]. supporting [LBR81]. suppression [CW85b]. Suprathermal [Wie83]. Surface [FOF88, Asa87, BE85, CL86b, For89, FW81, HCG^{+}85, LC88, MES80, MSB87, OvOH89, PJ83, SS88a, SS80, SS89]. surfaces [Keu86, ML83a, Sal86, TMCS85, War86]. surge [TGP80]. surgery [Dri88a, Dri88b]. surrounding [KH80, KGH80]. suspended [FP88b]. swarm [SBBY83]. swirling [KMA89]. switching [SS84b]. Symbolic [Bro89, Gla84a, SR85, KTT89, Lo85]. symmetric [CB89, CW81, GPS88, GDS88, GS88, Har83b, Kos84, KW85, MS81, Man85, Man87, NV89, Ret82, RN86, Yee87, vdV81]. symmetrical [Bra86]. Symmetry [DGKW84, Pat87, SN83, CW86, HBR82, MW87, NMG81]. Symposium [Ano80z]. synergetics [Zab81]. system [Abe89, CCCS81, EG86, Jar81, KP83b, MPPS84]. Systematic [FK89, Rob89b]. Systems [DGKW84, vdV81, AD87, Aiz85, Alz84, BL85, BCT89, BDJS86, CQ87, Cep83, CLS89, EFCS88, GDM83, Gus80b, Gus82, Hag80, HHS89, Har81, Har83b, HC81a, HMS89, Hen87, Hir89, Hoe83, HSSC83, HR86, Ixa80, JR81, KP85, KP86a, Kra80, Law85, Law89, Lub88, Mej86, MT84b, MS80, Moh84, NR85, Nol85, PT86, Pir80, RRS89, RM87, RSR86, Sad84, SGT88, Shi86, Spa85, TMCS85, TWM82, Tho87, Tur80, Uch83, VBSS89, Wol85, dFSS89].

Tables [ML85, ML87]. taking [Tak81]. Taming [AD87]. tangent [DS86a]. tangential [RL89, ZO83]. targets [Bon84, CW80, LBR81]. Targetted [DS89]. Tau [LO87, GTD89]. Taylor [BK87, Jon85, Ker88, MZ83, MML83, Try88, Zuf89]. tearing [HCH^{+}81]. Technical [Bas83]. Technique [Mun81, Adh81a, AF86, Duk84b, GPS88, JR81, KT86, LA83, LJ81, MEK86, Mat86, Mei89, Mel88, Mer86, MT86b, Nav83, NR80a, NE82, Rag82, RMS85, Sak81, SS86, SSC86, TL81, TF89]. techniques [BS86a, Cul83b, DA86, FLP80, Gla84a, GMT^{+}81, Hun86, KMRM88, KMA89, KS82, KG87, Ota88, SK83a]. temperature [OA85, PV83, RA87]. temperatures [Bir86]. tension [FOF88]. tensor [HW80]. term [Per87]. terms [Sch80a, SB88, Zal81]. terrestrial [KL80]. tesselation [RHOG88]. tessellation [TOO83]. test [EHD89, Mei89, Tre81, WH87]. testing [Kob85]. tests [GR82b, LV86]. textures [Bas84]. th [GS83]. Their [CP88a, AB81, Bar82, BV82, BL82, Sch80a, Yee87]. theorem [LBM87, MW84]. theorems [Dec84, Vic86]. theories [Pol87, TDhLcSb89]. Theory [Col81a, Fer80, AHP81, Bal85, BL88, BST85, CA85a, Chr88, DKD85, DKWD89, GS82, GS83, Gla84a, HC87, KKMP86, LO86, MC88a, ML83b, MM81b, OJJ88, Ota88, Pom88, RE87, Rok85, Rok89, TAJ^{+}81, Tre83, YHS86, BW88a, ZGB88]. thermal [BR80b, BDJS86, CL81b]. thermally [Chr88, GHK88]. thermodynamic [Kra80]. thermohydraulic [Ste81]. thick [LMM87, LM89a]. thin [AH88a, Hun89, Kam81, LBR81]. third [JY83]. third-order [JY83]. Thomas [CS84, FLP80, FRRT87, Kru82]. thousand [EG86]. Three [AB84, HR84a, Ohr81b, RM82, SS82b, ZMC83, AG81, AEBS86, AB85, Bas83, Ber88, Bue88, Cat88, CA85a, CA85b, CPJ89, CQ89, CBL81, Duk88, EH85, EG86, Eyr84, Eyr87, FP88b, FS87, Gla88b, Gre88, HRS80, HHS89, HK85a, HS86, HMJ86, HH86, HSA89, HGS88, KD88, KO86, KH80, KGH80, KHT87, KHTR89, LM86, MEK86, MP89a, Med86, MPPS84, MT84b, MNM85, PM89, RG88, Ros89, Sal86, SBC84, Sch80b, Sco85, SM84, SP87b, SN88, Tan85, Tan89, TOO83, WR84, WR86]. three-body [AEBS86, CQ89, Eyr84, Eyr87, MPPS84]. Three-dimensional [AB84, Ohr81b, RM82, SS82b, ZMC83, AG81, AB85, Bas83, Bue88, CA85a, CA85b, CPJ89, CBL81, Duk88, EH85, FP88b, FS87, Gla88b, Gre88, HHS89, HK85a, HMJ86, HH86, HSA89, HGS88, KD88, KH80, KGH80, KHT87, KHTR89, MEK86, MP89a, Med86, MT84b, MNM85, PM89, RG88, Ros89, SBC84, Sch80b, SM84, SN88, Tan85, Tan89, TOO83, WR84, WR86]. three-point [Ber88]. three-thousand [EG86]. thrust [Yam83]. thunderstorm [HNK89]. tidal [For83a, KG85, Pla81, Rag82, Tha80]. tides [LV86]. tight [Laf89, MJ81]. tight-binding [MJ81]. tilt [HSA89]. tilted [CW84a]. Time [Alz84, Den81, DGKW84, GS89, HG88, MK82, OA85, Soh87, Tho87, AG81, AF86, ADR84, Bai80, BK85b, BF86, CH88, CLF82, CFL86, CBG86, DQS87, For83b, FW81, FT84, Gan85, GHC^{+}86, Got88, Gus80b, HS86, Hu89, KK83, LCF83, Lar88, LW85, LM89b, MK80, Mor84, Neu81, NR84, Per81b, PC86, RA81, Roe86, SSM83, SGT88, SAR82, SH84c, Van84, Yor86, ZBP83, ZBP84]. time- [ADR84]. Time-Dependent [MK82, GS89, OA85, AF86, Bai80, BF86, CFL86, FW81, Gan85, HG88, Hu89, Lar88, MK80, PC86, Roe86, SAR82]. Time-filtering [Den81]. time-implicit [CBG86]. Time-independent [HG88]. time-integration [ZBP84]. Time-marching [Soh87]. Time-Reversal [DGKW84, Alz84]. time-step [LCF83, SH84c]. time-stepsize [ZBP83]. Toda [FFM82]. tokamak [DJT80, HPL80, NR84, ONM83, YNAS86, NB89]. tokamaks [CD80, GMT^{+}81, HG86, JPD86, Man86, Pet87, SB87b]. tomographic [Nol85]. tomography [SGT88]. tool [BK85b, BGM87b, KK83]. topics [RE87]. topological [Dri88b]. topology [DCA89]. toroidal [CLJ^{+}89, CHH^{+}86, CC87, GDM83, HHS89, HK85b, Jar81, Mer86, Sto86, WG86]. Toroidally [RN86]. Torque [CM82]. torus [Shu84]. total [BD83, YWH85]. trace [GWZ85b]. trace-biased [GWZ85b]. Tracing [SSF83]. tracking [BSS82a, CGM^{+}86, HL84, Löt82, SSG81, SH84b, YP88]. traction [Bel82]. trajectory [BC87b]. Transfer [MK82, Bon84, CDS83, Fuc89, FF86, GS87a, HK83, LM87b, Lar88, SC85]. transform [BK85b, CBLL86, GM81, GS85b, Kli83, LR80a, NM80, Pic83, CFR84]. transformation [Aiz85, HW81, Kra83, MFC^{+}85, PG88b]. Transformations [HH88, Tuc89b, GRP82]. transforms [Can81, SS88b, Tem83a, Tem83c]. Transient [Kam81, KH80, KGH80, CL86b, CHS88, HW84, JR85, Keu86, KG87]. transition [FFG87, Jon85, Pes81]. transitions [Bin85]. translating [WOZ84]. transmission [CCCS81, LBR81]. Transmittance [LM87a]. Transonic [Vij86, Boe82, CFG83, DH82, Dic83, EM81, Lyt87, SC82, SZH85]. Transport [KL80, NDP88, AOK81, AEG80, BCE88, BHS87, BST85, BF86, CRS88, CBG86, FC84, GS82, GS83, GF86, Hag87, HM80a, HC87, HPP^{+}82, Her89, HKA89, Jar81, JPD86, KW87, KPBRF83, LBR81, LMM87, Lar88, LM89a, MPA88, MEGH85, MD80, Mih85, MK86, MC85, PC82, PGBO87, Pet87, Rob89a, SP81, SBAM82, SS89, Smo84, SC86, SM89, SMR89, Tam81, WHW87, Yor86, ZBP84]. transversely [Ana89]. trapped [CD80]. trapped-electron [CD80]. travel [SGT88]. traveling [DW81]. treating [Pol87]. treatment [CQ87, HM89, HGM86, KP83b, LBR81, LPS85, MM81b, Phi86, Rai86a, Smi88, Spi88b, ZBP84]. treatments [TB82]. tree [Mak89]. trees [Hal89]. trial [CRS88]. triangular [Fre81, FOF88]. triangulated [OvOH89]. Tricomi [MMV81]. TRIDIF [Fre81]. trigonometric [Lyn84]. triple [JR82]. tritium [NR84]. Truncation [Chi88, Sch80a, ACM88, Hof82, ZMC83]. tsunami [KRW88]. turbulence [BM89, Bas83, But88, FFG87, KD88, KL89, NMG86, RRS89, Sco88, VM86, YP88]. Turbulent [MHK86, Set84, Ano82i, DS86c, Fer87, HP87, Hor87a, Lam81, O’R89]. Turkel [FT84, NN89]. turn [O’R84]. turning [CA85b, PA86, TL81]. TVB [CLS89]. TVD [WW89, YWH85, Yee87]. Two [Col81b, MG89, Rob89a, SK86, SW84, TS81, AG81, AM82, Ana89, And89, AF86, BST86, BS82, BF82, BR86, Cat88, CB84, CL83, CC84, CY85, CH84, DP86, Dyk87, Eas87, EG83, EB81, Fle83, For83b, For84, FB88, FFG87, Gar86, GGR82, Gau88, GMM81, GRP82, Gre85b, GS85b, GM88, GX89, HC87, HK86, HMJ86, Jac89, LA85, LS87, LG81, Lim80, LBS88, Löt82, Mah82, MEGH85, Mak89, MP84, Mei89, MSW87, Mru85, MTM81, OG86, PG88a, Qua81, QR89, Ram85b, Ram86, RM87, RH84, RH88a, RH87, Rok89, Sch82, SK80, Sco85, ST88, Ste84, TR86, Uch83, Wil82, WBL86, WC84, WOZ84, vDR89]. two- [AG81]. two-center [GS85b]. Two-dimensional [SK86, TS81, AM82, Ana89, And89, AF86, CL83, CC84, CY85, CH84, DP86, Eas87, Fle83, For84, FB88, FFG87, GGR82, Gau88, Gre85b, GX89, HC87, HK86, HMJ86, LA85, LG81, Lim80, LBS88, MP84, Mei89, MTM81, OG86, Qua81, Ram85b, Ram86, RM87, SK80, Uch83, WC84, vDR89]. two-factor [Cat88]. two-parameter [EG83]. Two-phase [Rob89a, SW84, BST86, CB84, EB81, LS87, Löt82, RH84, RH88a, TR86]. Two-point [MG89, Gar86, GM88, Jac89, QR89, ST88, WBL86]. two-potential [Mru85]. two-pressure [RH84, RH88a]. two-step [Dyk87, For83b, Mah82, MSW87]. two-valley [MEGH85]. type [BAF84, BDJS86, BB86a, DH89, Dru80, GS85b, Ioa82, IMT^{+}89, Kal84, iKI85, Lev82, Mel88, SS86, TNY85].

Uhlenbeck [FK89]. umbrella [Mez87]. Unavoidable [ERDF83]. Unbiased [Wag87, Wag88]. unbounded [Boy82, KP89a]. undercooled [SLO87]. underwater [FH81, SL85]. unfolding [HS88b]. Unified [Lyn82, Wit84]. uniform [Ana89, CFR84, Hof82, Mor81, MC85, Pan89, ROS85, SMR89]. Uniformly [HEOC87, Wik89]. Unigrid [HMR82]. unitary [KSD81, Shi86]. unsplit [BDS88]. Unstable [She87, GMM81, Hal88, Ker80, Str89, Vie88]. Unsteady [Ing84, SG88b, AFLS83, BE85, BS89b, CC84, Coo86, EM81, GGR89, GG88, Hua81, McL80, RS81]. unstructured [FS89]. update [RWC89]. Updating [BBG^{+}87, Mil89b]. upper [Bai80, GC89]. upstream [BP80, Dav80b, SP81, Vij86]. Upwind [Chr85, OC83, Str82, BW88b, Dav84, FS89, LC89, MA84, MV85]. Usage [MvdV81]. Use [DA86, KD80, LR80b, Pea81, SMR89, Wil80a, AG80, BR80b, BS89a, BM80b, CW86, Cul82, Gar83, Gre84b, HRT80, Hos86, MZ83, MM85b, Pic83, Shi81, SG86, Tur80]. useful [Afa87, Afa89]. Using [BP88, Abd87a, Abd87b, ABS82, Bar87, BSS82b, BB84a, BW81, Boe82, Bor86, Boy87b, CFR84, Cha86, CHH^{+}86, Cla87, CRS88, DS86a, DD89, Die83, Dri88b, DP86, Eck89, EHD89, EHM81, Fou84, GHC^{+}86, GGS82, GS85a, GHK88, Gla88b, GS88, Gus80b, Har81, Hun86, Ino83, Ino85, IMT^{+}89, KKMP86, Ker88, KTT89, LLL88, MT86a, Mih85, MK86, MT86b, Neu81, NMG81, NE82, Noh87, Ohr81a, Ohr81b, PP88a, Pes81, Pic86, PC81a, RS81, RM87, Rap85, Rei81, SH89, Spa85, Ta’87, Tan89, Ves82, Vij86, WG86, YH88a, YH88b]. utilization [CW85a].

V [Kil87]. vacuum [HCG^{+}85, McL80]. validated [OW85]. Validation [SG88a]. validity [GM85]. valley [MEGH85]. Value [Cou85, AM82, Asc80, Coo86, Dav80a, Dav83, Dav81, DD87, Gar86, GP80, HMR82, Jac89, KR84, Mey86, OS83a, QR89, RSR86, SB84a, Sid85, Smo82, Spi88b, Str80, VTB89, WBL86, Wol81, Zeb84]. values [BG85a, WC85a]. valves [MP80a, PM80]. Variable [CL80, LJ81, Alv89, Dar89, DO86, GT85, Got88, HL85, HK86, KK88, LO86, MIK88, Mos85, ZBP84]. variables [Abd87a, EG86, GGR82, Mih85, MK86, VL84, Van86]. variance [BB88, GRP82]. variance-reduction [GRP82]. variances [Mil89b]. variation [Whi87, YWH85]. Variational [ABDV81, CRS88, Sal86, Bar80, Ber83, CSW85, HK81, IA88, LW85, MM89]. variational-iterative [CSW85]. variations [And85, CX86, RA87]. various [Bal85]. varying [Hit80, WK89]. Vector [NOY85, ABS82, CM82, CM89, DQ85, FWHC86, HA89, KE87, McD80, Mül88, RN86, SW81, SMO85, Tut86, WR84, WR86]. Vectorized [Bor86, IMT^{+}89, CM86, Din86]. Vectorizing [Hor87b]. vectors [Roe81]. velocities [BM89]. velocity [Alv89, And83, And86, BM85, FF86, GGR82, GL85, Kaw87, Kli87, Spe87, SMR89, Tre83, Tuc89a]. version [And83, FWHC86]. versus [Boy86]. vertical [Lim80]. very [BJ84, CJW85, CW81, JR81, KS81]. vessels [Set84]. via [Dil85, Lev88]. vibration [Kob85]. vibration-rotation [Kob85]. vibronic [Law87]. vindication [GL85]. violating [BKP87]. visco [HT86]. visco-resistive [HT86]. viscoelastic [Keu86, Lon80]. viscosity [FOF88, GR84, Noh87, Wil80a]. viscous [And89, AD83, BR81, For85, GS89, GX89, KTH88, LG81, MSB87, MK80, O’B81, PM89, PL86, Qua81, QN86, RM87, Rig81, Ros89]. viscous-flow [KTH88]. Vlasov [Law85, Whe81, WMM86, ZGB88, ZBG88]. VLSI [HCM89]. VOF [HN81]. Vogelaere [Moh84]. voltage [MB80]. Volterra [EHM81, Pro82]. Volume [HN81, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano82d, Ano82e, Ano82f, Ano82g, Ano83e, Ano83f, Ano83g, Ano83h, Ano84d, Ano84e, Ano84f, Ano84g, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano88j, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Duk88, Lev88, MP89b, RS81, Vin89, WW89]. Voronoi [Aug84, AP85, Med86, RHOG88, TOO83]. Vortex [GC87, GH86, Leo80, And85, And86, BM87a, BM85, BL82, Cha88, CHS88, CBL81, ET88, GHK88, Gre85a, Ker88, Kra86, MP80a, Mei89, MTM81, Per85, Rob85, SG88a, SS85, Sum89, Tan89, Ten82, Try89, Zuf89]. vortex-in-cell [CBL81, Zuf89]. vortices [CH84, Jon85, MMR89, Riz88]. Vorticity [And89, Qua81, DQ83, FF86, GGR82, O’R84, SS89, Spe87, Tut86, WR84]. vorticity-stream [DQ83]. vorticity-vector [WR84]. vorticity-velocity [FF86, GGR82, Spe87]. Vries [AI80, AA89, HPBY86, MS87a, NS89, Slo88, TA84b, TA88].

walk [FC84, GS85a, GS87a]. walks [Cep83]. Wall [HG84, CHS88, NR84, SS80, Whi87]. Wall-compatibility [HG84]. wall-driven [CHS88]. Walsh [PC81a]. water [Aug84, CL86a, CS82, FN80, For83b, For84, For86, Gil81, Hen81, MM81a, MM81b, Mat86, Nav83, NN89, Ohr81a, Sch80b, Sho86, Wit84]. Wave [Lon80, SS85, Alv89, BBM88, BH87a, BFH86, BP80, CLJ^{+}89, CL83, CL86b, Chi88, DW81, EB81, For83a, Gol86b, HC81b, IMT^{+}89, KG85, KK86, Lev88, LA83, LJ81, MP84, MLB87, MSB87, Ohr81a, Ohr81b, SSC86, Sco88, SL85, Ste87b, SB84b, WMS88, Wei85]. Wave-induced [SS85]. wavefunction [Che84, Kob85]. wavefunctions [Hut84, MT84a]. waveguide [CDP80]. wavelike [GT85]. waves [BGT85, BPS85, BH87a, BL82, CG89, GM85, Hen81, KRW88, MS87a, MOFK88, Mat86, MNM85, Miy86, Mus81, SSC81, Sid85, Wit84]. wavy [Jon85]. way [Gol86b]. Weak [BGM87a, MJ87, RRS89]. weakly [KRW88, SBBY83]. weather [Cul83b, SOS89, Ste87a]. Weber [SGA81]. weighted [SP81]. weighting [Bar87]. welding [DKD85]. well [Hag80, WH87]. well-conditioned [Hag80]. well-known [WH87]. Wendroff [BLT^{+}81]. where [LO87]. which [BM84, OJJ88, SA81, WC85a]. while [Osg81]. wide [MLB87]. Wiener [Hal87]. wind [Kum82, Mel88]. wind-driven [Kum82]. wing [Ros87, SC82]. within [HNK89]. without [HG88]. wrong [Gre85a]. Wu [Fet84].

X [GKL^{+}87]. X-MP [GKL^{+}87]. X-MP/24 [GKL^{+}87].

Zakharov [PND83]. Zernike [HA85]. zero [Hew80, Kru82]. zero-inertia-electron-hybrid [Hew80]. zeroes [Afa89, ANAN84, Dye84]. zeros [BPD81, Dav86, IA85, Pie84]. zeroth [Wal84]. ZIP [Zal81]. zonal [Rai86a]. zone [Nor80]. zoned [BR86]. zoning [BS82]. Zwas [NN89].

References

Aoyagi:1989:PEC

[AA89] Akira Aoyagi and Kanji Abe. Parametric excitation of computational modes inherent to leap-frog scheme applied to the Korteweg–de Vries equation. Journal of Computational Physics, 83(2):447–462, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901307.

Anderson:1981:ACF

[AB81] David V. Anderson and Daniel C. Barnes. Accurate calculations of field-reversed axisymmetric equilibria and their MHD stability properties. Journal of Computational Physics, 42(2):288– 308, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902461.

Aydemir:1984:TDN

[AB84] A. Y. Aydemir and D. C. Barnes. Three-dimensional nonlinear incompressible MHD calculations. Journal of Computational Physics, 53(1):100–123, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918490055X.

Aydemir:1985:IAC

[AB85] A. Y. Aydemir and D. C. Barnes. An implicit algorithm for compressible three-dimensional magnetohydrodynamic calculations. Journal of Computational Physics, 59(1):108–119, May 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918590110X.

Abdallah:1987:NSI

[Abd87a] S. Abdallah. Numerical solutions for the incompressible Navier– Stokes equations in primitive variables using a non-staggered grid, II. Journal of Computational Physics, 70(1):193–202, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918790009X.

Abdallah:1987:NSP

[Abd87b] S. Abdallah. Numerical solutions for the pressure Poisson equation with Neumann boundary conditions using a non-staggered grid, I. Journal of Computational Physics, 70(1):182–192, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900088.

Arickx:1981:VDN

[ABDV81] F. Arickx, J. Broeckhove, E. Deumens, and P. Van Leuven. Variational discretization: A new algorithm for the generator coordinate method. Journal of Computational Physics, 39(2):272– 281, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901522.

Abe:1989:RGF

[Abe89] Takashi Abe. Rarefied gas flow analysis by direct simulation Monte Carlo method in body-fitted coordinate system. Journal of Computational Physics, 83(2):424–432, August 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901289.

Anderson:1982:CME

[ABS82] D. V. Anderson, J. Breazeal, and C. B. Sharp. Calculations of 3D mirror equilibria using a finite element vector potential representation. Journal of Computational Physics, 46(2):189–214, May 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900134.

Allen:1983:NEI

[ABW83] Richard C. Allen, Jr., W. Robert Boland, and G. Milton Wing. Numerical experiments involving Galerkin and collocation methods for linear integral equations of the first kind. Journal of Computational Physics, 49(3):465–477, March 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901419.

Alexander:1988:PSC

[ACM88] S. A. Alexander, R. L. Coldwell, and H. J. Monkhorst. Polyatomic SCF calculations with numerical orbitals. II. Methods to reduce integration and truncation error. Journal of Computational Physics, 76(2):263–280, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890143X.

Aubert:1983:SVF

[AD83] Xavier Aubert and Michel Deville. Steady viscous flows by compact differences in boundary-fitted coordinates. Journal of Computational Physics, 49(3):490–522, March 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901432.

Adams:1987:TES

[AD87] D. J. Adams and G. S. Dubey. Taming the Ewald sum in the computer simulation of charged systems. Journal of Computational Physics, 72(1):156–176, September 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900763.

Adams:1988:IFP

[Ada88] D. J. Adams. The implementation of fluid phase Monte Carlo on the DAP. Journal of Computational Physics, 75(1):138– 150, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901039.

Abarbanel:1986:ICR

[ADG86] Saul S. Abarbanel, Douglas L. Dwoyer, and David Gottlieb. Improving the convergence rate to steady state of parabolic ADI methods. Journal of Computational Physics, 67(1):1–18, November 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186901117.

Adhikari:1981:APT

[Adh81a] Sadhan K. Adhikari. Alternative to Padé technique for iterative solution of integral equations. Journal of Computational Physics, 43(2):382–393, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901303.

Adhikari:1981:ISH

[Adh81b] Sadhan K. Adhikari. Iterative solution of homogeneous integral equations. Journal of Computational Physics, 43(1):189–193, September 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901182.

Aziz:1982:CES

[ADK82] A. K. Aziz, M. R. Dorr, and R. B. Kellogg. Calculation of electromagnetic scattering by a perfect conductor. Journal of Computational Physics, 45(3):403–416, March 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901127.

Aleksandrov:1982:IRP

[ADKM82] L. Aleksandrov, M. Drenska, D. Karadjov, and P. Morozov. Inverse relativistic problem for the bound quark-antiquark states. Journal of Computational Physics, 45(2):291–299, February 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918290122X.

Ascoli:1987:DFS

[ADL87] E. P. Ascoli, D. S. Dandy, and L. G. Leal. On distortion functions for the strong constraint method of numerically generating orthogonal coordinate grids. Journal of Computational Physics, 72(2):513–519, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900994.

Axelrod:1984:ISC

[ADR84] Timothy S. Axelrod, Paul F. Dubois, and Clifford E. Rhoades, Jr. An implicit scheme for calculating time- and frequencydependent flux limited radiation diffusion in one dimension. Journal of Computational Physics, 54(2):205–220, May 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901141.

Alaylioglu:1986:SGS

[AEBS86] Ayse Alaylioglu, D. Eyre, M. Brannigan, and J. P. Svenne. Spline-Galerkin solution of integral equations for three-body scattering above break-up. Journal of Computational Physics, 62(2):383– 399, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918690135X.

Audenaerde:1980:STC

[AEG80] K. Audenaerde, G. A. Emmert, and M. Gordinier. SPUD-NUT: A transport code for neutral atoms in plasmas. Journal of Computational Physics, 34(2):268–284, February 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901096.

Arney:1986:TDM

[AF86] David C. Arney and Joseph E. Flaherty. A two-dimensional mesh moving technique for time-dependent partial differential equations. Journal of Computational Physics, 67(1):124–144, November 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918690118X.

Afanasiev:1987:CAE

[Afa87] G. N. Afanasiev. Closed analytical expressions for some useful sums and integrals involving Legendre functions. Journal of Computational Physics, 69(1):196–208, March 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901628.

Afanasiev:1989:CES

[Afa89] G. N. Afanasiev. Closed expressions for some useful integrals involving Legendre functions and sum rules for zeroes of Bessel functions. Journal of Computational Physics, 85(1):245–252, November 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189902064.

Aggarwal:1983:NOS

[AFLS83] S. K. Aggarwal, G. J. Fix, D. N. Lee, and W. A. Sirignano. Numerical optimization studies of axisymmetric unsteady sprays. Journal of Computational Physics, 50(1):101–115, April 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900438.

Axelsson:1980:UPC

[AG80] O. Axelsson and I. Gustafsson. On the use of preconditioned conjugate gradient methods for red-black ordered five-point difference schemes. Journal of Computational Physics, 35(2):284– 289, April 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180900893.

Abarbanel:1981:OTS

[AG81] Saul Abarbanel and David Gottlieb. Optimal time splitting for two- and three-dimensional Navier–Stokes equations with mixed derivatives. Journal of Computational Physics, 41(1):1–33, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900772.

Aitchison:1985:CHS

[AH85] Joyce M. Aitchison and S. D. Howison. Computation of Hele–Shaw flows with free boundaries. Journal of Computational Physics, 60(3):376–390, September 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900270.

Adams:1988:SMF

[AH88a] J. B. Adams and W. N. G. Hitchon. Solution of Master and Fokker–Planck equations by propagator methods, applied to Au/NaCl thin film nucleation. Journal of Computational Physics, 76(1):159–175, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188901350.

Anderson:1988:SNR

[AH88b] J. L. Anderson and D. W. Hobill. A study of nonlinear radiation damping by matching analytic and numerical solutions. Journal of Computational Physics, 75(2):283–299, April 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901131.

Attenberger:1987:SPC

[AHH87] S. E. Attenberger, W. A. Houlberg, and S. P. Hirshman. Some practical considerations involving spectral representations of 3D plasma equilibria. Journal of Computational Physics, 72(2):435– 448, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900921.

Amit:1981:ANT

[AHP81] R. Amit, C. A. Hall, and T. A. Porsching. An application of network’ theory to the solution of implicit Navier–Stokes difference equations. Journal of Computational Physics, 40(1):183– 201, March 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902060.

Abe:1980:FES

[AI80] Kanji Abe and Osamu Inoue. Fourier expansion solution of the Korteweg–de Vries equation. Journal of Computational Physics, 34(2):202–210, February 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901059.

Aizu:1985:CTI

[Aiz85] Ko Aizu. Canonical transformation invariance and linear multistep formula for integration of Hamiltonian systems. Journal of Computational Physics, 58(2):270–274, April 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901809.

Aki:1984:SPS

[Aki84] Keiiti Aki. Short period seismology. Journal of Computational Physics, 54(1):3–17, April 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184901360.

Alexander:1986:SBS

[Ale86] M. E. Alexander. Simulation of binary-single star and binarybinary scattering. Journal of Computational Physics, 64(1):195– 219, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900252.

Alpert:1981:IFC

[Alp81] Pinhas Alpert. Implicit filtering in conjunction with explicit filtering. Journal of Computational Physics, 44(1):212–219, November 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900474.

Alvin:1989:EFD

[Alv89] K. Alvin. An explicit finite-difference solution to the wave equation with variable velocity. Journal of Computational Physics, 84(1):257, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189901988.

Alzu:1984:TRI

[Alz84] Ko Alzu. Time-reversal invariance and linear multistep methods for integrating dynamical systems. Journal of Computational Physics, 55(1):154–156, July 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900202.

Abarbanel:1982:STD

[AM82] Saul S. Abarbanel and Earll M. Murman. Stability of two-dimensional hyperbolic initial boundary value problems for explicit and implicit schemes. Journal of Computational Physics, 48(2):160–167, November 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900444.

Anagnostopoulos:1989:NSL

[Ana89] Petros Anagnostopoulos. Numerical solution for laminar two-dimensional flow about a fixed and transversely oscillating cylinder in a uniform stream. Journal of Computational Physics, 85(2):434–456, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901587.

Aguilera-Navarro:1984:ZIS

[ANAN84] V. C. Aguilera-Navarro and M. C. K. Aguilera-Navarro. On the zeroes of infinite series. Journal of Computational Physics, 53(1):193–196, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184900615.

Andersen:1983:RVV

[And83] Hans C. Andersen. Rattle: a “velocity” version of the shake algorithm for molecular dynamics calculations. Journal of Computational Physics, 52(1):24–34, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900141.

Anderson:1985:VMF

[And85] Christopher R. Anderson. A vortex method for flows with slight density variations. Journal of Computational Physics, 61(3):417– 444, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900737.

Anderson:1986:MLC

[And86] Christopher R. Anderson. A method of local corrections for computing the velocity field due to a distribution of vortex blobs. Journal of Computational Physics, 62(1):111–123, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186901026.

Anderson:1989:VBC

[And89] Christopher R. Anderson. Vorticity boundary conditions and boundary vorticity generation for two-dimensional viscous incompressible flows. Journal of Computational Physics, 80(1):72– 97, January 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189900910.

Anonymous:1980:Aa

[Ano80a] Anonymous. Announcement. Journal of Computational Physics, 34(1):136, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180901187.

Anonymous:1980:Ab

[Ano80b] Anonymous. Announcement. Journal of Computational Physics, 34(3):435–436, March 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180900996.

Anonymous:1980:Ac

[Ano80c] Anonymous. Announcement. Journal of Computational Physics, 35(3):431, May 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180900807.

Anonymous:1980:Ad

[Ano80d] Anonymous. Announcement. Journal of Computational Physics, 36(1):133, June 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180901783.

Anonymous:1980:AIVa

[Ano80e] Anonymous. Author index for volume 34. Journal of Computational Physics, 34(3):439, March 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901011.

Anonymous:1980:AIVb

[Ano80f] Anonymous. Author index for volume 35. Journal of Computational Physics, 35(3):433, May 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180900820.

Anonymous:1980:AIVc

[Ano80g] Anonymous. Author index for volume 36. Journal of Computational Physics, 36(3):429, July 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901709.

Anonymous:1980:AIVd

[Ano80h] Anonymous. Author index for volume 37. Journal of Computational Physics, 37(3):442, October 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900480.

Anonymous:1980:AIVe

[Ano80i] Anonymous. Author index for volume 38. Journal of Computational Physics, 38(3):421, December 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901606.

Anonymous:1980:E

[Ano80j] Anonymous. Erratum. Journal of Computational Physics, 38(2):257, November 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180900571.

Anonymous:1980:LFAa

[Ano80k]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 34(1):137, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901199.

Anonymous:1980:LFAb

[Ano80l]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 34(2):285, February 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901102.

Anonymous:1980:LFAc

[Ano80m] Anonymous. List of forthcoming articles. Journal of Computational Physics, 34(3):437–438, March 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918090100X.

Anonymous:1980:LFAd

[Ano80n] Anonymous. List of forthcoming articles. Journal of Computational Physics, 35(1):146, March 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918090039X.

Anonymous:1980:LFAe

[Ano80o]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 35(2):290, April 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918090090X.

Anonymous:1980:LFAf

[Ano80p] Anonymous. List of forthcoming articles. Journal of Computational Physics, 35(3):432, May 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180900819.

Anonymous:1980:LFAg

[Ano80q]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 36(1):134, June 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901795.

Anonymous:1980:LFAh

[Ano80r]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 36(2):280, July 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901898.

Anonymous:1980:LFAi

[Ano80s] Anonymous. List of forthcoming articles. Journal of Computational Physics, 36(3):428, July 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901692.

Anonymous:1980:LFAj

[Ano80t]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 37(1):140, August 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900091.

Anonymous:1980:LFAk

[Ano80u] Anonymous. List of forthcoming articles. Journal of Computational Physics, 37(2):287, September 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900297.

Anonymous:1980:LFAl

[Ano80v]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 37(3):441, October 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900479.

Anonymous:1980:LFAm

[Ano80w] Anonymous. List of forthcoming articles. Journal of Computational Physics, 38(1):124, November 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900169.

Anonymous:1980:LFAn

[Ano80x] Anonymous. List of forthcoming articles. Journal of Computational Physics, 38(2):264, November 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900595.

Anonymous:1980:LFAo

[Ano80y] Anonymous. List of forthcoming articles. Journal of Computational Physics, 38(3):420, December 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918090159X.

Anonymous:1980:SCF

[Ano80z] Anonymous. Symposium on computers in flow predictions and fluid dynamics experiments. Journal of Computational Physics, 38(2):258–263, November 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180900583.

Anonymous:1981:Aa

[Ano81a] Anonymous. Announcement. Journal of Computational Physics, 43(2):399, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181901327.

Anonymous:1981:Ab

[Ano81b] Anonymous. Announcements. Journal of Computational Physics, 44(2):400, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181900619.

Anonymous:1981:AIVa

[Ano81c] Anonymous. Author index for volume 39. Journal of Computational Physics, 39(2):486, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901698.

Anonymous:1981:AIVb

[Ano81d] Anonymous. Author index for volume 40. Journal of Computational Physics, 40(2):528, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902291.

Anonymous:1981:AIVc

[Ano81e] Anonymous. Author index for volume 41. Journal of Computational Physics, 41(2):454, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901078.

Anonymous:1981:AIVd

[Ano81f] Anonymous. Author index for volume 42. Journal of Computational Physics, 42(2):414, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902564.

Anonymous:1981:AIVe

[Ano81g] Anonymous. Author index for volume 43. Journal of Computational Physics, 43(2):400, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901339.

Anonymous:1981:AIVf

[Ano81h] Anonymous. Author index for volume 44. Journal of Computational Physics, 44(2):401, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900620.

Anonymous:1981:E

[Ano81i] Anonymous. Erratum. Journal of Computational Physics, 39(1): 249, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901492.

Anonymous:1981:FGC

[Ano81j] Anonymous. Fourth GAMM conference on numerical methods in fluid mechanics: ENSTA, Paris, France, October 7–9, 1981. Journal of Computational Physics, 39(2):484, February 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901674.

Anonymous:1981:LFAa

[Ano81k]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 39(1):250, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901509.

Anonymous:1981:LFAb

[Ano81l]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 39(2):485, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901686.

Anonymous:1981:LFAc

[Ano81m]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 40(2):527, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190228X.

Anonymous:1981:LFAd

[Ano81n] Anonymous. List of forthcoming articles. Journal of Computational Physics, 41(1):231, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181900930.

Anonymous:1981:LFAe

[Ano81o]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 41(2):453, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901066.

Anonymous:1981:LFAf

[Ano81p] Anonymous. List of forthcoming articles. Journal of Computational Physics, 42(1):212, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181902400.

Anonymous:1981:LFAg

[Ano81q]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 42(2):413, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902552.

Anonymous:1981:LFAh

[Ano81r] Anonymous. List of forthcoming articles. Journal of Computational Physics, 43(1):194, September 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901194.

Anonymous:1982:Ab

[Ano82a] Anonymous. Announcement. Journal of Computational Physics, 46(3):475, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900298.

Anonymous:1982:Ac

[Ano82b] Anonymous. Announcement. Journal of Computational Physics, 48(1):157, October 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900419.

Anonymous:1982:Aa

[Ano82c] Anonymous. Announcements. Journal of Computational Physics, 45(2):303–304, February 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182901243.

Anonymous:1982:AIVa

[Ano82d] Anonymous. Author index for volume 45. Journal of Computational Physics, 45(3):469, March 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901152.

Anonymous:1982:AIVb

[Ano82e] Anonymous. Author index for volume 46. Journal of Computational Physics, 46(3):477, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900316.

Anonymous:1982:AIVc

[Ano82f] Anonymous. Author index for volume 47. Journal of Computational Physics, 47(3):498, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901000.

Anonymous:1982:AIVd

[Ano82g] Anonymous. Author index for volume 48. Journal of Computational Physics, 48(3):503, December 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900651.

Anonymous:1982:EF

[Ano82h] Anonymous. Editors’ foreword. Journal of Computational Physics, 48(3):313, December 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900547.

Anonymous:1982:ICN

[Ano82i] Anonymous. International conference on numerical methods in laminar and turbulent flow. Journal of Computational Physics, 46 (1):168, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900109.

Anonymous:1982:LFAa

[Ano82j]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 45(2):305, February 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901255.

Anonymous:1982:LFAb

[Ano82k]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 46(1):169, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900110.

Anonymous:1982:LFAc

[Ano82l] Anonymous. List of forthcoming articles. Journal of Computational Physics, 46(2):328, May 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900183.

Anonymous:1982:LFAd

[Ano82m]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 46(3):476, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900304.

Anonymous:1982:LFAe

[Ano82n] Anonymous. List of forthcoming articles. Journal of Computational Physics, 47(1):164, July 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900717.

Anonymous:1982:LFAf

[Ano82o]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 47(2):330, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900869.

Anonymous:1982:LFAg

[Ano82p] Anonymous. List of forthcoming articles. Journal of Computational Physics, 47(3):497, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900997.

Anonymous:1982:LFAh

[Ano82q]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 48(1):158, October 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900420.

Anonymous:1982:LFAi

[Ano82r] Anonymous. List of forthcoming articles. Journal of Computational Physics, 48(2):312, November 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900535.

Anonymous:1982:LFAj

[Ano82s] Anonymous. List of forthcoming articles. Journal of Computational Physics, 48(3):502, December 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918290064X.

Anonymous:1983:Ab

[Ano83a] Anonymous. Announcement. Journal of Computational Physics, 49(3):523, March 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183901444.

Anonymous:1983:Ac

[Ano83b] Anonymous. Announcement. Journal of Computational Physics, 50(1):178, April 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900487.

Anonymous:1983:Ad

[Ano83c] Anonymous. Announcement. Journal of Computational Physics, 50(3):513, June 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901109.

Anonymous:1983:Aa

[Ano83d] Anonymous. Announcements. Journal of Computational Physics, 49(2):354–355, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183901341.

Anonymous:1983:AIVa

[Ano83e] Anonymous. Author index for volume 49. Journal of Computational Physics, 49(3):525, March 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901468.

Anonymous:1983:AIVb

[Ano83f] Anonymous. Author index for volume 50. Journal of Computational Physics, 50(3):515, June 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901122.

Anonymous:1983:AIVc

[Ano83g] Anonymous. Author index for volume 51. Journal of Computational Physics, 51(3):537, September 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901705.

Anonymous:1983:AIVd

[Ano83h] Anonymous. Author index for volume 52. Journal of Computational Physics, 52(3):603, December 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900128.

Anonymous:1983:LFAa

[Ano83i]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 49(1):179, January 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901213.

Anonymous:1983:LFAb

[Ano83j]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 49(2):356, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901353.

Anonymous:1983:LFAc

[Ano83k]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 49(3):524, March 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901456.

Anonymous:1983:LFAd

[Ano83l]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 50(1):179, April 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900499.

Anonymous:1983:LFAe

[Ano83m] Anonymous. List of forthcoming articles. Journal of Computational Physics, 50(2):322, May 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900700.

Anonymous:1983:LFAf

[Ano83n]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 50(3):514, June 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901110.

Anonymous:1983:LFAg

[Ano83o] Anonymous. List of forthcoming articles. Journal of Computational Physics, 51(1):189, July 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900864.

Anonymous:1983:LFAh

[Ano83p]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 51(2):364, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901006.

Anonymous:1983:LFAi

[Ano83q] Anonymous. List of forthcoming articles. Journal of Computational Physics, 51(3):536, September 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901699.

Anonymous:1983:LFAj

[Ano83r]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 52(1):217, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900281.

Anonymous:1983:LFAk

[Ano83s] Anonymous. List of forthcoming articles. Journal of Computational Physics, 52(2):425, November 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900396.

Anonymous:1983:LFAl

[Ano83t] Anonymous. List of forthcoming articles. Journal of Computational Physics, 52(3):602, December 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900116.

Anonymous:1984:Aa

[Ano84a] Anonymous. Announcement. Journal of Computational Physics, 53(1):206–207, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184900640.

Anonymous:1984:Ab

[Ano84b] Anonymous. Announcement. Journal of Computational Physics, 55(1):173, July 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900238.

Anonymous:1984:Ac

[Ano84c] Anonymous. Announcements. Journal of Computational Physics, 55(3):515–516, September 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184900354.

Anonymous:1984:AIVa

[Ano84d] Anonymous. Author index for volume 53. Journal of Computational Physics, 53(3):536, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900780.

Anonymous:1984:AIVb

[Ano84e] Anonymous. Author index for volume 54. Journal of Computational Physics, 54(3):536, June 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901347.

Anonymous:1984:AIVc

[Ano84f] Anonymous. Author index for volume 55. Journal of Computational Physics, 55(3):518, September 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900378.

Anonymous:1984:AIVd

[Ano84g] Anonymous. Author index for volume 56. Journal of Computational Physics, 56(3):538, December 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918490113X.

Anonymous:1984:Ea

[Ano84h] Anonymous. Erratum. Journal of Computational Physics, 53(1): 205, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900639.

Anonymous:1984:Eb

[Ano84i] Anonymous. Erratum. Journal of Computational Physics, 54 (1):202, April 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918490144X.

Anonymous:1984:LFAa

[Ano84j]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 53(1):208, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900652.

Anonymous:1984:LFAb

[Ano84k]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 53(2):357, February 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900482.

Anonymous:1984:LFAc

[Ano84l]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 53(3):535, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900779.

Anonymous:1984:LFAd

[Ano84m]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 54(1):203, April 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901451.

Anonymous:1984:LFAe

[Ano84n] Anonymous. List of forthcoming articles. Journal of Computational Physics, 54(2):363, May 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184901220.

Anonymous:1984:LFAf

[Ano84o]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 54(3):535, June 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901335.

Anonymous:1984:LFAg

[Ano84p] Anonymous. List of forthcoming articles. Journal of Computational Physics, 55(1):174, July 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918490024X.

Anonymous:1984:LFAh

[Ano84q]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 55(2):346, August 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900123.

Anonymous:1984:LFAi

[Ano84r] Anonymous. List of forthcoming articles. Journal of Computational Physics, 55(3):517, September 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900366.

Anonymous:1984:LFAj

[Ano84s]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 56(1):173, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900901.

Anonymous:1984:LFAk

[Ano84t] Anonymous. List of forthcoming articles. Journal of Computational Physics, 56(2):361, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901025.

Anonymous:1984:LFAl

[Ano84u] Anonymous. List of forthcoming articles. Journal of Computational Physics, 56(3):537, December 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901128.

Anonymous:1985:Aa

[Ano85a] Anonymous. Announcement. Journal of Computational Physics, 57(3):495, February 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185901937.

Anonymous:1985:Ab

[Ano85b] Anonymous. Announcement. Journal of Computational Physics, 60(1):166, August 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185900257.

Anonymous:1985:AIVa

[Ano85c] Anonymous. Author index for volume 57. Journal of Computational Physics, 57(3):497, February 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901950.

Anonymous:1985:AIVb

[Ano85d] Anonymous. Author index for volume 58. Journal of Computational Physics, 58(3):433, May 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901743.

Anonymous:1985:AIVc

[Ano85e] Anonymous. Author index for volume 59. Journal of Computational Physics, 59(3):508, July 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185901299.

Anonymous:1985:AIVd

[Ano85f] Anonymous. Author index for volume 60. Journal of Computational Physics, 60(3):566, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900403.

Anonymous:1985:AIVe

[Ano85g] Anonymous. Author index for volume 61. Journal of Computational Physics, 61(3):525, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900816.

Anonymous:1985:LFAa

[Ano85h] Anonymous. List of forthcoming articles. Journal of Computational Physics, 57(1):155, January 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900579.

Anonymous:1985:LFAb

[Ano85i] Anonymous. List of forthcoming articles. Journal of Computational Physics, 57(2):326, January 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918590049X.

Anonymous:1985:LFAc

[Ano85j]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 57(3):496, February 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901949.

Anonymous:1985:LFAd

[Ano85k]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 58(1):155, March 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901639.

Anonymous:1985:LFAe

[Ano85l]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 58(2):282, April 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901822.

Anonymous:1985:LFAf

[Ano85m] Anonymous. List of forthcoming articles. Journal of Computational Physics, 58(3):432, May 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901731.

Anonymous:1985:LFAg

[Ano85n] Anonymous. List of forthcoming articles. Journal of Computational Physics, 59(1):176, May 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901159.

Anonymous:1985:LFAh

[Ano85o]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 59(2):351, June 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901524.

Anonymous:1985:LFAi

[Ano85p] Anonymous. List of forthcoming articles. Journal of Computational Physics, 59(3):507, July 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185901287.

Anonymous:1985:LFAj

[Ano85q]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 60(1):165, August 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900245.

Anonymous:1985:LFAk

[Ano85r] Anonymous. List of forthcoming articles. Journal of Computational Physics, 60(2):352, September 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900130.

Anonymous:1985:LFAl

[Ano85s] Anonymous. List of forthcoming articles. Journal of Computational Physics, 60(3):565, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900397.

Anonymous:1985:LFAm

[Ano85t]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 61(1):194, October 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900683.

Anonymous:1985:LFAn

[Ano85u] Anonymous. List of forthcoming articles. Journal of Computational Physics, 61(2):349, November 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900920.

Anonymous:1985:LFAo

[Ano85v] Anonymous. List of forthcoming articles. Journal of Computational Physics, 61(3):524, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900804.

Anonymous:1985:NPC

[Ano85w] Anonymous. NBS parallel computer benchmark collection. Journal of Computational Physics, 61(3):523, December 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900798.

Anonymous:1986:Aa

[Ano86a] Anonymous. Announcement. Journal of Computational Physics, 64(1):276, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900331.

Anonymous:1986:Ab

[Ano86b] Anonymous. Announcement. Journal of Computational Physics, 66(2):508, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900781.

Anonymous:1986:Ad

[Ano86c] Anonymous. Announcement. Journal of Computational Physics, 67(2):487, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186902755.

Anonymous:1986:Ac

[Ano86d] Anonymous. Announcements. Journal of Computational Physics, 67(1):237–238, November 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186901269.

Anonymous:1986:AIVa

[Ano86e] Anonymous. Author index for volume 62. Journal of Computational Physics, 62(2):505, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901439.

Anonymous:1986:AIVb

[Ano86f] Anonymous. Author index for volume 63. Journal of Computational Physics, 63(2):487, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918690207X.

Anonymous:1986:AIVc

[Ano86g] Anonymous. Author index for volume 64. Journal of Computational Physics, 64(2):516, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900495.

Anonymous:1986:AIVd

[Ano86h] Anonymous. Author index for volume 65. Journal of Computational Physics, 65(2):492, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902238.

Anonymous:1986:AIVe

[Ano86i] Anonymous. Author index for volume 66. Journal of Computational Physics, 66(2):510, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918690080X.

Anonymous:1986:AIVf

[Ano86j] Anonymous. Author index for volume 67. Journal of Computational Physics, 67(2):489, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902779.

Anonymous:1986:LFAa

[Ano86k]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 62(1):263, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901105.

Anonymous:1986:LFAb

[Ano86l]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 62(2):504, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901427.

Anonymous:1986:LFAc

[Ano86m]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 63(1):246, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900963.

Anonymous:1986:LFAd

[Ano86n]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 63(2):486, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902068.

Anonymous:1986:LFAe

[Ano86o] Anonymous. List of forthcoming articles. Journal of Computational Physics, 64(1):277, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900343.

Anonymous:1986:LFAf

[Ano86p]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 64(2):515, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900483.

Anonymous:1986:LFAg

[Ano86q] Anonymous. List of forthcoming articles. Journal of Computational Physics, 65(1):251, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918690015X.

Anonymous:1986:LFAh

[Ano86r]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 65(2):491, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902226.

Anonymous:1986:LFAi

[Ano86s] Anonymous. List of forthcoming articles. Journal of Computational Physics, 66(1):250, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900641.

Anonymous:1986:LFAj

[Ano86t]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 66(2):509, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900793.

Anonymous:1986:LFAk

[Ano86u] Anonymous. List of forthcoming articles. Journal of Computational Physics, 67(1):236, November 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901257.

Anonymous:1986:LFAl

[Ano86v] Anonymous. List of forthcoming articles. Journal of Computational Physics, 67(2):488, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902767.

Anonymous:1987:Aa

[Ano87a] Anonymous. Announcement. Journal of Computational Physics, 69(1):265, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187901677.

Anonymous:1987:Ab

[Ano87b] Anonymous. Announcement. Journal of Computational Physics, 70(2):526, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918790194X.

Anonymous:1987:Ac

[Ano87c] Anonymous. Announcement. Journal of Computational Physics, 72(1):275, September 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187900829.

Anonymous:1987:AIVa

[Ano87d] Anonymous. Author index for volume 68. Journal of Computational Physics, 68(2):507, February 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900714.

Anonymous:1987:AIVb

[Ano87e] Anonymous. Author index for volume 69. Journal of Computational Physics, 69(2):488, April 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901811.

Anonymous:1987:AIVc

[Ano87f] Anonymous. Author index for volume 70. Journal of Computational Physics, 70(2):528, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901963.

Anonymous:1987:AIVd

[Ano87g] Anonymous. Author index for volume 71. Journal of Computational Physics, 71(2):475, August 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900404.

Anonymous:1987:AIVe

[Ano87h] Anonymous. Author index for volume 72. Journal of Computational Physics, 72(2):521, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918790101X.

Anonymous:1987:AIVf

[Ano87i] Anonymous. Author index for volume 73. Journal of Computational Physics, 73(2):492, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901537.

Anonymous:1987:CPF

[Ano87j] Anonymous. Call for papers first national fluid dynamics congress. Journal of Computational Physics, 71(1):228–229, July 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900295.

Anonymous:1987:LFAa

[Ano87k]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 68(1):249, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900556.

Anonymous:1987:LFAb

[Ano87l]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 68(2):506, February 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900702.

Anonymous:1987:LFAc

[Ano87m]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 69(1):266, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901689.

Anonymous:1987:LFAd

[Ano87n]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 69(2):487, April 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918790180X.

Anonymous:1987:LFAe

[Ano87o] Anonymous. List of forthcoming articles. Journal of Computational Physics, 70(1):269, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187900155.

Anonymous:1987:LFAf

[Ano87p]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 70(2):527, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901951.

Anonymous:1987:LFAg

[Ano87q] Anonymous. List of forthcoming articles. Journal of Computational Physics, 71(1):230, July 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187900301.

Anonymous:1987:LFAh

[Ano87r]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 71(2):474, August 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900398.

Anonymous:1987:LFAi

[Ano87s] Anonymous. List of forthcoming articles. Journal of Computational Physics, 72(1):276, September 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900830.

Anonymous:1987:LFAj

[Ano87t]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 72(2):520, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901008.

Anonymous:1987:LFAk

[Ano87u] Anonymous. List of forthcoming articles. Journal of Computational Physics, 73(1):231, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918790115X.

Anonymous:1987:LFAl

[Ano87v] Anonymous. List of forthcoming articles. Journal of Computational Physics, 73(2):491, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901525.

Anonymous:1988:Aa

[Ano88a] Anonymous. Announcement. Journal of Computational Physics, 74(2):497, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188900927.

Anonymous:1988:Ad

[Ano88b] Anonymous. Announcement. Journal of Computational Physics, 79(2):487, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188900319.

Anonymous:1988:Ab

[Ano88c] Anonymous. Announcements. Journal of Computational Physics, 76(1):231, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901398.

Anonymous:1988:Ac

[Ano88d] Anonymous. Announcements. Journal of Computational Physics, 78(1):249, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188900472.

Anonymous:1988:AIVa

[Ano88e] Anonymous. Author index for volume 74. Journal of Computational Physics, 74(2):499, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900940.

Anonymous:1988:AIVb

[Ano88f] Anonymous. Author index for volume 75. Journal of Computational Physics, 75(2):506, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901271.

Anonymous:1988:AIVc

[Ano88g] Anonymous. Author index for volume 76. Journal of Computational Physics, 76(2):449, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901520.

Anonymous:1988:AIVd

[Ano88h] Anonymous. Author index for volume 77. Journal of Computational Physics, 77(2):550, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901866.

Anonymous:1988:AIVe

[Ano88i] Anonymous. Author index for volume 78. Journal of Computational Physics, 78(2):510, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900642.

Anonymous:1988:AIVf

[Ano88j] Anonymous. Author index for volume 79. Journal of Computational Physics, 79(2):489, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900332.

Anonymous:1988:LFAa

[Ano88k]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 74(1):263, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900800.

Anonymous:1988:LFAb

[Ano88l]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 74(2):498, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900939.

Anonymous:1988:LFAc

[Ano88m]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 75(1):251, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901118.

Anonymous:1988:LFAd

[Ano88n]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 75(2):505, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890126X.

Anonymous:1988:LFAe

[Ano88o] Anonymous. List of forthcoming articles. Journal of Computational Physics, 76(1):232, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188901404.

Anonymous:1988:LFAf

[Ano88p]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 76(2):448, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901519.

Anonymous:1988:LFAg

[Ano88q] Anonymous. List of forthcoming articles. Journal of Computational Physics, 77(1):282, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188901696.

Anonymous:1988:LFAh

[Ano88r]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 77(2):549, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901854.

Anonymous:1988:LFAi

[Ano88s] Anonymous. List of forthcoming articles. Journal of Computational Physics, 78(1):250, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900484.

Anonymous:1988:LFAj

[Ano88t]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 78(2):509, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900630.

Anonymous:1988:LFAk

[Ano88u] Anonymous. List of forthcoming articles. Journal of Computational Physics, 79(1):244, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900150.

Anonymous:1988:LFAl

[Ano88v] Anonymous. List of forthcoming articles. Journal of Computational Physics, 79(2):488, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900320.

Anonymous:1989:APAa

[Ano89a] Anonymous. Abstract of papers to appear in future issues. Journal of Computational Physics, 84(2):493–496, October 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902477.

Anonymous:1989:APAb

[Ano89b] Anonymous. Abstract of papers to appear in future issues. Journal of Computational Physics, 85(1):253–255, November 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189902076.

Anonymous:1989:Aa

[Ano89c] Anonymous. Announcement. Journal of Computational Physics, 81(1):236, March 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189900739.

Anonymous:1989:Ab

[Ano89d] Anonymous. Announcement. Journal of Computational Physics, 82(1):245, May 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189900466.

Anonymous:1989:Ac

[Ano89e] Anonymous. Announcement. Journal of Computational Physics, 84(2):492, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189902465.

Anonymous:1989:AIVa

[Ano89f] Anonymous. Author index for volume 80. Journal of Computational Physics, 80(2):513, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901174.

Anonymous:1989:AIVb

[Ano89g] Anonymous. Author index for volume 81. Journal of Computational Physics, 81(2):490, April 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902192.

Anonymous:1989:AIVc

[Ano89h] Anonymous. Author index for volume 82. Journal of Computational Physics, 82(2):502, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900624.

Anonymous:1989:AIVd

[Ano89i] Anonymous. Author index for volume 83. Journal of Computational Physics, 83(2):496, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901344.

Anonymous:1989:AIVe

[Ano89j] Anonymous. Author index for volume 84. Journal of Computational Physics, 84(2):497, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902489.

Anonymous:1989:AIVf

[Ano89k] Anonymous. Author index for volume 85. Journal of Computational Physics, 85(2):505, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901794.

Anonymous:1989:LFAa

[Ano89l]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 80(1):252, January 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900983.

Anonymous:1989:LFAb

[Ano89m]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 80(2):512, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901162.

Anonymous:1989:LFAc

[Ano89n]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 81(1):237, March 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900740.

Anonymous:1989:LFAd

[Ano89o]  Anonymous.  List of forthcoming articles.  Journal of

Computational Physics, 81(2):489, April 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902180.

Anonymous:1989:LFAe

[Ano89p] Anonymous. List of forthcoming articles. Journal of Computational Physics, 82(1):246, May 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189900478.

Anonymous:1989:LFAf

[Ano89q] Anonymous. List of forthcoming articles. Journal of Computational Physics, 82(2):499–501, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900612.

Anonymous:1989:LFAg

[Ano89r] Anonymous. List of forthcoming articles. Journal of Computational Physics, 83(1):246, July 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189902325.

Anonymous:1989:LFAh

[Ano89s] Anonymous. List of forthcoming articles. Journal of Computational Physics, 83(2):494–495, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901332.

Anonymous:1989:NA

[Ano89t] Anonymous. Notes to appear. Journal of Computational Physics, 85(2):504, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189901782.

Andrade:1981:SFP

[AOK81] A. Andrade, T. A. Oliphant, and T. Kammash. Solution of the Fokker–Planck transport equation by matrix factorization. Journal of Computational Physics, 42(2):367–381, August 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902503.

Augenbaum:1985:CVM

[AP85] Jeffrey M. Augenbaum and Charles S. Peskin. On the construction of the Voronoi mesh on a sphere. Journal of Computational Physics, 59(2):177–192, June 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901408.

Archer:1988:CSC

[AP88] B. J. Archer and G. A. Parker. A cross section code for diatom-spherical particle inelastic scattering. Journal of Computational Physics, 78(1):15–35, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900356.

Aregbesola:1981:NSL

[Are81] Y. A. S. Aregbesola. A numerical solution for the Laplace equation with normal derivative boundary condition. Journal of Computational Physics, 39(1):241–248, January 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901480.

Asaithambi:1987:CFS

[Asa87] N. S. Asaithambi. Computation of free-surface flows. Journal of Computational Physics, 73(2):380–394, December 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901434.

Ascher:1980:SBV

[Asc80] Uri Ascher. Solving boundary-value problems with a splinecollocation code. Journal of Computational Physics, 34(3):401– 413, March 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180900972.

Abe:1986:HOS

[ASIO86] Hirotade Abe, Natsuhiko Sakairi, Ryohei Itatani, and Hideo Okuda. High-order spline interpolations in the particle simulation. Journal of Computational Physics, 63(2):247–267, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186901932.

Adam:1982:ESC

[ASL82] J. C. Adam, A. Gourdin Serveniere, and A. B. Langdon. Electron sub-cycling in particle simulation of plasma. Journal of Computational Physics, 47(2):229–244, August 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900766.

Augenbaum:1984:LMS

[Aug84] Jeffrey M. Augenbaum. A Lagrangian method for the shallow water equations based on a Voronoi mesh-one dimensional results. Journal of Computational Physics, 53(2):240–265, February 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900408.

Ben-artzi:1989:GRP

[Ba89] Matania Ben-artzi. The generalized Riemann problem for reactive flows. Journal of Computational Physics, 81(1):70–101, March 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918990065X.

Baart:1981:CES

[Baa81] M. L. Baart. Computational experience with the spectral smoothing method for differentiating noisy data. Journal of Computational Physics, 42(1):141–151, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902370.

Ben-artzi:1986:AGR

[BaB86] Matania Ben-artzi and Amnon Birman. Application of the “Generalized Riemann Problem” method to 1-D compressible flows with material interfaces. Journal of Computational Physics, 65(1):170–178, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900100.

Ben-Artzi:1984:SOG

[BAF84] Matania Ben-Artzi and Joseph Falcovitz. A second-order Godunov-type scheme for compressible fluid dynamics. Journal of Computational Physics, 55(1):1–32, July 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900135.

Bailey:1980:NSC

[Bai80] G. J. Bailey. Numerical studies of coupled time-dependent diffusion equations for the Earth’s upper atmosphere: A method employing a self-diffusion coefficient. Journal of Computational Physics, 37(3):422–440, October 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900467.

Baker:1981:PFC

[Bak81] Timothy J. Baker. Potential flow calculation by the approximate factorisation method. Journal of Computational Physics, 42(1): 1–19, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902308.

Balarin:1985:AVA

[Bal85] M. Balarin. Accuracy of various approximations to the exponential integral in glow curve theory. Journal of Computational Physics, 57(1):26–42, January 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900518.

Barber:1980:OVA

[Bar80] Michael N. Barber. Optimal variational approximations to renormalization groups. II. Determination of optimal parameters. Journal of Computational Physics, 34(3):414–434, March 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900984.

Barnett:1982:CFE

[Bar82] A. R. Barnett. Continued-fraction evaluation of Coulomb functions F_{λ}(η, x), G_{λ}(η, x) and their derivatives. Journal of Computational Physics, 46(2):171–188, May 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900122.

Barlow:1987:ECU

[Bar87] Roger Barlow. Event classification using weighting methods. Journal of Computational Physics, 72(1):202–219, September 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900787.

Basdevant:1983:TID

[Bas83] C. Basdevant. Technical improvements for direct numerical simulation of homogeneous three-dimensional turbulence. Journal of Computational Physics, 50(2):209–214, May 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900645.

Basombrio:1984:FEF

[Bas84] Fernando G. Basombrio. Finite element formulation of radioisotope diffusion in metal grain textures. Journal of Computational Physics, 54(2):237–244, May 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901165.

Brackbill:1980:ENN

[BB80] J. U. Brackbill and D. C. Barnes. The effect of nonzero ∇ · B on the numerical solution of the magnetohydrodynamic equations. Journal of Computational Physics, 35(3):426–430, May 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900790.

Baker:1983:DNE

[BB83] George A. Baker, Jr. and John E. Brolley. Deconvolution of noisy experimental data. Journal of Computational Physics, 51(2):227– 240, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918390089X.

Bernhardt:1984:SEE

[BB84a] Paul A. Bernhardt and J. U. Brackbill. Solution of elliptic equations using fast Poisson solvers. Journal of Computational Physics, 53(3):382–394, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900676.

Booth:1984:CLS

[BB84b] Ian J. M. Booth and Andrew D. Booth. On a class of least-squares curve-fitting problems. Journal of Computational Physics, 53(1): 72–81, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900524.

Bieterman:1986:AML

[BB86a] M. Bieterman and I. Babuska. An adaptive method of lines with error control for parabolic equations of the reaction-diffusion type. Journal of Computational Physics, 63(1):33–66, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900835.

Booth:1986:ISS

[BB86b] Andrew D. Booth and Ian J. M. Booth. An interesting set of simultaneous, nonlinear equations. Journal of Computational Physics, 65(2):481–485, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902202.

Booth:1988:NPC

[BB88] Andrew D. Booth and Ian J. M. Booth. A note on the progressive calculation of the mean and variance. Journal of Computational Physics, 77(2):537–539, August 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901830.

Bauer:1980:NSN

[BBG80] F. Bauer, O. Betancourt, and P. Garabedian. Numerical studies of new stellarator concepts. Journal of Computational Physics, 35(3):341–355, May 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180900741.

Barbour:1987:UFL

[BBG^{+}87] I. M. Barbour, N.-E Behilil, P. E. Gibbs, M. Rafiq, K. J. M. Moriarty, and G. Schierholz. Updating fermions with the Lanczos method. Journal of Computational Physics, 68(1):227–236, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900532.

Buneman:1980:PCD

[BBGN80] O. Buneman, C. W. Barnes, J. C. Green, and D. E. Nielsen. Principles and capabilities of 3-D, E–M particle simulations. Journal of Computational Physics, 38(1):1–44, November 1, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900108.

Barry:1988:ABC

[BBM88] A. Barry, J. Bielak, and R. C. MacCamy. On absorbing boundary conditions for wave propagation. Journal of Computational Physics, 79(2):449–468, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900253.

Bennett:1987:AFD

[BC87a] John R. Bennett and Joan E. Campbell. Accuracy of a finitedifference method for computing lake currents. Journal of Computational Physics, 68(2):262–271, February 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918790057X.

Bennett:1987:ATC

[BC87b] John R. Bennett and Anne Hutchinson Clites. Accuracy of trajectory calculation in a finite-difference circulation model. Journal of Computational Physics, 68(2):272–282, February 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900581.

Berger:1989:LAM

[BC89] Marsha J. Berger and Phillip Colella. Local adaptive mesh refinement for shock hydrodynamics. Journal of Computational Physics, 82(1):64–84, May 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189900351.

Lawrence Livermore Laboratory Report No. UCRL-97196.

Barnes:1988:SMF

[BCE88] Michael S. Barnes, Tina J. Cotler, and Michael E. Elta. A staggered-mesh finite-difference numerical method for solving the transport equations in low pressure rf glow discharges. Journal of Computational Physics, 77(1):53–72, July 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901568.

Bell:1989:SOP

[BCG89] John B. Bell, Phillip Colella, and Harland M. Glaz. A secondorder projection method for the incompressible Navier–Stokes equations. Journal of Computational Physics, 85(2):257–283, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901514.

Bell:1989:HOG

[BCT89] John B. Bell, Phillip Colella, and John A. Trangenstein. Higher order Godunov methods for general systems of hyperbolic conservation laws. Journal of Computational Physics, 82(2):362– 397, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189900545.

Bramley:1982:CES

[BD82] J. S. Bramley and S. C. R. Dennis. The calculation of eigenvalues for the stationary perturbation of Poiseuille flow. Journal of Computational Physics, 47(2):179–198, August 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900730.

Brown:1983:NEP

[BD83] M. G. Brown and J. W. Dungey. Note economising plasma simulation by total neglect of the displacement current. Journal of Computational Physics, 52(1):205–208, October 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900256.

Bette:1986:BPC

[BDJS86] Srinivas Betté, Julio C. Diaz, William R. Jines, and Trond Steihaug. A block preconditioned conjugate gradient-type iterative solver for linear systems in thermal reservoir simulation. Journal of Computational Physics, 67(1):37–58, November 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901142.

Brandt:1980:MMS

[BDR80] Achi Brandt, J. E. Dendy, Jr., and Hans Ruppel. The multigrid method for semi-implicit hydrodynamics codes. Journal of Computational Physics, 34(3):348–370, March 1, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900947.

Barnett:1982:IBC

[BDR82] M. Barnett, R. T. Davis, and J. V. Rakich. Implicit boundary conditions for the solution of the parabolized Navier– Stokes equations for supersonic flows. Journal of Computational Physics, 48(2):168–181, November 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900456.

Bell:1988:UHO

[BDS88] John B. Bell, Clint N. Dawson, and Gregory R. Shubin. An unsplit, higher order Godunov method for scalar conservation laws in multiple dimensions. Journal of Computational Physics, 74(1): 1–24, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900654.

Bercovier:1980:FEM

[BE80] Michel Bercovier and Michael Engelman. A finite-element method for incompressible non-Newtonian flows. Journal of Computational Physics, 36(3):313–326, July 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901631.

Botta:1985:MSM

[BE85] Eugen F. F. Botta and Marcel H. M. Ellenbroek. A modified SOR method for the Poisson equation in unsteady free-surface flow calculations. Journal of Computational Physics, 60(1):119– 134, August 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900208.

Belward:1982:NST

[Bel82] J. A. Belward. Numerical solutions of the traction problem for a fibre-reinforced material by an integral-equation method. Journal of Computational Physics, 46(2):215–236, May 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900146.

Berryman:1983:CVB

[Ber83] James G. Berryman. Computing variational bounds for flow through random aggregates of spheres. Journal of Computational Physics, 52(1):142–162, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900219.

Bergers:1985:KMS

[Ber85a] Diethard Bergers. Kinetic model solution for axisymmetric flow by the method of discrete ordinates. Journal of Computational Physics, 57(2):285–302, January 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900464.

Beumee:1985:PFS

[BER85b] Johan G. B. Beumeé, Larry Eno, and Herschel Rabitz. Pointwise feature sensitivity analysis. Journal of Computational Physics, 57(2):318–325, January 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900488.

Berryman:1988:IIT

[Ber88] James G. Berryman. Interpolating and integrating threepoint correlation functions on a lattice. Journal of Computational Physics, 75(1):86–102, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901003.

Brackbill:1982:IME

[BF82] J. U. Brackbill and D. W. Forslund. An implicit method for electromagnetic plasma simulation in two dimensions. Journal of Computational Physics, 46(2):271–308, May 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918290016X.

Brooks:1986:IMC

[BF86] E. D. Brooks III and J. A. Fleck, Jr. An implicit Monte Carlo scheme for calculating time-dependent line transport. Journal of Computational Physics, 67(1):59–72, November 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901154.

Brannan:1986:HFN

[BFH86] J. R. Brannan, G. P. Forney, and R. F. Henrick. High-frequency numerical-analytic approximations for the separable acoustic wave equation. Journal of Computational Physics, 66(1):21–35, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900513.

Brandt:1985:ISM

[BFT85] Achi Brandt, Scott R. Fulton, and G. D. Taylor. Improved spectral multigrid methods for periodic elliptic problems. Journal of Computational Physics, 58(1):96–112, March 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901597.

Buland:1984:CFO

[BG84] Ray Buland and Freeman Gilbert. Computation of free oscillations of the Earth. Journal of Computational Physics, 54(1): 95–114, April 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901414.

Buendia:1985:CVC

[BG85a] E. Buendia and R. Guardiola. Critical values of the coupling constant of short range potentials. Journal of Computational Physics, 60(2):208–218, September 15, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900038.

Buendia:1985:NMS

[BG85b] E. Buendía and R. Guardiola. The Numerov method and singular potentials. Journal of Computational Physics, 60(3):561–564, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900385.

Brower:1987:WDP

[BGM87a] Richard C. Brower, Roscoe Giles, and K. J. M. Moriarty. Weak decay processes in lattice QCD. Journal of Computational Physics, 73(2):461–467, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901471.

Buendia:1987:IET

[BGM87b] E. Buendía, R. Guardiola, and M. M. Montoya. Integral equations: A tool to solve the Schrödinger equation. Journal of Computational Physics, 68(1):188–201, January 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900519.

Bayliss:1989:APS

[BGMM89] A. Bayliss, D. Gottlieb, B. J. Matkowsky, and M. Minkoff. An adaptive pseudo-spectral method for reaction diffusion problems. Journal of Computational Physics, 81(2):421–443, April 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189902155.

Brower:1989:CRG

[BGMT89] R. C. Brower, R. Giles, K. J. M. Moriarty, and P. Tamayo. Combining renormalization group and multigrid methods. Journal of Computational Physics, 80(2):472–479, February 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901113.

Bontoux:1980:AEB

[BGR80] Patrick Bontoux, Bernard Gilly, and Bernard Roux. Analysis of the effect of boundary conditions on numerical stability of solutions of Navier–Stokes equations. Journal of Computational Physics, 36(3):417–427, July 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901680.

Bayliss:1983:IMH

[BGT83] Alvin Bayliss, Charles I. Goldstein, and Eli Turkel. An iterative method for the Helmholtz equation. Journal of Computational Physics, 49(3):443–457, March 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901390.

Bayliss:1985:ACN

[BGT85] A. Bayliss, C. I. Goldstein, and E. Turkel. On accuracy conditions for the numerical computation of waves. Journal of Computational Physics, 59(3):396–404, July 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901196.

Booij:1987:POW

[BH87a] Nico Booij and Leo H. Holthuijsen. Propagation of ocean waves in discrete spectral wave models. Journal of Computational Physics, 68(2):307–326, February 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918790060X.

Byrne:1987:SOS

[BH87b] George D. Byrne and Alan C. Hindmarsh. Stiff ODE solvers: A review of current and coming attractions. Journal of Computational Physics, 70(1):1–62, May 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900015.

Betancourt:1983:CMS

[BHM^{+}83] O. Betancourt, F. Herrnegger, P. Merkel, J. Nührenberg, R. Gruber, and F. Troyon. Comparison of MHD stability results obtained with the BETA 3D and HERA 2D codes. Journal of Computational Physics, 52(1):187–197, October 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900232.

Beidler:1987:HMC

[BHS87] C. D. Beidler, W. N. G. Hitchon, and J. L. Shohet. ”hybrid” Monte Carlo simulation of ripple transport in stellarators. Journal of Computational Physics, 72(1):220–242, September 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900799.

Binder:1985:MCM

[Bin85] K. Binder. The Monte Carlo method for the study of phase transitions: A review of some recent progress. Journal of Computational Physics, 59(1):1–55, May 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901068.

Birnboim:1986:SST

[Bir86] A. Birnboim. Simultaneous solution of temperatures in plasmas with rapid equipartition rates. Journal of Computational Physics, 63(2):353–362, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901981.

Bonnerot:1981:CFE

[BJ81] R. Bonnerot and P. Jamet. A conservative finite element method for one-dimensional Stefan problems with appearing and disappearing phases. Journal of Computational Physics, 41(2):357– 388, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901017.

Bucher:1984:SVL

[BJ84] Ingrid Y. Bucher and Thomas L. Jordan. Solving very large elliptic problems on a supercomputer with solid state disk. Journal of Computational Physics, 55(2):340–345, August 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900111.

Belyaev:1985:FDA

[BK85a] V. B. Belyaev and O. I. Kartavtsev. Finite-dimensional approximation of the differential operator in problems of quantum mechanics. Journal of Computational Physics, 59(3):493–498, July 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185901263.

Bisseling:1985:FHT

[BK85b] Rob Bisseling and Ronnie Kosloff. The fast Hankel transform as a tool in the solution of the time dependent Schrödinger equation. Journal of Computational Physics, 59(1):136–151, May 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185901123.

Bolstad:1987:CAM

[BK87] J. H. Bolstad and H. B. Keller. Computation of anomalous modes in the Taylor experiment. Journal of Computational Physics, 69(1):230–251, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901641.

Bisseling:1988:OCG

[BK88a] R. H. Bisseling and R. Kosloff. Optimal choice of grid points in multidimensional pseudospectral Fourier methods. Journal of Computational Physics, 76(2):243–262, June 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901428.

Blaschak:1988:CSA

[BK88b] Jeffrey G. Blaschak and Gregory A. Kriegsmann. A comparative study of absorbing boundary conditions. Journal of Computational Physics, 77(1):109–139, July 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901593.

Biegler-Konig:1986:NNM

[BKH86] F. Biegler-König and Juergen Hinze. Nonrelativistic numerical MCSCF for atoms. Journal of Computational Physics, 67(2):290– 309, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186902640.

Barnes:1983:IPS

[BKLT83] D. C. Barnes, T. Kamimura, J.-N Leboeuf, and T. Tajima. Implicit particle simulation of magnetized plasmas. Journal of Computational Physics, 52(3):480–502, December 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900049.

Berezin:1987:CFD

[BKP87] Yu. A. Berezin, V. N. Khudick, and M. S. Pekker. Conservative finite-difference schemes for the Fokker–Planck equation not violating the law of an increasing entropy. Journal of Computational Physics, 69(1):163–174, March 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901604.

Burbea:1982:KWV

[BL82] Jacob Burbea and Michael Landau. The Kelvin waves in vortex dynamics and their stability. Journal of Computational Physics, 45(1):127–156, January 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901061.

Baum:1985:CSN

[BL85] Joseph D. Baum and Jay N. Levine. A critical study of numerical methods for the solution of nonlinear hyperbolic equations for resonance systems. Journal of Computational Physics, 58(1):1– 28, March 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901536.

Basden:1988:HON

[BL88] Bryan Basden and Robert R. Lucchese. High-order Newton– Cotes integration methods in scattering theory. Journal of Computational Physics, 77(2):524–536, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901829.

Blottner:1982:IBA

[Blo82] F. G. Blottner. Influence of boundary approximations and conditions on finite-difference solutions. Journal of Computational Physics, 48(2):246–269, November 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900493.

Brunel:1981:MPC

[BLT^{+}81] F. Brunel, J. N. Leboeuf, T. Tajima, J. M. Dawson, M. Makino, and T. Kamimura. Magnetohydrodynamic particle code: Lax– Wendroff algorithm with finer grid interpolations. Journal of Computational Physics, 43(2):268–288, October 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901236.

Birdsall:1980:PSH

[BM80a] Charles K. Birdsall and Neil Maron. Plasma self-heating and saturation due to numerical instabilities. Journal of Computational Physics, 36(1):1–19, June 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901710.

Briley:1980:SUL

[BM80b] W. R. Briley and H. McDonald. On the structure and use of linearized block implicit schemes. Journal of Computational Physics, 34(1):54–73, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901126.

Bridges:1984:DEP

[BM84] T. J. Bridges and P. J. Morris. Differential eigenvalue problems in which the parameter appears nonlinearly. Journal of Computational Physics, 55(3):437–460, September 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900329.

Beale:1985:HOA

[BM85] J. T. Beale and A. Majda. High order accurate vortex methods with explicit velocity kernels. Journal of Computational Physics, 58(2):188–208, April 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901767.

Baba:1987:HOA

[BM87a] Nobuhiro Baba and Hideaki Miyata. Higher-order accurate difference solutions of vortex generation from a circular cylinder in an oscillatory flow. Journal of Computational Physics, 69(2): 362–396, April 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901732.

Bayliss:1987:FRO

[BM87b] A. Bayliss and B. J. Matkowsky. Fronts, relaxation oscillations, and period doubling in solid fuel combustion. Journal of Computational Physics, 71(1):147–168, July 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900246.

Balachandar:1989:MEF

[BM89] S. Balachandar and M. R. Maxey. Methods for evaluating fluid velocities in spectral simulations of turbulence. Journal of Computational Physics, 83(1):96–125, July 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902246.

Briggs:1983:FMI

[BNS83] William L. Briggs, Alan C. Newell, and Talib Sarie. Focusing: A mechanism for instability of nonlinear finite difference equations. Journal of Computational Physics, 51(1):83–106, July 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900827.

Berger:1984:AMR

[BO84] Marsha J. Berger and Joseph Oliger. Adaptive mesh refinement for hyperbolic partial differential equations. Journal of Computational Physics, 53(3):484–512, March 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900731.

Boerstoel:1982:MAS

[Boe82] J. W. Boerstoel. A multigrid algorithm for steady transonic potential flows around aerofoils using Newton iteration. Journal of Computational Physics, 48(3):314–343, December 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900559.

Bogomolny:1988:MAS

[Bog88] Alexander Bogomolny. A method of approximating stiffness matrices for some pseudodifferential equations. Journal of Computational Physics, 74(1):85–93, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900708.

Bond:1984:ACM

[Bon84] D. J. Bond. Application of the Chebyshev method to radiative transfer calculations for laser heated targets. Journal of Computational Physics, 55(3):369–376, September 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900263.

Booth:1981:EMC

[Boo81] Thomas E. Booth. Exact Monte Carlo solution of elliptic partial differential equations. Journal of Computational Physics, 39(2):396–404, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901595.

Booth:1982:AO

[Boo82a] Andrew D. Booth. On the anharmonic oscillator. Journal of Computational Physics, 46(3):423–428, June 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900249.

Booth:1982:RMC

[Boo82b] Thomas E. Booth. Regional Monte Carlo solution of elliptic partial differential equations. Journal of Computational Physics, 47(2):281–290, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900791.

Boris:1986:VNN

[Bor86] Jay Boris. A vectorized ‘near neighbors’ algorithm of order N using a monotonic logical grid. Journal of Computational Physics, 66(1):1–20, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900501.

Bowen:1984:GDG

[Bow84] Jeffrey M. Bowen. Graphic displays of gravitational initial data. Journal of Computational Physics, 56(1):42–50, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900822.

Boyd:1982:OCC

[Boy82] John P. Boyd. The optimization of convergence for Chebyshev polynomial methods in an unbounded domain. Journal of Computational Physics, 45(1):43–79, January 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901024.

Boyd:1984:ACH

[Boy84] John P. Boyd. Asymptotic coefficients of Hermite function series. Journal of Computational Physics, 54(3):382–410, June 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184901244.

Boyd:1985:CCM

[Boy85] John P. Boyd. Complex coordinate methods for hydrodynamic instabilities and Sturm–Liouville eigenproblems with an interior singularity. Journal of Computational Physics, 57(3):454–471, February 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901901.

Boyd:1986:PSV

[Boy86] John P. Boyd. Polynomial series versus sinc expansions for functions with corner or endpoint singularities. Journal of Computational Physics, 64(1):266–270, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900318.

Boyd:1987:ORF

[Boy87a] John P. Boyd. Orthogonal rational functions on a semi-infinite interval. Journal of Computational Physics, 70(1):63–88, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900027.

Boyd:1987:SMU

[Boy87b] John P. Boyd. Spectral methods using rational basis functions on an infinite interval. Journal of Computational Physics, 69(1):112– 142, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901586.

Brown:1980:GWP

[BP80] Philip S. Brown, Jr. and Joseph P. Pandolfo. A gravitywave problem with the upstream difference method. Journal of Computational Physics, 37(2):141–150, September 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900170.

Borgers:1987:LFS

[BP87] Christoph Börgers and Charles S. Peskin. A Lagrangian fractional step method for the incompressible Navier–Stokes equations on a periodic domain. Journal of Computational Physics, 70(2):397– 438, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901896.

Boyer:1988:MDC

[BP88] L. L. Boyer and G. S. Pawley. Molecular dynamics of clusters of particles interacting with pairwise forces using a massively parallel computer. Journal of Computational Physics, 78(2):405– 423, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900575.

Branrers:1981:RAZ

[BPD81] M. Branrers, R. Piessens, and M. De Meue. Rational approximations for zeros of Bessel functions. Journal of Computational Physics, 42(2):403–405, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902539.

Bona:1985:NSM

[BPS85] J. L. Bona, W. G. Pritchard, and L. R. Scott. Numerical schemes for a model for nonlinear dispersive waves. Journal of Computational Physics, 60(2):167–186, September 15, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900014.

Barnhill:1980:ICS

[BR80a] Maurice V. Barnhill III and Craig Russ. Improving the convergence of solutions to the Fredholm integral equation of the first kind. Journal of Computational Physics, 37(2):268–273, September 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900261.

Bell:1980:UOE

[BR80b] G. E. Bell and S. I. M. Ritchie. The use of the odd-even hopscotch algorithm for a thermal shock problem. Journal of Computational Physics, 38(3):413–419, December 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901588.

Brandeis:1981:IMC

[BR81] Julius Brandeis and Josef Rom. Interactive method for computation of viscous flow with recirculation. Journal of Computational Physics, 40(2):396–410, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902187.

Brackbill:1986:FMA

[BR86] J. U. Brackbill and H. M. Ruppel. FLIP: A method for adaptively zoned, particle-in-cell calculations of fluid flows in two dimensions. Journal of Computational Physics, 65(2):314–343, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186902111.

Bramley:1984:NCE

[Bra84] J. S. Bramley. Note on the calculation of eigenvalues for the stationary perturbation of Poiseuille flow. Journal of Computational Physics, 53(3):524–529, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900755.

Bramley:1986:CES

[Bra86] J. S. Bramley. The calculation of eigenvalues for the stationary perturbation of symmetrical pipe Poiseuille flow. Journal of Computational Physics, 64(1):258–265, May 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900306.

Brackbill:1988:RIP

[Bra88] J. U. Brackbill. The ringing instability in particle-in-cell calculations of low-speed flow. Journal of Computational Physics, 75(2): 469–492, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901234.

Brown:1980:MEN

[Bro80] Philip S. Brown, Jr. More efficient numerical solution of diurnal kinetics problems. Journal of Computational Physics, 34(1):111– 123, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180901163.

Brown:1985:ISE

[Bro85] Philip S. Brown, Jr. An implicit scheme for efficient solution of the coalescence/collision-breakup equation. Journal of Computational Physics, 58(3):417–431, May 15, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918590172X.

Brooks:1989:SIM

[Bro89] Eugene D. Brooks III. Symbolic implicit Monte Carlo. Journal of Computational Physics, 83(2):433–446, August 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901290.

Bennett:1981:CRN

[BS81] John R. Bennett and David J. Schwab. Calculation of the rotational normal modes of oceans and lakes with general orthogonal coordinates. Journal of Computational Physics, 44(2):359–376, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181900577.

Brackbill:1982:AZS

[BS82] J. U. Brackbill and J. S. Saltzman. Adaptive zoning for singular problems in two dimensions. Journal of Computational Physics, 46(3):342–368, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900201.

Bell:1983:AGF

[BS83] J. B. Bell and G. R. Shubin. An adaptive grid finite difference method for conservation laws. Journal of Computational Physics, 52(3):569–591, December 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900086.

Brazier-Smith:1984:LSH

[BS84] P. R. Brazier-Smith. On the limitations of spherical harmonics for the solution of Laplace’s equation. Journal of Computational Physics, 54(3):524–530, June 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901311.

Baker:1986:BIT

[BS86a] G. R. Baker and M. J. Shelley. Boundary integral techniques for multi-connected domains. Journal of Computational Physics, 64(1):112–132, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900215.

Bateman:1986:DDA

[BS86b] Glenn Bateman and Robin G. Storer. Direct determination of axisymmetric magnetohydrodynamic equilibria in Hamada coordinates. Journal of Computational Physics, 64(1):161–176, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900239.

Boyd:1989:UMN

[BS89a] I. D. Boyd and J. P. W. Stark. On the use of the modified Nanbu direct simulation scheme. Journal of Computational Physics, 80(2):374–386, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990106X.

Briscolini:1989:DMM

[BS89b] M. Briscolini and P. Santangelo. Development of the mask method for incompressible unsteady flows. Journal of Computational Physics, 84(1):57–75, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901812.

Bell:1982:FIS

[BSS82a] J. B. Bell, G. R. Shubin, and J. M. Solomon. Fully implicit shock tracking. Journal of Computational Physics, 48(2):223– 245, November 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900481.

Bell:1982:SAG

[BSS82b] J. B. Bell, G. R. Shubin, and A. B. Stephens. A segmentation approach to grid generation using biharmonics. Journal of Computational Physics, 47(3):463–472, September 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918290095X.

Blom:1988:SMG

[BSSV88] J. G. Blom, J. M. Sanz-Serna, and J. G. Verwer. On simple moving grid methods for one-dimensional evolutionary partial differential equations. Journal of Computational Physics, 74(1):191– 213, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900769.

Bogy:1980:FDS

[BST80] D. B. Bogy, S. J. Shine, and F. E. Talke. Finite difference solution of the Cosserat fluid jet equations. Journal of Computational Physics, 38(3):294–326, December 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901515.

Boffi:1985:SNI

[BST85] V. C. Boffi, G. Spiga, and J. R. Thomas, Jr. Solution of a nonlinear integral equation arising in particle transport theory. Journal of Computational Physics, 59(1):96–107, May 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901093.

Bell:1986:MRN

[BST86] John B. Bell, Gregory R. Shubin, and John A. Trangenstein. A method for reducing numerical dispersion in two-phase black-oil reservoir simulation. Journal of Computational Physics, 65(1): 71–106, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900057.

Bayliss:1982:FFB

[BT82] Alvin Bayliss and Eli Turkel. Far field boundary conditions for compressible flows. Journal of Computational Physics, 48(2):182– 199, November 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900468.

Buell:1988:MFD

[Bue88] Jeffrey C. Buell. A mixed finite difference/Galerkin method for three-dimensional Rayleigh–Bénard Convection. Journal of Computational Physics, 75(1):54–72, March 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900988.

Buttke:1988:NSS

[But88] Thomas F. Buttke. A numerical study of superfluid turbulence in the self-induction approximation. Journal of Computational Physics, 76(2):301–326, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901453.

Botta:1982:LRM

[BV82] Eugen F. F. Botta and Arthur E. P. Veldman. On local relaxation methods and their application to convection-diffusion equations. Journal of Computational Physics, 48(1):127–149, October 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900390.

Bridges:1986:LED

[BV86] Thomas J. Bridges and Leonid N. Vaserstfin. The local evaluation of the derivative of a determinant. Journal of Computational Physics, 65(1):107–119, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900069.

Blue:1981:CEE

[BW81] James L. Blue and Charles L. Wilson. Calculating eigenvalues and eigenfunctions using an interior constraint. Journal of Computational Physics, 44(1):70–83, November 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900383.

Baines:1988:MFE

[BW88a] M. J. Baines and A. J. Wathen. Moving finite element methods for evolutionary problems. I. Theory. Journal of Computational Physics, 79(2):245–269, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900162.

Brio:1988:UDS

[BW88b] M. Brio and C. C. Wu. An upwind differencing scheme for the equations of ideal magnetohydrodynamics. Journal of Computational Physics, 75(2):400–422, April 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901209.

Beam:1982:SAN

[BWY82] R. M. Beam, R. F. Warming, and H. C. Yee. Stability analysis of numerical boundary conditions and implicit difference approximations for hyperbolic equations. Journal of Computational Physics, 48(2):200–222, November 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918290047X.

Bogomolov:1985:NSN

[BY85] Ya. L. Bogomolov and A. D. Yunakovsky. Numerical simulation of nonstationary processes in free electron lasers. Journal of Computational Physics, 58(1):80–95, March 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901585.

Chang:1985:NASa

[CA85a] Sin-Chung Chang and John J. Adamczyk. A new approach for solving the three-dimensional steady Euler equations. I. General theory. Journal of Computational Physics, 60(1):23–40, August 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900154.

Chang:1985:NASb

[CA85b] Sin-Chung Chang and John J. Adamczyk. A new approach for solving the three-dimensional steady Euler equations. II. Application to secondary flows in a turning channel. Journal of Computational Physics, 60(1):41–61, August 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900166.

Candel:1981:SCF

[Can81] Sébastien M. Candel. Simultaneous calculation of Fourier– Bessel transforms up to order N. Journal of Computational Physics, 44(2):243–261, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900504.

Carver:1980:PCM

[Car80] Michael B. Carver. Pseudo characteristic method of lines solution of the conservation equations. Journal of Computational Physics, 35(1):57–76, March 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900340.

Caramana:1989:CFD

[Car89] E. J. Caramana. Consistent finite difference operators that satisfy ∇ · ∇x = 0. Journal of Computational Physics, 82(2):492– 498, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189900600.

Catherall:1988:OAF

[Cat88] D. Catherall. The optimisation of approximate-factorisation schemes for solving elliptic partial differential equations in three dimensions, featuring a new two-factor scheme. Journal of Computational Physics, 78(1):138–159, September 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900411.

Chin:1980:NSC

[CB80] R. C. Y. Chin and R. L. Braun. Numerical solutions of chemically reacting flows in porous media. Journal of Computational Physics, 34(1):74–93, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901138.

Chang:1984:NCS

[CB84] Chiechun J. Chang and Robert A. Brown. Natural convection in steady solidification: Finite element analysis of a twophase Rayleigh–Bénard problem. Journal of Computational Physics, 53(1):1–27, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184900494.

Campbell:1989:CMR

[CB89] Edwin S. Campbell and David Belford. The construction of the Maxwell representation for a cylindrically symmetric spherical harmonic of arbitrary order. Journal of Computational Physics, 84(2):403–413, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902404.

Cohen:1980:OAP

[CBCF80] Bruce I. Cohen, Thomas A. Brengle, Davis B. Conley, and Robert P. Freis. An orbit averaged particle code. Journal of Computational Physics, 38(1):45–63, November 1, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918090011X.

Cranfill:1986:TIM

[CBG86] C. W. Cranfill, J. U. Brackbill, and S. R. Goldman. A time-implicit Monte Carlo collision algorithm for particlein-cell electron transport models. Journal of Computational Physics, 66(1):239–249, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918690063X.

Couet:1981:STD

[CBL81] Benoit Couët, Oscar Buneman, and Anthony Leonard. Simulation of three-dimensional incompressible flows with a vortexin-cell method. Journal of Computational Physics, 39(2):305– 328, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901546.

Candel:1986:DAF

[CBLL86] Sebastien M. Candel, Eric Boussarie, Jean-Marc Loesch, and Annie Lelarge. A dual algorithm for fast calculation of the H_{0}^{1}- transform. Journal of Computational Physics, 65(2):344–369, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186902123.

Chen:1984:FAN

[CC84] Ching-Jen Chen and Hamn-Ching Chen. Finite analytic numerical method for unsteady two-dimensional Navier–Stokes equations. Journal of Computational Physics, 53(2):209–226, February 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918490038X.

Cheng:1987:NNC

[CC87] C. Z. Cheng and M. S. Chance. NOVA: A nonvariational code for solving the MHD stability of axisymmetric toroidal plasmas. Journal of Computational Physics, 71(1):124–146, July 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900234.

Chance:1981:NSH

[CCCS81] E. M. Chance, G. M. Clore, A. R. Curtis, and E. P. Shephard. Numerical solution of the Hodgkin–Huxley equations in a moving coordinate system. Simulation of nerve impulse transmission over long distances. Journal of Computational Physics, 40(2):318– 326, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902138.

Crystal:1980:CST

[CD80] T. L. Crystal and J. Denavit. Computer simulation of trapped-electron modes in tokamaks. Journal of Computational Physics, 37(2):151–170, September 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900182.

Carpentier:1982:SEI

[CD82] M. P. Carpentier and A. F. Dos Santos. Solution of equations involving analytic functions. Journal of Computational Physics, 45(2):210–220, February 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901176.

Cook:1981:PCM

[CDH81] Thomas L. Cook, Ruth B. Demuth, and Francis H. Harlow. PIC calculations of multiphase flow. Journal of Computational Physics, 41(1):51–67, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900796.

Caron:1980:NSS

[CDP80] A. Caron, M. Dupuy, and Ch. Pichot. Numerical solution of singular integral equations. Application to the flanged plane waveguide. Journal of Computational Physics, 38(3):357–377, December 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180901540.

Cupini:1983:MCS

[CDS83] E. Cupini, A. De Matteis, and R. Simonini. Monte Carlo simulation of the charge-transfer reaction in a plasma. Journal of Computational Physics, 52(1):122–129, October 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900190.

Cupini:1985:AER

[CDS85] E. Cupini, A. De Matteis, and R. Simonini. Approximation error in the rate coefficients of neutral-ion interactions. Journal of Computational Physics, 57(1):77–89, January 15, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900531.

Ceperley:1983:SQS

[Cep83] D. Ceperley. The simulation of quantum systems with random walks: A new algorithm for charged systems. Journal of Computational Physics, 51(3):404–422, September 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901614.

Cohen:1982:SAO

[CF82] Bruce I. Cohen and Robert P. Freis. Stability and application of an orbit-averaged magneto-inductive particle code. Journal of Computational Physics, 45(3):367–373, March 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901097.

Cox:1983:LSF

[CFG83] C. L. Cox, G. J. Fix, and M. D. Gunzburger. A least squares finite element scheme for transonic flow around harmonically oscillating airfoils. Journal of Computational Physics, 51(3):387–403, September 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901602.

Cryer:1982:SLC

[CFH^{+}82] C. W. Cryer, P. M. Flanders, D. J. Hunt, S. F. Reddaway, and J. Stansbury. The solution of linear complementarity problems on an array processor. Journal of Computational Physics, 47(2):258– 280, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918290078X.

Coyle:1986:SME

[CFL86] J. Michael Coyle, Joseph E. Flaherty, and Raymond Ludwig. On the stability of mesh equidistribution strategies for time-dependent partial differential equations. Journal of Computational Physics, 62(1):26–39, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900987.

Cain:1984:DOF

[CFR84] A. B. Cain, J. H. Ferziger, and W. C. Reynolds. Discrete orthogonal function expansions for non-uniform grids using the Fast Fourier Transform. Journal of Computational Physics, 56(2):272– 286, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900962.

Cohen:1982:OAI

[CFT82] Bruce I. Cohen, Robert P. Freis, and Vincent Thomas. Orbit-averaged implicit particle codes. Journal of Computational Physics, 45(3):345–366, March 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901085.

Colella:1985:ESA

[CG85] Phillip Colella and Harland M. Glaz. Efficient solution algorithms for the Riemann problem for real gases. Journal of Computational Physics, 59(2):264–289, June 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901469.

Christie:1989:NSN

[CG89] Ian Christie and G. H. Ganser. A numerical study of nonlinear waves arising in a one-dimensional model of a fluidized bed. Journal of Computational Physics, 81(2):300–318, April 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902106.

Chin:1984:P

[CGH84] R. C. Y. Chin, Freeman Gilbert, and Gerald Hedstrom. Preface. Journal of Computational Physics, 54(1):1–2, April 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901359.

Chern:1986:FTG

[CGM^{+}86] I.-L Chern, J. Glimm, O. Mcbryan, B. Plohr, and S. Yaniv. Front tracking for gas dynamics. Journal of Computational Physics, 62(1):83–110, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186901014.

Choi:1984:APT

[CH84] Youngki Choi and Joseph A. C. Humphrey. Analytical prediction of two-dimensional potential flow due to fixed vortices in a rectangular domain. Journal of Computational Physics, 56(1):15– 27, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900809.

Compagner:1987:MLS

[CH87] A. Compagner and A. Hoogland. Maximum-length sequences, cellular automata, and random numbers. Journal of Computational Physics, 71(2):391–428, August 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900374.

Cloot:1988:GRF

[CH88] A. Cloot and B. M. Herbst. Grid resonances, focusing and Benjamin–Feir instabilities in leapfrog time discretizations. Journal of Computational Physics, 75(1):31–53, March 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900976.

Chapman:1981:FND

[Cha81] Milt Chapman. FRAM-nonlinear damping algorithms for the continuity equation. Journal of Computational Physics, 44(1):84– 103, November 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181900395.

Chang:1986:SEP

[Cha86] Sin-Chung Chang. Solution of elliptic PDEs by fast Poisson solvers using a local relaxation factor. Journal of Computational Physics, 67(1):91–123, November 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901178.

Chang:1988:RVM

[Cha88] Chien-Cheng Chang. Random vortex methods for the Navier– Stokes equations. Journal of Computational Physics, 76(2):281– 300, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901441.

Chernov:1984:AWN

[Che84] S. V. Chernov. Analytical wavefunction normalization procedure in KKR-method. Journal of Computational Physics, 53(1):173– 180, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900585.

Charlton:1986:NCU

[CHH^{+}86] L. A. Charlton, J. A. Holmes, H. R. Hicks, V. E. Lynch, and B. A. Carreras. Numerical calculations using the full MHD equations in toroidal geometry. Journal of Computational Physics, 63(1):107– 129, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900860.

Chiang:1988:TAE

[Chi88] Yi-Ling F. Chiang. Truncation and accumulated errors in wave propagation. Journal of Computational Physics, 79(2):353–372, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900216.

Chorin:1980:FAP

[Cho80] Alexandre Joel Chorin. Flame advection and propagation algorithms. Journal of Computational Physics, 35(1):1–11, March 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900303.

Chorin:1982:NEH

[Cho82] Alexandre Joel Chorin. Numerical estimates of Hausdorff dimension. Journal of Computational Physics, 46(3):390–396, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900225.

Chorin:1985:CS

[Cho85] Alexandre Joel Chorin. Curvature and solidification. Journal of Computational Physics, 57(3):472–490, February 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901913.

Christie:1985:UCF

[Chr85] I. Christie. Upwind compact finite difference schemes. Journal of Computational Physics, 59(3):353–368, July 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901160.

Christiansen:1987:IAD

[Chr87] J. P. Christiansen. Integrated analysis of data from JET. Journal of Computational Physics, 73(1):85–106, November 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901070.

Christodoulides:1988:RAE

[Chr88] C. Christodoulides. A rational approximation for the exponential appearing in the theory of thermally stimulated processes. Journal of Computational Physics, 79(2):469–472, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900265.

Choi:1988:RVS

[CHS88] Y. Choi, J. A. C. Humphrey, and F. S. Sherman. Random-vortex simulation of transient wall-driven flow in a rectangular enclosure. Journal of Computational Physics, 75(2):359–383, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901180.

Chin:1984:NMS

[CHT84] R. C. Y. Chin, Gerald Hedstrom, and Lewis Thigpen. Numerical methods in seismology. Journal of Computational Physics, 54(1): 18–56, April 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901372.

Chang:1986:EMS

[CJ86] Shih-Hung Chang and Gary M. Johnson. An embedding method for the steady Euler equations. Journal of Computational Physics, 63(1):191–200, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900914.

Cliffs:1985:NCV

[CJW85] K. A. Cliffs, C. P. Jackson, and K. H. Winters. Natural convection at very high Rayleigh numbers. Journal of Computational Physics, 60(1):155–160, August 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900221.

Chattaraj:1987:SAF

[CKD87] P. K. Chattaraj, S. Rao Koneru, and B. M. Deb. Stability analysis of finite difference schemes for quantum mechanical equations of motion. Journal of Computational Physics, 72(2):504– 512, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900982.

Chalmers:1987:ACQ

[CKR87] C. B. Chalmers, R. D. Kenway, and D. Roweth. Algorithms for calculating quark propagators on large lattices. Journal of Computational Physics, 70(2):500–520, June 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901926.

Clifton:1980:VPM

[CL80] Joseph M. Clifton and Robert A. Leacock. Variable phase method for the calculation of the scattering phase shift: The bound-state region. Journal of Computational Physics, 38(3):327–334, December 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180901527.

Chen:1981:NMS

[CL81a] Y. M. Chen and David T. S. Lee. Numerical methods for solving differential equations with inadequate data. Journal of Computational Physics, 42(2):238–256, August 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902424.

Chen:1981:NAR

[CL81b] Y. M. Chen and J. Q. Liu. A numerical algorithm for remote sensing of thermal conductivity. Journal of Computational Physics, 43(2):315–326, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190125X.

Chen:1983:NAS

[CL83] Y. M. Chen and J. Q. Liu. A numerical algorithm for solving inverse problems of two-dimensional wave equations. Journal of Computational Physics, 50(2):193–208, May 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900633.

Chen:1984:INA

[CL84] Y. M. Chen and J. Q. Liu. An iterative numerical algorithm for solving multi-parameter inverse problems of evolutional partial differential equations. Journal of Computational Physics, 53(3):429–442, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918490069X.

Calder:1986:MWB

[CL86a] A. C. Calder and J. G. Laframboise. Multiple-water-bag simulation of inhomogeneous plasma motion near an electrode. Journal of Computational Physics, 65(1):18–45, July 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900021.

Cheng:1986:EMT

[CL86b] S. I. Cheng and Yulin Lu. An Eulerian method for transient nonlinear free surface wave problems. Journal of Computational Physics, 62(2):429–440, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901385.

Cliffs:1986:CFE

[CL86c] K. A. Cliffs and D. A. Lever. A comparison of finite-element methods for solving flow past a sphere. Journal of Computational Physics, 62(2):321–330, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901300.

Chen:1987:EIG

[CL87] Y. M. Chen and M. S. Liu. Efficiency improvement of GPST inversion algorithm. Journal of Computational Physics, 72(2):372– 382, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918790088X.

Clark:1987:CMR

[Cla87] Bradley A. Clark. Computing multigroup radiation integrals using polylogarithm-based methods. Journal of Computational Physics, 70(2):311–329, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901859.

Cohen:1982:ITI

[CLF82] Bruce I. Cohen, A. Bruce Langdon, and A. Friedman. Implicit time integration for plasma simulation. Journal of Computational Physics, 46(1):15–38, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918290002X.

Cohen:1984:SSG

[CLF84] Bruce J. Cohen, A. Bruce Langdon, and Alex Friedman. Smoothing and spatial grid effects in implicit particle simulation. Journal of Computational Physics, 56(1):51–64, October 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900834.

Cohen:1989:POD

[CLHP89] Bruce I. Cohen, A. Bruce Langdon, Dennis W. Hewett, and Richard J. Procassini. Performance and optimization of direct implicit particle simulation. Journal of Computational Physics, 81(1):151–168, March 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900685.

Carreras:1989:FWC

[CLJ^{+}89] B. A. Carreras, V. E. Lynch, E. F. Jaeger, D. B. Batchelor, and J. S. Tolliver. Full-wave calculations in flux coordinates for toroidal geometry. Journal of Computational Physics, 85(2):502, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901666.

Chiccoli:1988:EGE

[CLM88] C. Chiccoli, S. Lorenzutta, and G. Maino. On the evaluation of generalized exponential integrals E_{ν}(x). Journal of Computational Physics, 78(2):278–287, October 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900502.

Cloutman:1987:CFL

[Clo87] Lawrence D. Cloutman. A convective flux limiter for non-Lagrangian computational fluid dynamics. Journal of Computational Physics, 73(2):349–363, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901410.

Cockburn:1989:TRK

[CLS89] Bernardo Cockburn, San-Yih Lin, and Chi-Wang Shu. TVB Runge–Kutta local projection discontinuous Galerkin finite element method for conservation laws III: One-dimensional systems. Journal of Computational Physics, 84(1):90–113, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901836.

Cullen:1980:AEE

[CM80] M. J. P. Cullen and K. W. Morton. Analysis of evolutionary error in finite element and other methods. Journal of Computational Physics, 34(2):245–267, February 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901084.

Campbell:1982:TAP

[CM82] Edwin S. Campbell and Mihaly Mezei. Torque algorithms: The permanent multipole and induced dipole vector contributions in a set of charge distributions. Journal of Computational Physics, 47(2):245–257, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900778.

Celmaster:1986:MVR

[CM86] William Celmaster and K. J. M. Moriarty. A method for vectorized random number generators. Journal of Computational Physics, 64(1):271–275, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918690032X.

Chang:1989:RBF

[CM89] Chad-Lyan Chang and Charles L. Merkle. The relation between flux vector splitting and parabolized schemes. Journal of Computational Physics, 80(2):344–361, February 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901046.

Camerlengo:1980:OBC

[CO80] Alejandro L. Camerlengo and James J. O’Brien. Open boundary conditions in rotating fluids. Journal of Computational Physics, 35(1):12–35, March 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900315.

Cole:1981:TMA

[Col81a] Lee A. Cole. On “Theory and method for accelerating the convergence of self-consistent electronic structure calculations”. Journal of Computational Physics, 41(1):200–202, May 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900887. See [Fer80].

Cole:1981:TSR

[Col81b] R. J. Cole. Two series representations of the integral ∫_{0}^{∞} exp[−s(ψ + y cos ψ − z sin ψ)] dψ. Journal of Computational Physics, 44(2):388–396, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900590.

Cooke:1986:OSU

[Coo86] C. H. Cooke. On operator splitting for unsteady boundary value problems. Journal of Computational Physics, 67(2):472–478, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918690272X.

Cornille:1982:PSN

[Cor82] P. Cornille. A pseudospectral scheme for the numerical calculation of shocks. Journal of Computational Physics, 47(1):146–159, July 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900699.

Coughran:1985:NBC

[Cou85] W. M. Coughran, Jr. On noncharacteristic boundary conditions for discrete hyperbolic initial-boundary-value problems. Journal of Computational Physics, 60(1):135–154, August 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918590021X.

Concus:1987:C

[CP87] P. Concus and W. Proskurowski. Corrigendum. Journal of Computational Physics, 73(2):490, December 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901513.

Constable:1988:SSS

[CP88a] C. G. Constable and R. L. Parker. Smoothing, splines and smoothing splines; their application in geomagnetism. Journal of Computational Physics, 78(2):493–508, October 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900629.

Csepes:1988:ERA

[CP88b] Z. Csépes and J. Pipek. An effective recursive algorithm for generating many-body Hugenholtz and Goldstone diagrams. Journal of Computational Physics, 77(1):1–17, July 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901532.

Chen:1989:SRA

[CPJ89] H. C. Chen, V. C. Patel, and S. Ju. Solutions of Reynoldsaveraged Navier–Stokes equations for three-dimensional incompressible flows. Journal of Computational Physics, 85(2):502– 503, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901678.

Canuto:1985:PMR

[CQ85] Claudio Canuto and Alfio Quarteroni. Preconditioned minimal residual methods for Chebyshev spectral calculations. Journal of Computational Physics, 60(2):315–337, September 15, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900105.

Canuto:1987:BTS

[CQ87] Claudio Canuto and Alfio Quarteroni. On the boundary treatment in spectral methods for hyperbolic systems. Journal of Computational Physics, 71(1):100–110, July 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900210.

Chin:1989:EEC

[CQ89] Yuan-Shun Chin and Chaoyu Qint. Explicit energy-conserving schemes for the three-body problem. Journal of Computational Physics, 83(2):485–493, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901320.

Cushman-Roisin:1984:ALS

[CR84] Benoit Cushman-Roisin. Analytical, linear stability criteria for the leap-frog, Dufort–Frankel method. Journal of Computational Physics, 53(2):227–239, February 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900391.

Colombo:1988:VDN

[CRS88] V. Colombo, P. Ravetto, and M. Sumini. Variational determination of the neutron integral transport equation eigenvalues using space asymptotic trial functions. Journal of Computational Physics, 77(2):318–329, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901714.

Chodura:1981:CME

[CS81a] R. Chodura and A. Schlüter. A 3D code for MHD equilibrium and stability. Journal of Computational Physics, 41(1):68–88, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900802.

Coope:1981:SEL

[CS81b] J. A. R. Coope and D. W. Sabo. Several eigenvectors of a large matrix. II. The generalized eigenvalue problem. Journal of Computational Physics, 44(1):20–30, November 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900358.

Chavent:1982:FEM

[CS82] G. Chavent and G. Salzano. A finite-element method for the 1-D water flooding problem with gravity. Journal of Computational Physics, 45(3):307–344, March 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901073.

Civan:1984:STF

[CS84] Faruk Civan and C. M. Sliepcevich. On the solution of the Thomas–Fermi equation by differential quadrature. Journal of Computational Physics, 56(2):343–348, November 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900998.

Caldwell:1985:NVI

[CSW85] J. Caldwell, R. Saunders, and P. Wanless. A note on variationaliterative schemes applied to Burgers’ equation. Journal of Computational Physics, 58(2):275–281, April 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901810.

Cullen:1982:UQF

[Cul82] M. J. P. Cullen. The use of quadratic finite element methods and irregular grids in the solution of hyperbolic problems. Journal of Computational Physics, 45(2):221–245, February 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901188.

Cullen:1983:ASL

[Cul83a] M. J. P. Cullen. Analysis of some low-order finite element schemes for the Navier–Stokes equations. Journal of Computational Physics, 51(2):273–290, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900931.

Cullen:1983:CPP

[Cul83b] M. J. P. Cullen. Current progress and prospects in numerical techniques for weather prediction models. Journal of Computational Physics, 50(1):1–37, April 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900402.

Cullen:1983:ESL

[Cul83c] M. J. P. Cullen. Experiments with some low-order finite element schemes for the Navier–Stokes equations. Journal of Computational Physics, 51(2):291–312, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900943.

Cullen:1989:IFD

[Cul89] M. J. P. Cullen. Implicit finite difference methods for modelling discontinuous atmospheric flows. Journal of Computational Physics, 81(2):319–348, April 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902118.

Curzon:1982:MS

[Cur82] A. E. Curzon. A method for solving J_{1}(x)Y_{1}(ρx)−J_{1}(ρx)Y_{1}(x) = 0. Journal of Computational Physics, 46(1):39–42, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900031.

Christiansen:1980:NML

[CW80] J. P. Christiansen and N. K. Winsor. A numerical model for laser targets. Journal of Computational Physics, 35(3):291–318, May 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900728.

Cullum:1981:CEV

[CW81] Jane Cullum and Ralph A. Willoughby. Computing eigenvalues of very large symmetric matrices—an implementation of a Lanczos algorithm with no reorthogonalization. Journal of Computational Physics, 44(2):329–358, December 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900565.

Chan:1982:ASM

[CW82] Kwing L. Chan and Charles L. Wolff. ADI on staggered mesh — a method for the calculation of compressible convection. Journal of Computational Physics, 47(1):109–129, July 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900675.

Cliffe:1984:NSC

[CW84a] K. A. Cliffe and K. H. Winters. A numerical study of the cusp catastrophe for Bénard convection in tilted cavities. Journal of Computational Physics, 54(3):531–534, June 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901323.

Colella:1984:PPM

[CW84b] Phillip Colella and Paul R. Woodward. The Piecewise Parabolic Method (PPM) for gas-dynamical simulations. Journal of Computational Physics, 54(1):174–201, April 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901438.

Castillo:1985:RRG

[CW85a] JoséE Castillo and Robert E. Wyatt. Recursive residue generation method for laser-molecule interaction: utilization of structured sparsity. Journal of Computational Physics, 59(1):120– 135, May 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901111.

Chapman:1985:NSA

[CW85b] M. Chapman and E. M. Waisman. A noise suppression algorithm for the numerical solution of Maxwell’s equations. Journal of Computational Physics, 58(1):44–58, March 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918590155X.

Cliffe:1986:USB

[CW86] K. A. Cliffe and K. H. Winters. The use of symmetry in bifurcation calculations and its application to the Bénard problem. Journal of Computational Physics, 67(2):310–326, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186902652.

Chang:1989:MAA

[CW89] Qianshun Chang and Goubin Wang. Multigrid and adaptive algorithm for solving the nonlinear Schrödinger equation. Journal of Computational Physics, 85(2):504, December 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901721.

Chen:1986:IMS

[CX86] Y. M. Chen and G. Q. Xie. An iterative method for simultaneous determination of bulk and shear moduli and density variations. Journal of Computational Physics, 62(1):143–163, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918690104X.

Chikhliwala:1985:AOM

[CY85] E. D. Chikhliwala and Y. C. Yortsos. Application of orthogonal mapping to some two-dimensional domains. Journal of Computational Physics, 57(3):391–402, February 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918590186X.

Delsanto:1986:UMC

[DA86] P. P. Delsanto and J. D. Alemar. Use of Monte Carlo techniques for complex angular momentum algebra calculations. Journal of Computational Physics, 64(2):510–514, June 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900471.

Danielopoulos:1980:ESC

[Dan80] Stylianos D. Danielopoulos. On evaluating strategies for the computation of DWBA integrals. Journal of Computational Physics, 36(2):271–279, July 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901886.

Daripa:1989:ACV

[Dar89] Prabir Daripa. On applications of a complex variable method in compressible flows. Journal of Computational Physics, 85(2): 503, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918990171X.

Davey:1980:NSD

[Dav80a] A. Davey. On the numerical solution of difficult boundary-value problems. Journal of Computational Physics, 35(1):36–47, March 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900327.

Davies:1980:PUD

[Dav80b] Huw C. Davies. A pseudo-upstream differencing scheme for advection. Journal of Computational Physics, 37(2):280–286, September 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180900285.

Davies:1981:IVA

[Dav81] Charles W. Davies. An initial value approach to the production of discrete orthogonal coordinates. Journal of Computational Physics, 39(1):164–178, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901431.

Davey:1983:AOM

[Dav83] A. Davey. An automatic orthonormalization method for solving stiff boundary-value problems. Journal of Computational Physics, 51(2):343–356, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900980.

Davis:1984:RBU

[Dav84] Stephen F. Davis. A rotationally biased upwind difference scheme for the Euler equations. Journal of Computational Physics, 56(1): 65–92, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900846.

Davies:1986:LZA

[Dav86] B. Davies. Locating the zeros of an analytic function. Journal of Computational Physics, 66(1):36–49, September 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900525.

Davies:1989:CPM

[Dav89] K. T. R. Davies. Complex-plane methods for evaluating integrals with highly oscillatory integrands. Journal of Computational Physics, 80(2):498–505, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901149.

Dukowicz:1989:GTG

[DCA89] John K. Dukowicz, Michael C. Cline, and Frank L. Addessio. A general topology Godunov method. Journal of Computational Physics, 82(1):29–63, May 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990034X.

Dorfi:1987:SAG

[DD87] E. A. Dorfi and L. O’C. Drury. Simple adaptive grids for 1D initial value problems. Journal of Computational Physics, 69(1):175– 195, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901616.

Demaret:1989:CPS

[DD89] P. Demaret and M. O. Deville. Chebyshev pseudospectral solution of the Stokes equations using finite element preconditioning. Journal of Computational Physics, 83(2):463–484, August 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901319.

Decyk:1984:EMC

[Dec84] Viktor K. Decyk. Energy and momentum conservation theorems for electrostatic simulations. Journal of Computational Physics, 56(3):461–468, December 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901074.

Denavit:1981:TFP

[Den81] J. Denavit. Time-filtering particle simulations with ω_{pe}Δt ≫ 1. Journal of Computational Physics, 42(2):337–366, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902497.

Dawson:1988:EAL

[DF88] Silvina Ponce Dawson and Constantino Ferro Fontán. Extension of the Ablowitz–Ladik method to the derivative nonlinear Schrödinger equation. Journal of Computational Physics, 76(1): 192–200, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901374.

Dose:1981:IAI

[DFG81] V. Dose, Th. Fauster, and H.-J. Gossmann. The inversion of autoconvolution integrals. Journal of Computational Physics, 41 (1):34–50, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181900784.

deForcrand:1985:OLQ

[dFLR85] Ph. de Forcrand, D. Lellouch, and C. Roiesnel. Optimizing a lattice QCD simulation program. Journal of Computational Physics, 59(2):324–330, June 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901494.

Delfour:1981:FDS

[DFP81] M. Delfour, M. Fortin, and G. Payr. Finite-difference solutions of a non-linear Schrödinger equation. Journal of Computational Physics, 44(2):277–288, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900528.

deFrutos:1989:SSS

[dFSS89] J. de Frutos and J. M. Sanz-Serna. Split-step spectral schemes for nonlinear Dirac systems. Journal of Computational Physics, 83(2):407–423, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901277.

Dewar:1988:HEM

[DG88] R. L. Dewar and H. J. Gardner. A harmonic expansion for the magnetic field of the helical solenoid. Journal of Computational Physics, 77(2):485–500, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901799.

Dongarra:1984:SSE

[DGKW84] J. J. Dongarra, J. R. Gabriel, D. D. Koelling, and J. H. Wilkinson. Solving the secular equation including spin orbit coupling for systems with inversion and time-reversal symmetry. Journal of Computational Physics, 54(2):278–288, May 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.netlib.org/utk/people/JackDongarra/PAPERS/
Inversion-and-Time-Reversal-Symmetry.pdf;   http://www.
sciencedirect.com/science/article/pii/0021999184901190.

Deconinck:1982:MMT

[DH82] Herman Deconinck and Charles Hirsch. A multigrid method for the transonic full potential equation discretized with finite elements on an arbitrary body fitted mesh. Journal of Computational Physics, 48(3):344–365, December 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900560.

Dunn:1984:NCO

[DH84] Mark H. Dunn and S. I. Hariharan. Numerical computations on one-dimensional inverse scattering problems. Journal of Computational Physics, 55(1):157–165, July 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900214.

Dennis:1989:CFD

[DH89] S. C. R. Dennis and J. D. Hudson. Compact h^{4} finite-difference approximations to operators of Navier–Stokes type. Journal of Computational Physics, 85(2):390–416, December 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901563.

Dick:1983:OME

[Dic83] E. Dick. An overrelaxation method for Euler equations in steady transonic flow. Journal of Computational Physics, 49(2):298– 309, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901286.

Dick:1988:FDS

[Dic88] E. Dick. A flux-difference splitting method for steady Euler equations. Journal of Computational Physics, 76(1):19–32, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901295.

Dierckx:1983:AED

[Die83] Paul Dierckx. An algorithm for experimental data deconvolution using spline functions. Journal of Computational Physics, 52(1):163–186, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900220.

Dilts:1985:CSH

[Dil85] Gary A. Dilts. Computation of spherical harmonic expansion coefficients via FFT’s. Journal of Computational Physics, 57(3):439–453, February 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185901895.

Ding:1986:IHB

[Din86] Hong-Qiang Ding. An improved heat bath method for vectorized processing. Journal of Computational Physics, 67(1):28–36, November 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901130.

DeLucia:1980:IMM

[DJT80] J. DeLucia, S. C. Jardin, and A. M. M. Todd. An iterative metric method for solving the inverse tokamak equilibrium problem. Journal of Computational Physics, 37(2):183–204, September 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900200.

Davis:1985:SSP

[DKD85] Michael Davis, Phiroze Kapadia, and John Dowden. Solution of a Stefan problem in the theory of laser welding by the method of lines. Journal of Computational Physics, 60(3):534–548, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918590035X.

Davis:1989:SPT

[DKWD89] Michael Davis, Phiroze Kapadia, Paul Whiting, and John Dowden. Solution of a problem in the theory of laser medicine by the method of lines. Journal of Computational Physics, 81(2): 406–420, April 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189902143.

Deville:1985:CPS

[DM85] M. Deville and E. Mund. Chebyshev pseudospectral solution of second-order elliptic equations with finite element preconditioning. Journal of Computational Physics, 60(3):517–533, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900348.

Demkowicz:1986:ACP

[DO86] L. Demkowicz and J. T. Oden. An adaptive characteristic Petrov– Galerkin finite element method for convection-dominated linear and nonlinear parabolic problems in one space variable. Journal of Computational Physics, 67(1):188–213, November 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918690121X.

Dvinsky:1986:NST

[DP86] A. S. Dvinsky and A. S. Popel. Numerical solution of twodimensional Stokes equations for flow with particles in a channel of arbitrary shape using boundary-conforming coordinates. Journal of Computational Physics, 67(1):73–90, November 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901166.

Dennis:1983:DSV

[DQ83] S. C. R. Dennis and L. Quartapelle. Direct solution of the vorticity-stream function ordinary differential equations by a Chebyshev approximation. Journal of Computational Physics, 52(3):448–463, December 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900025.

Dennis:1985:SAV

[DQ85] S. C. R. Dennis and L. Quartapelle. Spectral algorithms for vector elliptic equations in a spherical gap. Journal of Computational Physics, 61(2):218–241, November 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918590083X.

Donea:1987:ATD

[DQS87] J. Donea, L. Quartapelle, and V. Selmin. An analysis of time discretization in the finite element solution of hyperbolic problems. Journal of Computational Physics, 70(2):463–499, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187901914.

Delic:1985:SEE

[DR85] G. Delic and G. H. Rawitscher. Sturmian eigenvalue equations with a Chebyshev polynomial basis. Journal of Computational Physics, 57(2):188–209, January 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900427.

Dritschel:1988:CDS

[Dri88a] David G. Dritschel. Contour dynamics/surgery on the sphere. Journal of Computational Physics, 79(2):477–483, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900289.

Dritschel:1988:CST

[Dri88b] David G. Dritschel. Contour surgery: A topological reconnection scheme for extended integrations using contour dynamics. Journal of Computational Physics, 77(1):240–266, July 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901659.

Drury:1980:NSO

[Dru80] L. O’C Drury. Numerical solution of Orr–Sommerfeld-type equations. Journal of Computational Physics, 37(1):133–139, August 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918090008X.

Daripa:1986:EAG

[DS86a] Prabir K. Daripa and Lawrence Sirovich. Exact and approximate gas dynamics using the tangent gas. Journal of Computational Physics, 62(2):400–413, February 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901361.

Daripa:1986:IMS

[DS86b] Prabir K. Daripa and Lawrence Sirovich. An inverse method for subcritical flows. Journal of Computational Physics, 63(2):311– 328, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901968.

Degani:1986:CTS

[DS86c] David Degani and Lewis B. Schifft. Computation of turbulent supersonic flows around pointed bodies having crossflow separation. Journal of Computational Physics, 66(1):173–196, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900598.

Dreitlein:1989:TSM

[DS89] Joseph F. Dreitlein and George F. Sowers. Targetted stochastic matrix inversion. Journal of Computational Physics, 80(2):362– 373, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901058.

Dukowicz:1980:PFN

[Duk80] John K. Dukowicz. A particle-fluid numerical model for liquid sprays. Journal of Computational Physics, 35(2):229–253, April 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918090087X.

Dukowicz:1984:CRR

[Duk84a] John K. Dukowicz. Conservative rezoning (remapping) for general quadrilateral meshes. Journal of Computational Physics, 54(3): 411–424, June 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901256.

Dukowicz:1984:SAM

[Duk84b] John K. Dukowicz. A simplified adaptive mesh technique derived from the moving finite element method. Journal of Computational Physics, 56(2):324–342, November 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900986.

Dukowicz:1985:GNI

[Duk85] John K. Dukowicz. A general, non-iterative Riemann solver for Godunov’s method. Journal of Computational Physics, 61(1):119–137, October 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185900646.

Dukowicz:1988:EVC

[Duk88] John K. Dukowicz. Efficient volume computation for threedimensional hexahedral cells. Journal of Computational Physics, 74(2):493–496, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188900915.

Dunn:1981:IMC

[Dun81] William L. Dunn. Inverse Monte Carlo analysis. Journal of Computational Physics, 41(1):154–166, May 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900851.

Dunn:1985:BPA

[Dun85] William L. Dunn. The back-projection angular flux estimator. Journal of Computational Physics, 61(3):391–402, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900713.

Delic:1987:NSA

[DVW87] G. Delic, E. J. Janse Van Rensburg, and G. Welke. A non-self-adjoint general matrix eigenvalue problem. Journal of Computational Physics, 69(2):325–340, April 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901719.

Dendy:1981:NMC

[DW81] J. E. Dendy, Jr. and Burton Wendroff. A numerical method for the calculation of traveling wave solutions of a quench front problem. Journal of Computational Physics, 39(2):418–432, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181901613.

Dilber:1983:CSO

[DWD83] I. Dilber, J. M. Walsh, and J. Denavit. Convergence of stochastic orbit computations. Journal of Computational Physics, 52(3):545–568, December 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900074.

Dyer:1984:MMR

[Dye84] Charles C. Dyer. A modified Muller routine for finding the zeroes of a non-analytic complex function. Journal of Computational Physics, 53(3):530–534, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900767.

Dykhuizen:1987:SFD

[Dyk87] Ronald C. Dykhuizen. Stability of finite difference representations of partial differential equations — a two-step process. Journal of Computational Physics, 70(1):262–268, May 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900143.

Eastwood:1987:TDA

[Eas87] James W. Eastwood. The two-dimensional adiabatic relaxation method for MHD modelling. Journal of Computational Physics, 73(1):148–167, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901112.

Ermak:1980:NIL

[EB80] Donald L. Ermak and Helen Buckholz. Numerical integration of the Langevin equation: Monte Carlo simulation. Journal of Computational Physics, 35(2):169–182, April 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900844.

Eidelman:1981:NSN

[EB81] Shmuel Eidelman and Alexander Burcat. Numerical solution of a nonsteady blast wave propagation in two-phase (“separated flow”) reactive medium. Journal of Computational Physics, 39(2):456–472, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901649.

Ettouney:1983:FEM

[EB83] Hisham M. Ettouney and Robert A. Brown. Finite-element methods for steady solidification problems. Journal of Computational Physics, 49(1):118–150, January 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901171.

Eckert:1989:SDS

[Eck89] Michael Eckert. Solving the 1-, 2-, and 3-dimensional Schrödinger equation for multiminima potentials using the Numerov–Cooley method. an extrapolation formula for energy eigenvalues. Journal of Computational Physics, 82(1):147–160, May 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900399.

Editors:1982:EF

[Edi82] The Editors. Editors’ foreword. Journal of Computational Physics, 48(2):159, November 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900432.

Epele:1988:IMC

[EFCS88] L. N. Epele, H. Fanchiotti, C. A. García Canal, and S. J. Sciutto. Improvement of the Monte Carlo projector method for spin systems. Journal of Computational Physics, 78(1):79–93, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900381.

Eaton:1983:CCB

[EG83] Brian Eaton and Karl Gustafson. Calculation of critical branching points in two-parameter bifurcation problems. Journal of Computational Physics, 50(1):171–177, April 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900475.

Ecker:1986:SSN

[EG86] A. Ecker and D. H. E. Gross. A system of simultaneous non-linear equations in three-thousand variables. Journal of Computational Physics, 64(1):246–252, May 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900288.

Eastwood:1985:CSR

[EH85] J. W. Eastwood and K. I. Hopcraft. Comments on the stability of the RSF three-dimensional MHD plasma simulation algorithm. Journal of Computational Physics, 60(3):549–557, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900361.

Elowitz:1989:TMA

[EHD89] Mark Elowitz, Frank Hill, and Thomas L. Duvall, Jr. A test of a modified algorithm for computing spherical harmonic coefficients using an FFT. Journal of Computational Physics, 80(2):506– 511, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901150.

Evans:1981:ISV

[EHM81] G. A. Evans, J. Hyslop, and A. P. G. Morgan. Iterative solution of Volterra integral equations using Clenshaw–Curtis quadrature. Journal of Computational Physics, 40(1):64–76, March 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901996.

Ehrlich:1981:AHS

[Ehr81] Louis W. Ehrlich. An ad hoc SOR method. Journal of Computational Physics, 44(1):31–45, November 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190036X.

Eiseman:1982:CGP

[Eis82a] Peter R. Eiseman. Coordinate generation with precise controls over mesh properties. Journal of Computational Physics, 47(3):331–351, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900870.

Eiseman:1982:HLC

[Eis82b] Peter R. Eiseman. High level continuity for coordinate generation with precise controls. Journal of Computational Physics, 47(3):352–374, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900882.

Eggert:1987:SFC

[EJL87] Norman Eggert, Mary Jarratt, and John Lund. Sinc function computation of the eigenvalues of Sturm–Liouville problems. Journal of Computational Physics, 69(1):209–229, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918790163X.

Elsasser:1980:MMM

[EM80] Klaus Elsässer and Winfried Maasjost. Macroscopic and microscopic models for the relativistic beam-plasma instability: A comparison. Journal of Computational Physics, 36(1):71–92, June 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180901758.

Engquist:1981:NRB

[EM81] Bjorn Engquist and Andrew Majda. Numerical radiation boundary conditions for unsteady transonic flow. Journal of Computational Physics, 40(1):91–103, March 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902011.

Eleuterio:1982:ESA

[EM82] S. M. Eleuterio and R. V. Mendes. Equivalence and singularities: An application of computer algebra. Journal of Computational Physics, 48(1):150–156, October 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900407.

Emanuel:1981:MIC

[Ema81] George Emanuel. Multiple isotope cascade algorithm. Journal of Computational Physics, 44(2):318–328, December 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900553.

Eriksson:1985:CAA

[ER85] Lars E. Eriksson and Arthur Rizzi. Computer-aided analysis of the convergence to steady state of discrete approximations to the Euler equations. Journal of Computational Physics, 57(1):90–128, January 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900543.

Erber:1983:SRP

[ERDF83] T. Erber, T. M. Rynne, W. F. Darsow, and M. J. Frank. The simulation of random processes on digital computers: Unavoidable order. Journal of Computational Physics, 49(3):394–419, March 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901377.

Eydeland:1988:CMS

[ET88] Alexander Eydeland and Bruce Turkington. A computational method of solving free-boundary problems in vortex dynamics. Journal of Computational Physics, 78(1):194–214, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900447.

Eyre:1988:AMG

[EW88] D. Eyre and C. J. Wright. Adaptive mesh grading for finite element solutions of an integral equation in quantum scattering. Journal of Computational Physics, 74(1):73–84, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900691.

Eyre:1988:SCA

[EWR88] D. Eyre, C. J. Wright, and G. Reutert. Spline-collocation with adaptive mesh grading for solving the stochastic collection equation. Journal of Computational Physics, 78(2):288–304, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900514.

Eyre:1984:STB

[Eyr84] D. Eyre. Splines and a three-body separable expansion for scattering problems. Journal of Computational Physics, 56(1):149– 164, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900883.

Eyre:1987:STB

[Eyr87] D. Eyre. Solving three-body integral equations with blending functions. Journal of Computational Physics, 73(2):447–460, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918790146X.

Farnell:1980:SPE

[Far80] L. Farnell. Solution of Poisson equations on a nonuniform grid. Journal of Computational Physics, 35(3):408–425, May 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900789.

Foreman:1988:NSA

[FB88] M. G. G. Foreman and A. F. Bennett. A numerical stability analysis for the two-dimensional incompressible Euler equations. Journal of Computational Physics, 79(1):70–84, November 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900046.

Fleck:1984:RWP

[FC84] J. A. Fleck, Jr. and E. H. Canfield. A random walk procedure for improving the computational efficiency of the implicit Monte Carlo method for nonlinear radiation transport. Journal of Computational Physics, 54(3):508–523, June 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918490130X.

Friedman:1982:APM

[FD82] Morton B. Friedman and David R. Dellwo. Accelerated projection methods. Journal of Computational Physics, 45(1):108–126, January 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918290105X.

Friedman:1981:SIR

[FDS81] Alex Friedman, Jacques Denavit, and R. N. Sudan. Strong ion ring equilibria formed by injection and intrinsic stochasticity of orbits. Journal of Computational Physics, 44(1):104–133, November 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900401.

Ferreira:1980:TMA

[Fer80] Luiz G. Ferreira. Theory and method for accelerating the convergence of self-consistent electronic structure calculations. Journal of Computational Physics, 36(2):198–208, July 1, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901849. See comment [?].

Ferziger:1987:SIT

[Fer87] Joel H. Ferziger. Simulation of incompressible turbulent flows. Journal of Computational Physics, 69(1):1–48, March 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901549.

Ferm:1988:OBC

[Fer88] Lars Ferm. Open boundary conditions for stationary inviscid flow problems. Journal of Computational Physics, 78(1):94–113, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900393.

Fettis:1984:CWI

[Fet84] Henry F. Fettis. On the calculation of Wu’s integral. Journal of Computational Physics, 53(1):197–204, January 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900627.

Fusegi:1986:PFF

[FF86] Toru Fusegi and Bakhtier Farouk. Predictions of fluid flow and heat transfer problems by the vorticity-velocity formulation of the Navier–Stokes equations. Journal of Computational Physics, 65(1):227–243, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900136.

Fortin:1987:NST

[FFG87] A. Fortin, M. Fortin, and J. J. Gervais. A numerical simulation of the transition to turbulence in a two-dimensional flow. Journal of Computational Physics, 70(2):295–310, June 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901847.

Ferguson:1982:NNM

[FFM82] W. E. Ferguson, Jr., H. Flaschka, and D. W. McLaughlin. Nonlinear normal modes for the Toda chain. Journal of Computational Physics, 45(2):157–209, February 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901164.

Feit:1982:SSE

[FFS82] M. D. Feit, J. A. Fleck, Jr., and A. Steiger. Solution of the Schrödinger equation by a spectral method. Journal of Computational Physics, 47(3):412–433, September 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900912.

Flores:1981:GMA

[FH81] Jolen Flores and Maurice Holt. Glimm’s method applied to underwater explosions. Journal of Computational Physics, 44(2):377– 387, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181900589.

Foidl:1989:SGL

[FK89] Ch Foidl and P. Kasperkovitz. Systematic generation of linear graphs — check and extension of the list of Uhlenbeck and Ford. Journal of Computational Physics, 85(2):504, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901770.

Fletcher:1983:CFE

[Fle83] C. A. J. Fletcher. A comparison of finite element and finite difference solutions of the one- and two-dimensional Burgers’ equations. Journal of Computational Physics, 51(1):159–188, July 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900852.

Floryan:1985:CMB

[Flo85] J. M. Floryan. Conformal-mapping-based coordinate generation method for channel flows. Journal of Computational Physics, 58(2):229–245, April 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185901780.

Floryan:1986:CMB

[Flo86] J. M. Floryan. Conformal-mapping- based coordinate generation method for flows in periodic configurations. Journal of Computational Physics, 62(1):221–247, January 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901087.

Flagg:1980:INI

[FLP80] R. C. Flagg, C. D. Luning, and W. L. Perry. Implementation of new iterative techniques for solutions of Thomas– Fermi and Emden–Fowler equations. Journal of Computational Physics, 38(3):396–405, December 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901564.

Fernandez:1983:ACE

[FMC83] Francisco M. Fernández, Alejandro M. Mesón, and Eduardo A. Castro. Accurate calculation of the eigenvalues of the x^{2}+λx^{2}(1+ gx^{2}) potential. Journal of Computational Physics, 51(3):519–526, September 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901675.

Fairweather:1980:LAM

[FN80] G. Fairweather and I. M. Navon. A linear ADI method for the shallow-water equations. Journal of Computational Physics, 37(1):1–18, August 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180900017.

Fyfe:1988:STV

[FOF88] D. E. Fyfe, E. S. Oran, and M. J. Fritts. Surface tension and viscosity with Lagrangian hydrodynamics on a triangular mesh. Journal of Computational Physics, 76(2):349–384, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901477.

Fogelson:1984:MMN

[Fog84] Aaron L. Fogelson. A mathematical model and numerical method for studying platelet adhesion and aggregation during blood clotting. Journal of Computational Physics, 56(1):111–134, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918490086X.

Foreman:1983:AWE

[For83a] M. G. G. Foreman. An analysis of the “wave equation” model for finite element tidal computations. Journal of Computational Physics, 52(2):290–312, November 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900323.

Foreman:1983:ATS

[For83b] M. G. G. Foreman. An analysis of two-step time discretizations in the solution of the linearized shallow water equations. Journal of Computational Physics, 51(3):454–483, September 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918390164X.

Foreman:1984:TDD

[For84] M. G. G. Foreman. A two-dimensional dispersion analysis of selected methods for solving the linearized shallow water equations. Journal of Computational Physics, 56(2):287–323, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900974.

Fornberg:1985:SVF

[For85] Bengt Fornberg. Steady viscous flow past a circular cylinder up to Reynolds number 600. Journal of Computational Physics, 61(2):297–320, November 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900890.

Foreman:1986:AAB

[For86] M. G. G. Foreman. An accuracy analysis of boundary conditions for the forced shallow water equations. Journal of Computational Physics, 64(2):334–367, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900379.

Forbes:1989:ADF

[For89] Lawrence K. Forbes. An algorithm for 3-dimensional freesurface problems in hydrodynamics. Journal of Computational Physics, 82(2):330–347, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189900521.

Foulser:1984:MDT

[Fou84] R. W. S. Foulser. Mass diffusion through a moving interface using a moving finite element mesh. Journal of Computational Physics, 55(3):408–425, September 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900305.

Fettis:1982:MCO

[FP82] Henry E. Fettis and Robert L. Pexton. More on the calculation of oscillatory integrals. Journal of Computational Physics, 47(3):473–476, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900961.

Fauci:1988:CMA

[FP88a] Lisa J. Fauci and Charles S. Peskin. A computational model of aquatic animal locomotion. Journal of Computational Physics, 77(1):85–108, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188901581.

Fogelson:1988:FNM

[FP88b] Aaron L. Fogelson and Charles S. Peskin. A fast numerical method for solving the three-dimensional Stokes’ equations in the presence of suspended particles. Journal of Computational Physics, 79(1):50–69, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900034.

Friedman:1987:IRF

[FR87] M. Friedman and A. Rabinovitch. Improved regula falsi method for solving the Schrödinger equation with a piecewise constant potential. Journal of Computational Physics, 68(1):180–187, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900507.

Freeman:1981:TTM

[Fre81] John R. Freeman. TRIDIF, A triangular mesh diffusion code. Journal of Computational Physics, 41(1):142–153, May 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190084X.

Frenkel:1983:CESa

[Fre83a] A. Frenkel. A Chebyshev expansion of singular integral equations with a logarithmic kernel. Journal of Computational Physics, 51(2):326–334, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900967.

Frenkel:1983:CESb

[Fre83b] A. Frenkel. A Chebyshev expansion of singular integrodifferential equations with a ∂^{2} ln |s − t|/∂s∂t kernel. Journal of Computational Physics, 51(2):335–342, August 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900979.

Friedman:1987:TFE

[FRRT87] M. Friedman, A. Rabinovitch, Y. Rosenfeld, and R. Thieberger. Thomas–Fermi equation with non-spherical boundary conditions. Journal of Computational Physics, 70(2):284–294, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187901835.

Forsyth:1986:PCA

[FS86] P. A. Forsyth, Jr. and P. H. Sammon. Practical considerations for adaptive implicit methods in reservoir simulation. Journal of Computational Physics, 62(2):265–281, February 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901270.

Fong:1987:SIF

[FS87] Jefferson Fong and Lawrence Sirovich. Supersonic inviscid flow- a three-dimensional characteristics approach. Journal of Computational Physics, 68(2):378–392, February 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900635.

Fezoui:1989:CIU

[FS89] L. Fezoui and B. Stoufflet. A class of implicit upwind schemes for Euler simulations with unstructured meshes. Journal of Computational Physics, 84(1):174–206, September 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901873.

Friedman:1981:LHC

[FSD81] A. Friedman, R. N. Sudan, and J. Denavit. A linearized 3D hybrid code for stability studies of field-reversed ion rings. Journal of Computational Physics, 40(1):1–35, March 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901972.

Fulton:1984:GTT

[FT84] Scott R. Fulton and G. D. Taylor. On the Gottlieb–Turkel time filter for Chebyshev spectral methods. Journal of Computational Physics, 55(2):302–312, August 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900081.

Fuchs:1989:ASL

[Fuc89] Norman H. Fuchs. Approximate solutions for large transfer matrix problems. Journal of Computational Physics, 83(1):201–211, July 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189902283.

Frederiksen:1981:FEM

[FW81] C. S. Frederiksen and A. M. Watts. Finite-element method for time-dependent incompressible free surface flow. Journal of Computational Physics, 39(2):282–304, February 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901534.

Fyfe:1981:FES

[FWB^{+}81] D. Fyfe, A. Weiser, I. Bernstein, S. Eisenstat, and M. Schultz. A finite element solution of a reduced Fokker–Planck equation. Journal of Computational Physics, 42(2):327–336, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902485.

Fryxell:1986:IEH

[FWCW86] Bruce A. Fryxell, Paul R. Woodward, Phillip Colella, and Karl-Heinz Winkler. An implicit-explicit hybrid method for Lagrangian hydrodynamics. Journal of Computational Physics, 63(2):283–310, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186901956.

Friesner:1986:GVR

[FWHC86] Richard A. Friesner, Robert E. Wyatt, C. Hempel, and B. Criner. A generalized version of the recursive residue generation method for vector computers. Journal of Computational Physics, 64(1): 220–229, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900264.

Floryan:1987:SCM

[FZ87] J. M. Floryan and Charles Zemach. Schwarz–Christoffel mappings: A general approach. Journal of Computational Physics, 72(2):347–371, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187900878.

Floryan:1988:QRS

[FZ88] J. M. Floryan and Charles Zemach. Quadrature rules for singular integrals with application to Schwarz–Christoffel mappings. Journal of Computational Physics, 75(1):15–30, March 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900964.

Ganapol:1985:RTD

[Gan85] B. D. Ganapol. Reconstruction of the time-dependent monoenergetic neutron flux from moments. Journal of Computational Physics, 59(3):468–483, July 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185901238.

Gary:1978:BCH

[Gar78] John Gary. On boundary conditions for hyperbolic difference schemes. Journal of Computational Physics, 26(3):339– 351, March 1978. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0021999178900748. See comment

[Gus80a].

Garlick:1983:UDG

[Gar83] A. R. Garlick. The use of distorting grids and flux splitting to model axisymmetric adiabatic explosions. Journal of Computational Physics, 52(3):427–447, December 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900013.

Gartland:1986:HOF

[Gar86] Eugene C. Gartland, Jr. High order finite-difference methods for two-point boundary value problems with singular sources. Journal of Computational Physics, 64(2):473–489, June 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900458.

Gauthier:1988:SCM

[Gau88] Serge Gauthier. A spectral collocation method for twodimensional compressible convection. Journal of Computational Physics, 75(1):217–235, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188901088.

Ghoniem:1987:VSL

[GC87] Ahmed F. Ghoniem and Yves Cagnon. Vortex simulation of laminar recirculating flow. Journal of Computational Physics, 68(2):346–377, February 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900623.

Guo:1988:FPM

[GC88] Ben-Yu Guo and Wei-Ming Cao. The Fourier pseudospectral method with a restrain operator for the RLW equation. Journal of Computational Physics, 74(1):110–126, January 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900721.

Gent:1989:RGP

[GC89] Peter R. Gent and Mark A. Cane. A reduced gravity, primitive equation model of the upper equatorial ocean. Journal of Computational Physics, 81(2):444–480, April 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902167.

Gelinas:1981:MFE

[GDM81] R. J. Gelinas, S. K. Doss, and Keith Miller. The moving finite element method: Applications to general partial differential equations with multiple large gradients. Journal of Computational Physics, 40(1):202–249, March 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902072.

Grimm:1983:IMS

[GDM83] R. C. Grimm, R. L. Dewar, and J. Manickam. Ideal MHD stability calculations in axisymmetric toroidal coordinate systems. Journal of Computational Physics, 49(1):94–117, January 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918390116X.

Gardner:1988:FBE

[GDS88] H. J. Gardner, R. L. Dewar, and W. N-C Sy. The free-boundary equilibrium problem for helically symmetric plasmas. Journal of Computational Physics, 74(2):477–487, February 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900897.

Giannakopoulos:1988:DCG

[GE88] Antonios E. Giannakopoulos and Antonie J. Engel. Directional control in grid generation. Journal of Computational Physics, 74(2):422–439, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900861.

Greenbaum:1986:PGF

[GF86] A. Greenbaum and J. M. Ferguson. A Petrov–Galerkin finite element method for solving the neutron transport equation. Journal of Computational Physics, 64(1):97–111, May 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900203.

Guerra:1986:NMI

[GG86] Jaime Guerra and Bertil Gustafsson. A numerical method for incompressible and compressible flow problems with smooth solutions. Journal of Computational Physics, 63(2):377–397, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186902007.

Gottlieb:1988:ARS

[GG88] J. J. Gottlieb and C. P. T. Groth. Assessment of Riemann solvers for unsteady one-dimensional inviscid flows of perfect gases. Journal of Computational Physics, 78(2):437–458, October 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900599.

Gatski:1982:NST

[GGR82] T. B. Gatski, C. E. Grosch, and M. E. Rose. A numerical study of the two-dimensional Navier–Stokes equations in vorticity-velocity variables. Journal of Computational Physics, 48(1):1–22, October 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900328.

Gatski:1989:NSN

[GGR89] T. B. Gatski, C. E. Grosch, and M. E. Rose. The numerical solution of the Navier–Stokes equations for 3-dimensional, unsteady, incompressible flows by compact schemes. Journal of Computational Physics, 82(2):298–329, June 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990051X.

Ghia:1982:HRS

[GGS82] U. Ghia, K. N. Ghia, and C. T. Shin. High-Re solutions for incompressible flow using the Navier–Stokes equations and a multigrid method. Journal of Computational Physics, 48(3):387–411, December 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900584.

Gustafson:1986:VDC

[GH86] K. Gustafson and K. Halasi. Vortex dynamics of cavity flows. Journal of Computational Physics, 64(2):279–319, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900355.

Gustafson:1987:CFD

[GH87] K. Gustafson and K. Halasi. Cavity flow dynamics at higher Reynolds number and higher aspect ratio. Journal of Computational Physics, 70(2):271–283, June 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901823.

Garcia:1986:NMC

[GHC^{+}86] L. Garcia, H. R. Hicks, B. A. Carreras, L. A. Charlton, and J. A. Holmes. 3D nonlinear MHD calculations using implicit and explicit time integration schemes. Journal of Computational Physics, 65(2):253–272, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902081.

Ghoniem:1988:NST

[GHK88] Ahmed F. Ghoniem, Ghassem Heidarinejad, and Anantha Krishnan. Numerical simulation of a thermally stratified shear layer using the vortex element method. Journal of Computational Physics, 79(1):135–166, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900083.

Gilliland:1981:SSW

[Gil81] Ronald L. Gilliland. Solutions of the shallow water equations on a sphere. Journal of Computational Physics, 43(1):79–94, September 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901121.

Gruber:1987:CME

[GIT87] R. Gruber, R. Iacono, and F. Troyon. Computation of MHD equilibria by a quasi-inverse finite hybrid element approach. Journal of Computational Physics, 73(1):168–182, November 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901124.

Gary:1981:CSO

[GK81] John Gary and D. R. Kassoy. Computation of steady and oscillatory convection in saturated porous media. Journal of Computational Physics, 40(1):120–142, March 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902035.

Grimes:1987:SLD

[GKL^{+}87] Roger Grimes, Henry Krakauer, John Lewis, Horst Simon, and Su-Hai Wei. The solution of large dense generalized eigenvalue problems on the Cray X-MP/24 with SSD. Journal of Computational Physics, 69(2):471–481, April 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901781.

Gresho:1985:PVB

[GL85] P. M. Gresho and R. L. Lee. Partial vindication of the bilinear velocity, piecewise constant pressure element. Journal of Computational Physics, 60(1):161–164, August 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900233.

Gladd:1984:SMT

[Gla84a] N. T. Gladd. Symbolic manipulation techniques for plasma kinetic theory derivations. Journal of Computational Physics, 56(2):175–202, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184900913.

Glatzmaier:1984:NSS

[Gla84b] Gary A. Glatzmaier. Numerical simulations of stellar convective dynamos. I. the model and method. Journal of Computational Physics, 55(3):461–484, September 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900330.

Glaister:1988:ALRa

[Gla88a] P. Glaister. An approximate linearised Riemann solver for the Euler equations for real gases. Journal of Computational Physics, 74(2):382–408, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900848.

Glaister:1988:ALRb

[Gla88b] P. Glaister. An approximate linearised Riemann solver for the three-dimensional Euler equations for real gases using operator splitting. Journal of Computational Physics, 77(2):361–383, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918890174X.

Glasser:1989:MFE

[Gla89] Alan H. Glasser. A moving finite element model of the high density Z-pinch. Journal of Computational Physics, 85(1):159– 209, November 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189902040.

Gabutti:1981:NAD

[GM81] B. Gabutti and B. Minetti. A new application of the discrete Laguerre polynomials in the numerical evaluation of the Hankel transform of a strongly decreasing even function. Journal of Computational Physics, 42(2):277–287, August 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190245X.

Gingold:1982:KEB

[GM82] R. A. Gingold and J. J. Monaghan. Kernel estimates as a basis for general particle methods in hydrodynamics. Journal of Computational Physics, 46(3):429–453, June 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900250.

Goodman:1985:VME

[GM85] Jonathan Goodman and Andrew Majda. The validity of the modified equation for nonlinear shock waves. Journal of Computational Physics, 58(3):336–348, May 15, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901664.

Guerrero:1988:HOT

[GM88] Antonio L. Guerrero and Pablo Martin. Higher order two-point quasi-fractional approximations to the Bessel functions J_{0}(x) and J_{1}(x). Journal of Computational Physics, 77(1):276–281, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901684.

Glimm:1980:SRF

[GMM80] J. Glimm, D. Marchesin, and O. McBryan. Subgrid resolution of fluid discontinuities, II. Journal of Computational Physics, 37(3):336–354, October 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900418.

Glimm:1981:NMT

[GMM81] J. Glimm, D. Marchesin, and O. McBryan. A numerical method for two phase flow with an unstable interface. Journal of Computational Physics, 39(1):179–200, January 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901443.

Goldston:1981:NTC

[GMT^{+}81] R. J. Goldston, D. C. McCune, H. H. Towner, S. L. Davis, R. J. Hawryluk, and G. L. Schmidt. New techniques for calculating heat and particle source rates due to neutral beam injection in axisymmetric tokamaks. Journal of Computational Physics, 43(1):61–78, September 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190111X.

Goldman:1986:CPS

[Gol86a] Saul Goldman. The convergence properties and stochastic characteristics inherent in force-biased and in Metropolis Monte Carlo simulations on liquids. Journal of Computational Physics, 62(2):441–463, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901397.

Goldstein:1986:FEM

[Gol86b] C. I. Goldstein. Finite element methods applied to nearly oneway wave propagation. Journal of Computational Physics, 64(1): 56–81, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900185.

Gottlieb:1988:SNG

[Got88] James J. Gottlieb. Staggered and nonstaggered grids with variable node spacing and local time stepping for the random choice method. Journal of Computational Physics, 78(1):160– 177, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900423.

Grundy:1980:AHA

[GP80] R. E. Grundy and G. M. Phillips. Application of Hermite approximation to a boundary value problem. Journal of Computational Physics, 38(2):251–256, November 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918090056X.

Greenberg:1981:FAM

[GP81] J. B. Greenberg and C. Presser. A fully automatic method for predicting internal reacting flows. Journal of Computational Physics, 40(2):361–375, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902163.

Goedheer:1985:CFD

[GP85] W. J. Goedheer and J. H. H. M. Potters. A compact finite difference scheme on a non-equidistant mesh. Journal of Computational Physics, 61(2):269–279, November 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900865.

Goussis:1989:RIE

[GP89] Dimitrios A. Goussis and Arne J. Pearlstein. Removal of infinite eigenvalues in the generalized matrix eigenvalue problem. Journal of Computational Physics, 84(1):242–246, September 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901897.

Gambolati:1988:IIO

[GPS88] Giuseppe Gambolati, Giorgio Pini, and Flavio Sartoretto. An improved iterative optimization technique for the leftmost eigenpairs of large symmetric matrices. Journal of Computational Physics, 74(1):41–60, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900678.

Guardiola:1982:NISa

[GR82a] R. Guardiola and J. Ros. On the numerical integration of the Schrödinger equation in the finite-difference schemes. Journal of Computational Physics, 45(3):374–389, March 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901103.

Guardiola:1982:NISb

[GR82b] R. Guardiola and J. Ros. On the numerical integration of the Schrödinger equation: Numerical tests. Journal of Computational Physics, 45(3):390–402, March 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901115.

Giles:1984:NSS

[GR84] Michael B. Giles and Milton E. Rose. A numerical study of the steady scalar convective diffusion equation for small viscosity. Journal of Computational Physics, 56(3):513–529, December 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184901104.

Greengard:1987:FAP

[GR87] Leslie Greengard and Vladimir Rokhlin. A fast algorithm for particle simulations. Journal of Computational Physics, 73(2):325– 348, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901409. This paper is cred-

ited as the origin of the fast multipole method, with an O(N) algorithm. It was reprinted in the same journal, vol. 135, pp. 280–292, August 1997.

Grinstein:1983:MMA

[GRA83] F. F. Grinstein, H. Rabitz, and A. Askar. The multigrid method for accelerated solution of the discretized Schrödinger equation. Journal of Computational Physics, 51(3):423–443, September 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901626.

Greenspan:1984:CNM

[Gre84a] Donald Greenspan. Conservative numerical methods for ẍ = f(x). Journal of Computational Physics, 56(1):28–41, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900810.

Greenstadt:1984:UNF

[Gre84b] John Greenstadt. The use of nonrational functions to represent steep front solutions to partial differential equations. Journal of Computational Physics, 53(2):299–318, February 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900433.

Greengard:1985:CSV

[Gre85a] Claude Greengard. The core spreading vortex method approximates the wrong equation. Journal of Computational Physics, 61(2):345–348, November 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185900919.

Greywall:1985:SCT

[Gre85b] Mahesh S. Greywall. Streamwise computation of two-dimensional incompressible potential flows. Journal of Computational Physics, 59(2):224–231, June 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185901433.

Greywall:1988:SCT

[Gre88] Mahesh S. Greywall. Streamwise computation of threedimensional incompressible potential flows. Journal of Computational Physics, 78(1):178–193, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900435.

Grotzbach:1983:SRR

[Grö83] Günther Grötzbach. Spatial resolution requirements for direct numerical simulation of the Rayleigh–Bénard convection. Journal of Computational Physics, 49(2):241–264, February 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901250.

Gordon:1982:EVR

[GRP82] Heather L. Gordon, Stuart M. Rothstein, and Timothy R. Proctor. Efficient variance-reduction transformations for the simulation of a ratio of two means: application to quantum Monte Carlo simulations. Journal of Computational Physics, 47(3):375– 386, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900894.

Greenside:1989:CPN

[GRS89] H. S. Greenside, A. H. Reiman, and A. Salas. Convergence properties of a nonvariational 3D MHD equilibrium code. Journal of Computational Physics, 81(1):102–136, March 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900661.

Garcia:1982:MTT

[GS82] R. D. M. Garcia and C. E. Siewert. Multigroup transport theory with anisotropic scattering. Journal of Computational Physics, 46(2):237–270, May 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900158.

Garcia:1983:MMT

[GS83] R. D. M. Garcia and C. E. Siewert. Multislab multigroup transport theory with L th the order anisotropic scattering. Journal of Computational Physics, 50(2):181–192, May 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900621.

Ghoniem:1985:GFS

[GS85a] Ahmed F. Ghoniem and Frederick S. Sherman. Grid-free simulation of diffusion using random walk methods. Journal of Computational Physics, 61(1):1–37, October 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900580.

Grotendorst:1985:FTT

[GS85b] Johannes Grotendorst and E. Otto Steinborn. The Fourier transform of a two-center product of exponential-type functions and its efficient evaluation. Journal of Computational Physics, 61(2):195–217, November 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900828.

Giorla:1987:RWM

[GS87a] J. Giorla and R. Sentis. A random walk method for solving radiative transfer equations. Journal of Computational Physics, 70(1): 145–165, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900064.

Giovangigli:1987:CEL

[GS87b] V. Giovangigli and M. D. Smooke. Calculation of extinction limits for premixed laminar flames in a stagnation point flow. Journal of Computational Physics, 68(2):327–345, February 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900611.

Grimes:1988:NSL

[GS88] Roger G. Grimes and Horst D. Simon. New software for large dense symmetric generalized eigenvalue problems using secondary storage. Journal of Computational Physics, 77(1):270–275, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901672.

Goodrich:1989:TDV

[GS89] John W. Goodrich and W. Y. Sox. Time-dependent viscous incompressible Navier–Stokes equations: The finite difference Galerkin formulation and streamfunction algorithms. Journal of Computational Physics, 84(1):207–241, September 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901885.

Giles:1985:PSW

[GT85] M. B. Giles and W. T. Thompkins, Jr. Propagation and stability of wavelike solutions of finite difference equations with variable coefficients. Journal of Computational Physics, 58(3):349– 360, May 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901676.

Gardner:1989:MTS

[GTD89] David R. Gardner, Steven A. Trogdon, and Rod W. Douglass. A modified tau spectral method that eliminates spurious eigenvalues. Journal of Computational Physics, 80(1):137–167, January 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900934.

Gupta:1981:CNS

[Gup81] Murli M. Gupta. A comparison of numerical solutions of convective and divergence forms of the Navier–Stokes equations for the driven cavity problem. Journal of Computational Physics, 43(2):260–267, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901224.

Gupta:1984:FOP

[Gup84] Murli M. Gupta. A fourth-order Poisson solver. Journal of Computational Physics, 55(1):166–172, July 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900226.

Gustafsson:1980:CBC

[Gus80a] Bertil Gustafsson. Comment on “On boundary conditions for hyperbolic difference schemes”. Journal of Computational Physics, 34(1):108–110, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901151. See [Gar78].

Gustafsson:1980:NSH

[Gus80b] Bertil Gustafsson. Numerical solution of hyperbolic systems with different time scales using asymptotic expansions. Journal of Computational Physics, 36(2):209–235, July 1, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901850.

Gustafsson:1982:CNB

[Gus82] Bertil Gustafsson. The choice of numerical boundary conditions for hyperbolic systems. Journal of Computational Physics, 48(2):270–283, November 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918290050X.

Gustafsson:1980:FDM

[GW80] Bertil Gustafsson and Per Wahlund. Finite-difference methods for computing the steady flow about blunt bodies. Journal of Computational Physics, 36(3):327–346, July 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901643.

Glaz:1985:HOG

[GW85] Harland M. Glaz and Andrew B. Wardlaw. A high-order Godunov scheme for steady supersonic gas dynamics. Journal of Computational Physics, 58(2):157–187, April 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901755.

Guralnik:1985:FIM

[GWZ85a] Gerald Guralnik, Tony Warnock, and Charles Zemach. Factorization, invariant measure, and random selection of matrices in SU(n) and other groups. Journal of Computational Physics, 61(1):89–104, October 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185900622.

Guralnik:1985:RST

[GWZ85b] Gerald Guralnik, Tony Warnock, and Charles Zemach. Random selection from trace-biased invariant distributions of SU (3) matrices. Journal of Computational Physics, 61(1):105–118, October 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900634.

Guo:1989:SDM

[GX89] Ben-Yu Guo and Yeu-Shan Xiong. A spectral-difference method for two-dimensional viscous flow. Journal of Computational Physics, 84(2):259–278, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902337.

Harder:1985:SOZ

[HA85] J. M. Harder and A. R. Allnatt. Solution of Ornstein– Zernike integral equations for lattice gases. Journal of Computational Physics, 59(3):499–506, July 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901275.

Heron:1989:PCO

[HA89] A. Héron and J. C. Adam. Particle code optimization on vector computers. Journal of Computational Physics, 85(2):284–301, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901526.

Hagin:1980:CWC

[Hag80] Frank Hagin. On the construction of well-conditioned systems for Fredholm I problems by mesh adapting. Journal of Computational Physics, 36(2):154–169, July 1, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901813.

Hagin:1981:SNA

[Hag81] Frank Hagin. Some numerical approaches to solving onedimensional inverse problems. Journal of Computational Physics, 43(1):16–30, September 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181901091.

Hagstrom:1987:BCO

[Hag87] Thomas M. Hagstrom. Boundary conditions at outflow for a problem with transport and diffusion. Journal of Computational Physics, 69(1):69–80, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901562.

Haines:1985:GRF

[Hai85] E. M. Haines. Generation of random fields defined on a cluster of points when the random field has a given site-to-site correlation. Journal of Computational Physics, 60(3):353–375, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900269.

Hain:1987:PDC

[Hai87] Klaus H. Hain. The partial donor cell method. Journal of Computational Physics, 73(1):131–147, November 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901100.

Hald:1981:CFM

[Hal81] Ole H. Hald. Convergence of Fourier methods for Navier– Stokes equations. Journal of Computational Physics, 40(2):305– 317, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902126.

Hald:1987:AWI

[Hal87] Ole H. Hald. Approximation of Wiener integrals. Journal of Computational Physics, 69(2):460–470, April 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918790177X.

Hall:1988:IBC

[Hal88] M. S. Hall. The interaction between a compliant material and an unstable boundary layer flow. Journal of Computational Physics, 76(1):33–47, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901301.

Halton:1989:PRT

[Hal89] John H. Halton. Pseudo-random trees: Multiple independent sequence generators for parallel and branching computations. Journal of Computational Physics, 84(1):1–56, September 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901800.

Harada:1980:CSC

[Har80] I. Harada. Computation of strongly compressible rotating flows. Journal of Computational Physics, 38(3):335–356, December 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180901539.

Harding:1981:MCS

[Har81] M. P. Harding. Monte Carlo simulation of Ising systems using a data structure. Journal of Computational Physics, 44(2):227– 242, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181900498.

Harned:1982:QHS

[Har82] Douglas S. Harned. Quasineutral hybrid simulation of macroscopic plasma phenomena. Journal of Computational Physics, 47(3):452–462, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900948.

Harten:1983:HRS

[Har83a] Ami Harten. High resolution schemes for hyperbolic conservation laws. Journal of Computational Physics, 49(3):357–393, March 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901365.

Harten:1983:SFS

[Har83b] Amiram Harten. On the symmetric form of systems of conservation laws with entropy. Journal of Computational Physics, 49(1):151–164, January 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183901183.

Harten:1989:ESS

[Har89] Ami Harten. ENO schemes with subcell resolution. Journal of Computational Physics, 83(1):148–184, July 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990226X.

Hautot:1981:IR

[Hau81] André Hautot. Interaction λx^{2}/(1 + gx^{2}) revisited. Journal of Computational Physics, 39(1):72–93, January 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901376.

Hautot:1982:AGP

[Hau82] André Hautot. Application of generalized Padé approximants to the special function evaluation problem. Journal of Computational Physics, 47(3):477–488, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900973.

Hewett:1987:SDS

[HB87] Dennis W. Hewett and John K. Boyd. Streamlined Darwin simulation of nonneutral plasmas. Journal of Computational Physics, 70(1):166–181, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187900076.

Head:1982:APS

[HBR82] John D. Head, George Blyholder, and Fernando Ruette. An automatic procedure for the symmetry blocking of semi-empirical Hamiltonians. Journal of Computational Physics, 45(2):255–265, February 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182901206.

Haussling:1981:MGO

[HC81a] H. J. Haussling and R. M. Coleman. A method for generation of orthogonal and nearly orthogonal boundary-fitted coordinate systems. Journal of Computational Physics, 43(2):373–381, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181901297.

Huang:1981:HOA

[HC81b] Chi-Hua Huang and John H. Cushman. High order accurate, explicit, difference formulas for the classical wave equation. Journal of Computational Physics, 40(2):376–395, April 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902175.

Hasan:1987:TDF

[HC87] Mohammad Z. Hasan and Robert W. Conn. A two-dimensional finite element multigroup diffusion theory for neutral atom transport in plasmas. Journal of Computational Physics, 71(2):371– 390, August 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900362.

Hender:1985:CSE

[HCG^{+}85] T. C. Hender, B. A. Carreras, L. Garcia, J. A. Rome, and V. E. Lynch. The calculation of stellarator equilibria in vacuum flux surface coordinates. Journal of Computational Physics, 60(1): 76–96, August 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918590018X.

Hicks:1981:NCR

[HCH^{+}81] H. R. Hicks, B. Carreras, J. A. Holmes, D. K. Lee, and B. V. Waddell. 3D nonlinear calculations of resistive tearing modes. Journal of Computational Physics, 44(1):46–69, November 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900371.

Hicks:1985:RCR

[HCH85] H. R. Hicks, B. A. Carreras, and J. A. Holmes. The reliability of calculations with the RSF algorithm. Journal of Computational Physics, 60(3):558–560, September 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900373.

Hortensius:1989:VAP

[HCM89] Peter D. Hortensius, Howard C. Card, and Robert D. McLeod. A VLSI architecture for percolation simulation. Journal of Computational Physics, 84(1):76–89, September 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901824.

Horak:1985:NCM

[HD85] W. C. Horak and J. J. Dorning. A nodal coarse-mesh method for the efficient numerical solution of laminar flow problems. Journal of Computational Physics, 59(3):405–440, July 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901202.

Heinrichs:1988:LRS

[Hei88a] Wilhelm Heinrichs. Line relaxation for spectral multigrid methods. Journal of Computational Physics, 77(1):166–182, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901611.

Heinrichs:1988:MMC

[Hei88b] Wilhelm Heinrichs. Multigrid methods for combined finite difference and Fourier problems. Journal of Computational Physics, 78(2):424–436, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900587.

Henry:1981:RSS

[Hen81] R. F. Henry. Richardson–Sielecki schemes for the shallowwater equations, with applications to Kelvin waves. Journal of Computational Physics, 41(2):389–406, June 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901029.

Henshaw:1987:SNS

[Hen87] W. D. Henshaw. A scheme for the numerical solution of hyperbolic systems of conservation laws. Journal of Computational Physics, 68(1):25–47, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900428.

Harten:1987:UHO

[HEOC87] Ami Harten, Bjorn Engquist, Stanley Osher, and Sukumar R. Chakravarthy. Uniformly high order accurate essentially non-oscillatory schemes, III. Journal of Computational Physics, 71(2):231–303, August 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900313.

Hermeline:1989:DPM

[Her89] F. Hermeline. A deterministic particle method for transport diffusion equations: Application to the Fokker–Planck equation. Journal of Computational Physics, 82(1):122–146, May 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900387.

Hewett:1980:GMS

[Hew80] D. W. Hewett. A global method of solving the electron-field equations in a zero-inertia-electron-hybrid plasma simulation code. Journal of Computational Physics, 38(3):378–395, December 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180901552.

Hirschel:1984:WCC

[HG84] E. H. Hirschel and A. Groh. Wall-compatibility condition for the solution of the Navier–Stokes equations. Journal of Computational Physics, 53(2):346–350, February 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900469.

Helton:1986:IMP

[HG86]  F. J. Helton and J. M. Greene.  Ideal MHD proper-

ties for proposed noncircular tokamaks. Journal of Computational Physics, 66(2):458–468, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900756.

Holz:1988:TIT

[HG88] J. Holz and W. Glockle. Time-independent and time-dependent potential scattering without angular momentum decomposition. Journal of Computational Physics, 76(1):131–158, May 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901349.

Huh:1986:MRN

[HGM86] Kang Y. Huh, Michael W. Golay, and Vincent P. Manno. A method for reduction of numerical diffusion in the donor cell treatment of convection. Journal of Computational Physics, 63(1):201–221, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900926.

Hosoya:1988:EBE

[HGS88] M. Hosoya, E. Goto, and T. Soma. Extrapolated boundary element method for three-dimensional inductance calculation. Journal of Computational Physics, 74(1):94–109, January 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890071X.

Harten:1983:SAG

[HH83] Ami Harten and James M. Hyman. Self adjusting grid methods for one-dimensional hyperbolic conservation laws. Journal of Computational Physics, 50(2):235–269, May 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900669.

Handy:1985:ACS

[HH85] C. R. Handy and S. P. Hirshman. Accelerated convergence of the steepest descent method for magnetohydrodynamic equilibria. Journal of Computational Physics, 60(2):338–345, September 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900117.

Hirshman:1986:OTD

[HH86] S. P. Hirshman and J. T. Hogan. ORMEC: A threedimensional MHD spectral inverse equilibrium code. Journal of Computational Physics, 63(2):329–352, April 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918690197X.

Haj-Hariri:1988:TRO

[HH88] Hossein Haj-Hariri. Transformations reducing the order of the parameter in differential eigenvalue problems. Journal of Computational Physics, 77(2):472–484, August 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901787.

Hayot:1986:SPC

[HHN^{+}86] F. Hayot, H. J. Herrmann, J.-M Normand, P. Farthouat, and M. Mur. A special purpose computer for the electrical conductivity of disordered media. Journal of Computational Physics, 64(2): 380–388, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900392.

Harafuji:1989:CST

[HHS89] Kenji Harafuji, Takaya Hayashi, and Tetsuya Sato. Computational study of three-dimensional magnetohydrodynamic equilibria in toroidal helical systems. Journal of Computational Physics, 81(1):169–192, March 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900697.

Hirao:1989:IMS

[Hir89] K. Hirao. Iterative methods for the solution of large systems of linear equations. Journal of Computational Physics, 80(1):232– 242, January 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918990096X.

Hitzl:1980:IHL

[Hit80] Donald L. Hitzl. Implementing Hamilton’s law of varying action with shifted Legendre polynomials. Journal of Computational Physics, 38(2):185–211, November 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900522.

Hu:1981:DCD

[HK81] Chin-Kun Hu and P. Kleban. Direct calculation of the derivatives of the free energy for Ising models by a modified Kadanoff variational method. Journal of Computational Physics, 43(2):289– 314, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901248.

Hatziavramidis:1983:PSL

[HK83] Dimitri Hatziavramidis and Hwar-Ching Ku. Pseudospectral solutions of laminar heat transfer problems in pipelines. Journal of Computational Physics, 52(2):414–424, November 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900384.

Holodniok:1984:DAC

[HK84] Martin Holodniok and Milan Kubícek. DERPER-an algorithm for the continuation of periodic solutions in ordinary differential equations. Journal of Computational Physics, 55(2):254–267, August 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900056.

Harned:1985:SIM

[HK85a] Douglas S. Harned and W. Kerner. Semi-implicit method for three-dimensional compressible magnetohydrodynamic simulation. Journal of Computational Physics, 60(1):62–75, August 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900178.

Hastings:1985:CSC

[HK85b] D. E. Hastings and T. Kamimura. Calculation of the selfconsistent electric field in toroidal nonaxisymmetric devices. Journal of Computational Physics, 61(2):286–296, November 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900889.

Hatziavramidis:1986:PMS

[HK86] Dimitri Hatziavramidis and Hwar-Ching Ku. A pseudospectral method for the solution of the two-dimensional Navier– Stokes equations in the primitive variable formulation. Journal of Computational Physics, 67(2):361–371, December 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902688.

Hitchon:1989:ESC

[HKA89] W. N. G. Hitchon, D. J. Koch, and J. B. Adams. An efficient scheme for convection-dominated transport. Journal of Computational Physics, 83(1):79–95, July 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902234.

Hui:1984:NIC

[HL84] Bertram Hui and Martin Lampe. A nonlinear implicit code for relativistic electron beam tracking studies. Journal of Computational Physics, 55(2):328–339, August 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918490010X.

Hariharan:1985:ASV

[HL85] S. I. Hariharan and Harold C. Lester. Acoustic shocks in a variable area duct containing near sonic flows. Journal of Computational Physics, 58(1):134–145, March 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901615.

Hicks:1986:IPM

[HL86] H. R. Hicks and V. E. Lynch. An introduction to programming multiple-processor computers. Journal of Computational Physics, 63(1):140–156, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900884.

Hewett:1987:EDI

[HL87] Dennis W. Hewett and A. Bruce Langdon. Electromagnetic direct implicit plasma simulation. Journal of Computational Physics, 72(1):121–155, September 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900751.

Haldenwang:1984:CDS

[HLAD84] P. Haldenwang, G. Labrosse, S. Abboudi, and M. Deville. Chebyshev 3-D spectral and 2-D pseudospectral solvers for the Helmholtz equation. Journal of Computational Physics, 55(1): 115–128, July 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900184.

Halbleib:1980:AMC

[HM80a] J. A. Halbleib and J. E. Morel. Adjoint Monte Carlo electron transport in the continuous-slowing-down approximation. Journal of Computational Physics, 34(2):211–230, February 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901060.

Hitzl:1980:AEP

[HM80b] Donald L. Hitzl and Frederick H. Maltz. Adaptive estimation procedures for multi-parameter Monte Carlo computations. Journal of Computational Physics, 37(2):218–241, September 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900224.

Hariharan:1982:IEP

[HM82] S. I. Hariharan and R. C. MacCamy. Integral equation procedures for eddy current problems. Journal of Computational Physics, 45(1):80–99, January 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901036.

Harned:1989:ASI

[HM89] Douglas S. Harned and Zoran Mikic’. Accurate semi-implicit treatment of the Hall effect in magnetohydrodynamic computations. Journal of Computational Physics, 83(1):1–15, July 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189902209.

Han:1983:OBC

[HMI83] T. Y. Han, J. C. S. Meng, and G. E. Innis. An open boundary condition for incompressible stratified flows. Journal of Computational Physics, 49(2):276–297, February 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901274.

Herrnegger:1986:CTD

[HMJ86] F. Herrnegger, P. Merkel, and J. L. Johnson. Comparison of two-dimensional and three-dimensional MHD equilibrium and stability codes. Journal of Computational Physics, 66(2):445– 457, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900744.

Herbst:1985:NEN

[HMM85] B. M. Herbst, J. Ll Morris, and A. R. Mitchell. Numerical experience with the nonlinear Schrödinger equation. Journal of Computational Physics, 60(2):282–305, September 15, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900087.

Holland:1982:UMB

[HMR82] W. Holland, S. McCormick, and J. Ruge. Unigrid methods for boundary value problems with nonrectangular domains. Journal of Computational Physics, 48(3):412–422, December 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900596.

Hayot:1989:IPB

[HMS89] F. Hayot, M. Mandal, and P. Sadayappan. Implementation and performance of a binary lattice gas algorithm on parallel processor systems. Journal of Computational Physics, 80(2):277– 287, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901009.

Herbst:1985:SNS

[HMW85] B. M. Herbst, A. R. Mitchell, and J. A. C. Weideman. On the stability of the nonlinear Schrödinger equation. Journal of Computational Physics, 60(2):263–281, September 15, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900075.

Hrymak:1986:IMF

[HMW86] Andrew N. Hrymak, Gregory J. McRae, and Arthur W. Westerberg. An implementation of a moving finite element method. Journal of Computational Physics, 63(1):168–190, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900902.

Hirt:1980:ALC

[HN80] C. W. Hirt and B. D. Nichols. Adding limited compressibility to incompressible hydrocodes. Journal of Computational Physics, 34(3):390–400, March 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900960.

Hirt:1981:VFV

[HN81] C. W. Hirt and B. D. Nichols. Volume of fluid (VOF) method for the dynamics of free boundaries. Journal of Computational Physics, 39(1):201–225, January 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901455.

Hirao:1982:GDM

[HN82] K. Hirao and H. Nakatsuji. A generalization of the Davidson’s method to large nonsymmetric eigenvalue problems. Journal of Computational Physics, 45(2):246–254, February 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918290119X.

Hodgson:1988:NSS

[HN88] M. J. Hodgson and C. M. M. Nex. On “A new splitting to solve a large Hermitian eigenproblem”. Journal of Computational Physics, 79(2):484–485, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900290. See [Nex87].

Hager:1989:EDE

[HNK89] William W. Hager, John S. Nisbet, and John R. Kasha. The evolution and discharge of electric fields within a thunderstorm. Journal of Computational Physics, 82(1):193–217, May 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900429.

Hedstrom:1980:ECR

[HO80] G. W. Hedstrom and Albert Osterheld. The effect of cell Reynolds number on the computation of a boundary layer. Journal of Computational Physics, 37(3):399–421, October 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900455.

Hoek:1983:SCE

[Hoe83] Jaap Hoek. Strong coupling expansion of the generating functional for gauge systems on a lattice with arbitrary sources. Journal of Computational Physics, 49(2):265–275, February 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901262.

Hoek:1984:SCE

[Hoe84] Jaap Hoek. Strong coupling expansion of the SU(3) and U(3) effective actions. Journal of Computational Physics, 54(2):245– 259, May 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901177.

Hoffman:1982:RBT

[Hof82] Joe D. Hoffman. Relationship between the truncation errors of centered finite-difference approximations on uniform and nonuniform meshes. Journal of Computational Physics, 46(3):469–474, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900286.

Horiuti:1987:CCR

[Hor87a] Kiyosi Horiuti. Comparison of conservative and rotational forms in large eddy simulation of turbulent channel flow. Journal of Computational Physics, 71(2):343–370, August 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900350.

Horowitz:1987:VIR

[Hor87b] Eric J. Horowitz. Vectorizing the interpolation routines of particle-in-cell codes. Journal of Computational Physics, 68(1): 56–65, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900441.

Hosack:1986:UCM

[Hos86] John M. Hosack. The use of Čebyšev mixing to generate pseudorandom numbers. Journal of Computational Physics, 67(2):482– 486, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186902743.

Haworth:1987:MCS

[HP87] D. C. Haworth and S. B. Pope. Monte Carlo solutions of a joint PDF equation for turbulent flows in general orthogonal coordinates. Journal of Computational Physics, 72(2):311–346, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900866.

He-Ping:1986:FPM

[HPBY86] Ma He-Ping and Guo Ben-Yu. The Fourier pseudospectral method with a restrain operator for the Korteweg–de Vries equation. Journal of Computational Physics, 65(1):120–137, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900070.

Holmes:1980:EFC

[HPL80] J. A. Holmes, Y.-K. M. Peng, and S. J. Lynch. Evolution of flux-conserving tokamak equilibria with preprogrammed cross sections. Journal of Computational Physics, 36(1):35–54, June 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180901734.

Heifetz:1982:MCM

[HPP^{+}82] D. Heifetz, D. Post, M. Petravic, J. Weisheit, and G. Bateman. A Monte–Carlo model of neutral-particle transport in diverted plasmas. Journal of Computational Physics, 46(2):309–327, May 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900171.

Hausman:1984:TDS

[HR84a] Mark A. Hausman and William W. Roberts, Jr. Three dimensional, stratified gas flows past an obstacle. Journal of Computational Physics, 55(3):347–368, September 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900251.

Helstrom:1984:CCD

[HR84b] Carl W. Helstrom and Stephen O. Rice. Computation of counting distributions arising from a single-stage multiplicative process. Journal of Computational Physics, 54(2):289–324, May 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901207.

Howland:1986:CIA

[HR86] R. A. Howland and D. L. Richardson. Computer implementation of an algorithm for the quadratic analytical solution of Hamiltonian systems. Journal of Computational Physics, 67(1):19–27, November 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901129.

Haidvogel:1980:AES

[HRS80] D. B. Haidvogel, A. R. Robinson, and E. E. Schulman. The accuracy, efficiency, and stability of three numerical models with application to open ocean problems. Journal of Computational Physics, 34(1):1–53, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901114.

Hanggi:1980:EIS

[HRT80] P. Hänggi, F. Roesel, and D. Trautmann. Evaluation of infinite series by use of continued fraction expansions: A numerical study. Journal of Computational Physics, 37(2):242–258, September 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900236.

Harned:1986:SIM

[HS86] Douglas S. Harned and D. D. Schnack. Semi-implicit method for long time scale magnetohydrodynamic computations in three dimensions. Journal of Computational Physics, 65(1):57–70, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900045.

Hale:1988:OCD

[HS88a] Jack K. Hale and Natalia Sternberg. Onset of chaos in differential delay equations. Journal of Computational Physics, 77(1): 221–239, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901647.

Hall:1988:KMU

[HS88b] Peter Hall and Richard L. Smith. The kernel method for unfolding sphere size distributions. Journal of Computational Physics, 74(2):409–421, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890085X.

Horowitz:1989:QTD

[HSA89] Eric J. Horowitz, Dan E. Shumaker, and David V. Anderson. QN3D: A three-dimensional quasi-neutral hybrid particle-in-cell code with applications to the tilt mode instability in field reversed configurations. Journal of Computational Physics, 84(2):279– 310, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189902349.

Hoogland:1983:SPP

[HSSC83] A. Hoogland, J. Spaa, B. Selman, and A. Compagner. A specialpurpose processor for the Monte Carlo simulation of Ising spin systems. Journal of Computational Physics, 51(2):250–260, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900918.

Hoffmann:1988:IHL

[HSV88] K.-H Hoffmann, J. Sprekels, and A. Visintin. Identification of hysteresis loops. Journal of Computational Physics, 78(1):215– 230, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900459.

Hosking:1986:CVR

[HT86] R. J. Hosking and J. Tendys. Computation of visco-resistive MHD instabilities. Journal of Computational Physics, 66(2):274– 293, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900677.

Harten:1981:FOA

[HTE81] A. Harten and H. Tal-Ezer. On a fourth order accurate implicit finite difference scheme for hyperbolic conservation laws. II. Five-point schemes. Journal of Computational Physics, 41(2): 329–356, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901005.

Hu:1989:MIS

[Hu89] Y. Q. Hu. A multistep implicit scheme for time-dependent 2-dimensional magnetohydrodynamic flows. Journal of Computational Physics, 84(2):441–460, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902428.

Huang:1981:PUD

[Hua81] Lan Chieh Huang. Pseudo-unsteady difference schemes for discontinuous solutions of steady-state, one-dimensional fluid dynamics problems. Journal of Computational Physics, 42(1):195– 211, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902394.

Hunt:1986:NSE

[Hun86] Roland Hunt. The numerical solution of elliptic free boundary problems using multigrid techniques. Journal of Computational Physics, 65(2):448–461, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902172.

Hunt:1989:NSP

[Hun89] Roland Hunt. The numerical solution of parabolic free boundary problems arising from thin film flows. Journal of Computational Physics, 84(2):377–402, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902398.

Hutson:1984:QSM

[Hut84] Jeremy M. Hutson. A quadrature scheme for matrix elements between numerical wavefunctions. Journal of Computational Physics, 56(1):165–172, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900895.

Hicks:1980:NDT

[HW80] H. R. Hicks and J. W. Wooten. Numerical divergence of a tensor. Journal of Computational Physics, 36(3):411–416, July 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180901679.

Hunt:1981:CTC

[HW81] Roland Hunt and Graham Wilks. Continuous transformation computation of boundary layer equations between similarity regimes. Journal of Computational Physics, 40(2):478–490, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902230.

Hu:1984:FIC

[HW84] Y. Q. Hu and S. T. Wu. A full-implicit-continuous-Eulerian (FICE) scheme for multidimensional transient magnetohydrodynamic (MHD) flows. Journal of Computational Physics, 55(1): 33–64, July 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900147.

Ioakimidis:1985:MDL

[IA85] N. I. Ioakimidis and E. G. Anastasselou. A modification of the Delves–Lyness method for locating the zeros of analytic functions. Journal of Computational Physics, 59(3):490–492, July 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901251.

Itoh:1988:HCD

[IA88] Toshiaki Itoh and Kanji Abe. Hamiltonian-conserving discrete canonical equations based on variational difference quotients. Journal of Computational Physics, 76(1):85–102, May 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901325.

If:1987:SSS

[IBCS87] F. If, P. Berg, P. L. Christiansen, and O. Skovgaard. Split-step spectral method for nonlinear Schrödinger equation with absorbing boundaries. Journal of Computational Physics, 72(2):501– 503, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900970.

Ixaru:1980:CSS

[ICP80] L. Gr. Ixaru, M. I. Cristu, and M. S. Popa. Choosing step sizes for perturbative methods of solving the Schrödinger equation. Journal of Computational Physics, 36(2):170–181, July 1, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901825.

Issa:1986:CCI

[IGW86] R. I. Issa, A. D. Gosman, and A. P. Watkins. The computation of compressible and incompressible recirculating flows by a non-iterative implicit scheme. Journal of Computational Physics, 62(1):66–82, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901002.

Ingham:1981:NSP

[IHM81] D. B. Ingham, P. J. Heggs, and M. Manzoor. The numerical solution of plane potential problems by improved boundary integral equation methods. Journal of Computational Physics, 42(1): 77–98, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902333.

Kanatani:1985:EAS

[iKI85]  Ken ichi Kanatani and Osamu Ishikawa.  Error analy-

sis for the stereological estimation of sphere size distribution: Abel type integral equation. Journal of Computational Physics, 57(2):229–250, January 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900440.

Iyengar:1988:HOD

[IM88] Satteluri R. K. Iyengar and Ram Manohar. High order difference methods for heat equation in polar cylindrical coordinates. Journal of Computational Physics, 77(2):425–438, August 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901763.

Ishibashi:1989:VPW

[IMT^{+}89] Kenji Ishibashi, Yuki Miura, Hiroshi Takada, Takeji Sakae, Yuzuru Matsumoto, and Akira Katase. Vectorized partial wave calculation using a Cauchy-type propagation matrix. Journal of Computational Physics, 80(1):17–26, January 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900880.

Ingham:1984:US

[Ing84] D. B. Ingham. Unsteady separation. Journal of Computational Physics, 53(1):90–99, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900548.

Inoue:1983:GGC

[Ino83] Kenji Inoue. Grid generation for cascades using conformal mapping. Journal of Computational Physics, 52(1):130–141, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900207.

Inoue:1985:GGI

[Ino85] Kenji Inoue. Grid generation for inlet configurations using conformal mapping. Journal of Computational Physics, 58(1):146– 154, March 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901627.

Israeli:1981:ARB

[IO81] Moshe Israeli and Steven A. Orszag. Approximation of radiation boundary conditions. Journal of Computational Physics, 41(1): 115–135, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181900826.

Ioakimidis:1981:IAN

[Ioa81a] Nikolaos I. Ioakimidis. An iterative algorithm for the numerical solution of singular integral equations. Journal of Computational Physics, 43(1):164–176, September 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901157.

Ioakimidis:1981:RAC

[Ioa81b] Nikolaos I. Ioakimidis. A remark on the application of closed and semi-closed quadrature rules to the direct numerical solution of singular integral equations. Journal of Computational Physics, 42(2):396–402, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902527.

Ioakimidis:1982:NIF

[Ioa82] Nikolaos I. Ioakimidis. A natural interpolation formula for Cauchy-type singular integral equations with generalized kernels. Journal of Computational Physics, 48(1):117–126, October 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900389.

Ixaru:1987:NMM

[IR87] L. Gr Ixaru and M. Rizea. Numerov method maximally adapted to the Schrödinger equation. Journal of Computational Physics, 73(2):306–324, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901392.

Issa:1986:SID

[Iss86] R. I. Issa. Solution of the implicitly discretised fluid flow equations by operator-splitting. Journal of Computational Physics, 62(1): 40–65, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900999.

Ingham:1989:NIS

[IT89] D. B. Ingham and T. Tang. Numerical investigation into the steady flow past a rotating circular cylinder at low and intermediate Reynolds numbers. Journal of Computational Physics, 84(1):256–257, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901976.

Ixaru:1980:HCP

[Ixa80] L. Gr. Ixaru. A highly convergent perturbative method for the solution of systems of coupled equations arising from the Schrödinger equation. Journal of Computational Physics, 36(2):182–197, July 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180901837.

Ixaru:1987:NMS

[Ixa87] L. Gr Ixaru. The Numerov method and singular potentials. Journal of Computational Physics, 72(1):270–274, September 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900817.

J:1982:BBM

[J.82] J. E. Dendy J. Black box multigrid. Journal of Computational Physics, 48(3):366–386, December 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900572.

Jacobs:1989:PMT

[Jac89] S. J. Jacobs. A pseudospectral method for two-point boundary value problems. Journal of Computational Physics, 85(2):502, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901654.

Jardin:1981:SCS

[Jar81] S. C. Jardin. Self-consistent solutions of the plasma transport equations in an axisymmetric toroidal system. Journal of Computational Physics, 43(1):31–60, September 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901108.

Johnson:1986:MRC

[JDK86] John L. Johnson, Robert L. Dewar, and John Killeen. In memoriam: Raymond C. Grimm (1944–1984). Journal of Computational Physics, 66(2):251–254, October 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900653.

Johnson:1983:NDS

[JH83] S. H. Johnson and A. C. Hindmarsh. Numerical dynamic simulation of solid-fluid reactions in isothermal porous spheres. Journal of Computational Physics, 52(3):503–523, December 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900050.

Joyce:1989:SRE

[JHLS89] Glenn Joyce, Richard Hubbard, Martin Lampe, and Steven Slinker. SARLAC: A relativistic electron beam code. Journal of Computational Physics, 81(1):193–205, March 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900703.

Joyce:1986:SLP

[JL86] Glenn Joyce and Martin Lampe. SIMM1: A linearized particle code. Journal of Computational Physics, 63(2):398–415, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186902019.

Jones:1981:AOS

[Jon81] Everett Jones. An asymptotic outer solution applied to the Keller box method. Journal of Computational Physics, 40(2): 411–429, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902199.

Jones:1985:NMT

[Jon85] C. A. Jones. Numerical methods for the transition to wavy Taylor vortices. Journal of Computational Physics, 61(2):321–344, November 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900907.

Jardin:1986:DMT

[JPD86] S. C. Jardin, N. Pomphrey, and J. Delucia. Dynamic modeling of transport and positional control of tokamaks. Journal of Computational Physics, 66(2):481–507, October 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918690077X.

Jacobs:1981:MSC

[JR81] Laurence Jacobs and Claudio Rebbi. Multi-spin coding: A very efficient technique for Monte Carlo simulations of spin systems. Journal of Computational Physics, 41(1):203–210, May 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900899.

Jhaveri:1982:ETI

[JR82] B. S. Jhaveri and F. Rosenberger. Exact triple integrals of beam functions. Journal of Computational Physics, 45(2):300– 302, February 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182901231.

Jayaraman:1985:ECS

[JR85] Lalitha Jayaraman and G. Ramanaiah. Effect of couple stresses on transient MHD Poiseuille flow. Journal of Computational Physics, 60(3):478–488, September 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900324.

Jordan:1986:ENE

[JRS86] Kirk E. Jordan, Gerard R. Richter, and Ping Sheng. An efficient numerical evaluation of the Green’s function for the Helmholtz operator on periodic structures. Journal of Computational Physics, 63(1):222–235, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900938.

Jarzebski:1986:SHA

[JT86] Andrzej B. Jarzȩbski and Jan W. Thullie. A stable highly accurate ADI method for hyperbolic heat conduction equation. Journal of Computational Physics, 63(1):236–239, March 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918690094X.

Johnson:1988:MFE

[JWB88] I. W. Johnson, A. J. Wathen, and M. J. Baines. Moving finite element methods for evolutionary problems. II. Applications. Journal of Computational Physics, 79(2):270–297, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900174.

Jafroudi:1983:CSL

[JY83] Hamid Jafroudi and H. T. Yang. Computation of steady laminar flow over a circular cylinder with third-order boundary conditions. Journal of Computational Physics, 49(2):181–198, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901225.

Jafroudi:1986:SLF

[JY86] Hamid Jafroudi and H. T. Yang. Steady laminar forced convection from a circular cylinder. Journal of Computational Physics, 65(1):46–56, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900033.

Kaijser:1983:SIM

[Kai83] Thomas Kaijser. A simple inversion method for determining aerosol size distributions. Journal of Computational Physics, 52(1):80–104, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900177.

Kalamboukis:1984:LTA

[Kal84] T. Z. Kalamboukis. A Lanczos-type algorithm for the generalized eigenvalue problem Ax = λBx. Journal of Computational Physics, 53(1):82–89, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900536.

Kameari:1981:TEC

[Kam81] Akihisa Kameari. Transient eddy current analysis on thin conductors with arbitrary connections and shapes. Journal of Computational Physics, 42(1):124–140, July 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902369.

Kansa:1981:AMC

[Kan81] Edward J. Kansa. An algorithm for multidimensional combusting flow problems. Journal of Computational Physics, 42(1):152–194, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902382.

Kansa:1983:RLR

[Kan83] Edward J. Kansa. Reply to the letter of Rarnshaw and Dukowicz. Journal of Computational Physics, 49(2):351–353, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918390133X. See [RD83].

Kanatani:1984:APP

[Kan84]  Ken-Ichi Kanatani.  The accuracy and the preservation

property of the discrete mechanics. Journal of Computational Physics, 53(1):181–187, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900597.

Kawakami:1987:FEM

[Kaw87] Ichiro Kawakami. Finite element method with optimal nodal velocity. Journal of Computational Physics, 73(1):59–72, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187901057.

Klapper:1980:UCM

[KD80] Michael H. Klapper and David Debrota. Use of Caley–Menger determinants in the calculation of molecular structures. Journal of Computational Physics, 37(1):56–69, August 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900042.

Kaplan:1988:TDS

[KD88] H. Kaplan and N. Dinar. A three-dimensional stochastic model for concentration fluctuation statistics in isotropic homogeneous turbulence. Journal of Computational Physics, 79(2):317–335, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900198.

Kapitza:1987:DPS

[KE87] H. Kapitza and D. Eppel. A 3-D Poisson solver based on conjugate gradients compared to standard iterative methods and its performance on vector computers. Journal of Computational Physics, 68(2):474–484, February 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900672.

Kelmanson:1983:IEM

[Kel83] M. A. Kelmanson. An integral equation method for the solution of singular slow flow problems. Journal of Computational Physics, 51(1):139–158, July 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900840.

Kelmanson:1984:SNE

[Kel84] M. A. Kelmanson. Solution of nonlinear elliptic equations with boundary singularities by an integral equation method. Journal of Computational Physics, 56(2):244–258, November 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900949.

Kershaw:1980:PUP

[Ker80] David S. Kershaw. On the problem of unstable pivots in the incomplete LU-conjugate gradient method. Journal of Computational Physics, 38(1):114–123, November 1, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900157.

Kershaw:1981:DDE

[Ker81] David S. Kershaw. Differencing of the diffusion equation in Lagrangian hydrodynamic codes. Journal of Computational Physics, 39(2):375–395, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901583.

Ker:1988:SRT

[Ker88] Robert M. Ker. Simulation of Rayleigh–Taylor flows using vortex blobs. Journal of Computational Physics, 76(1):48–84, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901313.

Kerner:1989:LSC

[Ker89] W. Kerner. Large-scale complex eigenvalue problems. Journal of Computational Physics, 85(1):1–85, November 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902003.

Keunings:1986:AST

[Keu86] Roland Keunings. An algorithm for the simulation of transient viscoelastic flows with free surfaces. Journal of Computational Physics, 62(1):199–220, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901075.

Karageorghis:1987:MFS

[KF87] Andreas Karageorghis and Graeme Fairweather. The method of fundamental solutions for the numerical solution of the biharmonic equation. Journal of Computational Physics, 69(2):434– 459, April 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901768.

Kinnmark:1985:SAS

[KG85] Ingemar Pe Kinnmark and William G. Gray. Stability and accuracy of spatial approximations for wave equation tidal models. Journal of Computational Physics, 60(3):447–466, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900300.

Kreskovsky:1987:ALT

[KG87] J. P. Kreskovsky and H. L. Grubin. Application of LBI techniques to the solution of the transient, multidimensional semiconductor equations. Journal of Computational Physics, 68(2):420– 461, February 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900659.

Keller:1989:ENR

[KG89] Joseph B. Keller and Dan Givoli. Exact non-reflecting boundary conditions. Journal of Computational Physics, 82(1):172– 192, May 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189900417.

Krieg:1980:TTDb

[KGH80] R. Krieg, B. Göller, and G. Hailfinger. Transient, threedimensional potential flow problems and dynamic response of the surrounding structures. Part II: Simultaneous coupling between fluid and structural dynamics (computer code SING–S). Journal of Computational Physics, 34(2):164–183, February 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901035.

Krieg:1980:TTDa

[KH80] R. Krieg and G. Hailfinger. Transient, three-dimensional potential flow problems and dynamic response of the surrounding structures. Part I: Description of the fluid dynamics by a singularity method (computer code SING). Journal of Computational Physics, 34(2):139–163, February 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901023.

Ku:1984:CEM

[KH84] Hwar-Ching Ku and Dimitri Hatziavramidis. Chebyshev expansion methods for the solution of the extended Graetz problem. Journal of Computational Physics, 56(3):495–512, December 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184901098.

Kubicek:1987:ADP

[KH87] Milan Kubícek and Martin Holodniok. Algorithms for determination of period-doubling bifurcation points in ordinary differential equations. Journal of Computational Physics, 70(1):203–217, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900106.

Ku:1987:PMS

[KHT87] Hwar C. Ku, Richard S. Hirsh, and Thomas D. Taylor. A pseudospectral method for solution of the three-dimensional incompressible Navier–Stokes equations. Journal of Computational Physics, 70(2):439–462, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901902.

Ku:1989:PME

[KHTR89] Hwar C. Ku, Richard S. Hirsh, Thomas D. Taylor, and Allan P. Rosenberg. A pseudospectral matrix element method for solution of three-dimensional incompressible flows and its parallel implementation. Journal of Computational Physics, 83(2):260– 291, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901198.

Anonymous:1987:MKV

[Kil87] John Killeen. In memoriam Keith V. Roberts. Journal of Computational Physics, 73(1):vi + 1–4, November 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901021.

Kerner:1986:FES

[KJL86] W. Kerner, A. Jakoby, and K. Lerbinger. Finite-element semidiscretization of linearized compressible and resistive MHD. Journal of Computational Physics, 66(2):332–355, October 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900707.

Knorr:1980:FOP

[KJM80] G. Knorr, G. Joyce, and A. J. Marcus. Fourth-order Poisson solver for the simulation of bounded plasmas. Journal of Computational Physics, 38(2):227–236, November 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900546.

Karasalo:1981:SSO

[KK81] Ilkka Karasalo and John Kurylo. On solving the stiff ODE’s of the kinetics of chemically reacting gas flow. Journal of Computational Physics, 40(1):167–182, March 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902059.

Kosloff:1983:FMS

[KK83] D. Kosloff and R. Kosloff. A Fourier method solution for the time dependent Schrödinger equation as a tool in molecular dynamics. Journal of Computational Physics, 52(1):35–53, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900153.

Kosloff:1986:ABW

[KK86] R. Kosloff and D. Kosloff. Absorbing boundaries for wave propagation problems. Journal of Computational Physics, 63(2):363– 376, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901993.

Kobeissi:1988:NVS

[KK88] Hafez Kobeissi and Majida Kobeissi. A new variable step method for the numerical integration of the one-dimensional Schrödinger equation. Journal of Computational Physics, 77(2):501–512, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901805.

Katayama:1989:SIM

[KK89] Yasunao Katayama and Masato Koda. Scaling integration method for singular integrals and its application to BEM. Journal of Computational Physics, 83(2):292–302, August 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901204.

Kennedy:1986:PEM

[KKMP86] A. D. Kennedy, J. Kuti, S. Meyer, and B. J. Pendleton. Program for efficient Monte Carlo computations of quenched SU(3) lattice gauge theory using the quasi-heatbath method on a CDC CYBER 205 computer. Journal of Computational Physics, 64(1): 133–160, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900227.

Kirkegaard:1980:TTR

[KL80] P. Kirkegaard and L. Løvborg. Transport of terrestrial y radiation in plane semi-infinite. Journal of Computational Physics, 36(1): 20–34, June 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180901722.

Koniges:1989:PPR

[KL89] A. E. Koniges and C. E. Leith. Parallel processing of random number generation for Monte Carlo turbulence simulation. Journal of Computational Physics, 81(1):230–235, March 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900727.

Klimas:1983:NMB

[Kli83] Alexander J. Klimas. A numerical method based on the Fourier– Fourier transform approach for modeling 1-D electron plasma evolution. Journal of Computational Physics, 50(2):270–306, May 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900670.

Klimas:1987:MOV

[Kli87] Alexander J. Klimas. A method for overcoming the velocity space filamentation problem in collisionless plasma model solutions. Journal of Computational Physics, 68(1):202–226, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900520.

Knorr:1980:RSL

[KM80] G. Knorr and M. Mond. The representation of shocklike solutions in an Eulerian mesh. Journal of Computational Physics, 38(2):212–226, November 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900534.

Kim:1985:AFS

[KM85] J. Kim and P. Moin. Application of a fractional-step method to incompressible Navier–Stokes equations. Journal of Computational Physics, 59(2):308–323, June 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901482.

Khorrami:1989:ASC

[KMA89] Mehdi R. Khorrami, Mujeeb R. Malik, and Robert L. Ash. Application of spectral collocation techniques to the stability of swirling flows. Journal of Computational Physics, 81(1):206–229, March 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900715.

Kelly:1988:PEE

[KMRM88] D. W. Kelly, R. J. Mills, J. A. Reizes, and A. D. Miller. A posteriori error estimates in finite difference techniques. Journal of Computational Physics, 74(1):214–232, January 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900770.

Kim:1986:GCM

[KO86] S. Y. Kim and H. Okuda. Guiding center magnetostatic particle simulation model in three dimensions. Journal of Computational Physics, 65(1):215–226, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900124.

Kobeissi:1985:TDV

[Kob85] Hafez Kobeissi. On testing diatomic vibration-rotation wavefunction for high levels. Journal of Computational Physics, 61(3):351– 358, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900695.

Koren:1988:DCM

[Kor88] Barry Koren. Defect correction and multigrid for an efficient and accurate computation of airfoil flows. Journal of Computational Physics, 77(1):183–206, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901623.

Koren:1989:MDC

[Kor89] Barry Koren. Multigrid and defect correction for the steady Navier–Stokes equations. Journal of Computational Physics, 84(1):256, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189901964.

Kosugi:1984:MLD

[Kos84] Nobuhiro Kosugi. Modification of the Liu–Davidson method for obtaining one or simultaneously several eigensolutions of a large real-symmetric matrix. Journal of Computational Physics, 55(3):426–436, September 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900317.

Koshizuka:1988:IMM

[KOTK88] Seiichi Koshizuka, Yoshiaki Oka, Yasumasa Togo, and Shunsuke Kondo. Interpolating matrix method: A finite difference method for arbitrary arrangement of mesh points. Journal of Computational Physics, 75(2):444–468, April 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901222.

Kannan:1983:NMN

[KP83a] R. Kannan and W. Proskurowski. A numerical method for the nonlinear Neumann problem. Journal of Computational Physics, 52(1):105–121, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900189.

Kuo-Petravic:1983:NTA

[KP83b] Gioietta Kuo-Petravic. Numerical treatment of the axial singularity in a flux coordinate system for particle simulation. Journal of Computational Physics, 52(1):209–213, October 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900268.

Kaas-Petersen:1985:CQP

[KP85] Chr. Kaas-Petersen. Computation of quasi-periodic solutions of forced dissipative systems. Journal of Computational Physics, 58(3):395–408, May 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901706.

Kaas-Petersen:1986:CQP

[KP86a] Chr Kaas-Petersen. Computation of quasi-periodic solutions of forced dissipative systems II. Journal of Computational Physics, 64(2):433–442, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900422.

Korczak:1986:ISE

[KP86b] Karol Z. Korczak and Anthony T. Patera. An isoparametric spectral element method for solution of the Navier–Stokes equations in complex geometry. Journal of Computational Physics, 62(2):361– 382, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901348.

Karageorghis:1989:SCM

[KP89a] A. Karageorghis and T. N. Phillips. Spectral collocation methods for Stokes flow in contraction geometries and unbounded domains. Journal of Computational Physics, 80(2):314–330, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901022.

Karageorghis:1989:CSC

[KP89b] Andreas Karageorghis and Timothy N. Phillips. Chebyshev spectral collocation methods for laminar flow through a channel contraction. Journal of Computational Physics, 84(1):114–133, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901848.

Kuo-Petravic:1987:NDM

[KPB87] G. Kuo-Petravic and A. H. Boozer. Numerical determination of the magnetic field line Hamiltonian. Journal of Computational Physics, 73(1):107–124, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901082.

Kuo-Petravic:1983:NEM

[KPBRF83] Gioietta Kuo-Petravic, Allen H. Boozer, James A. Rome, and Ronald H. Fowler. Numerical evaluation of magnetic coordinates for particle transport studies in asymmetric plasmas. Journal of Computational Physics, 51(2):261–272, August 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918390092X.

Kerrell:1981:CEM

[KPD81] A. Mc Kerrell, C. Phillips, and L. M. Delves. Chebyshev expansion methods for the solution of elliptic partial differential equations. Journal of Computational Physics, 41(2):444–452, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181901054.

Kadalbajoo:1984:NSS

[KR84] Mohan K. Kadalbajoo and K. S. Raman. Numerical solution of singular boundary value problems by invariant imbedding. Journal of Computational Physics, 55(2):268–277, August 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900068.

Kadalbajoo:1986:MIB

[KR86] Mohan K. Kadalbajoo and Y. N. Reddy. The method of inner boundary condition: A new approach for solving singular perturbation problems. Journal of Computational Physics, 62(2):349– 360, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901336.

Kress:1989:SIR

[KR89] Michael E. Kress and Kurt S. Riedel. Semi-implicit reduced magnetohydrodynamics. Journal of Computational Physics, 83(1): 237–239, July 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189902301.

Kratky:1980:NBC

[Kra80] K. W. Kratky. New boundary conditions for computer experiments of thermodynamic systems. Journal of Computational Physics, 37(2):205–217, September 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900212.

Krappe:1983:EDF

[Kra83] H. J. Krappe. On the evaluation of double-folding, heavy-ion interaction potentials by Fourier transformation methods. Journal of Computational Physics, 50(3):499–504, June 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901080.

Krasny:1986:DPV

[Kra86] Robert Krasny. Desingularization of periodic vortex sheet rollup. Journal of Computational Physics, 65(2):292–313, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918690210X.

Kress:1988:BAA

[Kre88] Michael E. Kress. Back-averaging: An accelerated iterative method for simulating plasma diffusion. Journal of Computational Physics, 76(1):201–230, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901386.

Krutter:1982:NIT

[Kru82] Harry Krutter. Numerical integration of the Thomas–Fermi equation from zero to infinity. Journal of Computational Physics, 47(2):308–312, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900833.

Kim:1988:ORB

[KRW88] Kwang Y. Kim, Robert O. Reid, and Robert E. Whitaker. On an open radiational boundary condition for weakly dispersive tsunami waves. Journal of Computational Physics, 76(2):327– 348, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901465.

Kirkpatrick:1981:VFS

[KS81] Scott Kirkpatrick and Erich P. Stoll. A very fast shiftregister sequence random number generator. Journal of Computational Physics, 40(2):517–526, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902278.

Kratky:1982:CTS

[KS82] K. W. Kratky and W. Schreiner. Computational techniques for spherical boundary conditions. Journal of Computational Physics, 47(2):313–320, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900845.

Keyes:1987:FSS

[KS87] D. E. Keyes and M. D. Smooke. Flame sheet starting estimates for counterflow diffusion flame problems. Journal of Computational Physics, 73(2):267–288, December 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901379.

Kugelmass:1989:SAL

[KS89] Steven D. Kugelmass and Kenneth Steiglitz. A scalable architecture for lattice-gas simulations. Journal of Computational Physics, 84(2):311–325, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902350.

Kent:1981:CAS

[KSD81] R. D. Kent, M. Schlesinger, and G. W. F. Drake. Calculations of atomic spin-orbit matrix elements in the unitary group approach. Journal of Computational Physics, 40(2):430–439, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902205.

Kubicek:1982:ISD

[KSM82] Milan Kubícek, Ivan Stuchl, and Milos Marek. “Isolas” in solution diagrams. Journal of Computational Physics, 48(1):106– 116, October 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900377.

Kunhardt:1986:MCT

[KT86] E. E. Kunhardt and Y. Tzeng. Monte Carlo technique for simulating the evolution of an assembly of particles increasing in number. Journal of Computational Physics, 67(2):279–289, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186902639.

Kinney:1988:HCA

[KTH88] R. B. Kinney, M. E. Taslim, and S. C. Hung. A hybrid computational approach to multiple-body viscous-flow problems: Application to large-eddy breakup in a boundary layer. Journal of Computational Physics, 78(2):378–404, October 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900563.

Klink:1989:CCG

[KTT89] W. H. Klink and Tuong Ton-That. Calculation of Clebsch– Gordan and Racah coefficients using symbolic manipulation programs. Journal of Computational Physics, 80(2):453–471, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901101.

Kuiken:1988:FMH

[Kui88] H. K. Kuiken. Fin methods of higher order. Journal of Computational Physics, 76(1):1–18, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901283.

Kumagai:1982:NSW

[Kum82] Michio Kumagai. A numerical study of wind-driven circulation in rectangular cavities. Journal of Computational Physics, 47(1): 130–145, July 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900687.

Kunhardt:1985:DSP

[KW85] E. E. Kunhardt and P. F. Williams. Direct solution of Poisson’s equation in cylindrically symmetric geometry: A fast algorithm. Journal of Computational Physics, 57(3):403–414, February 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185901871.

Koelling:1986:IER

[KW86] D. D. Koelling and J. H. Wood. On the interpolation of eigenvalues and a resultant integration scheme. Journal of Computational Physics, 67(2):253–262, December 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902615.

Kunhardt:1987:TMA

[KW87] E. E. Kunhardt and C. Wu. Towards a more accurate flux corrected transport algorithm. Journal of Computational Physics, 68(1):127–150, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900489.

Liu:1983:ENT

[LA83] Ta-Jo Liu and R. C. Ackerberg. An efficient numerical technique for the solution of a nonlinear capillary wave problem. Journal of Computational Physics, 49(2):218–240, February 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901249.

LeQuere:1985:CNC

[LA85] P. Le Quere and T. Alziary De Roquefortt. Computation of natural convection in two-dimensional cavities with Chebyshev polynomials. Journal of Computational Physics, 57(2):210–228, January 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900439.

Labontf:1987:SME

[Lab87] Gilles Labontf. On the solution of matrix equations, example: Application to invariant equations. Journal of Computational Physics, 69(2):420–433, April 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901756.

Lafon:1989:FFT

[Laf89] Earl E. Lafon. A factorizable formulation of tight binding. Journal of Computational Physics, 83(1):185–200, July 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902271.

Lamb:1981:SSP

[Lam81] Robert G. Lamb. A scheme for simulating particle pair motions in turbulent fluid. Journal of Computational Physics, 39(2):329– 346, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901558.

Lanquart:1982:EAA

[Lan82] J.-P Lanquart. Error attenuation in Abel inversion. Journal of Computational Physics, 47(3):434–443, September 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900924.

Larsen:1988:GTA

[Lar88] Edward Larsen. A grey transport acceleration method far time-dependent radiative transfer problems. Journal of Computational Physics, 78(2):459–480, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900605.

Lavery:1988:NSS

[Lav88] John E. Lavery. Nonoscillatory solution of the steady-state inviscid Burgers’ equation by mathematical programming. Journal of Computational Physics, 79(2):436–448, December 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900241.

Lawson:1985:MSV

[Law85] William S. Lawson. A method for simulation of Vlasov systems with discontinuous distribution functions. Journal of Computational Physics, 61(1):51–63, October 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900609.

Lawley:1987:EAS

[Law87] K. P. Lawley. An efficient algorithm for solving the vibronic coupling problem. Journal of Computational Physics, 70(1):218–231, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900118.

Lawson:1988:ACD

[Law88] William S. Lawson. Artificial cooling due to quiet injection in bounded plasma particle simulations. Journal of Computational Physics, 77(2):330–339, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901726.

Lawson:1989:PSB

[Law89] William S. Lawson. Particle simulation of bounded 1D plasma systems. Journal of Computational Physics, 80(2):253–276, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189900995.

Lambrakos:1987:GPM

[LB87] S. G. Lambrakos and J. P. Boris. Geometric properties of the monotonic Lagrangian grid algorithm for near neighbor calculations. Journal of Computational Physics, 73(1):183–202, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187901136.

Lewis:1987:LTA

[LBM87] H. Ralph Lewis, Daniel C. Barnes, and Karl J. Melendez. The Liouville theorem and accurate plasma simulation. Journal of Computational Physics, 69(2):267–282, April 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901690.

Lambrakos:1989:MSA

[LBO^{+}89] S. G. Lambrakos, J. P. Boris, E. S. Oran, I. Chandrasekhar, and M. Nagumo. A modified shake algorithm for maintaining rigid bonds in molecular dynamics simulations of large molecules. Journal of Computational Physics, 85(2):473–486, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901605.

Lanteri:1981:ASM

[LBR81] H. Lanteri, R. Bindi, and P. Rostaing. Application of splitting-up method to the numerical treatment of transport equation. Analysis of the transmission of electrons through thin self-supporting metallic targets. Journal of Computational Physics, 39(1):22– 45, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901352.

Lonsdale:1988:NMA

[LBS88] G. Lonsdale, J. S. Bramley, and D. M. Sloan. A nonlinear multigrid algorithm and boundary-fitted coordinates for the solution of a two-dimensional flow in a branching channel. Journal of Computational Physics, 78(1):1–14, September 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900344.

Lu:1988:EBB

[LC88] Wen-Qiang Lu and Hsueh-Chia Chang. An extension of the biharmonic boundary integral method to free surface flow in channels. Journal of Computational Physics, 77(2):340–360, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901738.

Liang:1989:IUS

[LC89] Shen-Min Liang and Jyh-Jang Chan. An improved upwind scheme for the Euler equations. Journal of Computational Physics, 84(2):461–473, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990243X.

Lynch:1986:SEM

[LCC^{+}86] V. E. Lynch, B. A. Carreras, L. A. Charlton, L. Garcia, T. C. Hender, H. R. Hicks, and J. A. Holmes. Stellarator expansion methods for MHD equilibrium and stability calculations. Journal of Computational Physics, 66(2):411–444, October 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900732.

Langdon:1983:DIL

[LCF83] A. Bruce Langdon, Bruce I. Cohen, and Alex Friedman. Direct implicit large time-step particle simulation of plasmas. Journal of Computational Physics, 51(1):107–138, July 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900839.

Liewer:1989:GCA

[LD89] Paulett C. Liewer and Viktor K. Decyk. A general concurrent algorithm for plasma particle-in-cell simulation codes. Journal of Computational Physics, 85(2):302–322, December 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901538.

Lee:1987:GPS

[Lee87] W. W. Lee. Gyrokinetic particle simulation model. Journal of Computational Physics, 72(1):243–269, September 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900805.

Leonard:1980:VMF

[Leo80] A. Leonard. Vortex methods for flow simulation. Journal of Computational Physics, 37(3):289–335, October 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900406.

Leventhal:1982:OCI

[Lev82] Stephen H. Leventhal. An operator compact implicit method of exponential type. Journal of Computational Physics, 46(1):138– 165, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900080.

Leveque:1988:HRF

[Lev88] Randall J. Leveque. High resolution finite volume methods on arbitrary grids via wave propagation. Journal of Computational Physics, 78(1):36–63, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900368.

Lynch:1980:FES

[LG80] Daniel R. Lynch and William G. Gray. Finite element simulation of flow in deforming regions. Journal of Computational Physics, 36(2):135–153, July 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901801.

Leone:1981:FES

[LG81] John M. Leone, Jr. and Philip M. Gresho. Finite element simulations of steady, two-dimensional, viscous incompressible flow over a step. Journal of Computational Physics, 41(1):167–191, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900863.

Livne:1985:FDS

[LG85] E. Livne and A. Glasner. A finite difference scheme for the heat conduction equation. Journal of Computational Physics, 58(1): 59–66, March 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901561.

Li:1980:CPS

[Li80] C. P. Li. A computational procedure for supersonic flows governed by the parabolic Navier–Stokes equations. Journal of Computational Physics, 35(3):356–380, May 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900753.

Li:1989:NCS

[Li89] Zi-Cai Li. A nonconforming combination for solving elliptic problems with interfaces. Journal of Computational Physics, 80(2):288–313, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189901010.

Lim:1980:IEM

[Lim80] T. H. Lim. An integral equation method for the numerical solution of a two-dimensional vertical jet under gravity. Journal of Computational Physics, 35(1):90–99, March 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900364.

Lohar:1981:VMC

[LJ81] B. L. Lohar and P. C. Jain. Variable mesh cubic spline technique for N-wave solution of Burgers’ equation. Journal of Computational Physics, 39(2):433–442, February 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901625.

Ling:1985:PSE

[LJ85] K. M. Ling and S. C. Jardin. The Princeton Spectral Equilibrium Code: PSEC. Journal of Computational Physics, 58(3):300– 335, May 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901652.

Laitone:1981:ENI

[LL81] J. A. Laitone and E. V. Laitone. Errors in the numerical integration of a linearized equation for particle flows. Journal of Computational Physics, 43(2):394–398, October 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901315.

Lavergnat:1988:NBP

[LLL88] J. Lavergnat, Th. Lehner, and A. Y. Le Roux. Nonlinear beam plasma interaction using finite difference methods. Journal of Computational Physics, 79(2):298–316, December 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900186.

Larsen:1985:DMO

[LLPS85] E. W. Larsen, C. D. Levermore, G. C. Pomraning, and J. G. Sanderson. Discretization methods for one-dimensional Fokker– Planck operators. Journal of Computational Physics, 61(3):359– 390, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900701.

Laloyaux:1989:ROD

[LLVL89] Th. Laloyaux, Ph. Lambin, J.-P Vigneron, and A. A. Lucas. Resolution of the one-dimensional scattering problem by a finite element method. Journal of Computational Physics, 83(2):398– 406, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901265.

Loewenstein:1986:APH

[LM86] Michael Loewenstein and William G. Mathews. Adiabatic particle hydrodynamics in three dimensions. Journal of Computational Physics, 62(2):414–428, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901373.

Ladera:1987:TCA

[LM87a] Celso L. Ladera and Pablo Martin. Transmittance of a circular aperture by an integrable fractional-like approximation to J_{0}(x) function. Journal of Computational Physics, 73(2):481– 489, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901501.

Larsen:1987:AMC

[LM87b] Edward W. Larsen and Bertrand Mercer. Analysis of a Monte Carlo method for nonlinear radiative transfer. Journal of Computational Physics, 71(1):50–64, July 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900192.

Larsen:1989:ASN

[LM89a] Edward W. Larsen and J. E. Morel. Asymptotic solutions of numerical transport problems in optically thick, diffusive regimes II. Journal of Computational Physics, 83(1):212–236, July 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902295.

Lee:1989:MFE

[LM89b] Robert L. Lee and Niel K. Madsen. A mixed finite element formulation for Maxwell’s equations in the time domain. Journal of Computational Physics, 85(2):503, December 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990168X.

Larsen:1987:ASN

[LMM87] Edward W. Larsen, J. E. Morel, and Warren F. Miller, Jr. Asymptotic solutions of numerical transport problems in optically thick, diffusive regimes. Journal of Computational Physics, 69(2):283– 324, April 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901707.

Lotz:1987:NEC

[LNS87] W. Lotz, J. Nührenberg, and A. Schlüter. Neoclassical energy confinement in stellarators. Journal of Computational Physics, 73(1):73–84, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187901069.

Lo:1985:AMS

[Lo85] Lilian L. Lo. Asymptotic matching by the symbolic manipulator MACSYMA. Journal of Computational Physics, 61(1):38– 50, October 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900592.

Leacock:1986:AVT

[LO86] Robert A. Leacock and Patrick W. O’Connor. Actionvariable theory and classical frequencies. Journal of Computational Physics, 62(1):164–179, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901051.

Liu:1987:TMA

[LO87] K. M. Liu and E. L. Ortiz. Tau method approximation of differential eigenvalue problems where the spectral parameter enters nonlinearly. Journal of Computational Physics, 72(2):299– 310, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900854.

Longman:1980:WPV

[Lon80] I. M. Longman. Wave propagation in a viscoelastic solid. Journal of Computational Physics, 37(2):171–182, September 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900194.

Lonsdale:1988:SRN

[Lon88] Guy Lonsdale. Solution of a rotating Navier–Stokes problem by a nonlinear multigrid algorithm. Journal of Computational Physics, 74(1):177–190, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900757.

Lotstedt:1982:FTM

[Löt82] Per Lötstedt. A front tracking method applied to Burgers’ equation and two-phase porous flow. Journal of Computational Physics, 47(2):211–228, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900754.

Larsen:1983:SMC

[LP83] Jan Emil Larsen and J. Boiden Pedersen. A simple method for calculating diffusional recombination probabilities. Journal of Computational Physics, 49(3):458–464, March 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901407.

Lavergnat:1980:NAD

[LPS80] J. Lavergnat, J. Portes, and M. Sylvain. A new algorithm to determine the parameters of a sinusoidal signal. Journal of Computational Physics, 36(3):347–365, July 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901655.

Lynch:1985:FES

[LPS85] Daniel R. Lynch, Keith D. Paulsen, and John W. Strohbehn. Finite element solution of Maxwell’s equations for hyperthermia treatment planning. Journal of Computational Physics, 58(2): 246–269, April 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901792.

Lugannani:1980:EBF

[LR80a] Robert Lugannani and Stephen Rice. Expressions for the behavior of a Fourier transform near its singularities. Journal of Computational Physics, 37(2):259–263, September 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900248.

Lugannani:1980:UGC

[LR80b] Robert Lugannani and Stephen Rice. Use of Gaussian convergence factors in the numerical evaluation of slowly converging integrals. Journal of Computational Physics, 37(2):264–267, September 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918090025X.

Lugannani:1981:NEC

[LR81] R. Lugannani and S. O. Rice. Numerical evaluation of a class of integrals by integrating along a string of saddle points. Journal of Computational Physics, 39(2):473–480, February 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901650.

Lynch:1985:HCD

[LS85] Daniel R. Lynch and John M. Sullivan, Jr. Heat conservation in deforming element phase change simulation. Journal of Computational Physics, 57(2):303–317, January 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900476.

Lee:1987:ETP

[LS87] Tai-Yong Lee and John H. Seinfeld. Estimation of two-phase petroleum reservoir properties by regularization. Journal of Computational Physics, 69(2):397–419, April 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901744.

Lubachevsky:1988:EPS

[Lub88] Boris D. Lubachevsky. Efficient parallel simulations of dynamic Ising spin systems. Journal of Computational Physics, 75(1):103– 122, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901015.

Luchini:1987:AMF

[Luc87] Paolo Luchini. An adaptive-mesh finite-difference solution method for the Navier–Stokes equations. Journal of Computational Physics, 68(2):283–306, February 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900593.

Luppi:1987:OQS

[Lup87] Jussi Luppi. An optimized quadrature scheme for evaluating general higher-order phase integrals. Journal of Computational Physics, 71(1):111–123, July 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900222.

LeProvost:1986:STP

[LV86] C. Le Provost and P. Vincent. Some tests of precision for a finite element model of ocean tides. Journal of Computational Physics, 65(2):273–291, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902093.

Liu:1989:NFC

[LV89] Yen Liu and Marcel Vinokur. Nonequilibrium flow computations. I. An analysis of numerical formulations of conservation laws. Journal of Computational Physics, 83(2):373–397, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901253.

Liou:1989:SIF

[LvLS89] Meng-Sing Liou, Bran van Leer, and Jian-Shun Shuen. Splitting of inviscid fluxes for real gases. Journal of Computational Physics, 84(1):256, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901952.

Lee:1985:CVM

[LW85] Clarence E. Lee and Bruce C. Wilson. A conservative variational method for multicomponent concentration time dependent diffusion. Journal of Computational Physics, 60(1):1–22, August 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900142.

Leroy:1986:ERN

[LW86] J. P. Leroy and R. Wallace. Extension of the renormalized Numerov method for second-order differential eigenvalue equations. Journal of Computational Physics, 67(2):239–252, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186902603.

Levine:1989:EMM

[LW89] Zachary H. Levine and John W. Wilkins. An energyminimizing mesh for the Schrödinger equation. Journal of Computational Physics, 83(2):361–372, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901241.

Li:1989:NMC

[LWP89] Ming Li, Lawrence Wilets, and Robert J. Perry. Numerical methods in calculating boson and fermion loop corrections. Journal of Computational Physics, 85(2):457–472, December 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901599.

Lynch:1982:UAS

[Lyn82] Daniel R. Lynch. Unified approach to simulation on deforming elements with application to phase change problems. Journal of Computational Physics, 47(3):387–411, September 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900900.

Lyness:1984:CTF

[Lyn84] J. N. Lyness. The calculation of trigonometric Fourier coefficients. Journal of Computational Physics, 54(1):57–73, April 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901384.

Lytton:1987:SEE

[Lyt87] C. C. Lytton. Solution of the Euler equations for transonic flow over a lifting aerofoil-the Bernoulli formulation (Roe/Lytton Method). Journal of Computational Physics, 73(2):395–431, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901446.

McDonald:1984:HOU

[MA84] B. Edward McDonald and John Ambrosiano. High-order upwind flux correction methods for hyperbolic conservation laws. Journal of Computational Physics, 56(3):448–460, December 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901062.

Mackrodt:1981:LSL

[Mac81] W. C. Mackrodt. Lattice sums for long-range interactions involving charged defects in non-cubic ionic crystals. Journal of Computational Physics, 44(1):206–211, November 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900462.

Macaraeg:1986:MPF

[Mac86] M. G. Macaraeg. A mixed pseudospectral/finite difference method for the axisymmetric flow in a heated, rotating spherical shell. Journal of Computational Physics, 62(2):297–320, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186901294.

Macrossan:1989:EFM

[Mac89] M. N. Macrossan. The equilibrium flux method for the calculation of flows with non-equilibrium chemical reactions. Journal of Computational Physics, 80(1):204–231, January 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900958.

Mahaffy:1982:SET

[Mah82] John H. Mahaffy. A stability-enhancing two-step method for fluid flow calculations. Journal of Computational Physics, 46(3):329– 341, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900195.

Makino:1989:CTD

[Mak89] Junichiro Makino. Comparison of two different tree algorithms. Journal of Computational Physics, 85(2):504, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901733.

Mann:1985:SMF

[Man85] Patrick J. Mann. Some mixed finite element-finite difference methods for spherically symmetric relativistic collapse. Journal of Computational Physics, 58(3):377–394, May 15, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918590169X.

Manickam:1986:ASR

[Man86] J. Manickam. Accessibility of second regions of stability in tokamaks. Journal of Computational Physics, 66(2):324–331, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900690.

Mann:1987:NFE

[Man87] Patrick J. Mann. A new finite element method for spherically symmetric relativistic collapse. Journal of Computational Physics, 72(2):467–485, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900945.

Martins:1981:SCF

[Mar81] Pedro A. P. Martins. Spherical Coulomb functions: Recurrence relations and continued fractions. Journal of Computational Physics, 41(1):223–230, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181900929.

Marder:1987:MIG

[Mar87] Barry Marder. A method for incorporating Gauss’ law into electromagnetic PIC codes. Journal of Computational Physics, 68(1): 48–55, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918790043X.

Mason:1981:IMP

[Mas81] Rodney J. Mason. Implicit moment particle simulation of plasmas. Journal of Computational Physics, 41(2):233–244, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900942.

Mason:1983:IMP

[Mas83] Rodney J. Mason. Implicit moment PIC-hybrid simulation of collisional plasmas. Journal of Computational Physics, 51(3):484– 501, September 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901651.

Mason:1987:EFA

[Mas87] Rodney J. Mason. An electromagnetic field algorithm for 2D implicit plasma simulation. Journal of Computational Physics, 71(2):429–473, August 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187900386.

Mattioli:1986:IFE

[Mat86] F. Mattioli. An improvement of the finite-element seriesexpansion technique for linear water waves. Journal of Computational Physics, 66(1):50–61, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900537.

Mavriplis:1989:LPI

[Mav89] Cathy Mavriplis. Laguerre polynomials for infinite-domain spectral elements. Journal of Computational Physics, 80(2):480–488, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901125.

Miranker:1980:FRA

[MB80] W. L. Miranker and K. Bu. A functional relation and an acceleration procedure for calculating the voltage response of Josephson junctions. Journal of Computational Physics, 37(2):274–279, September 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180900273.

Morrow:1985:FCT

[MC85] R. Morrow and L. E. Cram. Flux-corrected transport and diffusion on a non-uniform mesh. Journal of Computational Physics, 57(1):129–136, January 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900555.

MacKinnon:1988:AMI

[MC88a] R. J. MacKinnon and G. F. Carey. Analysis of material interface discontinuities and superconvergent fluxes in finite difference theory. Journal of Computational Physics, 75(1):151–167, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901040.

Murray:1988:FEA

[MC88b] Paul Murray and G. F. Carey. Finite element analysis of diffusion with reaction at a moving boundary. Journal of Computational Physics, 74(2):440–455, February 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900873.

McDonald:1980:CMS

[McD80] B. E. McDonald. The Chebychev method for solving nonselfadjoint elliptic equations on a vector computer. Journal of Computational Physics, 35(2):147–168, April 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900832.

McDonald:1989:FCP

[McD89] B. E. McDonald. Flux-corrected pseudospectral method for scalar hyperbolic conservation laws. Journal of Computational Physics, 82(2):413–428, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900569.

McLaughlin:1980:NSU

[McL80] R. McLaughlin. The numerical solution of the unsteady expansion of a gas into a near vacuum. Journal of Computational Physics, 35(1):77–89, March 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900352.

McNamara:1987:SES

[McN87] Brendan McNamara. Supercomputer environments for science applications. Journal of Computational Physics, 73(1):41–58, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901045.

Mehlhorn:1980:DOS

[MD80] Thomas A. Mehlhorn and James J. Duderstadt. A discrete ordinates solution of the Fokker–Planck equation characterizing charged particle transport. Journal of Computational Physics, 38(1):86–106, November 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900133.

McMurchie:1981:CII

[MD81a] Larry E. McMurchie and Ernest R. Davidson. Calculation of integrals over ab initio pseudopotentials. Journal of Computational Physics, 44(2):289–301, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190053X.

Mercier:1981:MCH

[MD81b] Philippe Mercier and Michel Deville. A multidimensional compact higher-order scheme for 3-D Poisson’s equation. Journal of Computational Physics, 39(2):443–455, February 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901637.

Mullins:1985:CSS

[MD85] M. Mullins and M. A. Dokainish. Comparison of solution schemes for atomic simulation problems. Journal of Computational Physics, 61(2):242–260, November 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900841.

Miller:1986:GMD

[MD86] A. D. Miller and R. L. Dewar. Galerkin method for differential equations with regular singular points. Journal of Computational Physics, 66(2):356–390, October 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900719.

Matsuno:1988:AME

[MD88] Kenichi Matsuno and Harry A. Dwyer. Adaptive methods for elliptic grid generation. Journal of Computational Physics, 77 (1):40–52, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901556.

Medvedev:1986:ATD

[Med86] N. N. Medvedev. The algorithm for three-dimensional Voronoi polyhedra. Journal of Computational Physics, 67(1):223–229, November 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901233.

Mains:1985:FDN

[MEGH85] R. K. Mains, M. A. El-Gabaly, and G. I. Haddad. Finitedifference numerical methods for solving the energy-momentum transport equations in two-valley semiconductors. Journal of Computational Physics, 59(3):456–467, July 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901226.

Meiburg:1989:ITD

[Mei89] E. Meiburg. Incorporation and test of diffusion and strain effects in the two-dimensional vortex blob technique. Journal of Computational Physics, 82(1):85–93, May 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900363.

Mejia:1986:CCP

[Mej86] R. Mejia. CONKUB: A conversational path-follower for systems of nonlinear equations. Journal of Computational Physics, 63(1): 67–84, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900847.

Marshall:1986:GCT

[MEK86] Guillermo Marshall, Peter Eiseman, and John T. Kuo. A general collapsing technique for three-dimensional algebraic grid generation. Journal of Computational Physics, 62(1):180–198, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186901063.

Melezhik:1986:CAN

[Mel86] V. S. Melezhik. Continuous analog of Newton method in the multichannel scattering problem. Journal of Computational Physics, 65(1):1–17, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918690001X.

Melia:1988:NTS

[Mel88] Fulvio Melia. A new technique for solving Parker-type wind equations. Journal of Computational Physics, 74(2):488–492, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900903.

Merz:1980:DAE

[Mer80] Paul H. Merz. Determination of adsorption energy distribution by regularization and a characterization of certain adsorption isotherms. Journal of Computational Physics, 38(1):64–85, November 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180900121.

Merkel:1986:IET

[Mer86] Peter Merkel. An integral equation technique for the exterior and interior Neumann problem in toroidal regions. Journal of Computational Physics, 66(1):83–98, September 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900550.

Malik:1980:QBS

[MES80] David J. Malik, Joseph Eccles, and Don Secrest. On quantal bound state solutions and potential energy surface fitting. A comparison of the finite element, Numerov–Cooley and finite difference methods. Journal of Computational Physics, 38(2):157–184, November 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180900510.

Meyer:1981:HSF

[Mey81] Gunter H. Meyer. Hele–Shaw flow with a cusping free boundary. Journal of Computational Physics, 44(2):262–276, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900516.

Meyer:1986:COB

[Mey86] Gunter H. Meyer. Continuous orthonormalization for boundary value problems. Journal of Computational Physics, 62(1):248– 262, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901099.

Mezei:1981:SPP

[Mez81] Mihaly Mezei. On the selection of the particle to be perturbed in the Monte Carlo method. Journal of Computational Physics, 39(1):128–136, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901418.

Mezei:1987:AUS

[Mez87] Mihaly Mezei. Adaptive umbrella sampling: Self-consistent determination of the non-Boltzmann bias. Journal of Computational Physics, 68(1):237–248, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900544.

Meson:1985:SCO

[MFC^{+}85] Alejandro M. Mesón, Francisco M. Fernández, Eduardo A. Castro, Alberto Maltz, and Rodolfo Rodriguez. Shift of the coordinate origin in calculating resonances by dilatation transformation. Journal of Computational Physics, 57(3):491–494, February 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185901925.

Miller:1988:FPE

[MFW88] G. Miller, V. Faber, and A. B. White, Jr. Finding plasma equilibria with magnetic islands. Journal of Computational Physics, 79(2):417–435, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918890023X.

Miller:1989:REM

[MFW89] G. Miller, V. Faber, and A. B. White, Jr. Resistive evolution of magnetic fields in plasmas. Journal of Computational Physics, 85(2):323–341, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990154X.

Monaghan:1983:SSP

[MG83] J. J. Monaghan and R. A. Gingold. Shock simulation by the particle method SPH. Journal of Computational Physics, 52(2):374– 389, November 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183900360.

Martin:1989:TPQ

[MG89] Pablo Martin and Antonio Luis Guerrero. Two-point quasifractional approximations to the Bessel function J_{ν}(x) of fractional order. Journal of Computational Physics, 85(2):487–492, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901617.

Mcisaac:1986:CFE

[MGM86] K. Mcisaac, J. E. Gottschalk, and E. N. Maslen. Closed form expressions for an integral involving the Coulomb potential. Journal of Computational Physics, 67(2):479–481, December 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902731.

McRae:1982:NSA

[MGS82] Gregory J. McRae, William R. Goodin, and John H. Seinfeld. Numerical solution of the atmospheric diffusion equation for chemically reacting flows. Journal of Computational Physics, 45(1): 1–42, January 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182901012.

Ma:1989:NSO

[MH89] Yamin Ma and John R. Hiller. Numerical solution of the one-pair equation in the massive Schwinger model. Journal of Computational Physics, 82(1):229–240, May 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900442.

Maurer:1986:QAM

[MHK86] M. Maurer, A. Hayd, and H. J. Kaeppeler. Quasi-analytical method for solving nonlinear differential equations for turbulent self-confined magneto-plasma. Journal of Computational Physics, 66(1):151–172, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900586.

Memon:1981:MDC

[MHM81] M. K. Memon, R. W. Hockney, and S. K. Mitra. Molecular dynamics with constraints. Journal of Computational Physics, 43(2):345–356, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181901273.

Mihalas:1985:CRT

[Mih85] Dimitri Mihalas. The computation of radiation transport using Feautrier variables I. Static media. Journal of Computational Physics, 57(1):1–25, January 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900506.

Manohar:1988:HOD

[MIK88] R. Manohar, S. R. K. Iyengar, and U. A. Krishnaiah. High order difference methods for linear variable coefficient parabolic equation. Journal of Computational Physics, 77(2):513–523, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901817.

Milgram:1989:DPL

[Mil89a] M. S. Milgram. Does a point lie inside a polygon? Journal of Computational Physics, 84(1):134–144, September 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990185X.

Miller:1989:UMV

[Mil89b] Alan J. Miller. Updating means and variances. Journal of Computational Physics, 85(2):500–501, December 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901630.

Miyata:1986:FDS

[Miy86] Hideaki Miyata. Finite-difference simulation of breaking waves. Journal of Computational Physics, 65(1):179–214, July 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900112.

Modrak:1981:NCD

[MJ81] P. Modrak and B. Jasiński. On the numerical calculations of the density of states in tight-binding approximation. Journal of Computational Physics, 40(1):250–253, March 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902084.

Min:1987:WNI

[MJ87] Han Min and Yih Nen Jeng. The weak nonlinear instability of Euler explicit scheme for the convective equation. Journal of Computational Physics, 68(2):251–261, February 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900568.

Moin:1980:NST

[MK80] P. Moin and J. Kim. On the numerical solution of time-dependent viscous incompressible fluid flows involving solid boundaries. Journal of Computational Physics, 35(3):381–392, May 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900765.

Mihalas:1982:STD

[MK82] Dimitri Mihalas and Richard I. Klein. On the solution of the timedependent inertial-frame equation of radiative transfer in moving media to O(v/c). Journal of Computational Physics, 46(1):97– 137, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999182900079.

Marchuk:1983:CFE

[MK83] G. I. Marchuk and V. I. Kuzin. On the combination of finite element and splitting-up methods in the solution of parabolic equations. Journal of Computational Physics, 52(2):237–272, November 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918390030X.

Mihalas:1986:CRT

[MK86] Dimitri Mihalas and Paul B. Kunasz. The computation of radiation transport using Feautrier variables. II. Spectrum line formation in moving media. Journal of Computational Physics, 64 (1):1–26, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900161.

Margoliash:1983:CSFa

[ML83a] D. J. Margoliash and P. W. Langhoff. Computational studies of first-born scattering cross sections I. Spectral properties of Bethe surfaces. Journal of Computational Physics, 49(1):44–66, January 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901146.

Margoliash:1983:CSFb

[ML83b] D. J. Margoliash and P. W. Langhoff. Computational studies of first-born scattering cross sections II. Moment-theory approach. Journal of Computational Physics, 49(1):67–93, January 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901158.

Milgram:1985:TDF

[ML85] M. S. Milgram and H. C. Lee. Tables of divergent Feynman integrals in the axial and light-cone gauges. Journal of Computational Physics, 59(2):331–346, June 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901500.

Milgram:1987:TDF

[ML87] M. S. Milgram and H. C. Lee. Tables of divergent Feynman integrals. II. Light-cone gauge with the Mandelstam–Leibbrandt prescription. Journal of Computational Physics, 71(2):316–323, August 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900337.

Mary:1987:MWA

[MLB87] Donald F. St Mary, Ding Lee, and George Botseas. A modified wide angle parabolic wave equation. Journal of Computational Physics, 71(2):304–315, August 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900325.

Marshall:1981:CAR

[MM81a] Guillermo Marshall and Raúl Méndez. Computational aspects of the random choice method for shallow water equations. Journal of Computational Physics, 39(1):1–21, January 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901340.

Marshall:1981:NTN

[MM81b] Guillermo Marshall and Angel N. Mentndez. Numerical treatment of nonconservation forms of the equations of shallow water theory. Journal of Computational Physics, 44(1):167–188, November 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181900437.

Monkewitz:1981:ASO

[MM81c] Peter A. Monkewitz and Martin A. Monkewitz. The asymptotic solution of the Orr–Sommerfeld equation at large distance from a shear layer. Journal of Computational Physics, 41(2):245–255, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900954.

Meintjes:1985:PAC

[MM85a] Keith Meintjes and Alexander P. Morgan. Performance of algorithms for calculating the equilibrium composition of a mixture of gases. Journal of Computational Physics, 60(2):219–234, September 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918590004X.

Mikic:1985:UPB

[MM85b] Zoran Mikić and Edward C. Morse. The use of a preconditioned bi-conjugate gradient method for hybrid plasma stability analysis. Journal of Computational Physics, 61(1):154–185, October 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918590066X.

Martin:1988:NMF

[MM88] Carlos A. Martin and Gustavo Monti. Nonlinear-multiple-function simultaneous least squares fitting procedure. Journal of Computational Physics, 75(1):244–250, March 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901106.

Mandel:1989:MVM

[MM89] Jan Mandel and Steve McCormick. A multilevel variational method for Au = λBu on composite grids. Journal of Computational Physics, 80(2):442–452, February 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901095.

Moser:1983:SNM

[MML83] R. D. Moser, P. Moin, and A. Leonard. A spectral numerical method for the Navier–Stokes equations with applications to Taylor–Couette flow. Journal of Computational Physics, 52(3):524–544, December 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900062.

Moriarty:1989:DIS

[MMR89] K. J. M. Moriarty, Eric Myers, and Claudio Rebbi. Dynamical interactions of superconducting flux vortices. Journal of Computational Physics, 81(2):481–488, April 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902179.

Martinez:1986:NSN

[MMS86] M. C. Martinez, V. Muñoz, and V. Such. A numerical solution for the nonlinear diffusion equation of the electromagnetic field in ferromagnetic materials. Journal of Computational Physics, 65(2):432–447, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902160.

Maino:1981:CPC

[MMV81] G. Maino, E. Menapace, and A. Ventura. Computation of parabolic cylinder functions by means of a Tricomi expansion. Journal of Computational Physics, 40(2):294–304, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902114.

Miyata:1985:FDS

[MNM85] Hideaki Miyata, Shinichi Nishimura, and Akira Masuko. Finite difference simulation of nonlinear waves generated by ships of arbitrary three-dimensional configuration. Journal of Computational Physics, 60(3):391–436, September 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900282.

McCrory:1980:SMM

[MO80] Robert L. McCrory and Steven A. Orszag. Spectral methods for multi-dimensional diffusion problems. Journal of Computational Physics, 37(1):93–112, August 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900066.

Matsuo:1988:NSC

[MOFK88] H. Matsuo, Y. Ohya, K. Fujiwara, and H. Kudoh. Numerical simulation of cylindrically converging shock waves. Journal of Computational Physics, 75(2):384–399, April 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901192.

Mohamed:1984:VMA

[Moh84] J. L. Mohamed. De Vogelaere’s method for the automatic solution of systems of coupled homogeneous second-order differential equations. Journal of Computational Physics, 54(3):457– 467, June 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918490127X.

Meiron:1981:ANC

[MOI81] Daniel I. Meiron, Steven A. Orszag, and Moshe Israeli. Applications of numerical conformal mapping. Journal of Computational Physics, 40(2):345–360, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902151.

Monaghan:1981:NLP

[Mon81] J. J. Monaghan. A note on Larson’s particle method. Journal of Computational Physics, 44(2):397–399, December 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900607.

Monaghan:1985:ESI

[Mon85] J. J. Monaghan. Extrapolating B splines for interpolation. Journal of Computational Physics, 60(2):253–262, September 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900063.

Monaghan:1989:PPP

[Mon89] J. J. Monaghan. On the problem of penetration in particle methods. Journal of Computational Physics, 82(1):1–15, May 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900326.

Morrow:1981:NSH

[Mor81] R. Morrow. Numerical solution of hyperbolic equations for electron drift in strongly non-uniform electric fields. Journal of Computational Physics, 43(1):1–15, September 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190108X.

Morrow:1982:SCE

[Mor82] R. Morrow. Space-charge effects in high-density plasmas. Journal of Computational Physics, 46(3):454–461, June 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900262.

Morishima:1984:SET

[Mor84] Nobuhiro Morishima. Spectral estimation through cubic-spline approximation of a discrete time series. Journal of Computational Physics, 56(1):93–110, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900858.

Morgan:1989:DMP

[Mor89] Ronald B. Morgan. Davidson’s methods and preconditioning for generalized eigenvalue problems. Journal of Computational Physics, 85(2):504, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901769.

Mosher:1985:VNF

[Mos85] M. Calvin Mosher. A variable node finite element method. Journal of Computational Physics, 57(2):157–187, January 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900415.

Moss:1986:CIE

[Mos86] William F. Moss. Collocation for an integral equation arising in duct acoustics. Journal of Computational Physics, 64(2):443– 458, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900434.

McCracken:1980:VMB

[MP80a] M. F. McCracken and C. S. Peskin. A vortex method for blood flow through heart valves. Journal of Computational Physics, 35(2):183–205, April 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900856.

Morton:1980:GGM

[MP80b] K. W. Morton and A. K. Parrott. Generalised Galerkin methods for first-order hyperbolic equations. Journal of Computational Physics, 36(2):249–270, July 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901874.

Marshall:1984:RCM

[MP84] Guillermo Marshall and Bradley Plohr. A random choice method for two-dimensional steady supersonic shock wave diffraction problems. Journal of Computational Physics, 56(3):410–427, December 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901049.

McQueen:1989:TDC

[MP89a] David M. McQueen and Charles S. Peskin. A three-dimensional computational method for blood flow in the heart. II. contractile fibers. Journal of Computational Physics, 82(2):289–297, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900508.

Morton:1989:FVS

[MP89b] K. W. Morton and M. F. Paisley. A finite volume scheme with shock fitting for the steady Euler equations. Journal of Computational Physics, 80(1):168–203, January 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900946.

Mafe:1988:NAI

[MPA88] S. Mafé, J. Pellicer, and V. M. Aguilella. A numerical approach to ionic transport through charged membranes. Journal of Computational Physics, 75(1):1–14, March 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900952.

Melezhik:1984:NSS

[MPPS84] V. S. Melezhik, I. V. Puzynin, T. P. Puzynina, and L. N. Somov. Numerical solution of a system of integrodifferential equations arising from the quantum mechanical three-body problem with Coulomb interaction. Journal of Computational Physics, 54(2): 221–236, May 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901153.

Miller:1983:BQO

[MRH83] Robert N. Miller, Allan R. Robinson, and Dale B. Haidvogel. A baroclinic quasigeostrophic open ocean model. Journal of Computational Physics, 50(1):38–70, April 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900414.

Mrugala:1985:LDM

[Mru85] Felicja Mrugala. Log-derivative method for two-potential scattering problems. Journal of Computational Physics, 58(1):113– 133, March 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901603.

Mrugala:1987:HAS

[Mru87] Felicja Mrugala. Hybrid approximate solution–approximate potential approach to the solution of coupled equations for atom–molecule reactive scattering. Journal of Computational Physics, 68(2):393–419, February 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900647.

Mobley:1980:NGB

[MS80] Curtis D. Mobley and Robert J. Stewart. On the numerical generation of boundary-fitted orthogonal curvilinear coordinate systems. Journal of Computational Physics, 34(1):124–135, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180901175.

Maier:1981:NSS

[MS81] Maximilian R. Maier and Donald R. Smith. Numerical solution of a symmetric one-dimensional diode model. Journal of Computational Physics, 42(2):309–326, August 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902473.

Manohar:1983:SCH

[MS83] Ram P. Manohar and John W. Stephenson. Single cell high order difference methods for the Helmholtz equation. Journal of Computational Physics, 51(3):444–453, September 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901638.

Matsuda:1986:RMB

[MS86] Y. Matsuda and J. J. Stewart. A relativistic multiregion bounceaveraged Fokker–Planck code for mirror plasmas. Journal of Computational Physics, 66(1):197–217, September 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900604.

Maritz:1987:PWS

[MS87a] M. F. Maritz and S. W. Schoombie. Parasitic waves and solitons in the numerical solution of the Korteweg–de Vries and modified Korteweg–de Vries equation. Journal of Computational Physics, 73(2):244–266, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901367.

Morton:1987:CFL

[MS87b] K. W. Morton and P. K. Sweby. A comparison of flux limited difference methods and characteristic Galerkin methods for shock modelling. Journal of Computational Physics, 73(1):203–230, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901148.

Miyata:1987:DSV

[MSB87] Hideaki Miyata, Toru Sato, and Nobuhiro Baba. Difference solution of a viscous flow with free-surface wave about an advancing ship. Journal of Computational Physics, 72(2):393–421, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900908.

Mankofsky:1987:HSI

[MSD87] A. Mankofsky, R. N. Sudan, and J. Denavit. Hybrid simulation of ion beams in background plasma. Journal of Computational Physics, 70(1):89–116, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900039.

Meyer-Spasche:1980:CAF

[MSK80] Rita Meyer-Spasche and Herbert B. Keller. Computations of the axisymmetric flow between rotating cylinders. Journal of Computational Physics, 35(1):100–109, March 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900376.

Miller:1987:TSI

[MSW87] H. G. Miller, M. C. Steenkamp, and C. J. Wright. A two-step iterative block Lanczos algorithm for a dominant eigenspan. Journal of Computational Physics, 72(2):383–392, October 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900891.

Miao:1981:FIS

[MT81] Charles C. Miao and T. G. Theofanous. A fast ICE solution procedure for flows with largely invariant compressibility. Journal of Computational Physics, 40(1):254–261, March 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902096.

Mayers:1984:ERO

[MT84a] D. F. Mayers and D. M. Turner. The evaluation of relativistic oscillatory wavefunctions on a logarithmic grid. Journal of Computational Physics, 55(3):397–407, September 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900299.

Miki:1984:DDO

[MT84b] Kazuyoshi Miki and Toshiyuki Takagi. A domain decomposition and overlapping method for the generation of threedimensional boundary-fitted coordinate systems. Journal of Computational Physics, 53(2):319–330, February 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900445.

Mayers:1986:RIS

[MT86a] D. F. Mayers and D. M. Turner. Relativistic inelastic scattering using a logarithmic grid. Journal of Computational Physics, 62(2):331–339, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901312.

Miki:1986:NSP

[MT86b] Kazuyoshi Miki and Toshiyuki Takagi. Numerical solution of Poisson’s equation with arbitrarily shaped boundaries using a domain decomposition and overlapping technique. Journal of Computational Physics, 67(2):263–278, December 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902627.

Myers:1981:PSS

[MTM81] Robert B. Myers, Thomas D. Taylor, and John W. Murdock. Pseudo-spectral simulation of a two-dimensional vortex flow in a stratified, incompressible fluid. Journal of Computational Physics, 43(1):180–188, September 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901170.

Mulder:1985:MRE

[Mul85] Wim A. Mulder. Multigrid relaxation for the Euler equations. Journal of Computational Physics, 60(2):235–252, September 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900051.

Muller:1988:FVS

[Mül88] Ewald Müller. Flux-vector splitting for the Euler equations for real gases. Journal of Computational Physics, 79(1):227–230, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900137.

Mulder:1989:NMA

[Mul89] Wim A. Mulder. A new multigrid approach to convection problems. Journal of Computational Physics, 83(2):303–323, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901216.

Munro:1981:CCL

[Mun81] R. G. Munro. Criteria for choosing a linearized least-squares technique for the exponential model exp(A_{1} + A_{2}X). Journal of Computational Physics, 44(1):189–197, November 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900449.

Munz:1988:NDH

[Mun88] Claus-Dieter Munz. On the numerical dissipation of high resolution schemes for hyperbolic conservation laws. Journal of Computational Physics, 77(1):18–39, July 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901544.

Musher:1981:NSK

[Mus81] S. L. Musher. On numerical solution of kinetic equations for plasma waves. Journal of Computational Physics, 43(2):250– 259, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901212.

Masovic:1983:SOI

[MV83a] D. R. Masovíc and F. R. Vukajlovic. Spin-orbit interaction in metals, elementary semiconductors, and semiconductor compounds. Journal of Computational Physics, 50(3):505–512, June 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901092.

Milkhailov:1983:CPS

[MV83b] Mikhail D. Milkhailov and Nikolai L. Vulchanov. Computational procedure for Sturm–Liouville problems. Journal of Computational Physics, 50(3):323–336, June 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901018.

Mulder:1985:EIU

[MV85] Wim A. Mulder and Bram Van Leer. Experiments with implicit upwind methods for the Euler equations. Journal of Computational Physics, 59(2):232–246, June 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901445.

Meijerink:1981:GUI

[MvdV81] J. A. Meijerink and H. A. van der Vorst. Guidelines for the usage of incomplete decompositions in solving sets of linear equations as they occur in practical problems. Journal of Computational Physics, 44(1):134–155, November 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900413.

McDonald:1984:CLR

[MW84] B. E. McDonald and J. M. Witting. A conservation law related to Kelvin’s circulation theorem. Journal of Computational Physics, 56(2):237–243, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900937.

Madhavan:1987:MSI

[MW87] P. V. Madhavan and J. L. Written. Molecular symmetry in ab initio calculations. Journal of Computational Physics, 70(1):253– 261, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900131.

Menikoff:1980:MNC

[MZ80] Ralph Menikoff and Charles Zemach. Methods for numerical conformal mapping. Journal of Computational Physics, 36(3):366– 410, July 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180901667.

Meng-Zhao:1982:ISN

[MZ82] Qin Meng-Zhao. An implicit scheme for nonlinear evolution equations. Journal of Computational Physics, 48(1):57–71, October 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900353.

Menikoff:1983:RTI

[MZ83] Ralph Menikoff and Charles Zemach. Rayleigh–Taylor instability and the use of conformal maps for ideal fluid flow. Journal of Computational Physics, 51(1):28–64, July 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900803.

Malik:1985:SCM

[MZH85] M. R. Malik, T. A. Zang, and M. Y. Hussaini. A spectral collocation method for the Navier–Stokes equations. Journal of Computational Physics, 61(1):64–88, October 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900610.

Nataf:1989:OBC

[Nat89] F. Nataf. An open boundary condition for the computation of the steady incompressible Navier–Stokes equations. Journal of Computational Physics, 85(1):104–129, November 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902027.

Navon:1983:NGT

[Nav83] I. M. Navon. A Numerov–Galerkin technique applied to a finiteelement shallow-water equations model with enforced conservation of integral invariants and selective lumping. Journal of Computational Physics, 52(2):313–339, November 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900335.

Navarra:1987:AAM

[Nav87] A. Navarra. An application of the Arnoldi’s method to a geophysical fluid dynamics problem. Journal of Computational Physics, 69(1):143–162, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901598.

Nicolai:1989:CMD

[NB89] A. Nicolai and P. Börner. A computational model describing confinement and performance of circular and D-shaped Tokamak plasmas. Journal of Computational Physics, 80(1):98–136, January 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900922.

Nielsen:1989:AOP

[ND89] Dale E. Nielsen, Jr. and Adam T. Drobot. An analysis and optimization of the pseudo-current method. Journal of Computational Physics, 85(2):503, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901691.

Naouri:1988:SEE

[NDAL88] G. Naouri, C. Deutsch, O. Atabek, and R. Lefebvre. Successive embeddings of excited atomic dipoles in plasmas. Journal of Computational Physics, 77(1):73–84, July 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890157X.

Niclot:1988:DPS

[NDP88] B. Niclot, P. Degond, and F. Poupaud. Deterministic particle simulations of the Boltzmann Transport Equation of semiconductors. Journal of Computational Physics, 78(2):313–349, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900538.

Nichols:1982:NSE

[NE82] Nancy K. Nichols and Roland England. Numerical solution of an elastic boundary layer problem using a multiple shooting technique. Journal of Computational Physics, 46(3):369–389, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900213.

Neuman:1981:ELN

[Neu81] Shlomo P. Neuman. A Eulerian–Lagrangian numerical scheme for the dispersion-convection equation using conjugate space-time grids. Journal of Computational Physics, 41(2):270–294, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900978.

Neutsch:1983:OSD

[Neu83] Wolfram Neutsch. Optimal spherical designs and numerical integration on the sphere. Journal of Computational Physics, 51(2):313–325, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900955.

Nex:1987:NSS

[Nex87] C. M. M. Nex. A new splitting to solve a large Hermitian eigenproblem. Journal of Computational Physics, 70(1):138–144, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900052. See comment [?].

Ng:1981:SNS

[Ng81] K. Y. K. Ng. Solution of Navier–Stokes equations by goal programming. Journal of Computational Physics, 39(1):103–111, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918190139X.

Nachamkin:1980:FBT

[NM80] Jack Nachamkin and C. J. Maggiore. A Fourier Bessel transform method for efficiently calculating the magnetic field of solenoids. Journal of Computational Physics, 37(1):41–55, August 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900030.

Nevins:1981:PSU

[NMG81] W. M. Nevins, Y. Matsuda, and M. J. Gerver. Plasma simulations using inversion symmetry as a boundary condition. Journal of Computational Physics, 39(1):226–232, January 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901467.

Norton:1986:NMB

[NMG86] N. J. Norton, J. C. McWilliams, and P. R. Gent. A numerical model of the balance equations in a periodic domain and an example of balanced turbulence. Journal of Computational Physics, 67(2):439–471, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902718.

Neta:1989:ATZ

[NN89] B. Neta and I. M. Navon. Analysis of the Turkel–Zwas scheme for the shallow-water equations. Journal of Computational Physics, 81(2):277–299, April 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990209X.

Noh:1987:ECS

[Noh87] W. F. Noh. Errors for calculations of strong shocks using an artificial viscosity and an artificial heat flux. Journal of Computational Physics, 72(1):78–120, September 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918790074X.

Nolet:1985:SRI

[Nol85] Guust Nolet. Solving or resolving inadequate and noisy tomographic systems. Journal of Computational Physics, 61(3):463– 482, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900750.

Noras:1980:GSB

[Nor80] J. M. Noras. A general strategy for Brillouin zone sampling and integration. Journal of Computational Physics, 36(3):304– 312, July 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918090162X.

Nordstrom:1989:IOB

[Nor89] Jan Nordström. The influence of open boundary conditions on the convergence to steady state for the Navier–Stokes equations. Journal of Computational Physics, 85(1):210–244, November 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189902052.

Nishiguchi:1985:VCP

[NOY85] A. Nishiguchi, S. Orii, and T. Yabe. Vector calculation of particle code. Journal of Computational Physics, 61(3):519–522, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900786.

Nelson:1980:AEC

[NR80a] P. Nelson, Jr. and A. K. Ray. An automatic error-control technique for computation of eigenlengths. Journal of Computational Physics, 37(3):388–398, October 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900443.

Ng:1980:NSO

[NR80b] B. S. Ng and W. H. Reid. On the numerical solution of the Orr– Sommerfeld problem: Asymptotic initial conditions for shooting methods. Journal of Computational Physics, 38(3):275–293, December 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180901503.

Nicolai:1984:MDT

[NR84] A. Nicolai and D. Reiter. A multicode describing the time evolution of the atomic protium, deuterium, and tritium densities in the first wall of a tokamak reactor. Journal of Computational Physics, 55(1):129–153, July 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900196.

Ng:1985:CMM

[NR85] B. S. Ng and W. H. Reid. The compound matrix method for ordinary differential systems. Journal of Computational Physics, 58(2):209–228, April 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185901779.

Nouri:1989:CFP

[NS89] F. Z. Nouri and D. M. Sloan. A comparison of Fourier pseudospectral methods for the solution of the Korteweg–de Vries equation. Journal of Computational Physics, 83(2):324–344, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901228.

Neutsch:1985:EIH

[NSJ85] W. Neutsch, E. Schrüfer, and A. Jessner. Efficient integration on the hypersphere. Journal of Computational Physics, 59(1):167– 175, May 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901147.

Naitou:1980:IPL

[NTK80] H. Naitou, S. Tokuda, and T. Kamimura. Initial particle loadings for a nonuniform simulation plasma in a magnetic field. Journal of Computational Physics, 38(3):265–274, December 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901497.

Natarajan:1989:NIS

[NV89] Ramesh Natarajan and David Vanderbilt. A new iterative scheme for obtaining eigenvectors of large, real-symmetric matrices. Journal of Computational Physics, 82(1):218–228, May 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900430.

Nishiguchi:1982:FSF

[NY82] A. Nishiguchi and T. Yabe. Finite-sized fluid particle in a nonuniform moving grid. Journal of Computational Physics, 47(2):297– 302, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918290081X.

Nishiguchi:1983:SOF

[NY83] A. Nishiguchi and T. Yabe. Second-order fluid particle scheme. Journal of Computational Physics, 52(2):390–413, November 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900372.

Olsson:1985:TDS

[OA85] Jim O. Olsson and Lars L. Andersson. Time-dependent solution of pre-mixed laminar flames with a known temperature profile. Journal of Computational Physics, 59(3):369–387, July 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901172.

OBrien:1980:SBF

[O’B80] D. M. O’Brien. Spherical Bessel functions of large order. Journal of Computational Physics, 36(1):128–132, June 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901771.

OBrien:1981:CMI

[O’B81] V. O’Brien. Conformal mappings for internal viscous flow problems. Journal of Computational Physics, 44(1):220–226, November 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900486.

Osher:1983:USB

[OC83] Stanley Osher and Sukumar Chakravarthy. Upwind schemes and boundary conditions with applications to Euler equations in general geometries. Journal of Computational Physics, 50(3):447– 481, June 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901067.

Oza:1987:NSS

[OC87] Dipak H. Oza and Joseph Callaway. Numerical solution of the Schrödinger equation with a non-local exchange kernel. Journal of Computational Physics, 68(1):89–103, January 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900465.

Oey:1982:ISD

[Oey82] Li-Yauw Oey. Implicit schemes for differential equations. Journal of Computational Physics, 45(3):443–468, March 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901140.

Oskam:1982:GRS

[OF82] B. Oskam and J. M. J. Fray. General relaxation schemes in multigrid algorithms for higher-order singularity methods. Journal of Computational Physics, 48(3):423–440, December 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900602.

Olivier:1986:RCM

[OG86] H. Olivier and H. Grönig. The random choice method applied to two-dimensional shock focusing and diffraction. Journal of Computational Physics, 63(1):85–106, March 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900859.

Ohring:1981:NWW

[Ohr81a] Samuel Ohring. Nonlinear water wave generation using the method of lines. Journal of Computational Physics, 39(1):137– 163, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918190142X.

Ohring:1981:TDS

[Ohr81b] Samuel Ohring. Three-dimensional ship wave generation using an efficient finite difference scheme with double model linearization. Journal of Computational Physics, 41(1):89–114, May 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900814.

Ohring:1983:AMM

[Ohr83] Samuel Ohring. Application of the multigrid method to Poisson’s equation in boundary-fitted coordinates. Journal of Computational Physics, 50(2):307–315, May 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900682.

Ogawa:1987:MCF

[OI87] Satoru Ogawa and Tomiko Ishiguro. A method for computing flow fields around moving bodies. Journal of Computational Physics, 69(1):49–68, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187901550.

Olsen:1988:SLM

[OJJ88] Jeppe Olsen, Hans Jørgen Aa Jensen, and Poul Jørgensen. Solution of the large matrix equations which occur in response theory. Journal of Computational Physics, 74(2):265–282, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900812.

Obayashi:1986:ALF

[OK86] Shigeru Obayashi and Kunio Kuwahara. An approximate LU factorization method for the compressible Navier–Stokes equations. Journal of Computational Physics, 63(1):157–167, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900896.

Oliphant:1980:DQE

[Oli80] Thomas Oliphant. Dynamic and quasi-equilibrium Lagrangian MHD in 1-D. Journal of Computational Physics, 38(3):406–412, December 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180901576.

Otsuka:1983:BNB

[ONM83] M. Otsuka, M. Nagami, and T. Matsuda. Birth: A neutral beam deposition code for non-circular tokamak plasmas. Journal of Computational Physics, 52(2):219–236, November 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900293.

ORourke:1984:TFV

[O’R84] Peter J. O’Rourke. The turn function and vorticity method for numerical fluid dynamics. Journal of Computational Physics, 53(3):359–381, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184900664.

ORourke:1989:SPN

[O’R89] Peter J. O’Rourke. Statistical properties and numerical implementation of a model for droplet dispersion in a turbulent gas. Journal of Computational Physics, 83(2):345–360, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918990123X.

Orszag:1980:SMP

[Ors80] Steven A. Orszag. Spectral methods for problems in complex geometries. Journal of Computational Physics, 37(1):70–92, August 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900054.

Ogen:1983:CFE

[OS83a] G. Ogen and B. Schiff. Constrained finite elements for singular boundary value problems. Journal of Computational Physics, 51 (1):65–82, July 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183900815.

Olver:1983:ALF

[OS83b] F. W. J. Olver and J. M. Smith. Associated Legendre functions on the cut. Journal of Computational Physics, 51(3):502–518, September 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901663.

Oset:1985:ESS

[OS85] E. Oset and L. L. Salcedo. Error sources in the standard numerical integration of the Schrödinger equation: An improved method. Journal of Computational Physics, 57(3):361–377, February 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901846.

Osher:1988:FPC

[OS88] Stanley Osher and James A. Sethian. Fronts propagating with curvature-dependent speed: Algorithms based on Hamilton– Jacobi formulations. Journal of Computational Physics, 79(1):12– 49, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900022.

Osgood:1981:ACP

[Osg81] Charles F. Osgood. Approximating convolution products better than the DFT while keeping the FFT. Journal of Computational Physics, 42(2):382–395, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902515.

Otani:1988:AIC

[Ota88] Niels F. Otani. The Alfvén ion-cyclotron instability. Simulation theory and techniques. Journal of Computational Physics, 78(2):251–277, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188900496.

Oostendorp:1989:ITS

[OvOH89] Thom F. Oostendorp, Adriaan van Oosterom, and Geertjan Huiskamp. Interpolation on a triangulated 3D surface. Journal of Computational Physics, 80(2):331–343, February 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901034.

OSullivan:1985:CVB

[OW85] Finbarr O’Sullivan and Grace Wahba. A cross validated Bayesian retrieval algorithm for nonlinear remote sensing experiments. Journal of Computational Physics, 59(3):441–455, July 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901214.

Piessens:1986:ATP

[PA86] Robert Piessens and Shafique Ahmed. Approximation for the turning points of Bessel functions. Journal of Computational Physics, 64(1):253–257, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918690029X.

Poo:1989:CMD

[PA89] J. Y. Poo and N. Ashgriz. A computational method for determining curvatures. Journal of Computational Physics, 84(2):483– 491, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189902453.

Pajunen:1989:CPP

[Paj89] Petri Pajunen. Complex Prüfer phase function with applications to scattering. Journal of Computational Physics, 82(1): 16–28, May 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189900338.

Panaras:1989:SNU

[Pan89] Argyris G. Panaras. The spatially non-uniform convergence of the numerical solutions of flows. Journal of Computational Physics, 82(2):429–453, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900570.

Papantonopoulos:1984:CPC

[Pap84] E. Papantonopoulos. A computational procedure for calculating matrix elements and Clebsch–Gordan coefficients of simple Lie algebras. Journal of Computational Physics, 53(1):28–41, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900500.

Patera:1984:SEM

[Pat84] Anthony T. Patera. A spectral element method for fluid dynamics: Laminar flow in a channel expansion. Journal of Computational Physics, 54(3):468–488, June 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901281.

Patera:1986:FDP

[Pat86] Anthony T. Patera. Fast direct Poisson solvers for high-order finite element discretizations in rectangularly decomposable domains. Journal of Computational Physics, 65(2):474–480, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186902196.

Patel:1987:SFA

[Pat87] V. A. Patel. Symmetry of the flow around a circular cylinder. Journal of Computational Physics, 71(1):65–99, July 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900209.

Pathria:1989:PSS

[Pat89] D. Pathria. Pseudo-spectral solution of nonlinear Schrödinger equations. Journal of Computational Physics, 84(1):257, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918990199X.

Pierce:1981:GNS

[PC81a] T. H. Pierce and R. I. Cukier. Global nonlinear sensitivity analysis using Walsh functions. Journal of Computational Physics, 41(2):427–443, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901042.

Pulliam:1981:DFI

[PC81b] T. H. Pulliam and D. S. Chaussee. A diagonal form of an implicit approximate-factorization algorithm. Journal of Computational Physics, 39(2):347–363, February 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190156X.

Paolucci:1982:SEF

[PC82] Samuel Paolucci and Donald R. Chenoweth. Stability of the explicit finite differenced transport equation. Journal of Computational Physics, 47(3):489–496, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900985.

Poinsot:1986:IDC

[PC86] Thierry Poinsot and Sébastien M. Candel. The influence of differencing and CFL number on implicit time-dependent non-linear calculations. Journal of Computational Physics, 62(2):282–296, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901282.

Pearson:1981:USC

[Pea81] Carl E. Pearson. Use of streamline coordinates in the numerical solution of compressible flow problems. Journal of Computational Physics, 42(2):257–265, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902436.

Pearson:1983:APR

[Pea83] Robert B. Pearson. An algorithm for pseudo random number generation suitable for large scale integrations. Journal of Computational Physics, 49(3):478–489, March 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901420.

Pert:1981:ASC

[Per81a] G. J. Pert. Algorithms for the self-consistent generation of magnetic fields in plasmas. Journal of Computational Physics, 43(1):111–163, September 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181901145.

Pert:1981:EIT

[Per81b] G. J. Pert. Efficient integration of the time dependent collisional-radiative equations. Journal of Computational Physics, 39(2):251–271, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181901510.

Pert:1981:PCN

[Per81c] G. J. Pert. Physical constraints in numerical calculations of diffusion. Journal of Computational Physics, 42(1):20–52, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918190231X.

Pert:1983:QLR

[Per83] G. J. Pert. Quasi-Lagrangian rezoning of fluid codes maintaining an orthogonal mesh. Journal of Computational Physics, 49(1): 1–43, January 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901134.

Perlman:1985:AVM

[Per85] Mirta Perlman. On the accuracy of vortex methods. Journal of Computational Physics, 59(2):200–223, June 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901421.

Peranich:1987:FDS

[Per87] L. S. Peranich. A finite difference scheme for solving a nonlinear Schrödinger equation with a linear damping term. Journal of Computational Physics, 68(2):501–505, February 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900696.

Peskun:1981:GCT

[Pes81] P. H. Peskun. Guidelines for choosing the transition matrix in Monte Carlo methods using Markov chains. Journal of Computational Physics, 40(2):327–344, April 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190214X.

Petravic:1987:OGC

[Pet87] M. Petravic. Orthogonal grid construction for modeling of transport in tokamaks. Journal of Computational Physics, 73(1):125– 130, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901094.

Pakleza:1987:PCS

[PG87] J. Pakleza and D. Gentile. Phase-change simulation in a propagative medium subject to a heat flux disturbance. Journal of Computational Physics, 71(1):34–49, July 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900180.

Papatzacos:1988:IFP

[PG88a] Paul Papatzacos and Sven-Åke Gustafson. Incompressible flow in porous media with two moving boundaries. Journal of Computational Physics, 78(1):231–248, September 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900460.

Pearlstein:1988:ETC

[PG88b] Arne J. Pearlstein and Dimitrios A. Goussis. Efficient transformation of certain singular polynomial matrix eigenvalue problems. Journal of Computational Physics, 78(2):305–312, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900526.

Patnaik:1987:BIC

[PGBO87] G. Patnaik, R. H. Guirguis, J. P. Boris, and E. S. Oran. A barely implicit correction for flux-corrected transport. Journal of Computational Physics, 71(1):1–20, July 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900167.

Pickering:1988:PEH

[PH88] W. M. Pickering and P. J. Harley. Poisson’s equation, hexagonal grids and FFT methods: Periodic boundary conditions. Journal of Computational Physics, 75(1):236–243, March 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890109X.

Paul:1989:IRN

[PHD89] W. Paul, Dieter W. Heermann, and Rashmi C. Desai. Implementation of a random number generator in OCCAM. Journal of Computational Physics, 82(2):487–491, June 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900594.

Phillips:1984:NCE

[Phi84a] Timothy N. Phillips. Natural convection in an enclosed cavity. Journal of Computational Physics, 54(3):365–381, June 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184901232.

Phillips:1984:NSC

[Phi84b] Timothy N. Phillips. Numerical solution of a coupled pair of elliptic equations from solid state electronics. Journal of Computational Physics, 53(3):472–483, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918490072X.

Phillips:1986:NTB

[Phi86] Timothy N. Phillips. On the numerical treatment of boundary singularities in elliptic problems. Journal of Computational Physics, 64(2):459–472, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900446.

Pickering:1983:NFD

[Pic83] W. M. Pickering. Nonpolynomial finite difference schemes and the use of the fast Fourier transform. Journal of Computational Physics, 51(2):357–363, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900992.

Pickering:1986:SPE

[Pic86] W. M. Pickering. On the solution of Poisson’s equation on a regular hexagonal grid using FFT methods. Journal of Computational Physics, 64(2):320–333, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900367.

Piessens:1984:CSA

[Pie84] R. Piessens. Chebyshev series approximations for the zeros of the Bessel functions. Journal of Computational Physics, 53(1):188– 192, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900603.

Pike:1987:GAA

[Pik87] J. Pike. Grid adaptive algorithms for the solution of the Euler equations on irregular grids. Journal of Computational Physics, 71(1):194–223, July 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187900271.

Piran:1980:NCC

[Pir80] Tsvi Piran. Numerical codes for cylindrical general relativistic systems. Journal of Computational Physics, 35(2):254–283, April 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900881.

Prosperetti:1983:NMP

[PJ83] Andrea Prosperetti and Jeffrey W. Jacobs. A numerical method for potential flows with a free surface. Journal of Computational Physics, 51(3):365–386, September 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901596.

Pouagare:1986:SMM

[PL86] M. Pouagare and B. Lakshminarayana. A space-marching method for viscous incompressible internal flows. Journal of Computational Physics, 64(2):389–415, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900409.

Platzman:1981:SRC

[Pla81] George W. Platzman. Some response characteristics of finiteelement tidal models. Journal of Computational Physics, 40(1): 36–63, March 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901984.

Peskin:1980:MPH

[PM80] Charles S. Peskin and David M. McQueen. Modeling prosthetic heart valves for numerical analysis of blood flow in the heart. Journal of Computational Physics, 37(1):113–132, August 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900078.

Peskin:1989:TDC

[PM89] Charles S. Peskin and David M. McQueen. A three-dimensional computational method for blood flow in the heart I. Immersed elastic fibers in a viscous incompressible fluid. Journal of Computational Physics, 81(2):372–405, April 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902131.

Payne:1983:NSZ

[PND83] G. L. Payne, D. R. Nicholson, and R. M. Downie. Numerical solution of the Zakharov equations. Journal of Computational Physics, 50(3):482–498, June 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901079.

Polyzou:1987:NMT

[Pol87] W. N. Polyzou. Nonperturbative method for treating square root operators in relativistic quantum theories. Journal of Computational Physics, 70(1):117–137, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900040.

Pomentale:1983:PSD

[Pom83] Tomaso Pomentale. Periodic solutions of differential equations and one-parameter family of operators. Journal of Computational Physics, 51(3):527–535, September 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901687.

Pomraning:1988:IBC

[Pom88] G. C. Pomraning. Initial and boundary conditions for flux-limited diffusion theory. Journal of Computational Physics, 75(1):73– 85, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918890099X.

Passot:1988:HCF

[PP88a] T. Passot and A. Pouquet. Hyperviscosity for compressible flows using spectral methods. Journal of Computational Physics, 75(2): 300–313, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901143.

Percus:1988:LRC

[PP88b] Ora E. Percus and Jerome K. Percus. Long range correlations in linear congruential generators. Journal of Computational Physics, 77(1):267–269, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188901660.

Primorac:1989:CMI

[Pri89] Miljenko Primorac. Computation of molecular integrals over partially generalized Hermite–Gaussian functions. Journal of Computational Physics, 84(2):474–482, October 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902441.

Proskurowski:1981:NSB

[Pro81a] Włodzimierz Proskurowski. A note on solving the Buckley– Leverett equation in the presence of gravity. Journal of Computational Physics, 41(1):136–141, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900838.

Prosperetti:1981:NDN

[Pro81b] Andrea Prosperetti. On numerical differentiation on a nonuniform grid. Journal of Computational Physics, 39(2):481–483, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181901662.

Prosperetti:1982:MSC

[Pro82] Andrea Prosperetti. A method for the solution of a class of singular Volterra integro-differential equations. Journal of Computational Physics, 46(3):462–468, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900274.

Pearson:1983:FPM

[PRT83] Robert B. Pearson, John L. Richardson, and Doug Toussain. A fast processor for Monte–Carlo simulation. Journal of Computational Physics, 51(2):241–249, August 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900906.

Pruess:1988:SAC

[Pru88] Steven Pruess. On shooting algorithms for calculating Sturm– Liouville eigenvalues. Journal of Computational Physics, 75(2): 493–497, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901246.

Pinero:1989:BSE

[PRV89] Petra Pinero, Luis F. Rull, and Jose M. Velarde. Baxter solution to the 0 − Z equation near the critical point. Journal of Computational Physics, 85(2):504, December 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901745.

Pryce:1987:EEC

[Pry87] J. D. Pryce. Error estimation for a class of differential eigenproblems. Journal of Computational Physics, 69(1):252–257, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187901653.

Pares-Sierra:1989:FSD

[PSV89] Alejandro Pares-Sierra and Geoffrey K. Vallis. A fast semidirect method for the numerical solution of non-separable elliptic equations in irregular domains. Journal of Computational Physics, 82(2):398–412, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900557.

Pawley:1982:ILC

[PT82] G. S. Pawley and G. W. Thomas. The implementation of lattice calculations on the DAP. Journal of Computational Physics, 47(2):165–178, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900729.

Pajunen:1986:ECP

[PT86] Petri Pajunen and Juha Tienari. Efficient computation of the Prüfer phase function for determining eigenvalues of Sturm– Liouville systems. Journal of Computational Physics, 65(1):159– 169, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900094.

Pullin:1980:DSM

[Pul80] D. I. Pullin. Direct simulation methods for compressible inviscid ideal-gas flow. Journal of Computational Physics, 34(2):231– 244, February 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180901072.

Puoskari:1988:MCB

[Puo88] M. Puoskari. A method for computing Bessel function integrals. Journal of Computational Physics, 75(2):334–344, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901167.

Pritchard:1983:SME

[PV83] H. O. Pritchard and S. R. Vatsya. Stiffness of the master equation for low-temperature reaction rates. Journal of Computational Physics, 49(1):173–178, January 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901201.

Peraire:1987:ARC

[PVMZ87] J. Peraire, M. Vahdati, K. Morgan, and O. C. Zienkiewicz. Adaptive remeshing for compressible flow computations. Journal of Computational Physics, 72(2):449–466, October 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900933.

Perram:1985:SMH

[PW85] John W. Perram and M. S. Wertheim. Statistical mechanics of hard ellipsoids. I. Overlap algorithm and the contact function. Journal of Computational Physics, 58(3):409–416, May 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185901718.

Qin:1988:EEC

[Qin88] Chao-Yu Qin. An explicit energy-conserving numerical method for equations of the form d^{2}x/dt^{2} = f(x). Journal of Computational Physics, 79(2):473–476, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900277.

Quartapelle:1986:ICP

[QN86] L. Quartapelle and M. Napolitano. Integral conditions for the pressure in the computation of incompressible viscous flows. Journal of Computational Physics, 62(2):340–348, February 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901324.

Quick:1989:SSF

[QQM89] J. F. H. Quick, R. M. Quick, and H. G. Miller. A simple strategy for finding the low-lying solutions of the restricted nuclear Hartree–Fock equations. Journal of Computational Physics, 80(1):243–251, January 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900971.

Quartapelle:1989:NST

[QR89] L. Quartapelle and S. Rebay. Numerical solution of twopoint boundary value problems. Journal of Computational Physics, 84(1):255, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189901927.

Quartapelle:1981:VCC

[Qua81]  L. Quartapelle.  Vorticity conditioning in the computa-

tion of two-dimensional viscous flows. Journal of Computational Physics, 40(2):453–477, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902229.

Rai:1981:GET

[RA81] Man Mohan Rai and D. A. Anderson. Grid evolution in time asymptotic problems. Journal of Computational Physics, 43(2):327–344, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181901261.

Ramshaw:1985:IIS

[RA85] J. D. Ramshaw and A. A. Amsden. Improved iteration scheme for partial equilibrium flow. Journal of Computational Physics, 59(3):484–489, July 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918590124X.

Ramshaw:1987:ITV

[RA87] J. D. Ramshaw and A. A. Amsden. Inclusion of temperature variations in a numerical scheme for partial equilibrium flow. Journal of Computational Physics, 71(1):224–227, July 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900283.

Raggio:1982:KTA

[Rag82] Gabriel Raggio. On the Kantorovich technique applied to the tidal equations in elongated lakes. Journal of Computational Physics, 46(1):43–53, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900043.

Rai:1986:CTZ

[Rai86a] Man Mohan Rai. A conservative treatment of zonal boundaries for Euler equation calculations. Journal of Computational Physics, 62(2):472–503, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901415.

Rai:1986:RAP

[Rai86b] Man Mohan Rai. A relaxation approach to patched-grid calculations with the Euler equations. Journal of Computational Physics, 66(1):99–131, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900562.

Ramshaw:1983:MES

[Ram83] John D. Ramshaw. A method for enforcing the solenoidal condition on magnetic field in numerical calculations. Journal of Computational Physics, 52(3):592–596, December 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900098.

Ramsey:1985:LMS

[Ram85a] Gordon P. Ramsey. The Laguerre method for solving integrodifferential equations. Journal of Computational Physics, 60(1): 97–118, August 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900191.

Ramshaw:1985:CRA

[Ram85b] John D. Ramshaw. Conservative rezoning algorithm for generalized two-dimensional meshes. Journal of Computational Physics, 59(2):193–199, June 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918590141X.

Ramshaw:1986:SSO

[Ram86] John D. Ramshaw. Simplified second-order rezoning algorithm for generalized two-dimensional meshes. Journal of Computational Physics, 67(1):214–222, November 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901221.

Rapaport:1980:ESP

[Rap80a] D. C. Rapaport. The event scheduling problem in molecular dynamic simulation. Journal of Computational Physics, 34(2):184– 201, February 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180901047.

Rapoport:1980:SLO

[Rap80b] David Rapoport. Splitting of linear operators applied to the stationary Navier–Stokes equation. Journal of Computational Physics, 35(3):393–407, May 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999180900777.

Rapaport:1985:MDS

[Rap85] D. C. Rapaport. Molecular dynamics simulation using quaternions. Journal of Computational Physics, 60(2):306–314, September 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900099.

Ramshaw:1981:NMP

[RC81] John D. Ramshaw and Lawrence D. Cloutman. Numerical method for partial equilibrium flow. Journal of Computational Physics, 39(2):405–417, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901601.

Rogers:1987:DAM

[RCK87] Stuart E. Rogers, James L. C. Chang, and Dochan Kwak. A diagonal algorithm for the method of pseudocompressibility. Journal of Computational Physics, 73(2):364–379, December 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901422.

Ramshaw:1983:RKR

[RD83] John D. Ramshaw and John K. Dukowicz. Reply to Kansa’s remarks on the efficiency of ICE-like schemes for combustion problems. Journal of Computational Physics, 49(2):349–350, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901328. See reply [Kan83].

Rizzi:1987:STT

[RE87] Arthur Rizzi and Björn Engquist. Selected topics in the theory and practice of computational fluid dynamics. Journal of Computational Physics, 72(1):1–69, September 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900726.

Reitz:1981:ODC

[Rei81] Rolf D. Reitz. One-dimensional compressible gas dynamics calculations using the Boltzmann equation. Journal of Computational Physics, 42(1):108–123, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902357.

Rettrup:1982:IMC

[Ret82] Sten Rettrup. An iterative method for calculating several of the extreme eigensolutions of large real non-symmetric matrices. Journal of Computational Physics, 45(1):100–107, January 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182901048. See [?].

Reiman:1988:NST

[RG88] A. H. Reiman and H. S. Greenside. Numerical solution of three-dimensional magnetic differential equations. Journal of Computational Physics, 75(2):423–443, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901210.

Ransom:1984:HTP

[RH84] V. H. Ransom and D. L. Hicks. Hyperbolic two-pressure models for two-phase flow. Journal of Computational Physics, 53(1):124– 151, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900561.

Rawitscher:1987:AEI

[RH87] G. H. Rawitscher and E. S. Hirschorn. Accurate evaluation of an integral involving the product of two Bessel functions and a Gaussian. Journal of Computational Physics, 68(1):104–126, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900477.

Ransom:1988:HTP

[RH88a] V. H. Ransom and D. L. Hicks. Hyperbolic two-pressure models for two-phase flow revisited. Journal of Computational Physics, 75(2):498–504, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188901258.

Roberts:1988:HSG

[RH88b] William W. Roberts, Jr. and Mark A. Hausman. Hypersonic, stratified gas flows past an obstacle: Direct simulation Monte Carlo calculations. Journal of Computational Physics, 77(2):283– 317, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901702.

Riedinger:1988:ADV

[RHOG88] R. Riedinger, M. Habar, P. Oelhafen, and H. J. Güntherodt. About the Delaunay–Voronoi tesselation. Journal of Computational Physics, 74(1):61–72, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890068X.

Rigal:1981:ACV

[Rig81] Alain Rigal. Acceleration of the convergence in viscous flow computations. Journal of Computational Physics, 43(1):177–179, September 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901169.

Rigal:1987:SMC

[Rig87] Alain Rigal. SOR methods for coupled elliptic partial differential equations. Journal of Computational Physics, 71(1):181–193, July 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918790026X.

Rizzi:1988:MCV

[Riz88] Arthur Rizzi. Multi-cell vortices computed in large-scale difference solution to the incompressible Euler equations. Journal of Computational Physics, 77(1):207–220, July 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901635.

Ryskin:1983:OM

[RL83] G. Ryskin and L. G. Leal. Orthogonal mapping. Journal of Computational Physics, 50(1):71–100, April 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900426.

Redi:1984:PEC

[RL84] M. H. Redi and D. A. Larrabee. Plasma equilibrium calculations by line successive over relaxation. Journal of Computational Physics, 53(2):351–356, February 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900470.

Roth:1989:IMS

[RL89] M. Roth and J. Lemaire. An iterative method to solve the nonlinear Poisson’s equation in the case of plasma tangential discontinuities. Journal of Computational Physics, 84(1):256, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901940.

Rudman:1982:TDB

[RM82] S. Rudman and F. Marconi. Three-dimensional boundary conditions in supersonic flow. Journal of Computational Physics, 48(2):284–301, November 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900511.

Rangwalla:1987:NSV

[RM87] A. A. Rangwalla and B. R. Munson. Numerical solution for viscous flow for two-dimensional domains using orthogonal coordinate systems. Journal of Computational Physics, 70(2):373–396, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187901884.

Runca:1985:AFD

[RMS85] E. Runca, P. Melli, and F. Sardei. An analysis of a finite-difference and a Galerkin technique applied to the simulation of advection and diffusion of air pollutants from a line source. Journal of Computational Physics, 59(1):152–166, May 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901135.

Reusch:1986:TSP

[RN86] M. F. Reusch and G. H. Neilson. Toroidally symmetric polynomial multipole solutions of the vector Laplace equation. Journal of Computational Physics, 64(2):416–432, June 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900410.

Roberts:1985:ARV

[Rob85] Stephen Roberts. Accuracy of the random vortex method for a problem with non-smooth initial conditions. Journal of Computational Physics, 58(1):29–43, March 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901548.

Roberts:1989:TPF

[Rob89a] Lawrence J. Roberts. Two-phase flow in porous media with multicomponent transport: Formulation and higher order numerical solution. Journal of Computational Physics, 84(1):255, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901939.

Robinson:1989:SMC

[Rob89b] P. A. Robinson. Systematic methods for calculation of the dielectric properties of arbitrary plasmas. Journal of Computational Physics, 85(2):503, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901708.

Roe:1981:ARS

[Roe81] P. L. Roe. Approximate Riemann solvers, parameter vectors, and difference schemes. Journal of Computational Physics, 43(2):357– 372, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901285.

Roe:1986:DMN

[Roe86] Philip L. Roe. Discrete models for the numerical analysis of time-dependent multidimensional gas dynamics. Journal of Computational Physics, 63(2):458–476, April 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902044.

Rokhlin:1985:RSI

[Rok85] Vladimir Rokhlin. Rapid solution of integral equations of classical potential theory. Journal of Computational Physics, 60 (2):187–207, September 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185900026.

Rokhlin:1989:RSI

[Rok89] V. Rokhlin. Rapid solution of integral equations of scattering theory in two dimensions. Journal of Computational Physics, 84(1): 255, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901915.

Rome:1989:IMC

[Rom89] James A. Rome. An improved method of calculating flux coordinates. Journal of Computational Physics, 82(2):348–361, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900533.

Rose:1980:AEM

[Ros80] Milton E. Rose. An artificial energy method for calculating flows with shocks. Journal of Computational Physics, 38(1):107–113, November 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180900145.

Rose:1983:CFD

[Ros83] Milton E. Rose. Compact finite difference schemes for the Euler and Navier–Stokes equations. Journal of Computational Physics, 49(3):420–442, March 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901389.

Ramshaw:1985:PGS

[ROS85] J. D. Ramshaw, P. J. O’Rourke, and L. R. Stein. Pressure gradient scaling method for fluid flow with nearly uniform pressure. Journal of Computational Physics, 58(3):361–376, May 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185901688.

Ross:1987:RSW

[Ros87] David S. Ross. Redesign of a supercritical wing in the presence of an engine nacelle. Journal of Computational Physics, 73(2):233– 243, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901355.

Rosenfeld:1989:MST

[Ros89] Moshe Rosenfeld. A method for solving three-dimensional viscous incompressible flows over slender bodies. Journal of Computational Physics, 85(2):502, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901642.

Reardon:1989:GEC

[RQ89] A. C. Reardon and D. J. Quesnel. Growth of equilibrium clusters of Lennard-Jones molecules. Journal of Computational Physics, 83(1):240–245, July 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189902313.

Radicati:1989:IAS

[RRS89] G. Radicati, Y. Robert, and S. Succi. Iterative algorithms for the solution of nonsymmetric systems in the modelling of weak plasma turbulence. Journal of Computational Physics, 80(2):489– 497, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901137.

Rudy:1980:NOB

[RS80] David H. Rudy and John C. Strikwerda. A nonreflecting outflow boundary condition for subsonic Navier–Stokes calculations. Journal of Computational Physics, 36(1):55–70, June 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901746.

Ramos:1981:CPU

[RS81] J. I. Ramos and W. A. Sirignano. The calculation of the pressure in unsteady flows using a control-volume approach. Journal of Computational Physics, 41(1):211–216, May 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900905.

Ribando:1984:EMC

[RS84] R. J. Ribando and M. A. Shadday, Jr. The Ekman matching condition in a partially filled, rapidly rotating cylinder. Journal of Computational Physics, 53(2):266–288, February 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918490041X.

Rajaona:1985:MSA

[RS85] R. D. Rajaona and P. Sulmont. A method of spectral analysis applied to periodic and pseudoperiodic signals. Journal of Computational Physics, 61(1):186–193, October 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900671.

Reuven:1986:SAB

[RSR86] Yakir Reuven, Mitchell D. Smooke, and Herschel Rabitz. Sensitivity analysis of boundary value problems: application to nonlinear reaction-diffusion systems. Journal of Computational Physics, 64 (1):27–55, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900173.

Reddy:1988:FEA

[RV88a] B. D. Reddy and H. F. Voyé. Finite element analysis of the stability of fluid motions. Journal of Computational Physics, 79(1):92– 112, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918890006X.

Rothstein:1988:SED

[RV88b] Stuart M. Rothstein and Jan Vrbik. Statistical error of diffusion Monte Carlo. Journal of Computational Physics, 74(1):127– 142, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900733.

Rothberg:1989:EBP

[RWC89] R. H. Rothberg, W. F. Walker, and A. J. Chapman. Enhanced boundary pressure update for incompressible flow simulation. Journal of Computational Physics, 85(2):504, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901757.

Swift:1981:BCW

[SA81] Daniel W. Swift and John J. Ambrosiano. Boundary conditions which lead to excitation of instabilities in plasma simulations. Journal of Computational Physics, 44(2):302–317, December 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900541.

Sadiq:1984:NAM

[Sad84] Abdullah Sadiq. A new algorithm for the Monte Carlo simulation of spin-exchange kinetics of Ising systems. Journal of Computational Physics, 55(3):387–396, September 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900287.

Sakai:1981:ACT

[Sak81] Katsuhiro Sakai. An accelerated computing technique for steady fluid flows. Journal of Computational Physics, 40(2):509–516, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902266.

Saltzman:1986:VMG

[Sal86] Jeffrey Saltzman. Variational methods for generating meshes on surfaces in three dimensions. Journal of Computational Physics, 63(1):1–19, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900811.

Smith:1982:NMS

[SAR82] Leslie N. Smith, Stuart D. Augustin, and Herschel Rabitz. Numerical methods for solving time-dependent quantummechanical problems with applications. Journal of Computational Physics, 45(3):417–442, March 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182901139.

Schumann:1982:DSD

[SB82] U. Schumann and J. Benner. Direct solution of the discretized Poisson–Neumann problem on a domain composed of rectangles. Journal of Computational Physics, 46(1):1–14, August 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900018.

Shizgal:1984:DOM

[SB84a] B. Shizgal and R. Blackmore. A discrete ordinate method of solution of linear boundary value and eigenvalue problems. Journal of Computational Physics, 55(2):313–327, August 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900093.

Sweby:1984:CRS

[SB84b] P. K. Sweby and M. J. Baines. On convergence of Roe’s scheme for the general non-linear scalar wave equation. Journal of Computational Physics, 56(1):135–148, October 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900871.

Schneider:1987:MRH

[SB87a] W. Schneider and Glenn Bateman. Mapping from rectangular to harmonic representation. Journal of Computational Physics, 71(1):169–180, July 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999187900258.

Stotler:1987:SHP

[SB87b] D. P. Stotler and H. L. Berk. Solution of hot particle ballooning mode integral equation in Tokamaks. Journal of Computational Physics, 73(1):5–40, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901033.

Sharif:1988:AFD

[SB88] M. A. R. Sharif and Ahmed A. Busnaina. Assessment of finite difference approximations for the advection terms in the simulation of practical flow problems. Journal of Computational Physics, 74(1):143–176, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900745.

Shumaker:1982:NST

[SBAM82] D. E. Shumaker, J. K. Boyd, S. P. Auerbach, and B. McNamara. Numerical simulation of transport in a field-reversed mirror plasma. Journal of Computational Physics, 45(2):266–290, February 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182901218.

Segur:1983:AMF

[SBBY83] Pierre Segur, Marie-Claude Bordage, Jean-Pierre Balaguer, and Mohamed Yousfi. The application of a modified form of the S_{N} method to the calculation of swarm parameters of electrons in a weakly ionised equilibrium medium. Journal of Computational Physics, 50(1):116–137, April 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918390044X.

Schnack:1984:PAT

[SBC84] D. D. Schnack, D. C. Baxter, and E. J. Caramana. A pseudospectral algorithm for three-dimensional magnetohydrodynamic simulation. Journal of Computational Physics, 55(3):485–514, September 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900342.

Schnack:1987:SIM

[SBM^{+}87] D. D. Schnack, D. C. Barnes, Z. Mikic, Douglas S. Harned, and E. J. Caramana. Semi-implicit magnetohydrodynamic calculations. Journal of Computational Physics, 70(2):330–354, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187901860.

Stephens:1984:FDG

[SBSH84] A. B. Stephens, J. B. Bell, J. M. Solomon, and L. B. Hackerman. A finite difference Galerkin formulation for the incompressible Navier–Stokes equations. Journal of Computational Physics, 53(1):152–172, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900573.

Shield:1986:NCO

[SBT86] T. W. Shield, D. B. Bogy, and F. E. Talke. A numerical comparison of one-dimensional fluid jet models applied to drop-on-demand printing. Journal of Computational Physics, 67(2):327– 347, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186902664.

Steinhardt:1981:ASF

[SC81] Paul J. Steinhardt and P. Chaudhari. Airy stress function for atomic models. Journal of Computational Physics, 42(2):266– 276, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902448.

Shmilovich:1982:AMM

[SC82] Arvin Shmilovich and D. A. Caughey. Application of the multigrid method to calculations of transonic potential flow about wing–fuselage combinations. Journal of Computational Physics, 48(3):462–484, December 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900626.

Scharmer:1985:NAM

[SC85] G. B. Scharmer and M. Carlsson. A new approach to multilevel non-LTE radiative transfer problems. Journal of Computational Physics, 59(1):56–80, May 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918590107X.

Smolarkiewicz:1986:MPD

[SC86] Piotr K. Smolarkiewicz and Terry L. Clark. The multidimensional positive definite advection transport algorithm: Further development and applications. Journal of Computational Physics, 67(2):396–438, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902706.

Scanlon:1987:AFL

[Sca87] P. J. Scanlon. An alternative form for the Legendre polynomial expansion coefficients. Journal of Computational Physics, 69(2): 482–486, April 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901793.

Schmitt:1980:TET

[Sch80a] James A. Schmitt. Truncation error terms in the kinetic energy calculation in the HELP algorithm and their consequences. Journal of Computational Physics, 35(2):206–228, April 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900868.

Schumann:1980:FES

[Sch80b] U. Schumann. Fast elliptic solvers and three-dimensional fluidstructure interactions in a pressurized water reactor. Journal of Computational Physics, 36(1):93–127, June 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918090176X.

Schippers:1982:AMM

[Sch82] H. Schippers. Application of multigrid methods for integral equations to two problems from fluid dynamics. Journal of Computational Physics, 48(3):441–461, December 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900614.

Schiff:1988:FEE

[Sch88] B. Schiff. Finite element eigenvalues for the Laplacian over an L-shaped domain. Journal of Computational Physics, 76(2):233– 242, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901416.

Scott:1985:MGM

[Sco85] T. Scott. Multi-grid methods for oil reservoir simulation in two and three dimensions. Journal of Computational Physics, 59(2): 290–307, June 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901470.

Scott:1988:ISD

[Sco88] Bruce D. Scott. An implicit simulation of drift-wave turbulence in a sheared magnetic field. Journal of Computational Physics, 78(1):114–137, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890040X.

Shokoohi:1987:NID

[SE87] F. Shokoohi and H. G. Elrod. Numerical investigation of the disintegration of liquid jets. Journal of Computational Physics, 71(2):324–342, August 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900349.

Sellwood:1983:QSG

[Sel83] J. A. Sellwood. Quiet starts for galaxy simulations. Journal of Computational Physics, 50(3):337–359, June 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918390102X.

Sethian:1984:TCO

[Set84] James Sethian. Turbulent combustion in open and closed vessels. Journal of Computational Physics, 54(3):425–456, June 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184901268.

Singh:1986:SPC

[SG86] Shri Singh and Saul Goldman. Some practical considerations in the use of Metropolis and force-biased Monte Carlo simulations on liquids. Journal of Computational Physics, 62(2):464– 471, February 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901403.

Sethian:1988:VSV

[SG88a] J. A. Sethian and Ahmed F. Ghoniem. Validation study of vortex methods. Journal of Computational Physics, 74(2):283–317, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900824.

Soh:1988:USI

[SG88b] W. Y. Soh and John W. Goodrich. Unsteady solution of incompressible Navier–Stokes equations. Journal of Computational Physics, 79(1):113–134, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900071.

Schulten:1981:NAE

[SGA81] Z. Schulten, R. G. Gordon, and D. G. M. Anderson. A numerical algorithm for the evaluation of Weber parabolic cylinder functions U(a, x), V (a, x), and W (a, ±x). Journal of Computational Physics, 42(2):213–237, August 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902412.

Scales:1988:FSL

[SGT88] John A. Scales, Adam Gersztenkorn, and Sven Treitel. Fast I_{p} solution of large, sparse, linear systems: Application to seismic travel time tomography. Journal of Computational Physics, 75(2): 314–333, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901155.

Sharp:1984:PSM

[SH84a] H. Thomas Sharp and Wesley L. Harris. A pseudo-spectral method and parametric differentiation applied to the Falkner– Skan equation. Journal of Computational Physics, 55(3):377–386, September 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900275.

Sicilian:1984:ECS

[SH84b] J. M. Sicilian and C. W. Hirt. An efficient computation scheme for tracking contaminant concentrations in fluid flows. Journal of Computational Physics, 56(3):428–447, December 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901050.

Swindoll:1984:MTS

[SH84c] R. D. Swindoll and J. M. Haile. A multiple time-step method for molecular dynamics simulations of fluids of chain molecules. Journal of Computational Physics, 53(2):289–298, February 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900421.

Schultz:1989:SPP

[SH89] W. W. Schultz and S. W. Hong. Solution of potential problems using an overdetermined complex boundary integral method. Journal of Computational Physics, 84(2):414–440, October 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902416.

Sharp:1981:CNM

[Sha81] N. A. Sharp. Comparison of numerical methods for the integration of the black hole geodesic equations. Journal of Computational Physics, 41(2):295–308, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918190098X.

Shampine:1984:SAS

[Sha84] L. F. Shampine. Stiffness and the automatic selection of ODE codes. Journal of Computational Physics, 54(1):74–86, April 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184901396.

Schmieg:1982:EMA

[SHB82] C. Schmieg, P. C. Hägele, and L. M. Beck. An efficient matrix algorithm for the calculation of the gradient of the conformational energy of polymer chains. Journal of Computational Physics, 48(1):45–56, October 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900341.

Sheldon:1986:ASI

[She86] John W. Sheldon. Application of a semi-implicit finite difference scheme to heat generation in explosives. Journal of Computational Physics, 65(2):486–490, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902214.

Sherwood:1987:UFP

[She87] J. D. Sherwood. Unstable fronts in a porous medium. Journal of Computational Physics, 68(2):485–500, February 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900684.

Shestakov:1988:CFD

[She88] A. I. Shestakov. Comparison of finite differences and finite elements on a parabolic problem. Journal of Computational Physics, 79(1):231–243, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900149.

Shizgal:1981:GQP

[Shi81] B. Shizgal. A Gaussian quadrature procedure for use in the solution of the Boltzmann equation and related problems. Journal of Computational Physics, 41(2):309–328, June 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900991.

Shipsey:1986:NMU

[Shi86] Edward J. Shipsey. Numerical method for unitary systems. Journal of Computational Physics, 66(1):218–224, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900616.

Shestakov:1988:SDE

[SHK88] A. I. Shestakov, J. A. Harte, and D. S. Kershaw. Solution of the diffusion equation by finite elements in Lagrangian hydrodynamic codes. Journal of Computational Physics, 76(2):385– 413, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901489.

Shoucri:1983:NCD

[Sho83] Magdi M. Shoucri. Numerical calculations of discontinuities by shape preserving splines. Journal of Computational Physics, 49(2):334–341, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901304.

Shoucri:1986:NSS

[Sho86] Magdi M. Shoucri. Numerical solution of the shallow water equations. Journal of Computational Physics, 63(1):240–245, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900951.

Shumaker:1984:NCC

[Shu84] D. E. Shumaker. Numerical calculation of compact torus equilibria. Journal of Computational Physics, 53(3):456–471, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900718.

Shyy:1985:SFD

[Shy85] Wei Shyy. A study of finite difference approximations to steady-state, convection-dominated flow problems. Journal of Computational Physics, 57(3):415–438, February 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901883.

Sidi:1985:NEC

[Sid85] Avram Sidi. Numerical evaluation of a Cauchy principal value integral that arises in a problem involving the generation of instability waves. Journal of Computational Physics, 59(3):388– 395, July 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185901184.

Schmidt:1988:ALG

[SJ88] G. H. Schmidt and F. J. Jacobs. Adaptive local grid refinement and multi-grid in numerical reservoir simulation. Journal of Computational Physics, 77(1):140–165, July 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890160X.

Schnack:1980:NTD

[SK80] D. Schnack and J. Killeen. Nonlinear, two-dimensional magnetohydrodynamic calculations. Journal of Computational Physics, 35(1):110–145, March 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900388.

Schreiber:1983:DCF

[SK83a] R. Schreiber and H. B. Keller. Driven cavity flows by efficient numerical techniques. Journal of Computational Physics, 49(2):310– 333, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901298.

Schreiber:1983:SSD

[SK83b] R. Schreiber and H. B. Keller. Spurious solutions in driven cavity calculations. Journal of Computational Physics, 49(1):165– 172, January 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901195.

Smooke:1986:TDF

[SK86] M. D. Smooke and M. L. Koszykowski. Two-dimensional fully adaptive solutions of solid-solid alloying reactions. Journal of Computational Physics, 62(1):1–25, January 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900975.

Scandrett:1986:SPC

[SKA86] C. L. Scandrett, G. A. Kriegsmann, and J. D. Achenbach. Scattering of a pulse by a cavity in an elastic half-space. Journal of Computational Physics, 65(2):410–431, August 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902159.

Skeel:1981:APC

[Ske81] Robert D. Skeel. Accuracy of point cyclic reductions for Poisson’s equation. Journal of Computational Physics, 44(1):198–205, November 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181900450.

Skollermo:1982:BDS

[Skö82] Anders Sköllermo. A better difference scheme for the Laplace equation in cylindrical coordinates. Journal of Computational Physics, 47(1):160–163, July 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900705.

Shestakov:1982:NMN

[SKS82] A. I. Shestakov, J. Killeen, and D. D. Schnack. A numerical model for nonaxisymmetric MHD instabilities. Journal of Computational Physics, 46(1):69–96, August 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900067.

StMary:1985:AIF

[SL85] D. F. St. Mary and Ding Lee. Analysis of an implicit finite difference solution to an underwater wave propagation problem. Journal of Computational Physics, 57(3):378–390, February 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901858.

Sullivan:1987:FES

[SLO87] John M. Sullivan, Jr., Daniel R. Lynch, and Kevin O’Neill. Finite element simulation of planar instabilities during solidification of an undercooled melt. Journal of Computational Physics, 69(1): 81–111, March 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187901574.

Sloan:1988:MID

[Slo88] D. M. Sloan. On modulational instabilities in discretisations of the Korteweg–de Vries equation. Journal of Computational Physics, 79(1):167–183, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900095.

Shestakov:1984:FMT

[SM84] A. I. Shestakov and A. A. Mirin. A Fourier method for three-dimensional partial differential equations in periodic geometry. Application: HELIAC. Journal of Computational Physics, 56(3):469–494, December 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901086.

Sloan:1986:NIL

[SM86] D. M. Sloan and A. R. Mitchell. On nonlinear instabilities in leap-frog finite difference schemes. Journal of Computational Physics, 67(2):372–395, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918690269X.

Steinle:1989:IFC

[SM89] P. Steinle and R. Morrow. An implicit flux-corrected transport algorithm. Journal of Computational Physics, 80(1):61–71, January 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900909.

Smith:1981:SIP

[Smi81] Jeffrey B. Smith. Shape instabilities and pattern formation in solidification: A new method for numerical solution of the moving boundary problem. Journal of Computational Physics, 39(1):112– 127, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901406.

Smith:1988:GNP

[Smi88] B. L. Smith. A general numerical procedure for the treatment of moving interfaces in implicit continuous Eulerian (ICE) hydrodynamics. Journal of Computational Physics, 77(2):384–424, August 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901751.

Smooke:1982:SBS

[Smo82] Mitchell D. Smooke. Solution of burner-stabilized premixed laminar flames by boundary value methods. Journal of Computational Physics, 48(1):72–105, October 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900365.

Smolarkiewicz:1984:FMP

[Smo84] Piotr K. Smolarkiewicz. A fully multidimensional positive definite advection transport algorithm with small implicit diffusion. Journal of Computational Physics, 54(2):325–362, May 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901219.

Sullivan:1985:MDV

[SMO85] Francis Sullivan, Raymond D. Mountain, and Julie O’Connell. Molecular dynamics on vector computers. Journal of Computational Physics, 61(1):138–153, October 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900658.

Steinle:1989:UIE

[SMR89] P. Steinle, R. Morrow, and A. J. Roberts. Use of implicit and explicit flux-corrected transport algorithms in gas discharge problems involving non-uniform velocity fields. Journal of Computational Physics, 85(2):493–499, December 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901629.

Salvetti:1985:NME

[SMRS85] M. Salvetti, R. Montagnani, P. Riani, and O. Salvetti. A numerical method for evaluating matrix elements of Hermitian operators. Journal of Computational Physics, 60(3):467–477, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900312.

Scherz:1983:SAH

[SN83] U. Scherz and B. Nestler. Symmetry adapted harmonic oscillator functions. Journal of Computational Physics, 50(3):436– 446, June 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901055.

Strikwerda:1988:NMI

[SN88] John C. Strikwerda and Yvonne M. Nagel. A numerical method for the incompressible Navier–Stokes equations in three-dimensional cylindrical geometry. Journal of Computational Physics, 78(1):64–78, September 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890037X.

Shu:1988:EIE

[SO88] Chi-Wang Shu and Stanley Osher. Efficient implementation of essentially non-oscillatory shock-capturing schemes. Journal of Computational Physics, 77(2):439–471, August 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901775.

Shu:1989:EIE

[SO89] Chi-Wang Shu and Stanley Osher. Efficient implementation of essentially non-oscillatory shock-capturing schemes, II. Journal of Computational Physics, 83(1):32–78, July 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902222.

Soh:1987:TMS

[Soh87] W. Y. Soh. Time-marching solution of incompressible Navier– Stokes equations for internal flow. Journal of Computational Physics, 70(1):232–252, May 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918790012X.

Sonnenschein:1985:IAM

[Son85] Roland Sonnenschein. An improved algorithm for molecular dynamics simulation of rigid molecules. Journal of Computational Physics, 59(2):347–350, June 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901512.

Skamarock:1989:AGR

[SOS89] William Skamarock, Joseph Oliger, and Robert L. Street. Adaptive grid refinement for numerical weather prediction. Journal of Computational Physics, 80(1):27–60, January 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900892.

Shapiro:1981:AUW

[SP81] Allen Shapiro and George F. Pinder. Analysis of an upstream weighted collocation approximation to the transport equation. Journal of Computational Physics, 39(1):46–71, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181901364.

Sarkar:1987:CSP

[SP87a] P. K. Sarkar and M. A. Prasad. A comparative study of pseudo and quasi random sequences for the solution of integral equations. Journal of Computational Physics, 68(1):66–88, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999187900453.

Siggia:1987:RCM

[SP87b] Eric D. Siggia and Alain Pumir. A rapidly convergent method for the inversion of separable, positive, self-adjoint discrete elliptic operators in three or more dimensions. Journal of Computational Physics, 72(2):498–500, October 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900969.

Sparis:1985:MGB

[Spa85] Panagiotis Demetriou Sparis. A method for generating boundaryorthogonal curvilinear coordinate systems using the biharmonic equation. Journal of Computational Physics, 61(3):445–462, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900749.

Speziale:1987:AVV

[Spe87] Charles G. Speziale. On the advantages of the vorticity-velocity formulation of the equations of fluid dynamics. Journal of Computational Physics, 73(2):476–480, December 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901495.

Sartoretto:1989:ASI

[SPG89] Flavio Sartoretto, Giorgio Pini, and Giuseppe Gambolati. Accelerated simultaneous iterations for large finite element eigenproblems. Journal of Computational Physics, 81(1):53–69, March 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900648.

Spigler:1988:NSC

[Spi88a] Renato Spigler. Numerical simulation for certain stochastic ordinary differential equations. Journal of Computational Physics, 74(1):244–262, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188900794.

Spigler:1988:NTB

[Spi88b] Renato Spigler. Numerical treatment of a boundary-value problem for a certain singular parabolic partial differential equation. Journal of Computational Physics, 74(1):233–243, January 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900782.

Steinberg:1985:SMC

[SR85] Stanly Steinberg and Patrick J. Roache. Symbolic manipulation and computational fluid dynamics. Journal of Computational Physics, 57(2):251–284, January 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900452.

Silliman:1980:SHN

[SS80] William J. Silliman and L. E. Scriven. Separating how near a static contact line: Slip at a wall and shape of a free surface. Journal of Computational Physics, 34(3):287–313, March 1, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900911.

Saito:1981:SCF

[SS81] H. Saito and L. E. Scriven. Study of coating flow by the finite element method. Journal of Computational Physics, 42(1): 53–76, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902321.

Sanz-Serna:1982:EFD

[SS82a] J. M. Sanz-Serna. An explicit finite-difference scheme with exact conservation properties. Journal of Computational Physics, 47(2):199–210, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999182900742.

Stevens:1982:TDS

[SS82b] W. K. Stevens and C. H. Su. Three-dimensional steady stratified flows: A numerical approach. Journal of Computational Physics, 46(3):397–422, June 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900237.

Sack:1984:SAO

[SS84a] Ch Sack and H. Schamel. SUNION-an algorithm for onedimensional laser-plasma interaction. Journal of Computational Physics, 53(3):395–428, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184900688.

Stephens:1984:EDS

[SS84b] A. B. Stephens and G. R. Shubin. Exponentially derived switching schemes for inviscid flow. Journal of Computational Physics, 55(2):175–191, August 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900019.

Smith:1985:WIB

[SS85] P. A. Smith and P. K. Stansby. Wave-induced bed flows by a Lagrangian vortex scheme. Journal of Computational Physics, 60(3):489–516, September 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900336.

Sakai:1986:FCT

[SS86] Katsuhiro Sakai and Tamotsu Sekiya. A fast computing technique for diffusion-type equations. Journal of Computational Physics, 65(2):462–473, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902184.

Santosa:1988:HFP

[SS88a] Fadil Santosa and William W. Symes. High-frequency perturbational analysis of the surface point-source response of a layered fluid. Journal of Computational Physics, 74(2):318–381, February 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900836.

Schumann:1988:FFT

[SS88b] Ulrich Schumann and Roland A. Sweet. Fast Fourier transforms for direct solution of Poisson’s equation with staggered boundary conditions. Journal of Computational Physics, 75(1):123– 137, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901027.

Smith:1989:ESA

[SS89] P. A. Smith and P. K. Stansby. An efficient surface algorithm for random-particle simulation of vorticity and heat transport. Journal of Computational Physics, 81(2):349–371, April 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918990212X.

Sanz-Serna:1981:PGM

[SSC81] J. M. Sanz-Serna and I. Christie. Petrov–Galerkin methods for nonlinear dispersive waves. Journal of Computational Physics, 39(1):94–102, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181901388.

Sanz-Serna:1986:SAT

[SSC86] J. M. Sanz-Serna and I. Christie. A simple adaptive technique for nonlinear wave problems. Journal of Computational Physics, 67(2):348–360, December 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902676.

Sulem:1983:TCS

[SSF83] Catherine Sulem, Pierre-Louis Sulem, and Hélène Frisch. Tracing complex singularities with spectral methods. Journal of Computational Physics, 50(1):138–161, April 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900451.

Shubin:1981:SST

[SSG81] G. R. Shubin, A. B. Stephens, and H. M. Glaz. Steady shock tracking and Newton’s method applied to one-dimensional duct flow. Journal of Computational Physics, 39(2):364–374, February 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181901571.

Sanz-Serna:1983:MIT

[SSM83] J. M. Sanz-Serna and V. S. Manoranjan. A method for the integration in time of certain partial differential equations. Journal of Computational Physics, 52(2):273–289, November 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900311.

Similon:1984:GCD

[SSS^{+}84] P. Similon, J. E. Sedlak, D. Stotler, H. L. Berk, W. Horton, and D. Choi. Guiding-center dispersion function. Journal of Computational Physics, 54(2):260–277, May 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901189.

Shih:1983:BCS

[SSSR83] Tom I-P Shih, Gene E. Smith, George S. Springer, and Y. Rimon. Boundary conditions for the solution of compressible Navier–Stokes equations by an implicit factored method. Journal of Computational Physics, 52(1):54–79, October 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900165.

Smith:1988:CSR

[ST88] P. Smith and S. Torquato. Computer simulation results for the two-point probability function of composite media. Journal of Computational Physics, 76(1):176–191, May 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901362.

Salmon:1989:GAJ

[ST89a] Rick Salmon and Lynne D. Talley. Generalizations of Arakawa’s Jacobian. Journal of Computational Physics, 83(2):247–259, August 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901186.

Solomonoff:1989:GPP

[ST89b] A. Solomonoff and E. Turkel. Global properties of pseudospectral methods. Journal of Computational Physics, 81(2):239–276, April 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189902088.

Stewart:1981:FSM

[Ste81] H. Bruce Stewart. Fractional step methods for thermohydraulic calculation. Journal of Computational Physics, 40(1): 77–90, March 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918190200X.

Stephenson:1984:SCD

[Ste84] J. W. Stephenson. Single cell discretizations of order two and four for biharmonic problems. Journal of Computational Physics, 55 (1):65–80, July 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900159.

Steinhoff:1986:BMG

[Ste86] John Steinhoff. Blending method for grid generation. Journal of Computational Physics, 65(2):370–385, August 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902135.

Steppeler:1987:ECG

[Ste87a] J. Steppeler. Energy conserving Galerkin finite element schemes for the primitive equations of numerical weather prediction. Journal of Computational Physics, 69(1):258–264, March 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901665.

Sternberg:1987:SMN

[Ste87b] Natalia Sternberg. A spectral method for nonlinear wave equations. Journal of Computational Physics, 72(2):422–434, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918790091X.

Storchi:1984:AMC

[Sto84] P. R. M. Storchi. Application of the P_{n}-method to the calculation of the annular flux of gamma rays. Journal of Computational Physics, 55(1):81–97, July 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900160.

Storer:1986:NST

[Sto86] R. G. Storer. Numerical studies of toroidal resistive magnetohydrodynamic instabilities. Journal of Computational Physics, 66(2):294–323, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900689.

Strikwerda:1980:IBV

[Str80] John C. Strikwerda. Initial boundary value problems for the method of lines. Journal of Computational Physics, 34(1):94– 107, January 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918090114X.

Strikwerda:1982:UDF

[Str82] John C. Strikwerda. Upwind differencing, false scaling, and nonphysical solutions to the driven cavity problem. Journal of Computational Physics, 47(2):303–307, August 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900821.

Strain:1989:BIA

[Str89] John Strain. A boundary integral approach to unstable solidification. Journal of Computational Physics, 85(2):342–389, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901551.

Summers:1989:RVS

[Sum89] D. M. Summers. A random vortex simulation of Falkner–Skan boundary layer flow. Journal of Computational Physics, 85(1):86– 103, November 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189902015.

Storchi:1989:NMC

[SV89] P. Storchi and R. J. Van Der Linden. A numerical method for the calculation of the diffusion of high energy electrons in a heterogeneous medium. Journal of Computational Physics, 85(2):417– 433, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901575.

Steger:1981:FVS

[SW81] Joseph L. Steger and R. F. Warming. Flux vector splitting of the inviscid gasdynamic equations with application to finite-difference methods. Journal of Computational Physics, 40(2):263–293, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902102.

Stewart:1984:TPF

[SW84] H. Bruce Stewart and Burton Wendroff. Two-phase flow: Models and methods. Journal of Computational Physics, 56(3):363–409, December 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901037.

Sidorowich:1985:MSS

[SW85] J. J. Sidorowich and W. H. Wells. Multiple scattering in stratified inhomogeneous media. Journal of Computational Physics, 57(1):137–154, January 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900567.

Seldner:1988:AIL

[SW88] David Seldner and Thomas Westermann. Algorithms for interpolation and localization in irregular 2D meshes. Journal of Computational Physics, 79(1):1–11, November 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900010.

Streett:1985:SMM

[SZH85] Craig L. Streett, Thomas A. Zang, and M. Yousuff Hussaini. Spectral multigrid methods with applications to transonic potential flow. Journal of Computational Physics, 57(1):43–76, January 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918590052X.

Taha:1984:ANAb

[TA84a] Thiab R. Taha and Mark I. Ablowitz. Analytical and numerical aspects of certain nonlinear evolution equations. II.

Numerical, nonlinear Schrödinger equation. Journal of Computational Physics, 55(2):203–230, August 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900032.

Taha:1984:ANAc

[TA84b] Thiab R. Taha and Mark I. Ablowitz. Analytical and numerical aspects of certain nonlinear evolution equations. III. Numerical, Korteweg–de Vries equation. Journal of Computational Physics, 55(2):231–253, August 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900044.

Taha:1984:ANAa

[TA84c] Thiab R. Taha and Mark J. Ablowitz. Analytical and numerical aspects of certain nonlinear evolution equations. I. Analytical. Journal of Computational Physics, 55(2):192–202, August 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900020.

Taasan:1987:MME

[Ta’87] Shlomo Ta’asan. Multigrid method for the equilibrium equations of elasticity using a compact scheme. Journal of Computational Physics, 73(2):432–446, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901458.

Taha:1988:ANA

[TA88] Thiab R. Taha and Mark J. Ablowitz. Analytical and numerical aspects of certain nonlinear evolution equations IV. Numerical, modified Korteweg–de Vries equation. Journal of Computational Physics, 77(2):540–548, August 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901842.

Thomas:1981:CNM

[TAJ^{+}81] L. D. Thomas, M. H. Alexander, B. R. Johnson, W. A. Lester, Jr., J. C. Light, K. D. McLenithan, G. A. Parker, M. J. Redmon, T. G. Schmalz, D. Secrest, and R. B. Walker. Comparison of numerical methods for solving the second-order differential equations of molecular scattering theory. Journal of Computational Physics, 41(2):407–426, June 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181901030.

Takemitsu:1980:FDA

[Tak80] Nobumasa Takemitsu. On a finite-difference approximation for the steady-state Navier–Stokes equations. Journal of Computational Physics, 36(2):236–248, July 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901862.

Takada:1981:SME

[Tak81] Toshikazu Takada. A simple method to eliminate linearly dependent AO integrals in ab initio LCAO MO calculations by taking advantage of shell structure. Journal of Computational Physics, 42(1):99–107, July 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902345.

Takemitsu:1985:FDM

[Tak85] Nobumasa Takemitsu. Finite difference method to solve incompressible fluid flow. Journal of Computational Physics, 61(3):499– 518, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999185900774.

Takano:1986:ARC

[Tak86] Yasunari Takano. An application of the random choice method to reactive gas with many chemical species. Journal of Computational Physics, 67(1):173–187, November 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901208.

Talman:1980:ECE

[Tal80] James D. Talman. Eigenvalue correction estimates for the one-dimensional Schrödinger equation. Journal of Computational Physics, 37(1):19–40, August 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900029.

Tamor:1981:ACC

[Tam81] S. Tamor. ANTIC: A code for calculation of neutral transport in cylindrical plasmas. Journal of Computational Physics, 40(1):104–119, March 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181902023.

Tan:1985:AST

[Tan85] C. S. Tan. Accurate solution of three-dimensional Poisson’s equation in cylindrical coordinate by expansion in Chebyshev polynomials. Journal of Computational Physics, 59(1):81–95, May 30, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185901081.

Tan:1986:ACI

[Tan86] Roger C. E. Tan. Accelerating the convergence of an iterative method for derivatives of eigensystems. Journal of Computational Physics, 67(1):230–235, November 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901245.

Tanaka:1988:MIE

[Tan88] Motohiko Tanaka. Macroscale implicit electromagnetic particle simulation of magnetized plasmas. Journal of Computational Physics, 79(1):209–226, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900125.

Tan:1989:MDS

[Tan89] C. S. Tan. A multi-domain spectral computation of threedimensional laminar horseshoe vortex flow using incompressible Navier–Stokes equations. Journal of Computational Physics, 85(1):130–158, November 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189902039.

Thompkins:1982:BTI

[TB82] W. T. Thompkins, Jr. and R. H. Bush. Boundary treatments for implicit solutions to Euler and Navier–Stokes equations. Journal of Computational Physics, 48(2):302–311, November 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900523.

Thompson:1986:CBF

[TB86] I. J. Thompson and A. R. Barnett. Coulomb and Bessel functions of complex arguments and order. Journal of Computational Physics, 64(2):490–509, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918690046X.

Theodoropoulos:1989:LEM

[TB89] T. Theodoropoulos and G. C. Bergeles. A Laplacian equation method for numerical generation of boundary-fitted 3D orthogonal grids. Journal of Computational Physics, 82(2):269–288, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900491.

Thorlacius:1987:AIS

[TC87] A. E. Thorlacius and E. D. Cooper. An algorithm for integrating the Schrödinger equation. Journal of Computational Physics, 72(1):70–77, September 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900738.

Teran:1989:CNM

[TDhLcSb89] Luis Mier Y. Teran, Enrique Díaz-herrera, Marcelo Lozadacassou, and Rafaél Saavedra-barrera. A comparison of numerical methods for solving nonlinear integral equations found in liquid theories. Journal of Computational Physics, 84(2):326– 342, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189902362.

Tal-Ezer:1986:PLM

[TE86] Hillel Tal-Ezer. A pseudospectral Legendre method for hyperbolic equations with an improved stability condition. Journal of Computational Physics, 67(1):145–172, November 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901191.

Temperton:1980:FAD

[Tem80] Clive Temperton. On the FACR(l ) algorithm for the discrete Poisson equation. Journal of Computational Physics, 34(3):314– 329, March 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999180900923.

Temperton:1983:FMR

[Tem83a] Clive Temperton. Fast mixed-radix real Fourier transforms. Journal of Computational Physics, 52(2):340–350, November 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900347.

Temperton:1983:NNP

[Tem83b] Clive Temperton. Note a note on prime factor FFT algorithms. Journal of Computational Physics, 52(1):198–204, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900244.

Temperton:1983:SSM

[Tem83c] Clive Temperton. Self-sorting mixed-radix Fast Fourier Transforms. Journal of Computational Physics, 52(1):1–23, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918390013X.

Temperton:1985:ISS

[Tem85] Clive Temperton. Implementation of a self-sorting in-place prime factor FFT algorithm. Journal of Computational Physics, 58(3):283–299, May 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185901640.

Temperton:1988:C

[Tem88a] Clive Temperton. Corrigendum. Journal of Computational Physics, 79(2):486, December 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188900307.

Temperton:1988:NSM

[Tem88b] Clive Temperton. A new set of minimum-add small-n rotated DFT modules. Journal of Computational Physics, 75(1):190– 198, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901064.

Temperton:1988:SSP

[Tem88c] Clive Temperton. A self-sorting in-place prime factor real/halfcomplex FFT algorithm. Journal of Computational Physics, 75(1):199–216, March 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999188901076.

Temperton:1989:NSP

[Tem89] Clive Temperton. Nesting strategies for prime factor FFT algorithms. Journal of Computational Physics, 82(2):247–268, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918990048X.

Teng:1982:EVM

[Ten82] Zhen-Huan Teng. Elliptic-vortex method for incompressible flow at high Reynolds number. Journal of Computational Physics, 46(1):54–68, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900055.

Terras:1981:ASI

[Ter81a] Riho Terras. Algorithms for some integrals of Bessel functions and multivariate Gaussian integrals. Journal of Computational Physics, 41(1):192–199, May 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181900875.

Terras:1981:GEO

[Ter81b] Riho Terras. Generalized exponential operators in the continuation of the confluent hypergeometric functions. Journal of Computational Physics, 44(1):156–166, November 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900425.

Terras:1981:MAI

[Ter81c] Riho Terras. A Miller algorithm for an incomplete Bessel function. Journal of Computational Physics, 39(1):233–240, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181901479.

Thompson:1989:AMT

[TF89] M. C. Thompson and J. H. Ferziger. An adaptive multigrid technique for the incompressible Navier–Stokes equations. Journal of Computational Physics, 82(1):94–121, May 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900375.

Thacker:1980:MAC

[TGP80] W. C. Thacker, A. Gonzalez, and G. E. Putland. A method for automating the construction of irregular computational grids for storm surge forecast models. Journal of Computational Physics, 37(3):371–387, October 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900431.

Thacker:1980:GFD

[Tha80] W. C. Thacker. A geodesic finite-difference method for curved domains: Simulations of tidal motion on a sphere. Journal of Computational Physics, 37(3):355–370, October 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918090042X.

Thompson:1987:TDB

[Tho87] Kevin W. Thompson. Time dependent boundary conditions for hyperbolic systems. Journal of Computational Physics, 68(1):1– 24, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900416.

Tajima:1981:ABC

[TL81] T. Tajima and Y. C. Lee. Absorbing boundary condition and Budden turning point technique for electromagnetic plasma simulations. Journal of Computational Physics, 42(2):406–412, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902540.

Takle:1984:ALL

[TL84] E. S. Takle and J. M. Leone, Jr. Application of a log-linear element to a finite-element boundary-layer flow model. Journal of Computational Physics, 53(3):513–523, March 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900743.

Tajima:1980:MPC

[TLD80] T. Tajima, J. N. Leboeuf, and J. M. Dawson. A magnetohydrodynamic particle code with force free electrons for fluid simulations. Journal of Computational Physics, 38(2):237–250, November 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900558.

Tourigny:1988:IEP

[TM88] Y. Tourigny and J. Ll Morris. An investigation into the effect of product approximation in the numerical solution of the cubic nonlinear Schrödinger equation. Journal of Computational Physics, 76(1):103–130, May 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901337.

Takagi:1985:NGB

[TMCS85] Toshiyuki Takagi, Kazuyoshi Miki, Brian C. J. Chen, and William T. Sha. Numerical generation of boundary-fitted curvilinear coordinate systems for arbitrarily curved surfaces. Journal of Computational Physics, 58(1):67–79, March 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901573.

Takewaki:1985:CIP

[TNY85] H. Takewaki, A. Nishiguchi, and T. Yabe. Cubic interpolated pseudo-particle method (CIP) for solving hyperbolic-type equations. Journal of Computational Physics, 61(2):261–268, November 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900853.

Tobocman:1986:IIS

[Tob86] W. Tobocman. Iterative inverse scattering method employing Gram–Schmidt orthogonalization. Journal of Computational Physics, 64(1):230–245, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900276.

Tanemura:1983:NAT

[TOO83] Masaharu Tanemura, Tohru Ogawa, and Naofumi Ogita. A new algorithm for three-dimensional Voronoi tessellation. Journal of Computational Physics, 51(2):191–207, August 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900876.

Toxvaerd:1982:NAM

[Tox82] Søren Toxvaerd. A new algorithm for molecular dynamics calculations. Journal of Computational Physics, 47(3):444–451, September 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999182900936.

Toxvaerd:1983:ECM

[Tox83] Søren Toxvaerd. Energy conservation in molecular dynamics. Journal of Computational Physics, 52(1):214–216, October 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918390027X.

Trapp:1986:NIH

[TR86] John A. Trapp and Richard A. Riemke. A nearly-implicit hydrodynamic numerical scheme for two-phase flows. Journal of Computational Physics, 66(1):62–82, September 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900549.

Treve:1981:NTR

[Tre81] Yvain M. Treve. A numerical test of the reliability of Galerkin approximations to the solutions of the Navier–Stokes equation. Journal of Computational Physics, 41(1):217–222, May 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181900917.

Trefethen:1983:GVI

[Tre83] Lloyd N. Trefethen. Group velocity interpretation of the stability theory of Gustafsson, Kreiss, and Sundström. Journal of Computational Physics, 49(2):199–217, February 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901237.

Tryggvason:1988:NSR

[Try88] Grétar Tryggvason. Numerical simulations of the Rayleigh– Taylor instability. Journal of Computational Physics, 75(2):253– 282, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918890112X.

Tryggvason:1989:SVS

[Try89] Grétar Tryggvason. Simulation of vortex sheet roll-up by vortex methods. Journal of Computational Physics, 80(1):1–16, January 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189900879.

Takeuchi:1981:TDP

[TS81] K. Takeuchi and A. C. Spencer. Two-dimensional predictorcorrector method of characteristics for fluid-structure interactions. Journal of Computational Physics, 40(1):143–166, March 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902047.

Turnbull:1983:PRD

[TS83] A. D. Turnbull and R. G. Storer. A plasma resistive diffusion model. Journal of Computational Physics, 50(3):409–435, June 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901043.

Turnbull:1986:LML

[TST^{+}86] A. D. Turnbull, M. A. Secrétan, F. Troyon, S. Semenzato, and R. Gruber. β limits in H-mode-like discharges. Journal of Computational Physics, 66(2):391–410, October 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900720.

Tuckerman:1989:DFV

[Tuc89a] Laurette S. Tuckerman. Divergence-free velocity fields in nonperiodic geometries. Journal of Computational Physics, 80(2):403– 441, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901083.

Tuckerman:1989:TMB

[Tuc89b] Laurette S. Tuckerman. Transformations of matrices into banded form. Journal of Computational Physics, 84(2):360–376, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189902386.

Turkel:1980:PUH

[Tur80] Eli Turkel. On the practical use of high-order methods for hyperbolic systems. Journal of Computational Physics, 35(3):319–340, May 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199918090073X.

Turkel:1987:PMS

[Tur87] Eli Turkel. Preconditioned methods for solving the incompressible and low speed compressible equations. Journal of Computational Physics, 72(2):277–298, October 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900842.

Tutty:1986:VPV

[Tut86] O. R. Tutty. On vector potential-vorticity methods for incompressible flow problems. Journal of Computational Physics, 64(2):

368–379, June 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900380.

Thompson:1982:BFC

[TWM82] Joe F. Thompson, Zahir U. A. Warsi, and C. Wayne Mastin. Boundary-fitted coordinate systems for numerical solution of partial differential equations — a review. Journal of Computational Physics, 47(1):1–108, July 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900663.

Takewaki:1987:CIP

[TY87a] Hideaki Takewaki and Takashi Yabe. The cubic-interpolated pseudo particle (CIP) method: application to nonlinear and multi-dimensional hyperbolic equations. Journal of Computational Physics, 70(2):355–372, June 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901872.

Tang:1987:FSC

[TY87b] Tao Tang and Wei Yuan. The further study of a certain nonlinear integro-differential equation. Journal of Computational Physics, 72(2):486–497, October 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900957.

Uchikawa:1983:GBF

[Uch83] Sadao Uchikawa. Generation of boundary-fitted curvilinear coordinate systems for a two-dimensional axisymmetric flow problem. Journal of Computational Physics, 50(2):316–321, May 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900694.

VanRun:1984:OTS

[Van84] A. M. J. G. Van Run. An optimum time step length for convergence to steady-state solution in compressible-flow calculations. Journal of Computational Physics, 56(2):356–360, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184901013.

Vanka:1986:BIM

[Van86] S. P. Vanka. Block-implicit multigrid solution of Navier–Stokes equations in primitive variables. Journal of Computational Physics, 65(1):138–158, July 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900082.

Varah:1984:CML

[Var84] J. M. Varah. Computational methods in linear algebra. Journal of Computational Physics, 54(1):87–94, April 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901402.

Vatsya:1987:RAS

[Vat87] S. R. Vatsya. A representation and approximation of the solutions of hyperbolic differential equations. Journal of Computational Physics, 73(2):289–305, December 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901380.

Vatsya:1989:ESN

[Vat89] S. R. Vatsya. Existence of the solution of a nonlinear integrodifferential equation. Journal of Computational Physics, 82(1): 241–244, May 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189900454.

Verwer:1989:AMG

[VBSS89] J. G. Verwer, J. G. Blom, and J. M. Sanz-Serna. An adaptive moving grid method for one-dimensional systems of partial differential equations. Journal of Computational Physics, 82(2):454– 486, June 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189900582.

Vuik:1985:NSE

[VC85] C. Vuik and C. Cuvelier. Numerical solution of an etching problem. Journal of Computational Physics, 59(2):247–263, June 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185901457.

vanDommelen:1989:FAS

[vDR89] Leon van Dommelen and Elke A. Rundensteiner. Fast, adaptive summation of point forces in the two-dimensional Poisson equation. Journal of Computational Physics, 83(1):126–147, July 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189902258.

vanDommelen:1980:SGS

[vDS80] L. L. van Dommelen and S. F. Shen. The spontaneous generation of the singularity in a separating laminar boundary layer. Journal of Computational Physics, 38(2):125–140, November 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900492.

vanderVorst:1981:ISM

[vdV81] Henk A. van der Vorst. Iterative solution methods for certain sparse linear systems with a non-symmetric matrix arising from PDE-problems. Journal of Computational Physics, 44(1):1–19, November 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181900346.

Valeo:1987:NML

[VE87] Ernest J. Valeo and David C. Eder. Numerical modeling of lower hybrid heating and current drive. Journal of Computational Physics, 69(2):341–361, April 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901720.

Vesely:1982:AMC

[Ves82] Franz J. Vesely. Angular Monte Carlo integration using quaternion parameters: a spherical reference potential for CCl_{4}. Journal of Computational Physics, 47(2):291–296, August 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900808.

Varoglu:1980:FEI

[VF80] Erol Varoglu and W. D. Liam Finn. Finite elements incorporating characteristics for one-dimensional diffusion-convection equation. Journal of Computational Physics, 34(3):371–389, March 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900959.

Vichnevetsky:1986:ITC

[Vic86] R. Vichnevetsky. Invariance theorems concerning reflection at numerical boundaries. Journal of Computational Physics, 63(2): 268–282, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186901944.

Viecelli:1983:EDO

[Vie83] J. A. Viecelli. Exponential difference operator approximation for the sixth order Onsager equation. Journal of Computational Physics, 50(1):162–170, April 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900463.

Viecelli:1984:SOE

[Vie84] J. A. Viecelli. Solution of the Onsager equation in doubly connected regions. Journal of Computational Physics, 56(3):530–536, December 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901116.

Viecelli:1988:ELR

[Vie88] James A. Viecelli. Estimating limits on resolution and predictability in computations of hydrodynamically unstable flows. Journal of Computational Physics, 74(2):456–476, February 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900885.

Vijayasundaram:1986:TFS

[Vij86] G. Vijayasundaram. Transonic flow simulations using an upstream centered scheme of Godunov in finite elements. Journal of Computational Physics, 63(2):416–433, April 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902020.

Vinokur:1983:ODS

[Vin83] Marcel Vinokur. On one-dimensional stretching functions for finite-difference calculations. Journal of Computational Physics, 50(2):215–234, May 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900657.

Vinokur:1989:AFD

[Vin89] Marcel Vinokur. An analysis of finite-difference and finite-volume formulations of conservation laws. Journal of Computational Physics, 81(1):1–52, March 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999189900636.

VanDooren:1989:PDS

[VJ89] R. Van Dooren and H. Janssen. Period doubling solutions in the Duffing oscillator: A Galerkin approach. Journal of Computational Physics, 82(1):161–171, May 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900405.

Vanderbilt:1984:MCS

[VL84] David Vanderbilt and Steven G. Louie. A Monte Carlo simulated annealing approach to optimization over continuous variables. Journal of Computational Physics, 56(2):259–271, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900950.

Vandromme:1986:ACT

[VM86] D. Vandromme and H. Ha Minh. About the coupling of turbulence closure models with averaged Navier–Stokes equations. Journal of Computational Physics, 65(2):386–409, August 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186902147.

Vrbik:1986:QAD

[VR86] Jan Vrbik and Stuart M. Rothstein. Quadratic accuracy diffusion Monte Carlo. Journal of Computational Physics, 63(1):130– 139, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900872.

Vadillo:1986:SNN

[VSS86] F. Vadillo and J. M. Sanz-Serna. Studies in numerical nonlinear instability. II. A new look at u_{t} + uu_{x} = 0. Journal of Computational Physics, 66(1):225–238, September 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900628.

Vanajakshi:1989:BVP

[VTB89] T. C. Vanajakshi, Kevin W. Thompson, and David C. Black. Boundary value problems in magnetohydrodynamics (and fluid dynamics). I. Radiation boundary condition. Journal of Computational Physics, 84(2):343–359, October 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902374.

VanDooren:1982:CSS

[VV82] R. Van Dooren and J. Vlassenbroeck. Chebyshev series solution of the controlled Duffing oscillator. Journal of Computational Physics, 47(2):321–329, August 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900857.

Wagner:1987:UMC

[Wag87] Wolfgang Wagner. Unbiased Monte Carlo evaluation of certain functional integrals. Journal of Computational Physics, 71(1): 21–33, July 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900179.

Wagner:1988:UMS

[Wag88] Wolfgang Wagner. Unbiased multi-step estimators for the Monte Carlo evaluation of certain functional integrals. Journal of Computational Physics, 79(2):336–352, December 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900204.

Walmsley:1984:EEM

[Wal84] John L. Walmsley. On the efficient evaluation of modified Bessel functions of zeroth and first orders for arguments of the form x exp(iπ/4). Journal of Computational Physics, 56(2):349–355, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901001.

Wang:1983:NMS

[Wan83] J. C. T. Wang. On the numerical methods for the singular parabolic equations in fluid dynamics. Journal of Computational Physics, 52(3):464–479, December 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900037.

Warsi:1986:NGG

[War86] Z. U. A. Warsi. Numerical grid generation in arbitrary surfaces through a second-order differential-geometric model. Journal of Computational Physics, 64(1):82–96, May 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900197.

Wallace:1986:IME

[WBF86] Jon M. Wallace, Jeremiah U. Brackbill, and David W. Forslund. An implicit moment electromagnetic plasma simulation in cylindrical coordinates. Journal of Computational Physics, 63(2):434– 457, April 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186902032.

Witmer:1986:FSP

[WBL86] George Witmer, Vemuri Balakotaiah, and Dan Luss. Finding singular points of two-point boundary value problems. Journal of Computational Physics, 65(1):244–250, July 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900148.

Woodward:1984:NST

[WC84] Paul Woodward and Phillip Colella. The numerical simulation of two-dimensional fluid flow with strong shocks. Journal of Computational Physics, 54(1):115–173, April 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901426.

Wilson:1985:PMW

[WC85a] Katherine A. Wilson and Robert L. Coldwell. A pairing method which improves convergence in Monte–Carlo estimation of quantum mechanical expectation values. Journal of Computational Physics, 61(3):483–498, December 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900762.

Winters:1985:CPF

[WC85b] K. H. Winters and K. A. Cliffe. Convergence properties of the finite-element method for Bénard Convection in an infinite layer. Journal of Computational Physics, 60(2):346–351, September 15, 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999185900129.

Whealton:1988:AME

[WCR^{+}88] J. H. Whealton, G. L. Chen, R. J. Raridon, R. W. McGaffey, E. F. Jaeger, M. A. Bell, and D. J. Hoffman. A 3D analysis of Maxwell’s equations for cavities of arbitrary shape. Journal of Computational Physics, 75(1):168–189, March 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901052.

Winkler:1984:ESM

[WD84] Joab R. Winkler and J. Brian Davies. Elimination of spurious modes in finite element analysis. Journal of Computational Physics, 56(1):1–14, October 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184900792.

Weiss:1985:EWP

[Wei85] Isaac Weiss. Electromagnetic wave propagation in relativistic magnetized plasmas. Journal of Computational Physics, 61(3):403–416, December 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999185900725.

Wesseling:1988:CCM

[Wes88] P. Wesseling. Cell-centered multigrid for interface problems. Journal of Computational Physics, 79(1):85–91, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188900058.

Wysocki:1986:NMC

[WG86] F. Wysocki and R. C. Grimm. A new method for computing normal modes in axisymmetric toroidal geometry using a scalar form of ideal MHD. Journal of Computational Physics, 66(2):255– 273, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999186900665.

William:1987:SET

[WH87] G. D. Venneri William and William G. Hoover. Simple exact test for well-known molecular dynamics algorithms. Journal of Computational Physics, 73(2):468–475, December 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901483.

Whealton:1981:EVS

[Whe81] J. H. Whealton. Expeditious Vlasov solver for computation of ion extraction from a plasma. Journal of Computational Physics, 40(2):491–496, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999181902242.

Whitley:1987:PFP

[Whi87] Norman L. Whitley. Potential flow past a sinusoidal wall by direct variation. Journal of Computational Physics, 68(2):462– 473, February 1987. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999187900660.

Wienke:1987:MPT

[WHW87] B. R. Wienke, T. R. Hill, and P. P. Whalen. Multigroup particle transport in a moving material. Journal of Computational Physics, 72(1):177–201, September 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900775.

Wienke:1983:SEE

[Wie83] B. R. Wienke. Suprathermal electron energy deposition in plasmas with the Fokker–Planck method. Journal of Computational Physics, 51(2):208–226, August 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183900888.

Wigley:1988:EMS

[Wig88] Neil M. Wigley. An efficient method for subtracting off singularities at corners for Laplace’s equation. Journal of Computational Physics, 78(2):369–377, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900551.

Wikramaratna:1989:ANM

[Wik89] Roy S. Wikramaratna. ACORN — A new method for generating sequences of uniformly distributed pseudo-random numbers. Journal of Computational Physics, 83(1):16–31, July 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189902210.

Wilkins:1980:UAV

[Wil80a] Mark L. Wilkins. Use of artificial viscosity in multidimensional fluid dynamic calculations. Journal of Computational Physics, 36(3):281–303, July 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180901618.

Willenberg:1980:MRF

[Wil80b] H. J. Willenberg. Multigroup representation of fusion product orbits in a plasma column. Journal of Computational Physics, 34(3):330–347, March 1, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900935.

Williamson:1980:LSR

[Wil80c] J. H. Williamson. Low-storage Runge–Kutta schemes. Journal of Computational Physics, 35(1):48–56, March 15, 1980. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999180900339.

Williamson:1981:FCI

[Wil81] J. H. Williamson. Four cheap improvements to the particle-mesh code. Journal of Computational Physics, 41(2):256–269, June 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181900966.

Wilsdon:1982:AME

[Wil82] C. E. Wilsdon. An approximate method for evaluating the ratio of two complete elliptic integrals of the first kind. Journal of Computational Physics, 46(1):166–167, August 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900092.

Witting:1984:UME

[Wit84] James M. Witting. A unified model for the evolution nonlinear water waves. Journal of Computational Physics, 56(2):203–236, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900925.

Wu:1989:NSP

[WK89] C. Wu and E. E. Kunhardt. Numerical solution of Poisson’s equation for rapidly varying driving functions. Journal of Computational Physics, 84(1):247–254, September 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901903.

Wienke:1984:FSP

[WL84] B. R. Wienke and B. L. Lathrop. Fast scheme for photon-Maxwellian electron cross sections. Journal of Computational Physics, 53(2):331–345, February 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184900457.

Whealton:1986:FDD

[WMM86] J. H. Whealton, R. W. McGaffey, and P. S. Meszaros. A finite difference 3-D Poisson–Vlasov algorithm for ions extracted from a plasma. Journal of Computational Physics, 63(1):20–32, March 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999186900823.

Wei:1988:NCS

[WMS88] Musheng Wei, George Majda, and Walter Strauss. Numerical computation of the scattering frequencies for acoustic wave equations. Journal of Computational Physics, 75(2):345–358, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901179.

Wolniewicz:1981:NII

[Wol81] L. Wolniewicz. Numerical integration of an inhomogeneous boundary value problem. Journal of Computational Physics, 40(2):440–452, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999181902217.

Wolf:1985:AAD

[Wol85] Thomas Wolf. An analytic algorithm for decoupling and integrating systems of nonlinear partial differential equations. Journal of Computational Physics, 60(3):437–446, September 30, 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900294.

Wu:1984:SSS

[WOZ84] H. M. Wu, E. A. Overman II, and N. J. Zabusky. Steady-state solutions of the Euler equations in two dimensions: Rotating and translating V -states with limiting cases. I. Numerical algorithms and results. Journal of Computational Physics, 53(1):42–71, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999184900512.

Weidman:1983:MAM

[WP83] Scott Weidman and Roger A. Pielke. A more accurate method for the numerical solution of nonlinear partial differential equations. Journal of Computational Physics, 49(2):342–348, February 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183901316.

Wyman:1988:DMA

[WP88] Douglas R. Wyman and Michael S. Patterson. A discrete method for anisotropic angular sampling in Monte Carlo simulations. Journal of Computational Physics, 76(2):414–425, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901490.

Wyman:1989:SRA

[WPW89] Douglas R. Wyman, Michael S. Patterson, and Brian C. Wilson. Similarity relations for anisotropic scattering in Monte Carlo simulations of deeply penetrating neutral particles. Journal of Computational Physics, 81(1):137–150, March 1989. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189900673.

Wong:1984:EVV

[WR84] A. K. Wong and J. A. Reizes. An effective vorticity-vector potential formulation for the numerical solution of three-dimensional duct flow problems. Journal of Computational Physics, 55(1): 98–114, July 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900172.

Wong:1986:VPN

[WR86] A. K. Wong and J. A. Reizes. The vector potential in the numerical solution of three-dimensional fluid dynamics problems in multiply connected regions. Journal of Computational Physics, 62(1):124–142, January 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186901038.

Wormer:1982:CGM

[WVP82] P. E. S. Wormer, F. Visser, and J. Paldus. Conjugate gradient method for the solution of linear equations: Application to molecular electronic structure calculations. Journal of Computational Physics, 48(1):23–44, October 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918290033X.

Wang:1989:HRT

[WW89] J. C. T. Wang and G. F. Widhopf. A high-resolution TVD finite volume scheme for the Euler equations in conservation form. Journal of Computational Physics, 84(1):145–173, September 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901861.

Wooten:1981:IEO

[WWM^{+}81] J. W. Wooten, J. H. Whealton, D. H. McCollough, R. W. Mc-Gaffey, J. E. Akin, and L. J. Drooks. Ion extraction and optics in 3D. Journal of Computational Physics, 43(1):95–110, September 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181901133.

Yamamoto:1983:EAC

[Yam83] H. Yamamoto. An efficient algorithm for calculating thrust in high multiplicity reactions. Journal of Computational Physics, 52(3):597–601, December 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999183900104.

Yee:1987:CEI

[Yee87] H. C. Yee. Construction of explicit and implicit symmetric TVD schemes and their applications. Journal of Computational Physics, 68(1):151–179, January 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187900490.

Yiantsios:1988:NSE

[YH88a] Stergios G. Yiantsios and Brian G. Higgins. Numerical solution of eigenvalue problems using the compound matrix method. Journal of Computational Physics, 74(1):25–40, January 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900666.

Yidana:1988:FRM

[YH88b] P. V. A. Yidana and K. Hartt. The formation of a rational S matrix using phase shifts from elastic scattering. Journal of Computational Physics, 78(2):481–492, October 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900617.

Yakowitz:1986:ECT

[YHS86] S. Yakowitz, K. Hutter, and F. Szidarovszky. Elements of a computational theory for glaciers. Journal of Computational Physics, 66(1):132–150, September 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900574.

Yoshida:1986:NMT

[YNAS86] H. Yoshida, H. Ninomiya, M. Azumi, and S. Seki. Numerical method for tokamak equilibrium with outside limiter. Journal of Computational Physics, 63(2):477–485, April 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186902056.

Yorston:1986:FFT

[Yor86] Ronald M. Yorston. Free-flight time generation in the Monte Carlo simulation of carrier transport in semiconductors. Journal of Computational Physics, 64(1):177–194, May 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999186900240.

Yeung:1988:ATF

[YP88] P. K. Yeung and S. B. Pope. An algorithm for tracking fluid particles in numerical simulations of homogeneous turbulence. Journal of Computational Physics, 79(2):373–416, December 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188900228.

Young:1980:FPM

[YRT80] E. R. Young, P. G. Richards, and D. G. Torr. A flux preserving method of coupling first and second order equations to simulate the flow of plasma between the protonosphere and the ionosphere. Journal of Computational Physics, 38(2):141–156, November 15, 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180900509.

Yee:1985:ITV

[YWH85] H. C. Yee, R. F. Warming, and A. Harten. Implicit total variation diminishing (TVD) schemes for steady-state calculations. Journal of Computational Physics, 57(3):327–360, February 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185901834.

Zabusky:1981:CSM

[Zab81] Norman J. Zabusky. Computational synergetics and mathematical innovation. Journal of Computational Physics, 43(2):195– 249, October 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181901200.

Zalesak:1981:HOZ

[Zal81] Steven T. Zalesak. High order “ZIP” differencing of convective terms. Journal of Computational Physics, 40(2):497–508, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902254.

Zaki:1988:FECb

[ZBG88] S. I. Zaki, T. J. M. Boyd, and L. R. T. Gardner. A finite element code for the simulation of one-dimensional Vlasov plasmas. II. Applications. Journal of Computational Physics, 79(1):200–208, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900113.

Zlatev:1983:SRT

[ZBP83] Z. Zlatev, R. Berkowicz, and L. P. Prahm. Stability restrictions on time-stepsize for numerical integration of first-order partial differential equations. Journal of Computational Physics, 51(1): 1–27, July 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183900797.

Zlatev:1984:IVS

[ZBP84] Zahari Zlatev, Ruwim Berkowicz, and Lars P. Prahm. Implementation of a variable stepsize variable formula method in the timeintegration part of a code for treatment of long-range transport of air pollutants. Journal of Computational Physics, 55(2):278– 301, August 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199918490007X.

Zachary:1986:OAD

[ZC86] Andrew L. Zachary and Bruce I. Cohen. An orbit-averaged Darwin quasi-neutral hybrid code. Journal of Computational Physics, 66(2):469–480, October 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999186900768.

Zebib:1984:CMS

[Zeb84] Abdelfattah Zebib. A Chebyshev method for the solution of boundary value problems. Journal of Computational Physics, 53(3):443–455, March 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999184900706.

Zebib:1987:RSM

[Zeb87] Abdelfattah Zebib. Removal of spurious modes encountered in solving stability problems by spectral methods. Journal of Computational Physics, 70(2):521–525, June 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999187901938.

Zerah:1985:ENM

[Zer85] Gilles Zerah. An efficient Newton’s method for the numerical solution of fluid integral equations. Journal of Computational Physics, 61(2):280–285, November 1985. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999185900877.

Zaki:1988:FECa

[ZGB88] S. I. Zaki, L. R. T. Gardner, and T. J. M. Boyd. A finite element code for the simulation of one-dimensional Vlasov plasmas. I. Theory. Journal of Computational Physics, 79(1):184– 199, November 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188900101.

Zabolitzky:1988:MCS

[ZH88] John G. Zabolitzky and Hans J. Herrmann. Multitasking case study on the Cray-2: The Q2R cellular automaton. Journal of Computational Physics, 76(2):426–447, June 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999188901507.

Ziolkowski:1983:TDC

[ZMC83] Richard W. Ziolkowski, Niel K. Madsen, and Robert C. Carpenter. Three-dimensional computer modeling of electromagnetic fields: A global lookback lattice truncation scheme. Journal of Computational Physics, 50(3):360–408, June 1983. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999183901031.

Zabusky:1983:RCD

[ZO83] Norman J. Zabusky and Edward A. Overman II. Regularization of contour dynamical algorithms. I. tangential regularization. Journal of Computational Physics, 52(2):351–373, November 1983. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999183900359.

Zou:1988:CDE

[ZOWZ88] Q. Zou, E. A. Overman II, H.-M Wu, and N. J. Zabusky. Contour dynamics for the Euler equations: curvature controlled initial node placement and accuracy. Journal of Computational Physics, 78(2):350–368, October 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199918890054X.

Zufiria:1989:LAV

[Zuf89] Juan A. Zufiria. Linear analysis of the vortex-in-cell algorithm applied to Rayleigh–Taylor instability. Journal of Computational Physics, 80(2):387–402, February 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999189901071.

Zang:1982:SMM

[ZWH82] Thomas A. Zang, Yau Shu Wong, and M. Y. Hussaini. Spectral multigrid methods for elliptic equations. Journal of Computational Physics, 48(3):485–501, December 1982. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999182900638.

Zang:1984:SMM

[ZWH84] Thomas A. Zang, Yau Shu Wong, and M. Yousuff Hussaini. Spectral multigrid methods for elliptic equations II. Journal of Computational Physics, 54(3):489–507, June 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999184901293.